A Complete Bibliography of Publications in the Journal of Theoretical Biology: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

25 June 2024 Version 1.01

Title word cross-reference

2 [DLK75]. 2^{k} [Pie72]. 4 [RHA75]. 5 [Abd76, Ami74]. 6 [CS75b]. ^{+} [MHMS75, WD78]. ^{14} [Gar70]. ^{2+} [CT73a]. _{12} [GS77]. _{2} [FSST78, FC70, SP73a]. _{6} [JI78]. _{mem}^{2+} [CT73b]. A(G_{0}) [Eva79]. α [BCDK77, DM78, Jör75b, KF76, Lew72, Par70, UINH73, Zgi75]. β [HM71a, LRH77, VSW^{+}79]. d [DLK75, GS71a]. δ [SBR77b]. Δ^{31} [KBH76]. F_{1} [MHY77]. F_{3} [MHY77]. γ [BS78, Ros78b, Sam77]. i [Thi78]. k [Koi73]. n [Gin77a]. π [BJM70]. t [KV74, Sch72a]. θ [FSST78, SBR77b].

-2 [DLK75]. -amino [Jör75b]. -bond [BJM70]. -chymotrypsin [KF76]. -chymotrypsm [UINH73]. -controlled [Ble71]. -Dipalmitoyl [BCDK77]. -fibrous [Par70]. -fluorouracil [Abd76]. -glutamyl [Sam77]. -helical [DM78, Zgi75]. -helices [Lew72]. -lactam [HM71a]. -methyl [RHA75]. -methylimidazole [Ami74]. -Nucleic [BS79]. -pentene [KBH76]. -phospho [DLK75]. -phosphopropionic [DLK75]. -ray [Ros78b]. -ribose [GS71a]. -thioguanine [CS75b]. -turn [VSW^{+}79].

0Harris [SBC79].

1 [SBB^{+}74d]. 1-179 [Ano70v]. 10-oxy-7 [KGH79]. 10-tetrahydrobenzo [KGH79]. 131 [Ano75b]. 1f [CS77, LM74b]. 1S [Bar77a]. 1v [Bar77a].

2 [Smy79, AFB76, AFDL74, CLR77, MSG^{+}78, SBB^{+}74d, TSP78]. 2’-deacyl [Smy79]. 27 [Ano70d].

3 [NDF79, KGH79]. 3-diformylglycerol [KFH74]. 3-dihydroxy- [KBH76]. 3-diphosphoglycerate [DLK75]. 3-olo-propionic [DLK75]. 3’-terminal [NDF79]. 30 [Ano71x]. 31 [Ano71f]. 35 [LT72]. 371-498 [Ano76y]. 386 [CB79a].

425 [Ano71x]. 48 [Ano75b]. 497 [Ano70d]. 499 [Ano71f].

5 [Jor73a, JA75a, JA75b, KGYG70, NDF79, PGPP78]. 5’- [NDF79]. 5’-cyclic [JA75a, JA75b]. 5’-cyclicmonophosphate [Jor73a]. 5’-nucleotides [PGPP78]. 5’dGMP [LBCV76]. 5’monophosphate [LBCV76].

76 [CB79a].

8-dihydroxy-9 [KGH79].

9th [Ano75a].

A-T [SKS77]. A. [Gol76, Lon76]. aberration [KMSP74]. aberrations [CH73, CP79b, SP73a, SP75, Vig75]. abnormal [Whe75]. absence [Chu74b]. absolute [Bar72b, Bod78]. Absorption [Pap71, Cos75, SR72, Win70b, Win75]. Abstract [Woh78, AC74]. abstraction [vdSJ71]. abundance [Sma76, Web74]. abundances [Lan77]. accelerate [RC77]. acceleration [GG75a, GG75b]. Acceptor [KPU78, Maz78, CNS79, NOB72]. acclimatory [SR77]. accommodates [Ji76]. according [LB77, Tyr70]. account [FSFD70]. accumulation [Bat74, MHMS75]. Accuracy [SM71]. acetic [Ami74]. acetylated [Jör75b]. Acetylation [Jör75b, SC77]. acetylcholine [AFDL74, SG72]. acetylmuramyl [CDS77]. acetylselenocholine [CLR77]. acid [Ami74, AW70, BS79, Ber72, BH77, Bro74, BP79c, CA79, Chu78, CB77, CB79b, CB79a, CD72, DLK75, Dra74, DCMR74, Eng78, FAT77, Fit76, GKYG70, HT73, Hil78b, Jon75, KGYG70, KF76, Koc71b, Lew72, LM75a, LM75b, LT72, Mal79, Man75a, Mik75a, MBG73, Nov73, Paa78, PB78, PM76, PP79, Ril73, Ril74, VPBD78, Van76, Vog77]. acid-base [KF76]. acid-producing [Paa78]. acids [BSH75, DWS76, Gam75, GPB71, GS71a, GS71b, Hai79, HKAT77, IFN72, Jur77, KR76, Les70, MWM76, Ree73, RS78, Sam77, SG71a, Sum79b, TNG74, ZG75]. acquisition [Hew78]. Acrolein [Ala72]. across [AHH^{+}76, CW74, CSVR78, CA77, Dra71, Fle77, GCS72, HTSH72, KVB77, NCM75, NP79, RBP78, Smi73, SB79, TBP76, Vai71b, ZS72]. act [Har78]. ACTH [FY74]. Actin [Lak74, Lym74a, Mor77, GG75a, GG75b, GF77]. acting [Dra70]. Action [MCB77, Str76, BB77b, CS75b, DW71, DW75a, Dra74, Fis73, Gam75, GC76, GT77, GW70, Hil72b, Hul78, Lev74a, MMRE77, MR74b, NOB72, RAJM76, SBB^{+}74b, SK79, SP71, Tsu74, VFK75, Wal74a, WSMS77, Wer76]. actions [Nic76]. Activated [See71b, Sza77]. Activation [Hag76, CT73a, Cot73, GGS78, Got75, Lym74a, OK76c, Smi77, SW74, Wes74a, Har71]. activators [Zem73]. Active [BL75, MMT73, Smi77, AH79a, BJM70, Bla77b, Boh77, Bun78b, Bun78a, CKJ79, CB75b, CS75c, Dra71, Har73a, Hod71a, HS78a, KGYG70, Kan73, LT77, LM74a, MB78, OYT74, RSL79, RAH75, Seg70, TKO74, TD77a, Vai71a, Wun75a, Wun75b, Yu77]. active-site-directed [CB75b]. activité [MCB77]. activities [Kie73, OV78, Wel77b]. activity [Bod78, BJ72, Bra71, CR73, CT73b, Coy79, DB78, Des77a, DHK78, Ger75, GR73, HDP^{+}78, Hol71, HK74, HM79, Hun74, KDH77, Kre70, KL79, LA76, MB78, OSLM79, RB77, Sal74, Tuc77, Tuc79, Ven76, Wal73b, WMHG77]. actomyosin [Lym75a, Lym75b, SRML79]. actual [SBC79]. adaptability [SC75b]. Adaptation [Har70, Sin76, Sin78, CJ76b, Gar74, GC79, Joh76, Lei70, Shu79, Wil78, Wil75b]. adapters [Fer73]. adaptive [Kau76, Kur71, Leó76, Utt70]. adaptors [Les70]. Additional [War74]. Additive [WSSB77, MGB73]. addressed [Nie71]. adenine [MSG^{+}78, MMRE77]. adenosine [AL72, Jor73a, Kan73, Wal73b]. adenosine-3 [Jor73a]. adenylyl [dH76]. adequacy [Hae78]. ADH [DS70]. Adhesion [Mar75a, Mar75b, Ste75, Dol75, Gla79, Goo72b, GF77, Har76a, Mit79a, RMF78, WH72]. Adhesion-guided [Ste75]. adhesiveness [Mar79a]. adjacent [Buy73]. adjustment [CO75]. Adler [Rob78]. administration [Orr79]. admissible [Rou79]. admittance [Che79]. admitting [Les74]. adopting [IMMY79]. adrenergic [Smy72]. adrenocortical [FY74]. adsorption [NKB72, OK76b]. adsorptive [KL79]. adult [Sey77, Sho74b, Sho74a]. adults [NMK78]. advantage [Smi76a, Tay76, Tay79, Vin71, Wil78, Wis77]. advantages [DCMR74, Pul73, Wei74]. advection [OW73]. Aeroacoustics [Bra79]. affect [Wes74b]. affected [Doe71b]. affecting [Bre75b]. afferents [HA72]. affinity [CB75b, Coh70, KM75, LS77]. After [Nel77, LG76, Ord76, Ran76, TR78a, WY75]. After-effects [Nel77]. against [All77, BPF74]. Age [Bed74, BC78b, Abe79, AWD77, Dem76, Gur73, KP77b, MJ76, Rot75, SATT79, Van77, WC76]. age-dependent [Gur73]. age-specific [Rot75]. age-structured [Dem76]. aged [Gat78]. ageing [FY76, Goo77, Kan75, Kro74, SD74]. agents [HZ77, Jok74, Sum79b, TLJ76]. agglutination [Del74]. aggregates [Goo72b]. aggregation [ARW73, Ant74, AMR^{+}75a, AMR^{+}75b, CR71a, CR71b, DP76, DJ78, Dur73, GS78a, Gin73, Goe74, HKM79, JP79, KS70, KP72, Nan73, NS76b, PS78, Ram77, ST79]. aggression [Ber76c]. aggressive [Sou75]. Aging [RR79, DP77a, GY75, Nag78, SHG^{+}71]. agonist [AJ79, Sza77]. agreement [MG77a]. agriculture [All77]. Agrobacterium [Ker75]. aids [TR79]. al [LR74]. alamethicin [JW76]. alanine [CDS77]. alanyl [MPP71]. albumin [KKJ77, SUE76, Sin75]. algae [CP76, Hon73, Sch77a, Shu79]. algal [FG77, Nyh78]. Algebraic [GBH70, BBGR71, BBG70, HR71b, HLR73, HLR74]. algorithm [Ano75b, Moo77, Nin79, Par74a, Ste76b, Whi79]. algorithms [Jea79a, Lin75]. aliphatic [TBP78]. alkylation [IFN72]. all-female [Heu78]. allele [Eds75]. alleles [Gin77a]. allergic [Bac74]. Allosteric [BK77b, CB75a, AG78b, Bar77b, Bar77c, BW78a, Bat72, CB76, Chu74a, Chu74b, Chu78, Coh78, Eps71a, Eps71b, Gar76, GA77, Hua77, KKDY74, KDKY76a, KDKY76b, KDKY76c, Kur77, SFT75, Smi77, SOF72, WD73a, Woe70, Chu74b]. allowance [Eng78]. along [AP77, Dav77, Sim79]. alteration [Arc71]. altering [Tea76]. alternate [Ros76e, ZG75]. alternating [KP79, Sou75]. alternation [Koc74a]. Alternative [Goo73a, Whe75, Jak72b, Lai75, WB77b]. altitudes [Gün77]. altruistic [Cha78b]. ameboid [PB77]. American [Ano78h]. amidases [CDS77]. amines [RAH75]. Amino [BH77, Jon75, Mal79, Ree73, Ril73, SC77, AW70, BSH75, CA79, Chu78, CB77, CB79b, CB79a, DWS76, Dra74, Eng78, FAT77, Fit76, GPB71, HT73, HKAT77, Hil78b, Jör75b, Les70, Mik75a, MBG73, Nov73, PM76, PP79, RS78, Sam77, VPBD78]. aminoacetonitrile [LCC72]. Aminoacyl [BPP70, Rea77, FC76, Sav79, SJ79]. Aminoacyl-transfer [BPP70]. aminoacyl-tRNA [FC76, Sav79, SJ79]. aminoacyladenylates [BCLV76, VBCL76]. aminoacylation [Les78]. amoebae [Haw76, NS76b]. amoeboid [OF75, Sub70]. among [How72, Pfa71, Sey77, Sha77, SBP77, WW77]. AMP [Coh77, Coh78, DW75b, DW75a, Jor73a, JA75a, JA75b, Ols75a, Ols75b, RB77]. AMP-mediated [Ols75a, Ols75b]. amphetamine [HMS75]. amphetamines [DHK78]. amphibian [Mad77a, PM77]. Amphipathicity [DJ78]. amphoteric [CW74]. Amplification [DW71, CFV75]. amplifications [Doy78]. amplifying [AA76]. amplitude [LM74a, SP71, WB77b]. anaerobic [Hor74]. analog [Cri77, Dra74, Jor73a, MFS77, Rap77]. analogous [Orr79]. analogs [Bye78]. analogue [Sch79d, SK72]. analogues [GKYG70, KGYG70, LR76, Smi77, GKYG70]. analyses [Hjo70]. Analysing [CL70, GB77b, JBM^{+}78]. Analysis [AW70, BK77b, Gin72, HA72, IC79, LB77, Mac77d, MSF71, MK74, Nea72, NW73, Orl70, SH79, Sin75, SBR77b, Sta76, SW74, Ste70, SHC71, Van77, WSL79, AJ79, AH79b, AE79, Bak76, BL71, BSM77, BH79, BDS73, BG75, BHK78, BCLV76, Bun78b, CA79, CLR77, CB75a, CB76, Cho72, Cla76, Cla77b, Coo73, Cou78, CS75c, DLK75, DO76, DS73b, DD77, DST79, ES77, EMM79, FSST78, Fin76b, Fis77, Fit73, FK79, FC71, Gar76, GK73b, GKS78, Gri77, HM71a, HM71b, HM72, IMMY79, Jon71, JBM^{+}78, KOH70, KS71b, KBW75, KOWG70, KM75, Kit75, KWG79, KRH77, KDKY76a, KDKY76b, KDKY76c, Kur77, LH79a, Lav76, LBCV76, Lev72, Lin79b, Lin79c, Lub73, Lym74a, MHMS75, MWM76, Mac73, Mad76, MZ79, Măr76a, Mil74, MJ76, Mob73, MA71, NG77]. analysis [Nol73, Nos77, OO77, Ott74, PMC78, PTB74, Ran75, Ran77, Rap76, Rob78, RZMF75, Rot79, Sav70, Sch79d, Ses72, SH72b, SATT79, SB78a, STD77, SW78, SF79, SB78b, SOF72, Sue78, Sug75, Tho78, TV75, UF71, VBCL76, VS71, WH73a, WL76, YZI77, ZW70]. Analytic [Com79]. Analytical [KV74, Ros75a, Ros75b, Shi78a, DC77a, KH79a, Mar79b, Tay79]. analyzing [SG73]. ancestral [MBG73]. Androctonus [SBB^{+}74b]. anemia [VP77]. angiosperm [Sch79d]. angiosperms [BH78]. angiotensin [GK74]. angle [GPP77a, GPP77b, Hon71, LG76]. anhydrase [Wer76]. animal [BG74, CO75, CZ76, Iwa78, Joh76, Jon72, Kle72, Niv77, Par73a, Par75d, Par75e, Pre73, Sch73a, Smi74, Str71, SHC71, VtM79]. animals [Bak76, Mac78, Sey72]. anionic [WZ72]. anions [RBP78]. anisogamy [Bel78, Cha78a, Par78]. anisotropic [GL70, MJ76, MNR72]. anisotropy [Pap73a]. Announcemen [Ano70a, Ano75c]. Announcement [Ano70b, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano73a, Ano74a, Ano74b, Ano75d, Ano77a, Ano77b]. Announcements [Ano70c, Ano71e]. anode [RP76]. antagonistic [Sza77]. antibiotic [Hua71a, Hua71b, Hua73a]. antibiotics [HM71a, Hod71a, Kre71]. antibodies [Coh70, HO76, Ros76c]. Antibody [DLV77b, BY74, Bel70, Bel71a, Bel71b, BGKS76, CNS79, Coh70, Ebr75a, Moo71, Nov75, PO79, Teo73, WB77a]. anticodon [FIN70]. anticonvulsive [Vog77]. anticooperative [Har73a]. antidiuretic [DS70]. antigen [CNS79, Del75a, DLV77b, GGS78, GDA75, WB77a]. antigen-Antibody [DLV77b, CNS79, WB77a]. antigenic [AP70]. antigens [MM79]. antimycine [MCB77]. antipodes [See71a]. antiserum [GDA75]. antiviral [GR73]. antocatalytic [HSK72]. Antonelli [Ant74]. ants [Tay78]. any [MMS74, MG77a, TD76b]. aorta [MJ76, RZMF75]. apects [Tho77a]. apex [Rob78]. apical [Sac72, Tho77b]. apoferritin [Fis76]. apomorphine [Nic76]. apoplasm [Mol76]. apparatus [Pai72]. apparent [BP76, SKTC78]. appearance [BPF74, Fis73]. applicability [HLR73, HLR74, NRW72, NRW73, Rei78]. applicable [Tho77b]. Application [CNS79, Dun74, MRS77, Ste76a, Str77, Swa74, TV75, YS71b, Adl77a, Ain77a, AW70, Bak76, BB77b, BFM75, FPB72, Fri71, Fri76a, Gar70, GKS78, HKK76, IHI79, Jea79a, Kel77, KBW75, LH78b, LT77, MMS74, Man78, MHB79, MG77b, NB72, RCW73, Ste76b, VRM77, War70a, Yag76, Yoc74, Che79, Cra73, HH77b, MT78b]. applications [Gro78, HLR73, HLR74, Mik77, RK72, War74, Bre79]. applied [Ben78a, Dun74, HK76, JS79, Mac77a, Pur77b, RT77]. appraisal [VtM79]. approach [Adl70, AHH^{+}76, Che73a, EMM79, EGF77, Fin72, GL78, Got75, GR74, HTSH72, HS79b, Jak72b, Jea79b, Jon75, Kau74, LD73, Lau72, Lay78, MC73, Mob73, MGB73, PBV74, Per73, SM78, Sum79a, Tri75, VAC76, Vol72b, WL72, Wes75, WMHG77, Wu77, dLA76]. approaches [McC78, Nir75, Orl79b, Yos77]. appropriate [GR75]. Approximate [KH79a, ABR72]. approximation [KD76, OD73, Sav70, SH79, VB70, Wal74a, ZS72]. approximations [NRW72, NRW73, RRS72]. April [Ano70m, Ano71l, Ano72j, Ano73l, Ano74s, Ano75w, Ano76w, Ano77-27, Ano77-34, Ano78w, Ano78-30, Ano79w, Ano79-29]. Apter [Vis77]. aquatic [WO71]. aqueous [GP72, Hvi75]. ara [GR73, MMRE77]. ara-adenine [MMRE77]. ara-C [GR73]. arbitrarily [Boh77]. arbitrary [BBT70, MO73]. arbres [Hyv73]. archaeorelation [Mik75a]. architectural [HR79]. architecture [BRS79]. Area [Dia74, CSL75]. areal [Coo78a]. arginine [LSL77]. arising [Ain75, Cla72]. arm [Cli77]. aromatic [GPB71, HS76, KGH79]. arrangement [Fis76]. Arrays [SMN73, LH78a, NS76a]. Arrhenius [Han72, KRL71]. arsenic [Ros71]. ART [GBH70, BBGR71, BBG70, CK71, HR71b, HLR73, HLR74]. artefact [Sto76]. arterial [KGN76, MTL76, NP79, Zam76]. Arthropoda [KT72a]. article [Bro77]. artificial [AR75, CP70, KH79b, Lai75, Lin79a]. aryloxy [Gam75]. ascites [KBW75, WZ72, WNHS75]. asocial [BP70a, BP70b]. aspartate [Smi77]. aspartic [Lew72]. aspect [Ran73, RL76]. aspects [Blu76, BP79c, Bun78a, CT73a, CT73b, CC75d, DW75b, DW75a, DC77a, Jon72, KC78, Kem74, Mac79, Mag73, Mag74a, Mic79, Mil73, Nol76, OS74, Pap73c, PB78, Rei70, Ros73a, vDHD78, vHB76]. assay [NB72, NW73, Nos77]. assays [KNJ74]. assemblies [LBM70, Ros79a]. assembly [And77, Hae78, Ste75, WL72]. Assessment [CB77, Par74c, Bri78, Lei77, Woo78, HR79]. assimilation [Bat74, Chi76, Kow71]. assisted [DUK73a]. associated [Che79, Ede70b, Ede72, Naz77, Ros75c]. associates [KD76]. associating [KDKY76b, KDKY76c, Ste74]. association [Got72, Kur77, MRTG79, SKS77, Ste70, WD78]. Associative [Tre79]. Assortment [Mar70, McC78]. assumes [Tre76]. assuming [CR75b]. assumption [Shi79, Sim73b]. assumptions [BC77, RF73]. asymbiotic [SP73b]. Asymmetric [Ste72, Eas79, MM78b, See71a, See72]. Asymmetry [Tas70, Age72, BL75, KD72]. asymptotic [De 75, Goo70]. Asynchronous [WD73b]. atherosclerosis [Pap74]. atom [DKU77]. atomic [KD76]. atoms [TBP78]. ATP [AW75, BV70, Boy73, DeV76, GG75a, GG75b, HKAT77, Jon72, NR70]. ATPase [Lym74a, Lym75a, Lym75b, Yag77]. atropine [WSMS77]. attachment [Lew70a, Vig75]. attainable [AF77]. attempt [GKP75, Thi78]. Attempted [War70a]. attendance [SATT79]. attention [Coh70, Ste75]. attentional [GL75]. attractant [LS75, Ram77, SS72]. attraction [DM75]. attributed [Abe79]. attrition [BC78a, BCS78, NTR77]. auditory [GH75, HJ70]. August [Ano70n, Ano71q, Ano72p, Ano73p, Ano73t, Ano74w, Ano75x, Ano76v, Ano77x, Ano77-33, Ano78z, Ano78-31, Ano79x, Ano79-30]. australis [SBB^{+}74b]. Autocatalysis [Ede71a]. autocatalytic [CGM79]. autointerfering [MYF77]. autolytic [Tho71]. automata [Ros75e, Sug75]. Automated [BG75]. Automatic [Tem75]. Automaton [Lai77]. autonomous [Ram76]. autoradiographic [BG75]. autoradiography [SM78]. autoregulation [WKG73, Whe75]. Autowave [ZZ73]. availability [Cho77]. available [WO71]. avalanche [BM79]. average [Man78]. Averaging [LK76]. avian [SW78]. avoidance [Coh76, How74]. Avoiding [Ben78b]. Axes [McG70]. axon [AO72, AR74, Got75, Gri79, Has78a, Mad77b, RP73, SS70, Sta73b, Sta73c, Sta76, Str76]. axons [Lun74, RP76]. axoplasm [Doe71b]. axoplasmic [Ode76]. Azacyclopropenylidenimine [YMM76].

B [CS75b, GGS78, SBB^{+}74d, JFR76, MCB77, BD75b, GS77, JI78]. B-DNA [CS75b]. B. [Ano71x]. B/r [BD75b]. bacillaceae [Hod71a]. Bacillus [EGF77]. back [MR74a, RC77]. backbone [JW76, PM76, SG71a, TNG74, VBCL76]. background [Rei70]. bacteria [Bre75b, BC77, BC78b, BCY79, Coo74, GWZR77, HS78b, KS71b, Koc70, Koc71b, LK70, Mil70, OK76a, Pai76c, Ros75a, Ros75b, RGZW78, SSR74, WP72, Zar75]. Bacterial [LS75, Sim73a, Ada77, Ano70d, BC76, Cal76a, CDS77, CGM79, CBS75, Doe72, EJB^{+}78, Goo70, GD79, Har74, HDP^{+}78, Hit76, Lef78, LD75, MH74, NW73, OF77, Pre70, RS77, Rig79b, Ros73a, Rös76f, Seg76, Tho71, TD77b, dL77]. bacterio [Ott74]. bacteriological [Per73]. bacteriophage [EF76]. bacteriophages [HS70]. bacterium [BD75b]. bacteroids [Paa78]. balance [DH78, TC79]. band [Ros76b]. bands [Ciu77, KS71b, OK76a, Ros75a, Ros75b, SSR74]. Banks [Whi71]. barbiturates [PCC72]. barley [SBC79]. barrier [Jor73a, Jor73b]. Barriers [RHA75]. basal [Eco79, Rud79]. Base [FSG^{+}72, CP79a, DZOP71, Elt74, FIN70, IFN72, KF76, Les73, Les74, MSG^{+}78, McG79, PB78, PSK76, RE76, SK79, YC77]. based [BY74, BKB76, CE78, Coo78a, DS73a, Doy78, Eds75, Eng78, Fal77, Har73a, Hut70, Lan77, LH79b, MP78, Orl70, PB77, Utt77a, Van77, Yu77]. bases [Ber72, KD76, PPS75]. basic [CT73a, Con71b, JK72, Pap73b, Sco79b, Wes74b, And76a, BNT71]. basis [Arc71, Bel71b, Bun78b, Bun78a, Don72, LP75, Lev74a, Lew70a, Mak76b, Mat78, MMRE77, PL77, RB77, Rot73, Sac72, Seb79, Spe70]. batch [Pai76c]. Batesian [CC75a, CC75b, CC75c]. bats [HP77]. bay [RN76]. Bayesian [EMM79]. be [AF77, Ber77b, CJ76a, Man79b, TS78a, Wen77, Yoc78]. bearing [CW74, NP71]. beating [Lin79c]. bee [Ros78e]. beetle [Bra75, LBGK76, Niv77]. Behavior [AG78b, ON77, Bea79, Cha78c, Coh76, DC73, Doe72, GK71, HL70, JS79, LK77a, SIS78, SG75a, SG75b, Tag78, Tay78, TR78a, Wal70]. behaviour [AG75, Ale72, Bar76, BM79, BC73b, CO75, Cha78b, Dol75, Eij70, Gil73c, Har73b, HR79, Jon72, Kit75, KDKY76a, KDKY76b, KDKY76c, Kur77, MGNR78, OYT74, Par74c, Sch79c, See71c, Sin78, Tay72, Utt77a, Utt77b, Vin71]. behaviour-movement [Dol75]. behavioural [CL70, Lüc75, vdSJ71]. being [BBGR71]. Bellman [SBC79]. Belousov [Mur76, TIO77]. Belousov-Zhabotinsky [TIO77]. bend [LB71, LB77]. Bending [RS70, Can70]. benefit [Hea77]. benign [Kau71]. beryllium [JFR76]. bessel [Shi71]. between [Abe73, BW78b, Bec78b, Ber72, Blu77, BC78b, BCY79, Cas74, CW74, Cha78b, Che76, Che73b, Chu74b, CMSC77b, CC70, CB79b, CB79a, CP79b, Dar73b, Dar75, DW71, Der78, DUK73a, Dra70, DST79, FAT77, FJThH73, Gar76, Gau73, GD79, HM71a, Har73a, HS79a, HM79, KS77a, KVB77, KDKY76b, KDKY76c, LT75, Les70, LS78, MS71, MFS77, May74, MNR72, NDF79, Nic76, NP71, Nir75, Ott74, Pai76a, PR76, RE76, SYNY76, Sin73, SMN73, SBB^{+}74c, SBB^{+}74d, SBB^{+}76, SD75, SM77, UIN74, WNHS75, Wes75, WD73a, WD78]. Beyond [HS79b]. bi [YK71]. bi-ionic [YK71]. bias [SZ71]. biased [Man79b]. biases [GL75]. biconcave [Can70]. Bidirectional [Din74]. biexpone [CSL75]. Bifurcation [Has78a, Sib78, So79b, Tho77b, ZR73]. bilayer [Ohk70, OO76]. bilayers [MM78b, Sco74, TW76]. bile [MWM76]. billing [Vas74]. Bimodality [Gat78]. binary [AK78, FC76]. binder [Mag74b]. Binding [Bye78, CNS79, NOB72, AG78a, BNW78, Bar77b, BW78b, Bro74, BGKS76, CB76, Cho77, Chu74a, Chu78, GGS78, KO78, MP78, MLM77, Mor72, NWW79, Rak78b, Rap76, Rap77, Rap79, ST78b, Smy79, SOF72, Ste74, Via71, VSW^{+}79, WD73a, YC77]. Binocular [LBM70, BL71, Nel75]. Binomial [Pal70, Cou78, Sch79b, WB74a]. bioassays [ASS79]. Biochemical [Ano70d, Ede70a, Pre70, Sav70, BC73b, Cum75, Duc77, GK73b, Har73b, Has78b, Hul71, KWG79, Mac77a, Mac77c, OF77, Par75a, Pav71, Wal70, Yag76, Yca74]. biochemistry [Ano75a]. bioconvection [PWW76]. biocybernetic [KOWG70]. biodegradation [UI73]. biodynamics [Mit79b]. bioenergetics [Chr76, Tri75]. biogenesis [And77, Yoc77a]. biogenic [NB72]. Biol [Ano70d, Ano71f, Ano71x, Ano75b, CB79a]. biologic [Lei77]. Biological [Blu73, Der77, Fin72, Fre73, Ger78, May75, Nic73, RD77a, Tul71, Wic76, Abe79, Adl77a, Bes70, BC73a, Bol73, Bor74, CA77, Cil75b, Com75, Con74b, Dia76, DKU77, Ede71a, Els72, Fab76, GT77, Gla78, Gla75, Gou76, Gre76a, Hai79, HZ77, HM79, Ima76, IMMY79, JK72, KJ72, Kau76, Koc71a, Koc71b, KH79b, Ler78, Lew70b, LK76, Lin79b, Lin79c, LW79b, Man78, MN73, NPV77, Olo73, Oos73, RB77, RT77, RBP78, Sco74, SFT75, Sme76, Spe78, Sug75, The71, TBP76, Tho77a, Tri75, WA78, Yag76, You76]. biologically [Hod71a]. Biology [vdSJ71, Ber77a, CM70, DT74b, Rös76f, Van72, Yoc79]. biomathematics [Ano78h]. biomechanical [HP77]. biomembranes [GGM75]. biomolecules [Han78]. biophysical [BH78, CCY^{+}75]. biopolymer [Les77]. biopolymers [Sch72c]. Biosynthesis [Rus72, Rus74, AHH^{+}76, CBS75, GS77, Koc70, MKLKV76, PBV74]. biosynthetic [Sav79, SJ79]. biotin [Mag73, Mag74a]. biotins [Mag73, Mag74a]. Bird [SV75, Hjo70, TV75]. birds [AP77, Ale79, Bra79, Cal79, SK72, Wic76, Yor79]. Birfurcation [Mac77a]. bistable [See71c]. bisubstrate [Fal77]. bithorax [HGR79, Kig73]. bivalent [CNS79]. biwae [Hon73]. black [MP73, Sue78]. bladder [Civ70]. blakesleeanus [Coh76]. bleomycin [MMT73, MMT76]. blocking [TLJ76]. Blood [ZR73, BLP76, Can70, KOWG70, KS71c, LS72a, Mil73, Ned75, Nel78, Ols79, RK72, Win70b]. Board [Ano70e, Ano70f, Ano70g, Ano70h, Ano71g, Ano71h, Ano71i, Ano71j, Ano72f, Ano72g, Ano72h, Ano72i, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. bodies [Nir75]. Body [Cal76b, Gün77, Kle72, Blu77, Der78, Hon71, How72, Kle75a, Lei77, New79, SMN73, Syn79, WP71]. Bohr [Bla75b]. Boltzmann [Bar79b, Mar79b, She73b, Shi78a]. bolus [LTP78]. Bond [HPCH79, Ami74, BJM70, GPP77a, GPP77b, McC72a, TSP78, Zem73]. bonded [CP79a, OO77]. bonding [HK76]. bonds [CSVR78, MSG^{+}78, WP78]. Bone [Huk78, HC76, McC75b, UI73]. bonlinear [WB74b]. Boolean [OB73, Tho73b, Tho74b]. bordered [CRC77]. Bose [WA78]. both [Bat73, Wel77b]. Botts [Kat72]. botulinum [HL79]. bound [Cas74, Hea74, Wel77b]. boundary [CAA79, HOR73, Off71]. bounds [Ada76]. bovine [BS75]. Br [BHD74]. Brain [HCBL70, DWS76, KMS77, LBM70, Pit78, Rap78, Tre75]. branch [Hon71]. branchial [Jon71]. Branching [BRS79, BCH73, Hon71, Leo71, Zam76]. break [RP76]. breaking [Blu71]. breaks [KRL71, Res76, Res78]. breeding [Ran73]. bridge [BB73a, DUK73a, DUK73b, Koi73, Sch73b]. bridge-assisted [DUK73a]. Brillouin [BGK77]. broken [Sma76]. broken-stick [Sma76]. Brooding [Hea79]. broods [Heu78]. Brownian [GCS72, Man79b]. BUdR [Eld79a]. buffer [IMMY79]. buffers [WP71]. building [HKK76]. bulk [Civ70]. bundles [MMS74]. burrows [WK78]. burst [Wei74]. butterflies [KD79]. butterfly [Heu78]. buzz [BH78].

C [GR73, ABCT75, FSG^{+}72, Yoc77b, Gar70, GPP74, Jor73a, Jor73b, VSW^{+}79]. Ca [CT73a, CT73b]. cable [Koo77, Mad77b, Ols79, Shi71, SB75b]. cables [GF77]. Cachexia [Ste78]. calcitonin [War74]. Calcium [VSW^{+}79, Dea75, JBM^{+}78, RB77, SM71]. calcium-cyclic [RB77]. calculated [TS78a]. calculating [Bla75b, GRY^{+}72]. Calculation [FRG74, BM73, MSG^{+}78, PW79, See73a, Yoc77a, SG71a, TNG74]. Calculations [GCS72, Ami74, Bla74, Bla75a, DHK78, GPP73, GPP74, GS71a, GS71b, HK74, Jor73b, KGH79, LT72, MPD70, MPP70, MPP71, SG72, SB79]. calf [Tem75]. calibration [SR72]. caloric [Leb73]. camin [NSS74]. camin-sokal [NSS74]. Can [CJ76a, Wen77, Yoc78, TS78a]. canavanine [Ros78d]. cancellation [Whi79]. cancer [Blu74a, Blu70, BB73a, Bre79, Duc77, DMRD78, GL78, Har70, Orl77, Ste78]. cancers [Sum79a]. Canid [Tod70]. cannibalism [Bra75]. Canonical [TF74]. capacitance [Mar76b, WB74b]. capacitor [IMMY79]. capacity [CJ76a, Cra71, Cra73, Eld79b, Hae78, LH78b]. Capillary [Lev72, MT74a, Nos77]. Capillary-tissue [Lev72]. capping [Hew79, RMF78, Ros79b]. capsulatus [HWD76]. Capture [Pai76a, Pai72]. carbon [TBP78]. Carbonic [Wer76]. Carboxylated [Mag74a]. carbylases [FC70]. carcinogen [Ros71]. carcinogenesis [Bec72, HR75, NM72, Ros77b, Tar72]. carcinogenic [BJM70]. carcinogenicity [KGH79]. cardiac [CKJ79, HS77b, Jon71, MHMS75, MTL76]. cardiovascular [MTL76]. care [Shi78b]. Carlo [BHK78, Cla77b, NKB72, TR78a]. carotenoid [BY75]. carrier [BP73, BP74, Hua71a, Hua71b, Hua73a, LS71, LS72b, Ste76b, YSK^{+}72]. carrier-mediated [LS72b]. carrying [GK71]. cartilage [Kem79, WE75]. carve [CSL75]. cascade [WK73a]. Case [Mar72b, All77, Bec78a, Coo77, Cra71, Cra73, Hon78, Ish77, MWS77]. cases [GH76]. castrated [Sho74b, Sho74a]. catabolite [Hed71]. catalase [Kre70]. catalases [JD77]. catalysed [Ain74b, Ain74a, Ain75, BW76, Dug79, GG75a, GG75b, Kan73, KNJ74, Mal76]. catalysis [Ain77a, AH79a, Con79, Pog71, PA74, See71b, Tho74a]. catalyst [LT72]. catalytic [Ain75, LT72]. catalyzed [AG78b, Van76, Zem73, Zer75]. catastrophe [BK76, BK77a, GY75, GI77, PM77, VRM77]. catastrophic [Ain79]. catenanes [Dou72a]. catenary [She71]. cathodes [Alb71]. cation [HKM79, RL77]. cation-induced [HKM79]. cations [DST79, Got75, RBP78, SO79a]. causality [vdSJ71]. causative [GMR75]. caused [Har76a]. causing [Sch79a]. cavity [Nel78]. Cell [DG77, DK76, Ede71b, JR77, Ond79, RG78, Ala72, Alb77, And77, And79, Ano70d, ARW73, AMR^{+}75a, AMR^{+}75b, Bab77, Bar74c, Bec78a, Bec78b, Bec72, Bha77, Bla78b, Blu74a, BB73b, BKB76, Bre79, BM79, BK71, Buy73, Can70, CE78, Con71a, Cra77, DS73b, Dol75, Dra71, DP74, DC77c, Ehr79, FGP^{+}75, FD73, Fra70, Fri72, GL78, Goe74, Goo72b, Goo73a, GGSW72, Got72, GB77b, Gre77a, Gre76b, Gre77b, GF77, Gre78, Hal73, Har76a, Har75, Hea74, HO74, Hew79, Höf71, HMT79, Hua73b, IE76, IHI79, Jon72, Kem79, Kig73, KBW75, KP77b, KOWG70, KV77, KRH77, Kub70, LH78a, LS76, Ler78, LS77, Lin76, LSGFdlT^{+}75, LM74b, Mak76b, MF79, May74, McE78, Mei78, MGNR78, NG77, NW74, Ols75a, Ols75b, Ost77, Pap74, Pas76, Pat70, PW79]. cell [PWW76, Pow75, Pre70, PR77, RMF78, Ram76, Ran77, RSS^{+}73, RC77, Ros79b, Rot79, RP75, RMG76, Rud79, SH72a, Seb79, Sha73, Sha74, Sha77, SS79, Sim73a, SS73b, Ste75, TW77, TL74, Tho71, TD72, TS71b, VP77, WK75, Whi78, WT79, WRT79, Wig72, Wig75, WD73b, YZI77, ZW70, ZV77, vW76, Sha73]. cell-cell [Bab77, BM79]. cell-growth [Ala72]. Cell-sorting [RG78]. Cell-specific [Ond79]. cells [Alb75, Alb77, AB71, AHH^{+}76, AWD77, BW77, Bon78a, Che73b, CV76, CP79b, Cra73, DBS^{+}70, GM72, GL70, GCG^{+}70, Goo72b, GD79, GCS72, Has74, HO78, Hew78, HTSH72, HH77a, JA75a, JA75b, Kau71, Kra78, Lei70, LG71, Ler78, MYF77, MFS77, Mar75a, Mar75b, Mil73, Nan76, NMK78, Ohs77, Ond79, Pin72, Pur76b, Ran76, Res78, Rig79b, RG78, Rot77b, Rub77, SS72, Shi71, SM77, SR72, SD76, Tor76, Utt77a, Utt77b, WZ72, WH72, WY75, WNHS75, Whe75, Wil70a, WG70]. cellsim [SD76]. Cellular [Ram77, SG76, Wig75, Ale72, And76b, Bel71b, BC73b, CR71a, CR71b, CP72, DW75b, DW75a, Doe76, DH78, Ede72, EFH77, Erh76, GK72, GJ76, Har70, HS77a, Hon78, HH73, JP79, Kro74, LT77, LBM70, Lin71, Mar71a, Mar72a, Mit79b, Nan76, Nos77, OS71, OS74, Pur77b, RD77a, RG78, Sch70, Ste72, SD74, Wel77a, Woo72, Coh77]. cent [Sha74]. center [DeV76]. Central [Jon72, TS71a, BW78b, Ber77a, CM78, Ebr75b, GM71, WSMS77, Yoc74, Yoc78]. centre [MMT73, Ott74]. centriole [Har75, Wen77]. centromere [Ima75, Ima76]. cereal [LBGK76]. cerebral [Bar72b, Bra74, Lau72]. cerevisiae [KD72]. certain [Arc71, Che75, HK76, MS71, McN79, NRW72, NRW73, SG75a, SG75b]. Certainty [MHB79]. Chain [Dav71, Arc71, BPF74, BHD74, Bro74, Cla72, FP73, HC78, Jon75, Kem74, NPV77, RP77, So79b, Sto70, Sto71]. chain-terminating [Cla72]. chains [Bel77a, Boy79b, Den75c, GC73, GR74, GGM75, JDTF73, Nov73, RD77b, SB75a, Sim79, TV75]. chamber [Bak76]. change [Gin73, GKS78, KM75, KRL71, KDKY76a, Wes78, WD73a]. Changes [RP73, Ain79, Bec78a, Bla79, Cla77a, Doe71b, DH78, DC77c, GL73, Hol76, Lat75, LSB78, MJ71, SA70, SM77, Sta73b, Sta73c]. Changeux [KDKY76a]. changing [GB77b]. channel [Che79, SBB^{+}74a, Sta73c]. channels [Che79, CA77, KM75, dO79]. chaos [May75]. Character [CAA79, TD76a, HE73]. characteristic [Den75b]. characteristics [Coh70, GM71]. characterized [PR76]. characters [CAA79, Sla78]. charge [Ami74, Bla79, Buy73, KS77a, KF76, Lav71, MM78b, OV78, SO79a, TBP76]. charged [Buy73, DST79, Mar75a, Mar75b, Mar77, MW78, RBP78, SO79a]. charges [LS78, NM75]. Chart [Ses72]. Cheating [Ros78a]. check [CGM77]. Chemical [Ano71f, BN72, Boy79a, Bun78b, Bun78a, Cze71a, Cze71b, HOR73, LNHK71, LN71, McC71, Sch72c, Sto70, Sto71, AG78b, AG79, Ano72v, Blu75, GPP77a, GPP77b, GS71a, GS71b, HSK72, HS78b, Hyv72, Ish73, IM75, Ish77, Ji76, Jor74, Kei75a, Kei75b, KMS77, LD73, LRB71, Loe71, LCC72, Mac76, Man78, Mar72b, NM72, Per76b, PGPP78, PP79, SG71a, Sch79c, Seb79, See70, See71b, See71c, SIS78, She73a, Smo76, Str77, Sug75, TNG74, Vai71b, Vai71a, ZZ73]. chemically [HM79, KK78]. chemicals [Ord76]. chemico [Sug75]. chemico-physical [Sug75]. chemicomechanical [Kou73]. chemiosmosis [Bun78b]. chemiosmotic [Ji76, RS76]. chemistry [Mag73, Mag74a]. chemo [KS77b]. chemo-mechanical [KS77b]. chemoreception [WMHG77, vdSJ71]. chemoreceptor [Ros76d]. chemostat [TL74]. chemotaaris [Dol75]. Chemotactic [Ros76a, KS71b, OK76a, Ros75a, Ros75b, Ros75c, Ros76b, SSR74, SS72]. Chemotaxis [CR71a, Nan73, HS78b, KS71a, LS75, LK79, LD75, NW73, OF77, Ros73a, Rös76f, Seg76]. Chen [Fin73]. chi [SB78a]. chiasma [Mag74b]. chick [WE75]. chimaeric [Wes75, WG70]. chiralities [Har73b]. chirping [Sou75]. Chitin [GMR75]. Chlamydomonas [Mar70]. chloramphenicol [Coo74]. chloroleukemia [TR78a]. chlorophyll [Men76, Ott74]. chlorophylls [JFR76]. chloroplasts [DST79, Mar70, Ost77]. chloroquine [Che76]. Chlorpromazine [Kie73]. Choice [You76, Cha78c]. Cholesterol [Pap74, FK75, Pat70]. cholinergic [ABD^{+}72, CT73a]. Choosing [GR75]. chorismate [AH79a]. chromaffin [HKM79]. chromatic [Bri78]. chromatid [SP73a]. Chromatin [RD77b, Vig75, New79]. Chromatin-nuclear [Vig75]. chromatographic [MWM76]. chromosomal [CP79b, Mar70, SP75]. Chromosome [KMSP74, CH73, Cli77, Din74, Doy78, NMK78, Vig75]. chromosomes [And79, BBZ79, DCT73, Eds75, Ger75, Gra73, Hol72, Ima75, Ima76, Ima78, Leó75b]. chronic [HKK76]. chymotrypsin [CC70, KF76]. chymotrypsm [UINH73]. cilia [Bla74, Bla75a, Blu71, LB77, RS70]. ciliary [Bla77b, McK73, VRM77]. Ciliata [WJ74]. ciliated [Blu74c, McK73]. ciliates [RL76]. circadian [Ber76a, Cum75, DB78, Pav71, SB78a, TE76, Win70a, Win72]. circuit [SFN77, Sug75, Yag76]. circuits [Con74a, Kau74, Tho73b, Tho74b]. Circular [Boh77, AH79b, Bec78c, Dou72a, GDA75]. circular-linear [Bec78c]. circulation [KGN76, She71]. circulatory [KKW78]. cistron [CE77]. CL [KV74]. cladistics [NSS74]. clamp [AR74]. clamps [Sta76]. class [BM79, Che75, Mur75, WC75]. classes [MLM77, WW77]. classical [BT71a, Utt75a, Utt75b]. Classification [Gla75]. classifications [HE73, VtM79]. clearances [HS77b]. cleavage [CE78, Doe76, Eld79a, Gre77b]. cleft [MFS77, SM77]. clefts [MHMS75]. CLI [Sch72a]. clinic [BB73a]. clinic-laboratory [BB73a]. CLM [KV74]. CLMII [Sch72a]. cloacal [SK72]. clock [CZ76, Cum75, Hej75, Win70a]. clocks [Kau76, Pav71]. clonal [Bel70, Bel71a, Bel71b, Lew73, PO79]. clone [Wes75]. clones [AA78, Lei78]. close [Adl77a]. closed [Bla77a, Has78b, May73, Ula72]. clustered [Lev74a]. clustering [Pre73, Kir79b]. CNDO [AFB76, AFDL74, CLR77, GS71b, JFR76, MSG^{+}78, SG72, TSP78, TNG74]. CNDO/2 [AFB76, AFDL74, CLR77, MSG^{+}78, TSP78]. CNS [Mik75b, Mik78]. Co [GK72, KK78, KO78, Lav71, Mar72a, MN73, Rak77, Rak78a, RBP78, TT75, BW78c, BW78a, BW78b, Bla78b, CB75c, KKDY74, LR76, Lev74a, Mac77b, NKB72, Rak78b, Sch72c, ST78b, Ste76b, VFKL74, FC70, Mit79b, Par77, Sam71]. co-operation [TT75]. Co-operative [GK72, KK78, KO78, Lav71, Rak77, Rak78a, TT75, LR76, NKB72, Sch72c, ST78b, VFKL74]. Co-operativity [Mar72a, MN73, BW78c, BW78a, BW78b, CB75c, KK78, KKDY74, LR76, Lev74a, Rak78b]. co-ordinate [Bla78b, Mac77b]. Co-transport [RBP78, Ste76b]. coastlines [AP77]. cochlear [Nie71]. code [AFS^{+}79, Bar77d, BPP70, Ebr75b, Eng78, FC76, FIN70, Jur77, Klä75, Les70, Mik75a, Pap73c, Sch71, Tsu79]. code-words [BPP70]. coding [HH73, Wri77]. codings [MM79]. Codon [Hil78b, SHG^{+}71, Woe70, GRY^{+}72]. Codon-restriction [SHG^{+}71]. coefficient [HH77c]. Coefficients [OW78, Che75, Jor75a, LS76, LT77, Mag70, Orl75a, Orl75b]. coenobial [Hon73]. Coenzyme [Bre75a, Bro74]. coenzymes [UIN74]. Coexistence [AM76, Koc74a, Cro79, Eps79, Koc74b, PG79]. COFLM [KRH77]. cognitive [NB76, Tre77]. coherent [AK78]. cohesion [GJ76]. cohort [KRH77, ZV77]. cold [DP74]. cold-induced [DP74]. Cole [Alb74]. Cole-Moore [Alb74]. coli [BHD74, BD75b, Coo70, DT79, Lan78, LK77b, Man79a]. collagen [CA79, Huk78, McC75b, ST79]. collagen-like [CA79]. collagenic [RZMF75]. collagenous [HC76]. Collective [Kem74, Mar71a]. colligative [Ste70]. Colloid [Rei70, Rei79]. colloidal [Fis77]. colonies [WD73b]. colony [WC76]. Color [LSB78, Ric75]. columns [HMT79]. combinations [FP72]. Combinatorial [Cho72, Les74]. Combined [Ale79, MAM79, Ste70]. come [BP79b]. Cometabolism [HK77b]. coming [DD77]. comme [MP79]. Comment [APK75, Che73a, Gol76, Lak75, AFDL74, Bar72a, Ber77a, BHK74, CM72, Ebr75b, ER74, Fin73, Fre73, Hjo70, May73, Ros79a, Smi78b]. commentary [Ste75]. Comments [Goe74, HSC73, LR74, Ano73g, Ant74, Bro77, NRW73]. Commitment [KH75c]. common [AH79a, Whi79]. common-factor [Whi79]. commonly [PMN73]. commune [KP72]. communication [Bar72b, Joh76, KT72a, Sch73a, SV75, Sou75]. communities [CH76, Gil74, LK77a, PT79, RT75, TP79]. Community [PG79, Has79]. compact [Lin79a]. compaction [Hai79]. comparable [KDKY76c]. Comparative [WC75, Che73a, Fin72, JFR76, MT73, TP79, WSMS77]. comparing [Hut70]. Comparison [BH79, NO74, BP74, Jor73b, KDKY76b, Sho74b, Sho74a, SS73b, Tem75, Ran76]. comparisons [CB79b, CB79a, KJ72]. compartment [AW70, BD70, Cut78, Jac77, MMS74, MK74, Rot79]. compartmental [KV77, Măr76a, SM76]. compensation [Sil76]. competence [EGF77]. competing [TB72]. Competition [Hin77, LT75, CM72, Cza73, EH70, Gil74, JR77, PS70, Ros73c, Smi76a]. Competitive [Eps79, Koc74b, Lad77, Gro78, Har73a, Ros73b, TLJ76, TKS79]. competitors [AM76]. complement [PW79, PTB74]. Complementarity [May74, ABCT75, Jak72b, Jak72a, Jak73, Mar73, Mar71b, RE76]. complementary [FSG^{+}72, LT75]. Complete [Sou71]. Complex [Van76, Bar77c, CB76, HGR79, Kig73, LSL77, Naz77, New70, SBB^{+}75, SD75, Dou72b]. complexes [Bas77, Hea74, SKTC78]. Complexity [Fuj77, McC77, Cas78, CPC78, Mac79, Mar72a, SH76, SH77, Wic79, Woe71]. Complexity-stability [Fuj77]. component [Ala72, Ros78d]. components [Abe79, DC77c, GK72, IFN72, Oli73]. comportement [MCB77]. composed [KO78]. composition [CB77, CB79b, CB79a, Elt74, HT73, Hil78b, Les74, PSK76, Sin76, Sou71]. compositions [BH77, FAT77, Les73]. compound [JD77, KFH74, KBH76]. compounds [Bel77b, GPP74, HKAT77, LD73]. compression [MCM79]. comprising [CNS79, Tho78]. Computation [GP72, GY79, WB75, YGJ78]. Computed [Gin71a]. Computer [BR75, CS75a, GR78, HMT79, HH77b, Jor72, KR76, OMF72, Ran75, Ran77, RZMF75, RG78, Sal74, ZW70, AA78, Cam77, MFS77, Mar72b, NS76b, PS78, PCT76, RD77a, RD77b, SK72, SV75, SF79, Ste75, SD76]. concentration [BK71, Buy73, HR71a, Jaf74, LCH78, Ott74, RWD76, Smi73, vHB79]. concentrations [Atk71, Cal79]. concentrative [WP71]. concentric [FC71]. concept [Cut78, IMMY79, Lak75, Mik75a, Mik72, Mit79b, Ohk73, Shi77b, Tho73a, Tyr70]. conception [Sat74]. concepts [Bak76, RV79]. conceptual [EWRA76]. Concerning [BW78c, BW78b, Boy79a, Hul78, Lew76, Pap74]. concordance [Vol72b]. condensation [WA78]. condition [Den75a, DC77c]. Conditional [Sco79b, Sco79c, Ish73]. conditional-open [Ish73]. conditioning [Gro71, Kur71, Utt75a, Utt75b]. Conditions [BNW78, Fis73, And74a, BH70, CAA79, CK71, KP79, Koc74b, LH79a, SM76, Whi79]. conductance [Alb74, BHK74, BHK78, CS77, EM78, Gil73b, Got75, GKS78, Kle72, RSS^{+}73, Sta73b, Sta73c]. conductances [Ana73]. Conductile [Mur74b]. conduction [Che79, Gri77, Gri79, Mad77b]. confidence [CBPT78, Fag78]. configurations [MO73]. conflict [Man75d, MSK79]. conflicts [Smi74]. Conformation [GK73a, HMS75, KFH74, KBH76, LSL77, Loe71, Bel77b, FIN74, GK74, GGM75, KH75a, KH75b, KGH79, MPD70, MPP70, MPP71, MME74, Par70, PCC72, War70a, War74, WTR75]. Conformational [AJ79, BCLV76, DKU77, HR79, KOH70, Kie73, OK76c, RAH75, SC75b, VBCL76, WH73a, AFDL74, BS79, CLR77, DLK75, Dar77a, FRG74, GPP77a, GPP77b, GS71a, GS71b, HC78, Ji76, Jor73a, KDKY76a, Lat75, LBCV76, MMRE77, PGPP78, PP79, SG71a, SG72, Sch72c, TNG74, WB75, WD73a, BCDK77]. conformations [AFB76, GGM75, KR76, Kum75b, PM76]. conformon [Vol72a, Zgi75]. conformons [KG73, Kem74, KG74]. congenital [MHB79]. conical [Kel77]. conifer [CRC77]. Conjectural [Gün77]. conjugation [DT79, TD77b]. connected [Mar78b, McM75, OYT74]. connections [Sch77b]. connectivity [KDH77, LA76]. Consecutive [SG75b, EFH77, GK71, KNJ74]. consequence [KRL71, Nov73]. Consequences [CR75a, CR75b, DMRD78, Adl77b, Bro74, CKJ79, Got72, Jea79b, Wil70b]. conservation [LR74, Per76b, Sto70, Sto71]. conservative [BY74]. conserved [Nov73, Par75b, Par75c, VS71]. consideration [Fle75, Odu77]. considerations [BPF74, BH70, Din74, DP74, FP72, Gou76, Has79, Joh76, KD72, Man75a, SO74, SFT75, SB71, Tay78, WL71, WD73b]. considering [Hel78]. consistent [Bas71, BM73]. consisting [HH77a]. constant [Cro75, DM75, Koc74b, NR71, Shi79, Sim73b, SB75b, ZS72]. constants [BW78b, BJM70, CD72, Dar77a, DCMR74, GPP73, GPP74, GPP78, RS70]. constitutive [Rig79b]. constraints [Dou72a, Dou72b, Kil72, Sej77]. constructing [MBG73]. construction [BP79a, Cal76a, Par74b]. contact [Adl74, Adl77a, Adl77b, Bab77, BC73b, Dol75, RL77, Rob77, Rob78, RG78]. containing [Ain74a, Bas73, GDA75, KH79b, OD73, War70a]. contemporary [FAT77]. content [Bal73, DMRD78, RHR79, Yoc77b]. context [HS75]. Contingency [Pie72]. Continual [Sch73a]. continumn [OF75]. Continuous [Orr79, Sch72a, WRT79, AW70, CR73, GK73b, OB73, RS77, TS78b]. continuously [Par77]. continuum [Ode76, Win76]. contractile [Coo73, HS70]. contracting [Mil78a]. contraction [Ash72, Che75, Dav73, Dra70, DBU76, GG77a, GG77b, HS70, Hat73, Hil70, IC79, KS77b, Kom71, KZ79, Kou73, LRS79, MPHG76, MG77a, Opl72, She70, SY75, SYNY76, SW74, TT75, War70b, Whi71]. contractions [BLP76]. Contradicting [Den75c]. contrast [Fis77, GL75]. Contribution [DH78, GC73, APA79, GT77, Jea79a, BP79c, PB78]. contributions [Rap79]. Control [And74a, Moy70, NPV77, AG75, Bha77, Bla78b, Bon78a, CV76, CM75, Coh77, Coh78, Con71b, CZ76, DB71, Eld73, FGP^{+}75, Fri73, GGS78, GK73b, Har70, HLC75, Jun77, Kau74, Kit75, KWG79, Lew70a, Man79a, Mar71a, Mar72a, Maz78, MO76, Mei78, Mit79a, Mit79b, MA71, Nol76, Ond79, Osb74, Pai76c, RB77, Sco79b, Sco79c, SS79, SG76, Tho73b, Tho74b, TS79, VL78, VRM77, Wal74b, Wat72, WD72, WB73, WKG73, Whe75, Woo70, WD73b, dO79, vW76]. controllability [KGN76]. controlled [Ble71, Dun74, Lev74b, Par75e, TS78b, Wal70]. controls [And79, Kem79, Koc70]. controversy [Ros78e]. Convection [WR77, Coh76]. Convective [Nor71, Sey72]. convenience [SFT75]. conventional [LS71]. Convergence [vDHD78, GZ76]. convergence-divergence [GZ76]. conversion [KS77b, Kou73, Tri75]. cooling [Kle72]. Cooper [BC77]. Cooper-Helmstetter [BC77]. cooperative [Bla79, Tha73, WSF76]. cooperativity [Bar77b, Fri79]. coordination [McK73]. Copper [BS79]. copying [Olo73]. cord [Odu77]. corn [SS73a]. cornea [Fri76a, Hod71b]. corneal [Fri71, Fri73, Fri74, San74]. Coronary [LTP78]. correcting [Orl79b]. correction [WB74b, WB77b, ZG75]. correlate [Kat79]. correlated [VSW^{+}79]. correlating [Kie73]. Correlation [FAT77, Jon75, Nic76, SBR77a]. Correlations [HM79, Sha77, OSLM79, Per76a]. Correlative [Pat70]. correlatives [Lav77]. correspondence [Nel77]. cortex [Bar72b, Bra74, Lau72, PSM70]. cortical [BFM75]. cost [Ben78b, BS78, Hea77, Jon71, Man75b, TD76b, Wel77a, Zah77]. costs [Han79]. could [Edw78]. count [Fri72]. counter [Win75]. countercurrent [Abe73]. coupled [BS76, BC73b, DB78, KNJ74, Lin79c, MWS77, Mik77, MT78b, Ses74, VB70]. coupling [BJM70, Chr76, Dea75, GPP73, GPP74, GPP78, Gre76a, Kei75a, Kei75b, NCM75, RS76, Sto71]. course [GKS78, Pur77b]. courtship [Itz79]. covariance [OW78, PG79]. Coverage [RMG76]. covered [Alb71]. coverslips [WH72]. creating [CE77]. creation [Ano78h, Zgi75]. crick [Les74, Edw78, SKS77]. criteria [BP76, Coo78a, De 75, Măr76a]. criterion [Bar74a, Coo74, LT77, LS71, dS75]. Critical [JBM^{+}78, Ano79h, BNW78, DO76, Esh78, Mad76, Met75, Naz77]. criticism [CS71, PM77]. Critique [Fri73, Har76a, Hil73, HK77b, Pra72]. Cross [VD78, Jon75, Jor75a, Ros76c, Sch73b]. cross-bridge [Sch73b]. cross-reactivity [Ros76c]. crosslinking [Sum79b]. crossover [Chi76]. Cruciform [vHB76, Man75a]. Cryoenzymology [Fin76a]. crystal [Har74, Huk78]. crystallographic [SBR77b]. crystals [HS70, IC79, McC75b]. culture [KH75c, Pai76c, Per73, TL74]. cultured [NMK78, SD76]. cultures [AHH^{+}76, Gre76b, HC73, HTSH72, JWS76, PWW76, WR77]. Current [Bas74, Fon77, Pur76a, Fri73, GKP75, HR79, JRP74, NR71, RP73, Win75]. Current-voltage [Bas74]. currents [AR74]. curvature [Hew77]. curve [BSM77, CSL75, Cri77, Eas79, HH77a, MA78, Sha74]. curves [BW78a, Cra71, Cra73, FSFD70, HS77b, Kie71, KRH77, MSG^{+}78, Mil70, SP73a, TR79]. cusp [VRM77]. cut [DBS^{+}70]. cut-and-patch [DBS^{+}70]. cuticles [Swa74]. cyanide [JY72, Loe71, YMM76]. cyclase [dH76]. cycle [AG75, Ain75, Alb77, And79, BS78, Gar70, Goo70, GB77b, Gre77a, Ish77, KBW75, KOWG70, LSGFdlT^{+}75, Pas76, Rot77b, Sam77, Sch79c, Sim73a, TW77, Tor79, TS78b, TS79, Wal70, WT79, WRT79]. cycles [Leó75a, May75, Rot77a, SB78b]. Cyclic [Don72, WSF76, Ano71f, BM78, Coh77, Coh78, Cze71a, Cze71b, DW75b, DW75a, Jor73a, JA75a, JA75b, Ols75a, Ols75b, RB77, Seg79, FT74]. cyclical [MC73]. cyclicmonophosphate [Jor73a]. cycling [HH77a, RC77, dC75]. cyclostat [FG77]. cylinder [Adl77a, Lev72]. cylinders [FC71, MNR72, NS76a, SMN73]. Cylindrical [HS70, BM73, Har74, IC79, KS77a, MMS74]. cysteine [War70a]. cystine [War70a]. Cytarabine [GR73]. Cytoarchitectural [Sco79c]. cytochrome [ABCT75, Che73b, FSG^{+}72, MV75, Mit76, Yoc77b]. cytochromes [MCB77]. cytoethology [Ale72]. cytokinetic [Sha73, Sha74, Sha77]. Cytological [Moo71]. cytomembrane [Ada77]. cytoplasmic [Got72, MPHG76, Ord76]. Czerlinski [Ano71f].

D [GS71b]. D-ribose [GS71b]. daily [Sho74b]. damage [Mil70, Ord76]. damped [SSZ^{+}72]. daring [Dou72a]. Darwinian [Pap73b, Pap73c, Pap73d, Wil75a]. Darwinism [HS79b]. data [Ain75, Bak76, Bar79a, BNT71, BT71a, BT71b, BGKS76, CB76, CB77, GP72, HLR73, HLR74, HM72, Koc71a, Koc71b, Mag70, MA78, MGB73, PMN73, Sea76, SBR77b, TW77]. date [Fit76]. David [Ano75b]. deacyl [Smy79]. deacylation [CC70]. deadlock [DD77]. death [Iwa78, Sha73]. debate [Jak72a]. decay [CSL75, Moy70, Spe78]. December [Ano70s, Ano71u, Ano72r, Ano73r, Ano75y, Ano76u, Ano77y, Ano77-30, Ano78u, Ano78-28, Ano79-28, Ano79-33]. deciding [CMSC77b]. decision [Kur71]. Decisions [Gro78, Man75c]. decomposition [BT70, Met75, SKTC78]. decussate [Rob77]. Deducing [MK76, Wil70b]. deduction [Fit76]. Deductive [JR77]. deep [Wig73]. defects [VFKL74]. defence [Cil75a, Esh78, McC72a]. defense [BBGR71, SH77]. deficiency [Sre74, Vog77]. define [Coo74]. defined [Cra71, Cra73]. definition [Atl74, BW78c, Chr76, Cut78, HL70, Pap73b]. deformation [IC79, MJ76]. degeneracy [Ebr75b, FIN70, Hil78b, Klä75]. degeneration [Web70]. Degradation [Ran76, Ege79, FY74, KKJ77, Rig79a]. degrading [SS72]. degree [Bar77b, Bar77a, Bar77c, MSK79, Whi79]. dehydrogenase [AH79a, Pal70, WPTF77]. dehydrogenases [FH71a, UIN74]. delay [BC76, CGM77, Lad76, MT74b, Nie71, TL74]. delayed [CR75a, CR75b, Hel78]. delineation [BKB76]. demon [McC71]. Denaturation [BK76, BK77a]. dendrogram [MGB73]. dendrograms [WS78]. denning [WL71]. dense [Fri76b]. densities [OV78]. densitometry [NW73]. Density [AF77, De 75, Erh76, Gin77a, Gra73, Koc74a, MM78b, RP77, SA70, VSW^{+}79, WO71, Has78c]. density-dependent [De 75, Gin77a]. deoxyguanosine [LBCV76]. deoxyribonucleic [Man75a]. Dependence [Smi73, Tha73, GS78b, GPP77a, GPP77b, Lef78, Mit75, PGPP78, vHB79]. dependency [Cal79, Koc72]. dependent [AF77, BB77a, CB75b, CD72, De 75, DCMR74, Dun74, Gin77a, Gur73, Han72, Has78c, Koc74a, Mei78, Off74, RMG76, Sta76]. depletion [DVM79, WT78]. depolarization [Bas77, KM75]. deposition [RP75, RMG76]. depth [Nel77]. depurination [IFN72]. derangement [Whe75]. Derepression [TS79]. Derivation [BGKS76, Cho76, Koi73, MMS74, TD77a, Sta79a]. derivatives [GKYG70, KGYG70]. derived [Fin76b, Leb73, MHY77, NDF79, WW77, Yoc78]. deriving [Ain74b, Ain74a, Woo70]. dermestid [Bra75]. describe [AG78a, Ain79, Gra73, TD77a]. described [CJ76a]. describing [MMS74]. Description [Hon71, Hon78, Bar79b, Bor74, CJ76b, Fal77, Fri74, Jic72, JK72, KBW75, Rot77a, Smo76, dKJS76]. descriptions [Ros76e]. descriptive [NW74]. desensitization [CT73b]. Design [Cal76a, HS78a, Sum79b, Gar70, Lat75, MK76, Pre73, Smi78a, WK78]. designs [Dug79]. desmids [LH78a]. desorption [Got75]. destruction [PMC74]. detailed [Abr71]. detecting [Ano75b, Par74a]. Detection [Lei78, Vin73, Eng78, ER74, HE73, HC77, LS72a, Les77, Sea76, Vin71]. determinable [Ros72a]. determinant [Wig71]. Determinants [McM75]. Determination [Nor75b, RT70, Thi78, Ano70d, BW78a, Cho77, Coo70, Dar77a, Fri72, IE76, Kig73, LS76, Mei78, MG77b, NG77, Pre70, ST78b, SM76, SR72, SH72c, Wes73, Zar75]. determine [Zer75]. determined [BBG70, FCSS78]. Determining [KP79, CKJ79, GT77, GKS78, HSK72, Jon72, KP77b, Lew76, Mag71, Wes78]. determinism [You76]. Deterministic [PR77, Kro74, SS73b, SG75a, SG75b]. Development [LH78a, Bat73, BR75, Coh77, Coh78, Com77a, CFS78, DB71, EGF77, Fis77, KP79, LW76, LSW77, LW79b, PM77, PF78, Sch77b, SCDC77, SD77, Sta70, SHG^{+}71]. Developmental [Jea79a, Lin71, Lin75, GL75, Her70, Kac76, Sib78]. developments [Fon77]. Deviations [Koc72]. device [Bri78, Hjo70]. di-valent [DST79]. diagnosis [Hue77]. diagrams [BK77a]. dialytic [Bar77c]. diameters [Bon78b, SB75b]. diatoms [Ros73b]. dichroism [Bas73, Boh77]. Dictyostelium [Cot73, Coh78, Coh77, Dur73, PS78, Ram76]. Did [BP79b]. dielectric [IHI79, NB72]. dielectrophoresis [CP72]. Dielectrophoretic [PC72]. Diet [Par73b, EWRA76]. Difference [Woo78, CB78, Dia76, Hel78, May75, Ros72b]. difference-differential [Ros72b]. differences [Fle77, Har76a, Pfa71, SBR77a, Sre74, Via70]. different [Bri74, Coo74, Har78, MLM77, PGPP78, Ste70, WW77]. differential [HC75a, HC75b, MT74b, Mit79a, Oon76, Ros72b, SO74, Ste75]. differentiated [Eld73]. differentiating [Lap78]. Differentiation [Kau71, Ede72, HH73, JR77, Jun77, NG77, Par75a, Ste72, TS71b, WK73b, vW76]. differentiative [PM77]. differing [Rot77b]. differon [NG77]. diffraction [TW76]. Diffusion [CG73, Hew78, JWS76, NR70, Poo75, Pop79, Str77, SB73, ASS79, CD75, CCY^{+}75, Cha72, Che73b, Cla77a, Dun74, Ede71b, EH73, GDA75, Gur73, HSK72, HTSH72, HK77a, Hua73b, HH77c, Jor75a, JS79, KD72, LL70, LS76, Lin76, MMS74, MPH72, MH74, McE78, Mur75, Nor71, Off71, Off74, OK76c, Ses74, Sto75, SHC71, WR77, WK78, You78]. diffusion-controlled [Dun74]. Diffusional [EH79, Nor75a, MAM79]. Diffusive [SAT72, STA73a, BS77, HSC73, Jor77, MM78a]. diffusivity [EH73]. diformylglycerol [KFH74]. digger [BGD79]. digital [MA71, SV75, Koc71b]. dihydrofolate [SBR77a]. dihydrofolate-reductase [SBR77a]. dihydroxy [KBH76, KGH79]. dilemma [Nov75]. diluting [Orl79a]. dilution [HS77b, Nor75b]. dimension [BP70b]. Dimensional [Fit73, Bar74a, BBGR71, BBG70, CH72, CK71, Gil72, GBH70, Hon78, JW76, MO75a, MO75b, OMF72, Par77, Pur76a, San74, ZR73, Zar75]. dimensions [Blu77, BP70a]. dimer [JY72, Loe71, YMM76]. dimerization [MT76]. dimers [CHH^{+}76, Dou72a]. dimorphism [Kno74, Man75b, PBS72]. Dioxetanes [Cil75b]. Dipalmitoyl [BCDK77]. dipeptide [HR79]. dipeptides [MPD70]. diphosphoglycerate [DLK75]. diploid [AA78, HC73, KH75c]. Diploidy [LW79a]. Dipolar [KOH70]. dipole [KD76, RRS72, You74]. diptera [Ger75]. Direct [San74, Jor73b]. directed [CB75b, Smi77]. directing [Ril73]. Directional [CK77]. Dirichlet [Hon78]. disappearance [Moy70]. disc [BB73b, Ran75, Ran77]. discocyte [BKB76]. discoideum [Coh78, Cot73, Coh77, Dur73, PS78, Ram76]. discontinuous [AG78b]. discovery [Dia74]. Discrete [Bas71, GRDC72, Bed74, NM75, Sch73a, Sil76, Zei77, SO79a]. discriminant [CA79]. Discrimination [BP76, Gro70, PO79]. discs [Ran77]. Discussion [LT77, Hvi75, Wic79]. disease [Fab76, FP72, GMR75, HKK76]. diseases [Med72, Sum79a]. disk [MHMS75]. dislocations [HS70, Har74, IC79]. dispersif [MP79]. dispersing [BP70a, BP70b]. dispersion [Boh77, CP79b, RM73]. dispirations [IC79]. displays [Hjo70, Moy70]. dissimilar [NOB72]. Dissipative [CR78, SJ72, Ede70b, Mar72b, Mat78]. dissociable [Bat72]. dissociating [KDKY76b, KDKY76c]. dissociation [CD72, DCMR74, Eas79, Kur77, dL77]. dissociation-association [Kur77]. distance [KZ79, Tha73, WNHS75]. distinguishing [Har73a, MS71]. distortion [Alb73, Cli77, Miz77]. distributed [BRB78, Par77, SH72a, Van77, ZZ73]. distributed-parameter [Van77]. Distribution [DC77c, SCDC77, WY75, AWD77, BW77, Bed74, BD70, BC78b, BGKS76, CC73b, CBPT78, Fle75, Gra73, HS77b, HH73, KP77b, KKJ77, MYF77, Mag70, Măr76a, MO75a, MO75b, MHY77, Oos70, OMF72, Pal70, PP79, Rom76, She73b, Sim79, Sta79a, TS78a, Tho77a, WT79, Wig71, WD78]. distribution-free [CBPT78]. distributions [BP70b, Lap79, LS72a, Sla78, TR78b]. diurnal [KV77]. divalent [Got75, SO79a]. divergence [GZ76]. diversities [Hut70]. Diversity [Rou79, BY74, BD74, Con72, CM72, Dem77, Ebr75a, KGL75, Lax78, LH79b, PS70, Van74b, Web74]. dividing [Rig79b]. division [Bha77, BC73b, CE78, Dol75, Gre78, Kem79, Pre70, Ran75, Ran77, RC77, TS78b, TS79, WD73b]. division-synchronization [RC77]. divison [Ano70d]. DNA [Ama74, BP79a, BY74, Bas73, Bas77, Bec72, Bec78c, Ber77b, Bla77a, Ble71, Bra70, BC77, CS75b, CE77, Coo74, DZOP71, Den72, Doy78, Edw78, EJB^{+}78, Elt74, Fri76b, FIN74, Gal78, GM72, GRY^{+}72, Got72, Hit76, JO77, KK73, KP77b, Kir79a, Kir79b, Kum74, Les73, Lew70a, LK70, LK77b, MSG^{+}78, Mey73, Moo74, MMT76, MTM77, PPS75, PSK76, Ran76, RHR79, Res76, Res78, RP77, Sch71, Sha78, SBM72, SBB^{+}74c, SBB^{+}74d, SK79, TS78a, TS79, Vig75, VFKL74, VFK75, vHB76]. DNA-containing [Bas73]. DNA-division [TS79]. DNA-like [FIN74]. DNA-protein [JO77]. DNA-steroid [Got72]. Do [DBS^{+}70, McC75b, Yoc79, AP77, CB79b, CB79a]. dodecyl [Mad76]. Does [CPC78, BL74]. dogma [Ber77a, Ebr75b, Yoc74, Yoc78]. Domain [Kil70, Ciu77]. Domains [Dia76, Hon78, Lau72, NWW79, Pai76b]. Dominance [BB79, Ber76c, CC75c]. Donnan [Shi77a, Shi77b]. Donor [KPU78, GM72]. Donor-Acceptor [KPU78]. don’t [Dal79]. dopamine [Nic76]. dormancy [Fri72]. dosage [Gau73]. dose [Cra71, Cra73, Nos77, TR79]. dose-response [Nos77, TR79]. dose-survival [Cra71, Cra73]. Double [MLM77, Sch72b, Sch74, BW78a, CP79a, Cut72, Des77b, Edw78, FIN74, NDF79, PR76, Res76, Res78, Ril73, RWD76, SM78]. double-layer [PR76]. Double-site [MLM77]. double-strand [Res76, Res78]. double-stranded [FIN74, NDF79, Ril73]. doublet [BP79a]. down [HP77, Sin78, Ste76b]. downstream [ZR73]. DPL [BCDK77]. Dr [CS71]. drags [SB73]. Dreitlein [Ros75d]. Dreitlein-Smoes [Ros75d]. drift [Ale79]. driven [Ada77, BL75, Seg70]. drop [RLB74]. droplet [AK76]. droplets [Kob77, WSL79]. dropout [GH76]. Drosophila [Win72, Kig73, MHY77, NG77, Ran75, Ran77]. drug [AHH^{+}76, BD70, CT73a, CT73b, DW75a, GC76, GT77, Ger78, GKS78, HK76, HTSH72, KH75a, KH75b, KMS77, TLJ76, VAA71, Vog77, WFW75]. drug-induced [KMS77]. drug-metabolizing [VAA71]. drug-receptor [CT73a, CT73b, HK76, KH75a, KH75b, WFW75]. drugs [HPMH76, Pur77a, Pur77b, Smi78a]. dry [Bra71]. Dual [Sch71]. duality [AMR^{+}75a, AMR^{+}75b]. ducks [Fed77]. due [GPB71, HOR73, IC79, Klä75, Miz77]. Duetting [Wic76]. duplication [Org77, Wen77]. duration [KP79, RP73, SP73a, Sch74]. during [BLP76, Coh70, CZ76, Cos75, DB71, DH78, Dou72b, DP74, DT79, Hod71a, LG71, Lew73, MW78, Nan76, OG74, Par78, SA70, STD77]. dye [Bas77]. dye-DNA [Bas77]. Dynamic [Hol72, Kum75a, Moo74, TS77, BL75, Coo78b, HKK76, JR77, KS74, OS71, OS74, Wal74b, Win76, Sav70]. dynamical [GG77b, Mil78a, OYT74, Ros75e]. Dynamics [Alb77, ABC^{+}70, HLC75, HC73, Nyh78, RP75, TL74, TP79, BSM77, Coh77, DC77a, DKU77, Ede72, FY74, FY76, GK72, Gla75, Goo70, Gre77b, Gro71, HB70, HCBL70, HSS79, Kau74, KS77b, Mit75, NG76, Nol76, Ros72a, Ros72b, SLK76, Sch73b, Sil76, TC79, TF74, WB77a, Woo72, Chu78]. dystrophic [DC77c].

E. [Ano70d, LK77b]. E.H.T. [LBCV76]. Eadie [CB76]. earlier [Yca74]. earliest [Sag73]. early [And76b, BR75, CR71a, CR71b]. eccrine [RT77]. Echinocyte [Bec78a]. Ecolog [Ano78-33]. Ecological [RBM76, BV75, Bra77, HB70, HM72, NGP78, PT79, Ros77a, RL76, SJ72, Sme76, TM74]. Ecology [Van72, Bak76, JR77, Niv77, Pie72, Ste76a]. economics [DP77b]. Ecosystem [AC74, Hae78, CP70, Cza73, EWRA76, Fin76b, Han76, Kri77, Mat78, TS77, VS71]. Ecosystems [Pan78, Bat74, CPC78, Han73, Han79, Kro77, Lad76, LBGK76, May73, Maz78, Mob73, MK74, MG77b, Ula72, WO71, Wu77]. ecoystems [Lev77]. edema [Fri74, Rap78]. edge [LR74]. Editor [HA75, Lon76]. Editorial [Dan70, Dan75, Ano70e, Ano70f, Ano70g, Ano70h, Ano71g, Ano71h, Ano71i, Ano71j, Ano72f, Ano72g, Ano72h, Ano72i, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. Effect [Bla75b, Bla79, EH73, FC70, Hil78a, LA76, Lym74b, MM78b, SZ71, AG75, BC76, Chu74b, Cli77, CMV76, DMRD78, DST79, Ehr79, Fri76a, GT77, Ger78, GH76, Hay74, IE76, LCH78, LG76, LT72, Mac76, Mar76b, MO75a, MO75b, MAM79, MJ76, Miz77, NM75, Oos75, Orl79a, SP73a, Sho74b, SKTC78, Tan78, TKS79, TL74, Wei72, WC76, WL76, YC77]. effectiveness [WSMS77]. effectors [NOB72, Smi77, Wat72]. Effects [Hon71, KS77a, KT72b, LK79, Mic79, NM78, See73a, WB74a, BW76, DS70, EH79, Gau73, GR74, Gün77, HSC73, HC75a, HC75b, HSK72, Hvi75, Jor74, Kil70, LP75, Lav71, Leb73, Les74, LST74b, Mor72, Nel77, Nel78, OK76c, Pfa71, Poo75, Rak77, Rak78a, Sch73a, SAT72, STA73a, TSP78, TD77a, WB73, WK78, RBP78]. effet [MP79]. efffector [GL78]. efffector-target [GL78]. Efficiency [McC75a, Leo71, PBV74, RL76]. efficient [Par74b, WB75]. egg [JRP74]. Ehrlich [KBW75, WZ72, WNHS75]. EHT [BCLV76, GS71b]. Einstein [WA78]. either [MB78]. ejection [Nol76]. elastic [MJ76, RS70, RK72]. elasticity [Mak76b, RV79, Sto75]. electric [CMV76, Mad77b, OK76c, Rob70, Shi71, SM77, Yag76]. Electrical [Fle77, MFS77, Gor78, NM75, Sal74, SFN77]. electrically [Buy73]. electrodiffusion [ABR72]. electroencephalogram [SSZ^{+}72]. electrogenic [CKJ79]. electrokinetic [The71, WZ72]. electrolyte [CW74, LCH78, Mar77, MW78]. electrolytes [Abe73, Jac77, Vai71b]. electrolytic [Jor75a]. electromechanical [McK73]. electron [AAQ71, AL72, Arc71, BBGR71, CSVR78, Con79, CK71, DUK73a, DUK73b, DKU77, FH71b, Fri76b, GBH70, Kem74, KPU78, San74, Sre74]. Electroneutral [ABR72]. Electronic [HZ77, HH73, DUK73a, HMS75, JY72, Jor73a, KOH70, LM75a, LM75b, MO73, Mag73, MSG^{+}78, MT73, Pai72]. electrons [BC73a, KVB77, Sco79a]. electrophoresis [And79, Mad76]. electrophoretic [GKS78]. electrophysiological [Sco79c]. electropotentials [Gin72]. Electrostatic [ABD^{+}72, LS78, NCM75, NP71, She70, BM73, CW74, Dea75, Eps71a, Eps71b, HSC73, Ohs77, SAT72, STA73a]. element [CV76]. elemental [BFM75, Kow71]. Elementary [Lüc75, Cha77, CR75a]. elements [WE75]. elimination [Bar77c, BKWT76, BB77a, BRB78, Den75c]. Elongation [GWZR77, Lev74b]. elution [BC78b]. embrace [MMT76]. embryonic [BR75, CE78, GR78, WE75]. emergence [VLB78]. emission [RO74]. empirical [Hae78, KHK79, Kum75b, Mof73]. encephalitogenicity [Wes78]. encoding [Les77]. end [Ada76, GKP75, GR74, KT72b, Mac77c, Sou71, WB74b, WB77b]. end-plate [Ada76, GKP75, KT72b, Sou71, WB74b, WB77b]. endocardial [Nel78]. endocellular [vW76]. endocrine [Lei77]. endocytic [DP77b]. endogenous [JRP74]. endometrial [BB73a]. endonucleases [SK79]. endorphins [Kat79]. endospore [GD79]. endosymbiotic [Rei76]. endothelial [GCS72, Rub77]. endothermy [Ste73]. endothial [NP79]. endpoint [Mur74b]. Energetic [Wei74, Chr76, GG77b, PP79]. energetics [KD79]. energies [HMS75, Mor72]. Energy [AL72, BS78, Fon74, SKS77, See73b, AFDL74, APA79, Bel77a, BM73, BCDK77, Can70, Con79, DS73a, Dav77, DUK73a, DUK73b, DH78, FRG74, Gar76, Gin71a, GC73, Gre76a, Hag76, Han79, HR79, Hor74, Jor73a, Kou73, Lak74, LRS79, MV75, May73, McC72a, McC72b, Men76, MV76, OK76c, OO77, Pai72, Pai76a, RS76, Rei79, RL76, SB77, See73a, She73a, Sme76, Sto70, Sto71, Sun77, Tri75, Ula72, WB75, Wel77a, YMM76]. energy-yielding [Hor74]. engulfment [AMR^{+}75a, AMR^{+}75b, GR78]. Enhancement [Hol71, GL75]. enrichment [AE79]. ensemble [Kau74]. entered [Jak73, Mar73]. enteric [Pai76c]. enthalpy [Bel77b]. entrainment [Wev72]. Entropy [Fox71, CGM77, Jea78, Rya72]. entry [Gol76, Jak72a, Sto76]. Enumerating [Lew71]. environment [Bry73, Her70, Jok74, Mou71, NG76, SD75, WP71]. Environmental [NGP78, Sch70, TSP78, Yod78, EH70, Fre70, Koc74b, Pie74]. environments [CB74]. enyzme [KDKY76a]. Enzymatic [BKWT76, Sch76b, And74a, KDKY76b, KDKY76c, SH79, SB76, SBP77]. Enzyme [ER78, Ain74b, Ain74a, Ain75, Ain77b, AG78a, Ain79, AG78b, And74b, Ano71f, BW76, Bar77a, Blu76, BP74, CB76, Cho72, Chu74b, CBS75, Con79, CBPT78, CD72, Cze71a, Cze71b, DS73a, Dar73a, Dar77a, Dar77b, DC73, Dug79, DCMR74, Dun74, EH73, Fin76a, FPB72, GC76, GK71, Han72, Har71, Har73a, Hea74, HH71, Hil78a, HK74, JI78, Kan73, Kat72, KVB77, KNJ74, KDKY76c, KL79, Mal76, MO76, MB78, OD73, Rak77, Rak78b, Rak78a, RF73, See71b, SS72, Ses72, SY75, SYNY76, SG75a, SG75b, Smi78a, SD75, SKTC78, SD76, TKO74, Tha73, TE76, Tys79, Van76, Wal74b, Wal74a, WB73, WSF76, Whi79, Zer75]. enzyme-catalysed [BW76, Dug79]. enzyme-catalyzed [Zer75]. enzyme-containing [Ain74a, OD73]. enzyme-enzyme [Hil78a]. enzyme-ligand [SKTC78]. enzyme-one [SG75a]. Enzyme-substrate [ER78, SD75]. Enzymes [Fer73, BW78a, Bye78, CB75b, Chu74a, Chu74b, Chu78, Fri79, GA77, HSC73, Har78, HC78, KS72, KKDY74, KDKY76a, KDKY76b, KDKY76c, Kur77, LR76, MLM77, Mor72, MMT73, SAT72, STA73a, Smi77, Sto71, Tho71, UIN74, VAA71]. enzymet [Sin73]. enzymic [Ain77a, Dar73b, Dar75, Dar77a, Des77b, EFH77, EH79, KDKY76b, Lym74b, MB78, Olo73, UINH73]. epidemic [SBC79]. epidermal [HMT79]. epidermis [HH77b, HH77c, Mit79a]. epigenetic [HS79b]. epilepsy [AC77]. Epileptogenesis [WJ78]. epithelia [Cut72, Gor78, KHP77, SFN77]. Epithelial [Jac77, Ana73, Fri74, MT78b, NR71, Shi71]. epithelium [AWD77]. epoxy [KGH79]. epoxy-metabolites [KGH79]. equation [Gin77b, Kat72, Lym74b, Mar79b, Off71, Off74, Oon76, RO74, Shi78a, SF79]. Equations [CS75c, RP77, Ain74b, Ain74a, ABR72, Bar77a, Bar77c, BS77, CNS79, Cho76, Chu74a, Dia76, Gur73, Hel78, Koi73, LR74, MMS74, Mar71a, May75, MT74b, MR74b, Mur75, Ros76b, SO74, SYNY76, Shi79, ST76b, Ste76b, TD77a, VB70, Whi79]. equilibration [KS71c]. equilibria [Bla75b, LM74a, MSF71, MS71, TD76a, Yod78]. Equilibrium [KGL75, Rot75, TE76, Ain77a, And74a, And74b, AF77, Che79, Chu74a, Chu74b, Chu78, Cro79, Dar73a, Dar73b, Dar75, Gin77a, GA77, Has78b, Ish73, IM75, Ish77, KDKY76b, KDKY76c, LW79b, Mag70, MSG^{+}78, NP71, NTR77, Ril79, RF73, TLJ76, The71, TS77, WB73, Wig75, SB71]. Equivalence [MS71, Kub76, LW76]. equivalent [GB77a, SFN77, Sug75, Yoc77c]. Errata [Ano71k, Ano75j, Ano77i]. Erratum [Ano70i, Ano73g, Ano74i, Ano74j, Ano77j, Ano78g, Ano79g, ST78a]. Error [GI77, GY75, PMN73, WK73a, vHB79]. errors [BY74]. erythrocyte [GDA75, LCH78, Sir70, SC77]. erythrocytes [Goo78]. erythropoiesis [MA71]. escape [HLC75]. Escherichia [BHD74, BD75b, Coo70, DT79, Lan78, Man79a]. especially [KJ72]. ESS [Mac78]. essence [Ler78]. established [HZ77]. Establishment [HMT79, Paa78]. Estabrook [Vin74]. ester [Zem73]. esterase [Hol71]. esteratic [MMT73]. esters [LM75a, LM75b, Rea77, Smy79]. estimated [GH76]. estimates [Naz77]. estimating [Dug79, Hut73, Rom76, Sch79b, Wun75a, Wun75b]. Estimation [HT73, HE73, HS77b, KS74, PMN73, Sch74, SM71, TR78b]. estrogens [MKLKV76]. etiology [DP77a, Pap74, Web70]. eucaryotic [PCT76]. Euglena [BJ72]. eukariotic [CV76]. eukaryote [GM72, Hal73]. eukaryotes [DH79, Rea75a, Rea75b, Sub77, TS79]. Eukaryotic [Eds75, Bon78a, DCT73, GD79, NDF79, Ond79, Ost77]. eusocial [Cha78c]. evaluate [Bar74a]. Evaluation [CBPT78, EH73, Koc71a, Koc71b, Rap79, SM77, VB70, Cza73, Mad76, MTL76, NGP77, ZS72]. even [Gat78]. even-aged [Gat78]. event [Zei77]. events [CL70, Got72, WJ78]. ever [HH71]. Evidence [Ima75, Kub70, Ohk73, Shi77b, Ima76, Lak75, Sco79c]. evoked [GKS78]. Evolution [Car78, Cas78, Cha78c, CP70, Hal71, Har73b, Hor74, Kam74, Koz79, Mat78, Nyb75, Ses71, SP73b, Woe71, And76b, ABCT75, BP79a, BB79, Bar77d, Bel78, Boy79a, Bro77, CAA79, Cha78b, Eds75, Fed77, Fre73, GS77, Gla78, GY75, GI77, Has78c, Hea79, HS79b, IM78, Jur77, KT72a, Kno74, KSR74, LW79a, Loe71, Mac75, Man75d, MG72, McC77, Nag76, Ost77, Pap73c, Pap73d, PBS72, Par74c, Par78, Pit78, PSK76, RHR79, RS76, Rei78, RNS71, Sam71, SH76, Sch71, Sco79a, Smi74, Ste73, TS71a, Tre76, Van72, VPBD78, Wel77a, Wic78, Wic79, Wil70b, Yoc79, CC75b, CC75c]. Evolutionarily [BGD79, Has78c, BY75]. evolutionarily-static [BY75]. Evolutionary [Con74a, KC78, Ril79, RS78, Bas71, BH77, Bre79, Goc78, HT73, Has78d, Hin77, HSS79, MT78a, Mey73, Pap73b, Pap73c, Pap73d, Pin70, Ros78a, WSSB77, WW77]. evolving [GS78c, Lav76, Thi78]. exact [Mar79b, Shi79]. exactly [Pan78]. examination [BC77, McE78, NWW79, RRS72]. examined [Gri77]. example [BK76, BK77a, Lap79, Lev74b, SJ72, Sou75]. examples [Hil78a, Ros79b, TR78a]. excess [Sre74]. exchange [ARW73, AMR^{+}75a, AMR^{+}75b, Cha72, Cli77, CS75c, Dar73a, Dar75, FC71, Goe74, Lev72, MF79, Mil73, RL77, WB73, Win75, Yag76, Zei77]. exchange-type [AMR^{+}75a, AMR^{+}75b]. exchangeable [Koc71a]. excitability [GS78b, Kob77, KMS77]. excitable [AK76, AR75, CJ76b, Dur73, Gri77, KK78, KOR^{+}77, MC72, OK76b, SM77, VG79, Wil70a]. excitation [Cla77b, CJ76a, CT79, Cos75, Dav73, Lau72, MT79, Pai72, RP76, Sha78, WT78]. excitations [Pai76a, Pai76b]. excitatory [Cla76]. Excited [DZOP71]. Exciton [AK78, Mar78a]. exclusion [Dun74]. excretion [Ond79]. exhibiting [Tag78]. exhibits [MM78a]. Existence [Blu71, Ros76b, BNW78, CPC78, Hyv72, Mur75, Sch70]. Exocytosis [Dea75]. expanding [Rob78]. expansion [LG76, MT74b]. expected [KP79]. experiment [Sho74b, Sho74a]. Experimental [Dug79, BW76, BGKS76, CSL75, CGM77, DVM79, HR71b, KJ72, MG77a, Sto76]. experiments [BT71b, CM75, CP70, Eij70, Gar70, JvS77, RM73, Res73, She71, Pap73c, Pap73d]. explain [Bac74, GGSW72, RBSS70]. explained [Huk78]. explaining [LB72a]. explanation [Can70, GG75a, GG75b, SJ72, Sre74, Wes73, BMS77]. explanatory [Van72]. Explicit [Fra74a, NM72, NM78]. Exploitation [Lev77]. Exploration [GC76, Kat79]. expo [Eas79]. expo-exponential [Eas79]. exponential [Abe79, AG78a, Ain79, Eas79, GR75, Nie71]. exposure [Cra77, Lan78, NB72]. expression [And77, AP70, DB71]. expressions [KV74]. Extended [Cou78, SG71a, GS71a]. extension [AG78a, Ain79, DBS^{+}70, HO74, NG77, OF75, OD73, Ran77, SUE76]. extent [Hut73]. external [GL73]. externalities [Fre73, Tul71]. extinction [Heu78, SM79]. extracellular [CBS75, WJ78]. extracts [MWM76]. extraneous [CSVR78]. extraordinary [HL79]. extrusion [Lew70b]. eye [LF77].

F [Vin74]. face [LS77]. facilitate [Les77]. Facilitated [KD72, Sto75, Bun78b, CG75, Vai71a]. fact [MP73]. factor [AA76, Erh76, Rus72, Whi79]. Factorability [Bar77b, BW78b]. factors [GKS78, Jon71, Sre74, TBP78, Tho74a]. failure [Fit76, Ros78c]. Fairy [ST76c]. familiarity [Bat73]. families [BH77]. Fang [CS71]. far [GA77]. far-from-equilibrium [GA77]. fast [HH77a, Paa78]. fast-growing [Paa78]. fate [CH73]. fatty [Bro74, Hai79, Van76]. fault [Hil73]. favor [CPC78]. feasible [AF77]. feature [LSB78]. features [Bri79, CK77, DLV77b]. febrifugine [Che76]. February [Ano70q, Ano71s, Ano72o, Ano73o, Ano74t, Ano75v, Ano76-28, Ano77-28, Ano77-32, Ano78v, Ano78y, Ano79v, Ano79-27]. fecundability [GH76]. feedback [FT74, HS77a, JK72, KJ72, Tho78, TF73, TFB74, Wal74b]. feedback-repression [FT74]. Feedforward [Sav79, SJ79]. Feedforwards [Ros78c]. feeding [KD79, Pap73d, Par75e]. female [Heu78, MHY77, PBS72, Sey77]. Ferritin [Fis76]. ferrohemoglobin [Miz77]. fertility [Gup75]. fertilization [Nag76]. fetal [KS71c, Spe70]. fiber [Alb74]. Fibonacci [Rob78]. fibres [Mor78, SP71, Wig73]. fibril [Huk78]. Fibrillar [RL77, Don72]. fibrillogenesis [ST79]. fibroblast [AA78]. fibroblasts [KH75c, Ler79]. fibrous [Par70, Swa74]. fidelity [PB78, Sch78, BP79c]. Field [Oos73, Bak76, CE78, CMV76, Jea79a, Oos75, Shi79, Sim73b, SM77, ZS72]. fields [CR75a, CR75b, Dur73, OK76c, Rob70]. fighting [Par74c]. filament [Che75, GG77a, GG77b, Hil70, LB77, Mit75, Sch73b]. Filamentous [Ols75a, Ols75b]. filaments [Dou75]. file [KH79b]. filling [Adl75]. films [NRW72, NRW73]. filtration [MT74a, Ned75]. final [Coh71]. fine [Adl77a]. finished [Cas74]. finite [GR74, Joh79, KH75c, Koo77, LM74a, Wu77]. finite-time [Wu77]. fireflies [SC75a]. firefly [CS75a]. Firing [CT79, Oon76, CG73, GRDC72]. First [Oon76, Tre75, BP79b, Fle72, WT79].

First-occupancy [Tre75]. First-order [Oon76, Fle72]. fish [Jon71, TC79, Wei74]. Fisher [Ben78a]. fishes [Itz79]. fissioning [Tod70]. fit [Bon78b, BT71a]. Fitness [Wit78, WP71, Cas78, CGM77, Orl79a, Orl79b]. fits [TW77]. fitter [Wil78]. Fitting [MA78]. five [AW70]. five-compartment [AW70]. fix [PW79]. fixation [KMSP74, Paa78, Rei76, SP73b]. fixed [LS75, WZ72]. flagella [Ada77, Bla74, Bla75a, Blu71, Cal76a, Dou75, LB71, Lub73, NR70]. Flagellar [Har74, Lev74b, BJ72, CH72, HM71b, Kel77, LB72b, SH72b]. flagellate [Kel77]. flagellum [Har74]. flamingos [Fed77]. Flash [SC75a, CS75a]. flat [Shi71]. flavins [GPP70b]. Flexibility [LBCV76, RAH75]. Flicker [Hol76]. flies [PMC78]. floating [dH76]. flocking [Pul73, Vin71]. flocks [Vin73]. flocs [MH74]. Flow [IMMY79, RK72, App72, Gin71b, Hil72a, JP79, KP77b, KH79a, Kle72, LS72a, LTP78, Mar77, May73, Mik77, Nor75b, Ols79, PF78, Pur76a, RL76, RWD76, SB79, TS71a, Ula72, Win70b, ZR73]. flowers [Wis77]. flowing [BKWT76, BB77a, BRB78]. flows [BLP76, Fin76b, MWS77, Mik77, MT78b, Seg70, Sme76]. fluctuating [Eij70, MPH72]. fluctuation [Lav71, Oos73, Oos75]. Fluctuations [dC75, dC77b, Ber78, Cla77b, HR71a, HH71, NGP77, NGP78, Ven76]. fluid [Alf74, Blu74c, Fin76a, Gre78, Mar79a, RK72, Wie79]. fluid-mechanical [Gre78]. Fluidity [Cla77a]. Fluidity-induced [Cla77a]. fluids [WP71]. Fluorescence [MV76, Bas77, Ott74]. fluorouracil [Abd76]. Flux [Par75a, KWG79, Smi73, Ste76b]. fluxes [Fri76a, KH79b, LT77, Vai71b, Vai71a]. fly [AP77]. folding [Sim79]. folic [Vog77]. follicle [CS73]. follicular [CD75]. follow [GH76]. follow-up [GH76]. followed [ZV77]. following [BD75b, Cra77, Sin78]. food [JDTF73, RT70, Ros73c, SB75a, So79b]. foodweb [Mac79]. foodwebs [Con72]. foot [CM75]. Foraging [Tay78, TV75]. Force [Mor78, Gin71a, Hil70, Mil78a, PC72, PR76]. Force-velocity [Mor78]. forced [Lin79b]. Forces [Doe71a, Sch79a, Coo73, Dra70, DST79, GJ76, HKM79, Isr73, Lav71, MNR72, Nir75, PN73, She70, SMN73]. foreword [HA75]. forks [Dou72b]. form [Hon71, OW78, RB77, SP71, Yu77]. Formal [Win70b, Atl74]. formaldehyde [VFK75]. formalism [GG77a, HB70, KZ79]. formalization [Tho73b, Tho74b]. formation [Bec78a, Bec78c, CS73, CM75, DP76, DG75, Del75a, Del75b, Ede72, GD79, Hol79, Hon73, JD77, Jor74, Lew72, PPS75, RRB76, SSR74, See71a, VFKL74, Win73, Zem73]. formed [AK76]. forming [Les74, VSW^{+}79]. Formose [CSIW72]. forms [Dar77a, Dou72b, GB77a, KDKY76b, KDKY76c, LS71, MB78, NWW79, Sch76a, ZG75]. formula [Hut70, OW78, Orl79b]. Formulation [LBGK76, LT77, TB72, KK78]. formycin [MME74]. formyl [KBH76]. forth [Mar75c]. FORTRAN [MA78]. foundation [Lüc75]. Foundations [CFS78, HLR73, HLR74]. founder [WC76]. Four [SK79, Blu77, CR75b, Edw78]. four-dimensions [Blu77]. four-point [CR75b]. Four-stranded [SK79, Edw78]. fraction [KRH77, RP73]. Fractional [Dan76]. fractionation [BC78b]. fragment [MRTG79]. fragmented [Bec78c]. fragments [KK73]. frame [BPF74]. framework [Cam77]. Free [Bro77, Goc78, APA79, Bel77a, BM73, Con77, CBPT78, DMRD78, FRG74, Gar76, GC73, HWD76, OK76c, She73a, Sun77, Wel77a]. freezing [Oli73]. frequencies [Eds75, EF76, GRY^{+}72, TS78a, TD77b, vHB79]. frequency [CA79, MHY77, Ove77, Pai76a, RB77, Wil75b]. Frobenius [Jea79a]. frog [HH77b, HH77c]. Frontispiece [WSB76]. fruiting [Wal79]. Function [Leó75b, Ain79, Bar77a, BDS73, BBT70, BT70, Cam77, Car78, Dea75, DP74, Eas79, Fin76b, GPP77a, GPP77b, Gil74, Goo73b, Hai79, KK73, KTS71, LHC76, McN79, MTL76, Mit76, MG77b, MMT73, Pap74, PGPP78, RP73, VLB78]. Functional [Gar74, Kri77, OB73, Sug75, ABCT75, BR79, CDS77, HM79, Kat79, Mur74b, Odu77, RRB76, Sum79b]. functionally [Sza77, Yoc77c]. functioning [Blu76, IP71, KVB77]. functions [Ano71x, Bar77b, CB75a, GA77, Gra73, HCBL70, Jea79a, Mag70, Nea72, Nic73, NTR77, NM78, Org77, SUE76, SYNY76, Wei71, Wig75]. Fundamental [Ros73a, DT74b]. fungal [Res78]. fungi [KP72]. fungus [Win73]. Further [BPF74, Bri79, For75, Ima76, NRW73, Zah77, Gup75, Lym75a, Lym75b, Orl75c, SG71c, Wel77b]. fusion [Bel77b, Bre79, DH79, Dol75, Has74, Nel75, Par78, PA71, Spe76]. futile [SB78b].

G [SP73a]. G. [Ano71f, Vin74]. GABA [AJ79, SC75b]. GABA-sensitive [SC75b]. galactolipids [Ege79]. galactosylglycerols [Ege79]. game [HSS79, NTR77, Rei78]. Games [HS79a, Ros78a, Smi74]. Gamete [Man75b, Kno74, PBS72]. gametes [Par78]. gamma [SH72a]. gap [BB73a]. gas [JWS76]. gas-permeable [JWS76]. gases [WK78]. gastritis [HKK76]. gating [LS78]. Gaussian [Mof70]. gel [Doe71b, Mad76, PB77]. gel-to-sol [PB77]. gels [GDA75]. Gene [JA75a, JA75b, Org77, BM78, DB71, Gau73, Ger75, Goo73b, KC78, Kau74, Pow75]. Gene-duplication [Org77]. Genealogy [Nov73]. General [BFM75, Fri71, Lym74b, TB72, Che75, Doy78, LT72, MO73, MMS74, MA78, Mik72, MPHG76, Ros78c, Sum79a, Tyr70, Vas74, BT71b, SO74, VBCL76, WD73b]. general-purpose [MO73]. generality [Fle72, Kub70]. generalization [Sta79b]. Generalizations [PN73]. Generalized [BD70, Mar73, BC78a, Cza73, DT74a, GL79, Jak72b, Jak72a, Jak73, Mar71b, MHC79, NSS74, OV78, She73b]. generated [Dou72b, Rom76]. generates [Nin79]. generating [DW75a, WE75]. generation [AM75, Fri72, Jor72, LK70, Sch76a, Wic79, Dou72a]. generations [BT70, Leó75a]. Generative [Hae78]. generic [Ros76b]. Genes [Rap71, Ger75, Mar70, PSK76, Yoc79]. Genetic [Llo77, NG77, TF73, AFS^{+}79, BNT71, BT71a, BT71b, Bar77d, Eng78, FT74, GC79, GS78c, Jur77, Ker75, Les70, Lew70a, Mik75a, Mil78b, Mit79a, Moo74, Pap73c, Rea77, RCW73, RE76, Sch71, Tho73b, Tho74b, Tsu79]. Genetical [FT74, Mic79, TFB74, Cha77]. genetics [CC75a, CC75b, CC75c, Cha78a, Lay78]. genome [Oli77]. genomes [Bec72, Bec78c, FSST78]. geometric [Abr71, BD74, Bon78b, Nin79, Pre73, Sch77b, SG76]. Geometrical [Sch73b, WL72]. Geometry [ARW73, Goe74, Ham71, Ano70d, Blu77, Har74, JY72, Lei77, Pre70]. Geotaxis [KP77a, WJ74]. geotropic [Sil76]. Germ [Ber77b, IE76]. germination [CGM79, Cot73, GB75, HJ75, Hag76, Lef78]. Giant [MT79, Gri79, Has78a, RP76, SS70, Sta73b, Sta73c, Str76]. Gibbs [Ciu77]. gill [Ols79]. Gingival [Vog77, Alf74]. given [MBG73, Mur75, Yoc77c]. gland [RT77]. Glansdorff [dS75]. gliding [Man79b]. Global [De 75, HH77c, Fuj77, Ros78c, So79b]. Globality [Nel75]. globular [Cri74, Cri75, CK77, GY79, MO75a, MO75b, OMF72, PP79, YGJ78]. globule [Lew76]. glomerular [MT74a]. Glu [LT72]. gluconeogenesis [Ste78]. glucose [Atk71, HPCH79]. glutamate [WPTF77]. glutamic [Lew72]. glutamyl [BS78, Sam77]. glyco [RRB76]. glycolytic [Tor79]. glycoprotein [Gin73]. glycoproteins [SD74]. glycosidic [Jor73a, Jor73b]. glycosyl [GPP77a, GPP77b]. glycosylated [Bea79]. glycyladenylate [BCLV76]. Goel [LR74, Ant74]. goldfish [Eij70, Sch77b]. Gompertz [BSM77]. gonadal [MHY77, Sho74b]. graded [Cla76, Cla77b]. gradient [Doe71a, HS75, LS75, MP78, Mei78, Pie74, Seg70]. gradients [Gor78, Hew78]. grain [Fri72]. grain-count [Fri72]. grammars [Hae78, Lin71]. grana [DST79]. granule [HKM79, Utt77a, Utt77b]. granules [DM75]. graph [HPCH79, Whi79]. graph-theoretical [Whi79]. graphes [Hyv73]. Graphic [Kat72, Hjo70]. Graphical [TR79, Bar77b, BW78c]. graphs [SB76, SBP77]. gravitational [Rei79]. gravity [DC77c, Sil76, WJ74]. grazing [Kri77, NM78]. green [MR74a, See73b]. gregarious [CO75]. Griffith [CMSC77b]. grooming [Sey77]. gross [Woe73]. Ground [MSG^{+}78]. group [Der77, DCMR74, KA74, Rap79, Sum79b, WP78]. groups [CW74, DUK73b, FH71a, HM79, Jör75b, NCM75, NP71, PR76]. growing [Koc70, Paa78, Rig79b, RD77b]. Growth [AH79b, Jea78, RMG77, Ala72, Alb75, AB71, Bla78b, Blu74a, BHD74, BCY79, BC76, Coh71, Coo70, De 75, Fra70, Goo70, Gre77a, Gre76b, GL73, Hel78, HC73, HS77a, Jea79a, Kem79, KH75c, Koc74a, Kra78, Kub70, Lev74b, Lew73, LSGFdlT^{+}75, LW79b, Nyh78, OG74, Osb74, Pai76c, PD74, Par73a, Par75d, Pas76, Ran77, RGZW78, Ros73b, RP77, Sha73, SG76, SE79, Sin76, Spe70, Tho71, TR78a, WKG73, Whe75, ZZ72, dKJS76]. guard [DS73b]. guide [CB78]. guided [Ste75]. gustatory [HA72]. gyro [Mac76]. gyro-chemical [Mac76].

H [Ano71f, Vin74, GPP74, TSP78, WD78]. H-bond [TSP78]. H. [Vin74]. Haase [Spe70]. Habitat [Bry73]. haematins [Kre70]. haemoglobin [NG77]. haemolysis [Goo78]. Haemophilus [Ran76]. Haldane [TB72]. hallucinogenic [DHK78]. hallucinogens [Nic76]. halving [Fri72]. hamiltonian [Woo70]. handedness [OG74]. Handicap [Zah77, DO76, Esh78, Smi76b, Smi78b, Zah75]. hang [HP77]. Hapten [DG75]. harbor [Shi78b]. harmful [Jok74]. harvest [IPH^{+}79]. having [BP79a, DUK73a]. HCN [MT76]. H⋯ [MSG^{+}78]. head [Fle75, Leb73, Ran75]. headed [SY75, SYNY76]. heart [FSG^{+}72, MHB79, Nel78, Sal74, Tor76, WD72]. heart-cavity [Nel78]. heart-rate [WD72]. Heat [Cos75, RT77, Str71, Bak76, Bra70, Bra71, Fle75, Kle72, Sey72, BG74]. Heaviside [Shi71]. heavy [DMRD78, NB72]. hector [SBB^{+}74b]. height [CSL75]. helical [DDTG73, Den75b, DM78, MMT76, Zgi75]. helically [Swa74]. helically-wound [Swa74]. helices [Edw78, Lew72, Moo74]. helix [Edw78, Les74, Man75a, vHB76]. helix-Cruciform [vHB76, Man75a]. helix-forming [Les74]. Helmstetter [BC77]. hemeprotein [Cas71]. hemodynamics [NLC79]. hemoglobin [Bea79, Bla75b, UF71, Wei72, YS71b]. hemolysis [Mak76a]. hemolytic [DG75, Del75a, Del75b]. Henri [Cze71b, Via70]. heparan [CV76]. Hepatic [BRB78]. herbicidal [Gam75]. herbicides [GS78c]. herd [Ham71]. hermaphroditism [Hea77, Hea79]. herpesvirus [Bec78c]. Hessian [BW78b]. Heterogeneity [WZ72, Elt74, Fri76a]. heterogeneous [CB74, EH73, EH79, LK70, Sha73, Sha74, Sha77, Str77, TkC74]. heterotrophic [dKJS76]. heuristic [Bla78b]. hexagonal [War70a, War74]. Hierarchical [SR77, Jea79b, Par74b]. hierarchies [BB79, NPV77]. hierarchy [HC76]. high [Gün77, LS77, McC72a, RB77]. Higher [BBZ79, DCT73, Dou72a, Eld79b, KP72, LR76, Sat74, Sch77a]. hiking [MP79]. hill [BW78b, KZ79, Cho76, SYNY76]. hippocampus [Utt77a, Utt77b]. histamine [Bac74, FRG74, Kum75b, RHA75]. histone [SBM72, SBB^{+}74c, SBB^{+}74d]. histones [Lew70a, Ril74, Tem75, WW77]. histopathology [Fri74]. histories [Leó76]. history [Has78c, Van74a]. Hitch [MP79]. Hitch-hiking [MP79]. Hodgkin [Has78a, MR74b, SS70, ST76b]. Hofstee [CB76]. holometabolous [Rot77a]. Home [Coo78a, Coo78b]. homeostasis [Con71a, Rei70, Rei79]. Homeostatic [Goo72a, Bar72a, Goo73a, Woo70]. homeothermism [Cal76b, Gün77]. homeotherms [Øri74]. hominid [Pit78]. homogeneous [ZS72]. homologies [Les77, WW77]. homologous [Hol72, NO74, VPBD78]. homologs [JFR76, VSW^{+}79]. homology [Eng78, HT73, Mal79]. homotetrameric [Kel71]. honesty [Zah77]. honey [Ros78e]. hormonal [VD78]. Hormone [EFH77, AFB76, DW71, DS70, GK73a, Hil72b, Sac75, dH76]. hormones [Cil75a, Hol79, Hul78, Sho74b]. horopter [Nel77]. host [GM72, Hal73, Man75d]. hot [YS71a]. H − [MSG^{+}78]. Hückel [GKYG70, GS71a, KGYG70, SG71a]. Human [SSZ^{+}72, AA78, Can70, CS71, FSG^{+}72, Fle75, Gra73, Har77, Jok74, JBM^{+}78, KKW78, KH75c, Kit75, LCH78, LS77, MWM76, MTL76, MJ76, Pin72, Sha77]. humans [Rei70, Rei79, Ros78b]. humoral [Win76]. Huxley [Has78a, MR74b, SS70, ST76b]. hybrid [Ash72, DBU76]. hybridization [Lan77]. hydra [CM75]. Hydrated [Sco79a]. hydration [Fri73]. hydraulic [DC77a]. hydro [Hil72b]. hydro-osmotic [Hil72b]. Hydrocarbon [AA76, GGM75]. Hydrocarbon-nitrosamine [AA76]. hydrocarbons [BJM70, KGH79, VAA71]. Hydrodynamic [Bla74, Bla75a, HM71b, Kum74, Naz77, Sub70, EM78, KP77a, PST74]. hydrodynamical [CRC77]. Hydrodynamics [Wie79, Bla78a]. Hydrogen [WP78, Ami74, CP79a, CSVR78, HK76, JY72, Loe71, Poo75, YMM76]. hydrolysis [GG75a, GG75b, SRML79]. hydrolytic [Wer76]. hydrophobic [Bel77a, GC73, WL76]. hydrostatic [Gor78]. hydroxylation [TBP78]. hydroxyprolyl [MPP70]. hymenoptera [Ben78a, Mac78]. hyperbolic [CB75a, Koc72, Wal74b]. hyperplasia [Vog77]. hypoglycemic [Adl70]. hypolipidemic [Gam75]. Hypotheses [GF77, FP73, Ji76]. Hypothesis [Hed71, And79, Boy79a, BKB76, BS75, Dol75, GY75, Got72, GR73, Har76a, Hea74, HP77, Leó75b, LB72a, Lew76, Luc79, McK73, MGB73, MPHG76, Nag78, NM72, Ond79, Pap74, PMC74, Shi78b, Ste75, Sto70, Sto71, Teo73, Vig75, Woe70, Yoc74, Yu77, dO79]. hypothesized [Les73]. Hypothetical [Fed77, SH72b]. hypotonic [BB73b, BKB76]. hypotonic-sphere [BB73b]. hysteresis [Bar79b, Ede70a]. Hysteretic [Kur77, KDKY76a, KDKY76b, KDKY76c].

I-albumin [Sin75]. ial [CSL75]. ideal [HL70]. Identification [HPF78]. identity [CB77, CB78]. idiotypic [Seg79]. IgG [PW79]. IgM [Del75b]. ignoring [GH76]. II [BS79, And76b, Bar79a, BT71a, Bel71a, BK76, Bla75a, BHK78, BCLV76, Bun78a, CT73b, CC75b, Chu74b, Coh78, CR75b, Cra73, Cri75, DBS^{+}70, Dan76, DW75a, DG75, Den75c, DLV77b, Dou72b, EMM79, FT74, FY76, GK74, GPP77b, GY79, GKYG70, Goo77, GS71b, GG77b, GGM75, Har73a, Hua71b, KJ72, Kir79b, KDKY76b, Lad77, Mag74a, MO75a, MO75b, MA78, Mik77, Pai76a, Pai76b, Pap73c, Ran76, Rus74, Sha74, SYNY76, SG71c, Sin78, SG75b, SBB^{+}74d, SBB^{+}74b, SBR77b, Sta73c, Sto71, WT79]. III [AK78, BT71b, Bel71b, BK77a, CC75c, dC75, Chu78, Del75a, Hua73a, Koc71a, KDKY76c, MT78b, Pap73d, PGPP78, SG71a, Sav70, Sha77, She71, Sin73, SBB^{+}74c, WRT79]. illuminant [Bri78, Bri79]. illuminant-invariant [Bri78, Bri79]. illusions [LG76, Nel77, Nin79]. illustrated [LR76]. illustrations [vdSJ71]. imagery [BG75]. images [CK71]. imaginal [Ran77]. imagination [Tre77]. imagined [Ste75]. immiscible [KVB77]. immobilized [HSC73]. Immortal [Ler79, SS79]. immune [Coh70, DLV77a, DLV77b, JAS76, Man75c, Nos77, Nov76, Ros76d, Teo73, Wal73a, Win76]. immunity [Erh76, For75]. Immunodiffusion [GDA75]. immunogenetic [Woh78]. immunogenicity [HZ71]. Immunoglobulin [Nov75, NG77, Nov73]. immunoglobulins [ST77]. Immunological [HZ71, Bel71b, CFV75, Del74, Rap71]. immunology [Bre79, GL78]. immunoprotective [Med72]. immunoregulation [SR76]. impairment [Arc71]. imperfect [Har74]. impermeable [JWS76]. implanted [SLK76]. Implication [WD78, Har77]. Implications [Bal72, Dow73, Man75d, For75, GS78c, Jea79a, LR76, SG71c, Zuc73]. Importance [WMHG77, How74, MP73, Res78]. Impossibility [Hyv72]. improved [AAQ71, Bar79a, NSS74]. Improvement [Fis77]. impulse [Doe71b, Nol73]. Impurity [GR74]. IN1 [Ano76y]. inactivated [New79]. Inactivation [Gil76a, Sum79a]. inactive [Ger75]. Inbreeding [BBT70, Ano71x, Ben78b, Mic79, Wei71]. including [Tor70]. inclusion [RO74]. inclusive [Orl79a, Orl79b]. incoherence [Pra74, Rus73]. incompatibility [KP72]. incomplete [Olo73]. Incorporation [Seg76, FN76]. increase [Abe79, SH76]. Independent [Kre70, Edw78, Eva79, Har73a, Koc74a, MPH72, MA78, Nea72]. indeterminacy [Els72]. Index [Ano70j, Ano74k, Ano74l, Ano75k, Ano76i, Ano76j, Ano76k, CB78, GB77a, LH79b, NS76a, Øri74, SZ71, Sta79b, Woo78]. indicator [HS77b, Nor75b]. indicator-dilution [Nor75b]. indices [Rou79, Sch72a, Sch72b]. Indirect [Leb73, KKDY74]. individual [BB79, BBG70, Bli77, MWM76, Tre76]. indoles [CHH^{+}76]. induced [BS75, Cla77a, DVM79, DP74, Goo78, HKM79, JP79, KMS77, LM74a, Par75a, Rig79a, SP73a, Sil76, YZI77, You74, Zei77]. inducible [Sin73]. Induction [VAA71, WK73b, Coh70, Ker75, Luc79, WN79]. inductive [AG75]. ineffectiveness [CM78]. inequalities [OD71]. infection [MYF77, SBC79]. infectious [Med72]. inference [Fag78]. Infinite [Mar78b, KH75c, PN73]. inflammatory [GMR75, LK79]. Influence [Vai71b, Vai71a, Bal73, Hew77, KHP77, Lat75, LS72b, MGNR78, MB78, PPS75, SATT79, TBP76, Win75]. influenced [OK76c, Rig79a]. Influences [Kir79a]. influenzae [Ran76]. Information [Ebr75a, Erh76, Fag78, Rya72, Wic78, And77, Bal73, Ber77a, Ebr75b, Eds75, Fra74b, GZ76, HGR79, Hew78, Hol79, Mac77b, Mad77a, Mik75b, Mur74b, Orl70, RHR79, RCW73, Ros76e, RBM76, Sch71, Sie72, Sug75, TS71a, Tor70, Tsu79, Vol79, Wic79, Wri77, Yoc74, Yoc77a, Yoc77b, Yoc78]. information-theoretical [Wic79]. Informational [FSST78, RL76]. informon [Utt70, Utt75a, Utt75b, Utt77a, Utt77b]. infrastructure [Wel77a]. infusion [AW70]. inheritance [Dow73]. inhibiting [Ond79]. Inhibition [LR76, Ain75, BM79, BC73b, BC76, CB75b, CR75a, CR75b, CT79, Del75b, Dol75, Har71, Har73a, KS72, Mac77c, Rak77, Rak78b, Rak78a, RG78, Sav79, SJ79, STA73a, VD78, DG75, Del75a]. inhibitor [Bar79a, Rak78a]. inhibitors [Har78, Smi78a]. inhibitory [SC75b]. Inhomogeneous [Lap79, GN75]. initial [Ain75, CSL75, CB76, HH77b, Mar75c, Ros75a, Ros75b, Zer75]. initial-rate [CB76]. initial-value [Ros75a, Ros75b]. Initiation [KS70, Ble71, Bro74, Coo74, LB77, Tor70]. Initio [GPP78, HR79]. injected [Nor75a, TV72]. injections [Sho74b]. injury [DBS^{+}70, NP79]. inner [Ciu75, LS77]. inquiry [Spe76]. insect [Dia74, LBGK76, MWH73, Mit79a]. insecta [WC75]. insecticide [Com77a]. insight [Orl75c]. insolubilized [SAT72, STA73a]. inspection [Lai77]. Instabilities [Blu75, Ede70b, BN72, LNHK71, LN71]. Instability [OS71, SS72, SIS78, BD75a, KS70, Mar75c, PWW76, SG76]. installation [KDKY76b, KDKY76c]. instance [Ove77]. Instantaneous [KWG79]. Insulin [YM78, Atk71]. intact [BKWT76]. Integral [WB77b]. Integrated [Win70a, Gra73, SF79]. Integrating [SR72]. integration [Chr76, GM72, MR74b]. intensity [Tsu74]. Inter [TT75, Lin79c, Tha73]. inter-coupled [Lin79c]. inter-enzyme [Tha73]. Inter-tropomyosin [TT75]. interacting [CNS79, Mur75, NOB72, SKT79, VAC76]. interacting-population [Mur75]. Interaction [Abe73, Har78, CW74, DHK78, DH78, Gar76, Gin71a, HKAT77, HK74, JO77, JA75a, JA75b, KA74, KH75a, KH75b, Les70, Les78, LS78, Man79a, Mor72, Ohs77, Ord76, PM76, PR76, SBM72, SBB^{+}74c, SBB^{+}74d, SBB^{+}76, TD72, UIN74, VB78, WFW75]. Interactions [Ber72, CS75b, Lei70, SO79a, APA79, Ano75a, BS79, Bat72, CT73a, CT73b, Chu74a, Chu74b, Chu78, Ehr79, Eps71a, Eps71b, FJThH73, GP72, Hil78a, KS77a, KD76, KKDY74, LL70, Lev77, Lin71, LST74b, NB72, Nov73, PPS75, PB78, RT75, RRS72, Rot75, SFT75, SG73, Tar72, Vas74, WNHS75]. Interactive [And76a, And76b, BHK78, Got75, Sta73b]. intercalated [MHMS75]. intercellular [Har76a, KH79a]. intercistronic [FM72]. interconversions [PB77]. interface [KVB77, OO76]. interfacial [DH78, Sue78]. interfilament [Hil70, Lub73]. intermediary [Wel77a]. intermediate [Bye78, Dar77a, Ji76]. intermediates [Ain74a, Cil75b, SJ79]. Intermolecular [GPP70b, GPB71, GPP70a, MRTG79, Nir75, RRS72]. Internal [Bel77b, Doe71a, EH73, ON77, RHA75, WP71, WW77]. international [Ano75a]. Interplay [EFH77]. Interpretation [CB78, MJ71, Rap77, WSF76, Ain75, BW78a, BT71b, Cra73, Cro79, Jaf74, KNJ74, MP78, RLS73, SP73a, Sug75, TR79, The71]. interpreted [Bar79b]. interrelations [DLV77b]. Interspecific [CM72, PS70]. interspike [TR78b]. interstitial [KKJ77]. intertidal [JS79]. intestinal [Win70b, Win75, Zaj77]. intestine [TD77a]. Intracellular [Sum79a, Wig72, Wig73, Gla79, Kle75b, KKJ77, Wal73a]. Intragenic [Cla72]. intramolecular [Bec78c, CC70]. Intravascular [Rei79]. Intrinsic [Naz77]. intruder [CM70]. invadability [RV79]. Invariance [PMN73]. invariant [Bri78, Bri79]. invention [All77]. Inversion [TD76a, Dea72, DT74a, IM78]. invertebrates [Kow71]. Investigation [Atk71, Sim79, Fon77, HR71b]. investigations [MPD70]. invoking [Orl75a, Orl75b]. involved [HKM79, MT73, Wen77]. Involvement [Nic76, Hol79, MB78, WJ78]. involving [Ain74a, Arc71, BM78, Dar77a, Har73b, Isr73, KNJ74, Mar75c, Res76]. invulnerability [RV79]. iodine [GW70]. iodo [Cil75a]. Ion [BK71, Buy73, CA77, FL74, MW78, Abe73, BJ72, Coy79, EM78, Gil76b, Hua71a, Hua71b, Hua73a, KOH70, Smi73, TBP76, Wig71, dLA76]. ion-movements [BJ72]. ion-translocation [TBP76]. Ionic [VG79, DVM79, Hai79, Jor74, Jor75a, MMS74, NP71, Off74, OK76b, Oos73, Smi73, YK71]. ionising [Cra77]. ionizable [CW74, NP71, PR76]. ionized [MB78]. ionizing [NB72, Res78]. ionophore [SBB^{+}75]. ions [App72, KVB77, Nor75a, TV72, Wer76, ZS72, dLA76]. iontophoresis [SB73]. iontophoretic [Pur77a]. iontophoretically [Nor75a, Pur77b, TV72]. iron [DeV76, Miz77]. iron-sulfur [DeV76]. irradiation [Leb73, Ran76, VFKL74]. irregular [Sea76]. irreversible [Rak77, Rak78b, Rak78a, Str77, Tyr70]. Ising [GR74, YS71b]. iso [Gin71b]. iso-osmotic [Gin71b]. isocline [SG73]. isoenzyme [Bat72]. isoenzymes [Pal70]. Isolated [KOR^{+}77, AK76, PMC78]. isomer [CMSC77a]. isomerase [Ano71f, Cze71a]. isomers [Kam74]. isometric [Mit75, Hil70]. isometrically [Mil78a]. isopotential [HC78]. isothermal [HJ75, Hag76]. isothiocyanic [LM75a, LM75b]. isotonic [BB73b]. isotonic-disc [BB73b]. Isotope [Ain77a, Yag76, BW76, CS75c, Dar73a, Dar75, FC71]. isotope-exchange [Dar75]. isotopes [NB72]. isotopic [Fle72]. isotropic [GCG^{+}70]. issues [BW78b]. item [Ano76l]. Iterative [Gil72, MGB73]. IV [Ano71f, Cze71a, Del75b, Hil78a, Koc71b, Kur77, Lau72, MPP70, MA78, MRTG79, TNG74].

J [Ano70d, Ano71f, Ano71x, Ano75b]. J. [CB79a]. jacket [HC76]. Jacob [BN72]. Jahn [Miz77, Wei72]. January [Ano70v, Ano71n, Ano72k, Ano73j, Ano74n, Ano76r, Ano77m, Ano77s, Ano78i, Ano78o, Ano79j, Ano79p]. Jerne [Nov76]. Jerusalem [Ano75a]. Jones [Jor73b]. July [Ano70k, Ano71m, Ano72l, Ano73k, Ano74r, Ano75p, Ano76z, Ano76n, Ano77l, Ano77r, Ano78m, Ano78s, Ano79k, Ano79q]. junction [HL79, KHP77, KH79a, SBB^{+}75]. junctional [Shi71]. June [Ano71t, Ano72s, Ano73v, Ano74x, Ano75t, Ano76t, Ano77z, Ano77-31, Ano78-27, Ano78-32, Ano79z, Ano79-32].

K-12 [DT79]. K-permeability [KM75]. Kacser [Kac76]. Kalanchoë [KJ72]. Karyotype [IM78]. Karyotypic [Tod70]. katydid [Sou75]. Kauffman [Lav77]. ketosis [Adl70]. kidney [Abe73, BB77a, DMRD78, Lin79a, WN79]. Kin [Joh79, Orl77, Cha77, Mic79, Orl75a, Orl75b, Orl75c, OW78]. kin-selection [Cha77]. kinase [Kan73, SD76, Wel77b]. kinds [Bri74, RS78]. Kinematics [SE79]. Kinetic [CS75c, Des77a, Dun74, FC71, GK71, Hua71a, Hua71b, Hua73a, KKDY74, MG72, WFW75, vHB76, Alb74, BW76, Blo77, CKJ79, dC75, dC77b, CBPT78, Dar73b, Dar75, DS74, Dug79, EJB^{+}78, EH73, GL78, Han72, Koc71a, Koc71b, KZ79, KDKY76a, KDKY76b, KDKY76c, Kur77, Lym74a, Man79a, MR74a, MA78, MAM79, Ned75, PTB74, RZMF75, Ros75d, Ses72, Sha77, SG75a, SG75b, Sta76, ST76b, TD77a, Vas74, Via71, WPTF77, Whi79, YZI77, DLV77b]. Kinetics [Dar73a, DWS76, HJ75, Har71, Har73a, HK77a, LK70, Rak77, Rak78b, Rak78a, RF73, Sam71, ARW73, Ant74, AMR^{+}75a, AMR^{+}75b, Bar77a, BP73, BP74, CB75b, CD72, Dar77b, DW75b, DP76, DG77, DG75, Del75a, Del75b, DCMR74, EH79, Fal77, FPB72, Fle72, Gar70, GC76, Goe74, GGSW72, Got75, HM71a, Han78, Hil78a, HC78, JP79, JBM^{+}78, KBW75, KRH77, KNJ74, Lan77, LRB71, Lev72, LS72b, Lin76, LSGFdlT^{+}75, McE78, Mor72, MM78b, Nea72, OD73, Pop79, Rig79a, RBP78, RMG77, SH72a, Seg76, SD77, Sta73c, Ste76b, ST76c, SD76, SM71, TE76, TBP76, Via70, WSF76, Whi78, Win70b, DW75a, Kur77]. kinetics-Cyclic [WSF76]. kinship [MSK79]. KNF [BW78a]. knowledge [You76]. known [Gau73, War74]. Krebs [Gar70]. krogh [Lev72].

L [HS75, Lin75, BCDK77, CDS77, Lew72, Jea79a]. l- [BCDK77]. L-systems [HS75, Lin75, Jea79a]. L. [SS73a]. label [Fri76b, Sch72b, Sch74, SM78]. labeled [SZ71, Sha74, TW77]. labeling [BHD74, JvS77, RP77, Sch72a, ZV77]. labelled [HH77a, KRH77]. labelling [CB75b]. laboratory [BB73a, MG77b]. lac [Ege79, JO77, Man79a, SB71, BB77b]. lac-operon [Ege79]. lactam [HM71a]. lactate [Dal79]. lactic [Pal70]. Lactose [Lan78, Ege79]. lag [Cas72, Mac77c, ST79]. Lagrange [TM74]. lags [BV75, Maz76]. lake [NS78]. lambda [EJB^{+}78]. lamella [San74]. lamellar [Mil79]. lamina [Rud79]. lampbrush [Leó75b]. lander [LSB78]. lands [Kri77]. language [Ros78e, SD76]. languages [Lin71]. Laplace [KV74, Orr79, SH72a]. large [Kau74, KHK79, McM75, MJ76, Nin79, Sam77]. large-neutral [Sam77]. Latency [Har76b, WB74a]. latent [RP76]. lateral [CR75a, CR75b, Spe78]. Latin [Ano78h]. lattice [MO75a, MO75b, OMF72, TD72]. lattices [AK78]. law [Kle72, Lav77, MHC79, Sav70, Spe70]. laws [Jok74, TB72]. layer [Des77b, NR71, PF78, PR76, SS72]. layered [PN73]. layers [LS72b, PWW76, TD77a, You74]. leading [Bec78c]. leaf [LBGK76]. leak [Arc71]. leaky [KHP77, KH79a]. learned [GZ76]. Learning [Eij70, App72, Con74a, Mro73, Pfa73, Tre77]. least [BH79, LL79, PMN73, SB77]. least-squares [SB77]. leaves [Par77, SS73a]. lecithin [BCDK77, FK75, TW76]. lecithin-cholesterol-water [FK75]. left [Nol76]. leg [Ran77]. lemma [Koc71a]. length [HPF78, Hon71, Kra78, Mit75, Naz77, SZ71]. lengths [Den75b]. Lennard [Jor73b]. Lennard-Jones [Jor73b]. lesion [Cra77]. lethality [Cra71, Cra73]. lethals [MHY77]. letter [Gol76, Lon76]. leukemogenesis [Ros78b]. leukocyte [LK79]. leukocytes [Ros76a]. level [Ale72, ABCT75, Leb73, MV75, War70b]. levels [BBZ79, Bea79, DW75b, DW75a, MHB79, NB76, Rig79b, Sho74b, Sho74a, SC77]. LH [Sho74b, Sho74a]. Liapunov [Gil74]. lichen [AH79b]. Life [BH70, Leó76, CSIW72, CC75d, Duc77, Goo73b, Has78c, Lei78, Mik75a, Rot77a]. lifetime [SD75]. Lifshitz [NRW72, NRW73, PN73]. ligand [APA79, Ano75a, Bla79, CNS79, CB76, Cho77, Chu78, KO78, MSF71, MS71, NOB72, Rap79, ST78b, SKTC78, Ste74, Via71, WD73a, Chu74a]. ligand-binding [CB76]. ligand-triggered [Bla79]. ligands [AG78a, Cho76, Han78, Hew77, MLM77, NKB72, Rap76, SOF72, TB72]. ligases [FC76]. Light [You74, Cri77, Fon77, IE76, Men76, Pap71, Sam71, WP72, You76, dS75]. like [CA79, FIN74, Hon71]. limb [Mad77a, WE75]. Limit [Leó75a, AG75, Dun74, Ish77, Jon71, Sch79c, TS78b, Wal70]. limitations [Ebr75a]. limited [Coh71, Lei78, MSK79, Nyh78, Ros73c]. limiting [Ana73, CC70, Coh71, RS78, Van76, Wil75a]. limits [Con74b, CBPT78, Fag78, Sch70, Wev72]. Limulus [CR75a, CR75b, LF77]. Line [LG76, Ber77b, Nie71]. Linear [Bun78b, JS79, ST79, BW78a, Bas73, Bec78c, BH79, BD70, Bun78a, dC77b, CB76, Cza73, GK73b, GN76a, Han72, Hyv72, Kem79, Koo77, Kub70, Lin79b, Mar76b, MWS77, Mik77, MJ71, MK74, NGP77, Ols79, OS74, Per76a, PL77, RS77, Sch72c, See70, See71b, SG71b, VSW^{+}79, Wal74b, WD73b, vdD74]. linearity [Ada76]. linearization [SM79]. lines [SS79]. linkage [Jor73a, Jor73b]. linked [BBT70, FH71a, KK73, WD78]. lipid [DST79, MM78b, Ohk70, Sco74, SFT75, Sue78]. lipid-protein [SFT75]. lipids [Hai79]. lipophilic [WMHG77]. lipoproteins [SA70]. liquid [Huk78]. liquids [KVB77]. lived [Orr79]. liver [Bar78, Bar79a, BKWT76, BB77a, BR79, DMRD78, FY76, Gar70, Poo75]. living [Age72, Koz79, McC71, MRTG79, Tas70, Wig71, ZW70]. lizard [BSM77]. loan [Con79]. Local [BB73a, Fuj77, Par75b, Par75c, Fis73, LS76, NP79, STU78]. localization [Cla72, GK72, Ima75, Ima76, LK77b, Woo72]. Localized [SO74, FP72, Ji76, TE76, Ven76]. locally [Van73]. location [WK78]. loci [BBT70]. locked [AG79]. locomotion [PB77, RMF78, WJ72]. locus [Ano71x, CC75a, HL79, TFB74, Wei71]. logic [MHB79]. Logical [Kur71, Tho78, GK73b, Sug75, VL78]. logistic [Wit78]. Loligo [Str76]. London [DM75]. Long [Nir75, WNHS75, HKM79, KS77a, MGNR78, NR71]. Long-distance [WNHS75]. long-range [HKM79, KS77a]. long-time [MGNR78]. longitudinal [Don72]. Loop [GM72]. loops [RB77, Tho78]. loss [BG74, Cra71, Cra73, Goo72a, NMK78, Ste78, Str71, WT77]. Lotka [Gro78, HB70, Jor77, So79b]. Low [ABCT75, Leb73, Mar79a, SA70, Shi71, YMM76]. lower [Ada76, Bli77]. LSD [Nic76]. lumbar [Odu77]. lumen [CD75]. lumped [She71]. lung [AA76, HDP^{+}78]. Lyapunov [dS75]. lymphocyte [DP76, GGS78, GR73]. lymphocytes [CS71]. lymphoma [IHI79]. lysine [LSL77]. lysine-triphosphate-arginine [LSL77]. lysozyme [LT72, Tho74a].

M [Ano75b]. Mach [Ciu77]. machinery [GY75, GI77]. machines [BNT71, BT71a, BT71b, Lai75, MN73, McC71, Vis77]. macro [Gin77b]. macro-equation [Gin77b]. macromodel [BB73b]. macromolecular [And76a, And76b, Oos75, Sim73a, Sin76, Woe71]. macromolecules [Eps79, Lew71, NP79, PB79, RRB76, WD73a]. macronuclear [McC78]. Macroscopic [HL70, Bun78b, Bun78a, Cha72, LT77, Lay78, Nir75]. magnesium [Wal73b]. magnetic [BJM70, Rob70, Yor79]. Magnolia [BY75]. Maintenance [Con71a, Heu78, LB72a]. major [KC78]. maker [Tor76]. malaria [Pow75]. male [Dal79, PBS72]. male-female [PBS72]. malformations [MHB79]. malignant [Bha77, Kau71, TR78a]. malonamide [Goo78]. malonamide-induced [Goo78]. Malthusian [FJThH73]. mammal [WK78]. Mammalian [NLC79, Alb75, Alb77, AHH^{+}76, BP79a, Bec72, DB78, DBS^{+}70, Goo73a, Goo78, HC73, HTSH72, HMT79, Ima75, Ima76, JA75a, JA75b, RBSS70, Ros78d, Sha73, Sha74, Sha77]. mammals [Dal79, Ima78, Wun75a, Wun75b]. man [DS70, KTS71]. management [Com77b, Com79]. manifestation [Ber77b]. manifestations [Bar77b, BW78c, BFM75, KKDY74, KDKY76a]. manipulation [Mil78b]. manoeuvrability [How74]. many [PN73]. many-layered [PN73]. map [Hej75]. Mapping [SBR77a]. March [Ano70u, Ano71w, Ano72t, Ano73u, Ano74m, Ano75o, Ano76p, Ano77o, Ano77u, Ano78l, Ano78r, Ano79o, Ano79u]. Margalef [Mat78]. margin [Olo73]. Marginal [Han76, Ish73]. marginal- [Ish73]. marginotomy [Goo77, Olo73]. marine [Fal77, RL76]. markers [Hit76, TD77b]. Markov [Boy79b, NPV77, TV75]. Markovian [CR73]. Mars [LSB78]. masking [GH75, WP78]. Mass [AW75, BB77b, Kra78, May73, Ula72, Blu77, Der78, Lei77, Per76b, RG78, SR72]. mast [Wal79]. Matching [Bar79a, RCW73]. Mate [Zah75]. materials [CJ77, CP76, NB72, PC76]. maternal [KS71c]. Mathematical [AA78, Bel70, Bel71a, Bel71b, DLV77a, DLV77b, EGF77, KBW75, KP77b, KTS71, Lau72, MA71, Nos77, Pfa73, Rot77a, Ber76a, BB73a, BKB76, BMS77, BLP76, CE78, EH70, Fis73, Fra70, GKS78, Hat73, HLR73, HLR74, IE76, LBGK76, Mac77d, MT74a, MG77a, NMK78, Niv77, Nol73, Ode76, Ott74, Rap78, Sin75, Smo76, SR76, WSF76, Wil70b, JK72]. Mathematics [HLR73, HLR74, Kil72]. mating [Har77]. matrices [Whi78]. Matrix [Lav76, Kib79, Tre79, UI73]. Matter [Sme76, Bat74]. maturation [BSH75]. Mauve [Rus72]. maximal [Duc77]. Maximizing [Coh71, Klä75]. maximum [Den75a, Gil73b, Jon71, MBG73]. Maxwell [McC71]. May [Ano70p, Ano71p, Ano72n, Ano73q, Ano74q, Ano75m, Ano76s, Ano77n, Ano77t, Ano78n, Ano78t, Ano79n, Ano79t, AF77, Ber76b, Ber77b, Jon71, Man79b]. mays [GL73, SS73a]. mean [BH79, MTL76, RLS73, Sch74]. meaning [Kle75a]. means [CMSC77b, Ril73, ZV77]. measure [Lax78]. measured [RLB74]. Measurement [LS72a, Bar72b, Com75, MG77b]. measurements [Hay74, Mur74b, RLS73, Ste70, Via71]. Measures [Dem76, Fin76b, Fag78, LD75]. measuring [RP77, SHC71]. Mechanical [Lun74, GPP74, GG77a, GG77b, Gre78, HR79, JY72, KG74, KS77b, KD76, LST74b, MK76, Mof70, Sue78, Woe73]. Mechanics [Kel77, APK75, DS73b, HB70, HO78, KD79, LST74a, OF75, SS73a]. Mechanism [Gam75, JAS76, Sub77, WL72, Adl70, AM75, Bec72, Bla77a, CS75b, Coh70, Con71b, CM78, DBS^{+}70, Dar73a, Dar77a, Dra74, GMR75, GGSW72, GW70, GN76b, HO76, HLC75, HZ71, Hol79, JD77, Kro74, LRS79, Luc79, LM74b, MV75, Mar75c, MC73, Med72, Men76, Mik75b, Moo74, Ond79, OF77, Paa78, PWW76, Pog71, Ros78c, Sch77a, SRML79, SBM72, SBB^{+}74c, SBB^{+}74d, SBB^{+}74b, SBB^{+}76, Sta70, SK79, Sun77, TT75, TS71b, UINH73, Wal74a, Wal73b, War70b, Wer76, Wes74a, Wes78, WK73a, Yag77, vW76].

Mechanisms [RLB74, WB73, Zgi75, Ain74b, Ain74a, Ain77a, Bac74, Bar72a, Bec78a, BP79c, CB76, Fin76a, Goo72a, Goo73a, HS70, Kel71, Mit76, Ols75a, Ols75b, PB78, RF73, Sco79b, Sco79c, Smy72, Tea76, WN79, Whe75, WB74a, Zei77]. mechanistic [Ain75]. mechanochemistry [Opl72]. media [Dur73, Fis77, PN73, Str77]. mediated [Cro79, LS72b, Ols75a, Ols75b, Sum79a]. mediation [Coh76, KMS77]. medicine [Hue77]. medium [Bla77b, Ciu75, DUK73a, DUK73b, MC72]. megagametophyte [Sch79d]. meiosis [Hol72, TD76b]. meiotic [Sub77]. melanin [MP73]. melanized [PMC74]. melanocyte [WG70]. melograph [Hjo70]. melting [Dra71]. Membrane [Ano79h, DP77b, Gin73, LHC76, OV78, PA71, SD74, WB74b, You73, AK76, Alb71, And79, AO72, ABR72, AR75, BL75, Bec78b, Bla79, Blu75, Bon78b, BC78b, BHK74, BHK78, BK71, Buy73, Chr76, Cla76, Cla77b, CS77, CJ76a, CS75c, Dow70, Dra71, DP74, DM78, EFH77, FK75, Fri76a, GS78b, Gil73a, Gil73c, Gil73b, Gil76b, Gil76a, Gin71a, GK71, Gou76, GKS78, Gri77, Gri79, GW70, HC75a, HC75b, Hea74, HO74, Hew77, Hew78, HPMH76, Höf71, Hol76, Hua73b, KFH74, KVB77, KK78, Kil70, Koo77, Lak75, LS76, Lev74a, Lew76, LS77, LST74b, Mad76, Mak76b, Mar77, MW78, Mar76b, Mik77, MR74b, Nag78, Nea72, NCM75, NM75, OK76b, Ohk70, Ohk73, Pap74, Ros79b, RWD76, Rud79, ST76a, SO79a, SIS78, Ses71, Shi78a, Sim73b]. membrane [SS73b, Smi73, Spe76, SH72c, Sue78, Tag78, Vig75, WJ78, Wie79, You74, ZS72, dLA76]. membrane-chemical [SIS78]. membrane-covered [Alb71]. Membrane-soluble [dLA76]. membranes [Alb74, Ana73, AHH^{+}76, AR75, Bas71, Bas74, Bun78b, Bun78a, CC73a, Che79, CA77, Cla77a, CJ76b, CMV76, DP76, DC73, DST79, Fle77, Gre76a, Hai79, HKM79, HTSH72, KH79b, LM74b, LS78, MV76, MT78b, OK76c, Ohk70, OO76, Pat70, RSS^{+}73, RBP78, Sco74, Shi71, SFT75, Spe78, TBP76, TD72, VG79, Vai71b, dLA76]. Memory [PAZ79, App72, Con77, GL75, LP75, Mik72, Mik75b, Mik78, PL77, RRB76, STU78, Tor70]. Menten [Cho76, HM71a, McE78, Pop79, Sta70, SF79, Wal74a]. meristems [Bar76]. mesenchyme [Web70]. mesomorphic [Rot73]. mesophyll [Par77]. messages [Ril73]. Messenger [ABCT75, Abr71, Bal72, Bal73, Blu74b, FSG^{+}72, Fit76, Klä75, Rig79a, ST77, SG71c, Sin73, VNB77]. Metabolic [Kle75a, Wel77a, BW77, Bod78, Eco79, Har78, Has74, Hor74, New72, Par74b, Par75b, Par75c, TBP78, Wat72, Wun75a, Wun75b]. metabolism [Bak76, Blu77, CJ77, CP76, Der78, EFH77, HPMH76, HPCH79, KKW78, Koc71b, LR76, Mac77d, MGNR78, Mil78b, PC76, Poo75, Pop79, RZMF75, Ros78d, SA70, SB78b, TD72]. metabolites [KGH79]. metabolizing [VAA71]. metal [Ano75a, Wer76]. metal-ligand [Ano75a]. metalloenzymes [Wer76]. metaphor [Mac73, Mac75]. metastable [Wig75]. metastatic [SLK76]. metazoa [CJ77, Rud79]. meteorites [Ree73]. methane [HWD76]. methanol [Ami74]. method [AE79, Bar77c, BH79, CSL75, Eng78, Fri72, GRY^{+}72, GKYG70, Har73a, HSK72, HM72, KGYG70, KHK79, Koo77, Kum75b, MWS77, Mik77, MT78b, MBG73, MHB79, Nol73, SB77, SH72a, Ses72, SR72, TNG74, LTP78]. Methodological [VtM79]. methodology [Cza73]. Methods [LL79, Utt77a, BW76, Com79, DDTG73, Den75c, GC76, Gil72, HR79, JBM^{+}78, KP77b, LK76, Nor75b, PMN73, SG73, VL78, WB75]. methyl [RHA75]. methylation [SK79]. methylimidazole [Ami74]. Methylococcus [HWD76]. metrics [RR79]. MgATP [SRML79]. mice [WG70]. Michaelis [Cho76, Cro75, Cze71b, HM71a, McE78, Nea72, Pop79, Sta70, SF79, Via70, Wal74a]. micro [Bra70, Bra71, CH72, Lap78, ON77, PC76, Sin76, Sin78, dKJS76]. micro-organisms [Bra70, Bra71, CH72, Lap78, ON77, PC76, Sin76, Sin78, dKJS76]. microbes [CJ77]. Microbial [JDTF73, Ber78, VAC76]. microelectrode [Alb73]. Microelectrodes [Pur76b]. Microevolution [Kur71]. Microfluorometry [ES77, EMM79, KP77b]. micrographs [FH71b, San74]. micronucleus [FCSS78]. microphonics [Yos77]. microscope [CK71]. microscopic [Nir75]. microscopical [Ano79h]. microscopy [AAQ71, BBGR71, Fri76b, GBH70]. microsomal [VAA71]. microthermal [Leb73]. Microtubule [Ate73]. microtubules [Hea74]. microwave [Leb73]. Migrating [ZS72, AP77, Ale79]. migration [AK78, BHK78, CSVR78, Com77a, Han78, Jor74, MM78b, NW74, Nos77, RG78, Ros75c]. mimicry [CC75a, CC75b, CC75c]. MINDO [KGH79]. MINDO/3 [KGH79]. mineral [McC75b]. minerals [CSIW72]. miniature [KT72b]. Minimal [PO79, HPF78, SBR77a, DP76]. minimization [SB77]. minimum [Can70]. minor [SATT79]. Mispair [ZG75]. missing [Moo77]. Mitchison [Rob78]. mitochondria [FY74, FY76, Mar70, Ost77]. mitochondrial [Arc71, Mey73, Oli77, Sto70, Sto71]. mitochondries [MCB77]. mitochrondria [You73]. mitoses [TW77]. mitosis [Eld79a, FGP^{+}75, HH77a, KRH77, Ond79, Sha74]. mitosis-inhibiting [Ond79]. Mitotic [WKG73, Whe75, Bar79a, Con71a, Con71b, Doe71a, Goo72a, KW75, Ond79, SZ71, SG76]. mixed [Cas74, Goo72b]. mixing [Bel77a, GC73, Lew73]. MO [TSP78]. mobile [Gin73]. Mobility [Hua73b]. Model [Coo70, Gin71b, Goo70, IP71, KS71a, LB71, LB72b, Ran77, Sou75, Woo70, Abr71, Adl74, Adl75, Adl77a, Ain77b, AG78a, Ain79, Alb75, Ama74, Ami74, AH79a, AC77, AR75, ASS79, Bar74b, Bar78, Bar79a, BY74, BRB78, Bat73, Bea79, Bec78c, Bel70, Bel71a, Bel71b, Bel77b, Ber78, Bes70, Bha77, BD70, Bla77b, Bla75b, Bla78b, BB73a, Blu75, Bon78a, BMS77, Bra75, BHK74, BCDK77, BH78, BC73b, BLP76, Cal73, Cam77, CR73, CC73a, Cas72, CP79a, CE78, CT73a, CRC77, CKJ79, CV76, Cla77a, CS73, CGM79, Coh70, Coh78, CBS75, CE77, Con79, CZ76, Cri77, Cum75, CG75, Cut72, Cza73, DBS^{+}70, DS73a, Dan76, DW75b, DW75a, DO76, DG77, Dea72, DT74a, DC77a, Des77a, DLV77a, DLV77b, DHK78, Dow70, Doy78, Dra71, DS74, DT79]. model [DM78, Ede70a, Ede71b, Eld73, Eld79a, EWRA76, Els72, EJB^{+}78, ER74, FD73, Fis73, FSFD70, FK75, Fox78, Fra70, Fre70, FC76, FM72, Fuj77, FY74, GS78b, GPP74, GM72, GB75, Goo77, GL79, Got75, GG77a, GG77b, GR75, GL73, GR74, Gri79, Gup75, HK76, HL79, HJ70, Har73b, HZ71, Has74, Has78a, Has78b, Hat73, HR71a, HC77, HDP^{+}78, Hej75, HO74, Hew79, HKAT77, Hil70, HPMH76, HS77a, Höf71, HK74, HKK76, HH77c, Huk78, Hun74, Hut73, IPH^{+}79, IE76, IHI79, Jac77, JRP74, Ji76, JK72, JA75a, JA75b, Kac76, KFH74, KBH76, Kan75, KS72, KJ72, KS74, KKW78, KG74, Ker75, KPU78, KH75a, KH75b, Kig73, KH75c, Klä75, KV77, Kou73, Kre71, KP72, Kum74, Kum75a, KDKY76a, LH78a, LH78b, Lad76, LD73, Lan78, LS72a, LBGK76, Lei77]. model [Les78, Lev72, Lin79a, LSGFdlT^{+}75, LR74, LB77, Mac77a, Mac77c, MO73, MP78, Mad77b, Mak76a, MPH72, Man79a, MR74a, Mar72a, Mar78b, MT74a, MF79, Maz73, Maz78, McC72b, MCM79, McN79, MCO77, Mil79, MM78a, MRTG79, Mit75, Mit79a, Moo71, Mor72, MG77a, MMT73, MT79, Mur76, Nag76, NSS74, NOB72, NMK78, NB76, NG77, Niv77, Nol73, NS76b, NR71, Odu77, OO76, Ohs77, Ols75a, Ols75b, Ols79, Orl75a, Orl75b, Ove77, Pai76c, PST74, Pan78, Par77, Par73a, PB77, Pol73, PL77, Pow75, PT75, PTB74, PR77, RRB76, RSL79, RAJM76, Ran73, Rap78, RBSS70, Res76, Ric75, RS77, Rig79b, Rob77, Rob78, RZMF75, RC77, Ros75d, Ros72b, Ros73b, Ros73c, Rot79, Rub77, SS70, ST76a, Sch79b, SFN77, Sco74, Sco79b, See71b, See71a, See73a, See73b, Seg76]. model [Sey77, Sha73, Sha74, Sha77, SCDC77, Shi71, SY75, Sho74b, Sho74a, Shu79, Sib78, Sim73a, Sin75, Sma76, Sou71, Sta73b, Sta73c, Sta76, ST76b, SB75b, Str76, SR76, Sub70, SM71, Sza77, Tag78, TSP78, TKO74, TT75, Tay79, TE76, TC79, Tho77b, TIO77, Tor70, Tre75, Tre78, Tre76, Tuc79, TS79, Van77, Vas74, WB77a, WD72, WSMS77, WT77, WPTF77, WJ78, WE75, WD73a, Wil70b, WTR75, Wil75b, WJ74, Win76, Woe73, WD73b, Wun75a, Wun75b, YS71b, ZV77, dH76]. modeling [MWS77, MWH73, RN76]. Modelling [Bat72, GS78a, NS78, Dea75, HH77b, Lin79b, Lin79c, PM77, SR77]. Models [Gri77, Kow71, RMF78, Wal79, AA78, Alb74, Atk71, APK75, BN72, BM78, BW78a, Bar74c, BNT71, Bed74, Ber76a, Bol73, BP76, Bra77, BHK78, CC75a, Cha78b, dC75, Che75, CR78, CMSC77b, CG73, CB74, Cra77, CP72, De 75, DP76, Elt74, EH70, Fis76, FPB72, Fit73, Gar76, GH75, GRDC72, Has78c, Her70, HS75, HC78, HTSH72, JR77, JW76, JS79, KC78, Kib79, KA74, KZ79, KKDY74, LH79a, Lai77, LA76, LK76, Llo77, MS71, MG72, MHY77, Maz76, Mil74, MK74, NPV77, NGP77, NM78, OW73, OB73, Pfa73, PT79, Ran76, RD77a, Rig79a, RD77b, She71, SS73b, SG75a, SG75b, SM79, Ste76a, SW74, TW77, TFB74, Tre77, Tuc77, TS77, VD78, VB78, VAA71, WSF76, WT79, WRT79, WFW75, Wit78, ZW70, vdD74]. modern [Fed77]. Modes [MTM77, Mar71a]. modifiable [Bri74]. Modification [GW70, Fri72, Kel71]. modifications [Doy78]. modified [BS78, CC73b, HM79, KZ79, Lin79a, SB75b]. modifier [Kat72, Lym74b, WB73]. modulating [Ros79a]. Modulation [Lin79b, Lin79c, Rib79, Sch78]. moieties [Ano75b, Par74a, Par75b, Par75c]. mold [KS70, Nan76]. molds [CR71a, CR71b]. Molecular [GK74, GPP73, GGM75, JI78, LL70, LT72, Mag70, MPD70, MPP70, MPP71, Smy72, UIN74, YC77, Age72, ABCT75, Bac74, Ber77a, Bla77a, Bre75a, BCDK77, BP79c, DS73a, FIN74, FN76, GK73a, Gil73a, Gil73c, Gil73b, Gil76b, Gil76a, GKYG70, GG75a, GG75b, GW70, IFN72, Jor73b, KGYG70, Kan73, KDH77, KHK79, KF76, KS77b, Lai75, LCC72, Mac75, MV75, McC72b, MME74, MRTG79, Mit76, MGB73, MMT76, PB78, PCC72, RNS71, Rot73, See73a, SBB^{+}74b, SBB^{+}74a, SBB^{+}75, SBB^{+}76, Smy79, SD75, STD77, Sun77, TS71b, UINH73, Vas74, War70b, WB75, Wic78, WTR75, Win76, Wri77, Yoc79, Sug75]. Molecule [Ros75c, Bla77a, LST74b, MSG^{+}78, Mof70, MMT73, Smi78a, Smy79, VFKL74]. molecule-membrane [LST74b]. molecules [And79, Dav77, DUK73a, Jor72, KS77a, KDH77, KHK79, Kum75a, MO75a, MO75b, NOB72, OMF72, PW79, WH73a]. moment [CSL75, Cou78]. moments [RS70, WT79]. mona [Hjo70]. Monitoring [Kle75b]. monkey [Sch77b]. monkeys [Sey77]. Mono [See71c, DST79]. Mono- [See71c, DST79]. monoclonal [WK75]. monocultures [Gat78]. Monod [BN72, KDKY76a]. Monod-Jacob [BN72]. Monolayers [OO76, Bec78b]. monomers [Ste70]. monopole [RRS72]. monosaccharide [Höf71]. monosaccharides [KD72]. monotonic [Pie74]. monovalent [ZS72]. Monte [BHK78, Cla77b, NKB72, TR78a]. Moonlight [TJ75]. Moore [Alb74]. Morales [Kat72]. morphine [GC76]. morphogen [MP78, Mei78]. Morphogenesis [Mar72b, BL74, CZ76, Ede71b, Hol79, LH78b, Lüc75, NG77, Tar72, Tho71]. morphogenetic [CR78, Hej75, LH79a, Mar75c, Ros79b, UI73]. morphological [Lap78]. morphology [Ano70d, Nan73, Pre70]. morphometry [BCH73]. morphostasis [Tar72]. mortality [Abe79, FSFD70]. mortalization [SS79]. mosaic [Hut73, MHY77, WTR75, Ric75]. Mössbauer [CMSC77a, CMSC77b]. motile [AW75, Doe72, KP77a]. motility [LK79, LD75, Seb79]. Motion [Rob70, DS73b, Don72, GCS72, Kel77, Man79b, Sch73b]. motions [STD77]. motoneurone [Zuc73]. moulds [LH78b, TJ75]. Movement [Pai76b, Bla78b, Civ70, Dol75, Gre78, Har74, How72, LG71, Ros79b, Sub70]. movements [Bla74, Bla75a, BJ72, GR78, RG78, VRM77]. mRNA [Car78, Lan77, Lew70a, NDF79, PCT76, Sin73]. MS [FSST78]. Mucus [Lew76]. Multi [BP74, Cra77, Vas74, CH76, Gal78, GR75, Gri79, HS75, IHI79, Maz73, Maz76, Van76, Wal74b]. multi-enzyme [Van76, Wal74b]. multi-exponential [GR75]. multi-gradient [HS75]. Multi-lesion [Cra77]. multi-membrane [Gri79]. multi-palindromic [Gal78]. multi-pool [Maz76]. multi-prey [CH76]. Multi-random [Vas74]. Multi-site [BP74]. multi-stage [Maz73]. multi-stratified [IHI79]. multicellular [Lin75, Pal70, Ste75, TS71b, WK75]. multicompartment [HH77c]. multicomponent [DD77, LCH78, MC72]. multienzyme [HK77a]. Multifit [MA78]. Multigroup [VAC76]. multilevel [Koz79]. multiphase [RLS73].

Multiple [Pol73, BN72, Cho72, DWS76, Dou72b, Ede70a, Eld73, HOR73, KMSP74, PR76, Ses74, Sha77, Wes74a]. multistable [Nic73]. multistate [BK77a, CAA79]. multivalency [Ehr79]. multivalent [CNS79, GGS78, Mor72]. Multivariate [SW78, Van74a]. multivibration [Ano72v]. murine [WK75]. muscimol [AJ79]. Muscle [LRS79, Ash72, CT73a, CKJ79, Che75, Coo73, Des77a, DBU76, GG77a, GG77b, Hat73, Hay74, HPCH79, Koi73, KS77b, Kom71, KZ79, Kou73, MHMS75, MMS74, McC72b, Mil78a, Mit75, MG77a, Mor78, Mur74a, RSL79, RAJM76, She70, SRML79, SW74, TKO74, TT75, War70b, Whi71, Wig73, WD78]. muscles [GG75a, GG75b]. Muscular [MPHG76, Hil70, Opl72, SY75, SYNY76]. mushrooms [TJ75]. mutability [Oli77]. mutagen [Abd76]. mutagenic [CS75b, FCSS78]. mutant [Cza73]. mutase [AH79a]. mutase-prephenate [AH79a]. mutation [PSK76, Ril73]. Mutations [Hin78, Vol79, Moo77, RCW73, Sim79]. mutator [PSK76]. Mutual [Bal73]. mutualistic [TP79, VB78]. MWC [BW78a]. Mycoplasma [Man79b]. mycoses [Med72]. myelin [Mil79, Wes74b]. myelinated [Mad77b, MJ71, RP76]. myeloma [WK75]. myocardial [MFS77]. myofilaments [Dra70]. myoglobin [Mur74a, War74]. myosin [GG75a, GG75b, Lym74a, Mor77, RSL79, SY75, SYNY76, Wal73b]. myosin-catalysed [GG75a, GG75b].

N [Cal73, DeV76, Gol76, Jör75b, Lon76, CDS77, Jor73a, Jor73b, KBH76, MSG^{+}78]. N-2 [DeV76]. N-acetylmuramyl-l-alanine [CDS77]. N-formyl-1 [KBH76]. N-terminal [Jör75b]. N. [Ant74]. NAD [UIN74]. NADH [UIN74]. naked [Kob77]. narcissus [WTR75]. nascent [Cas74]. Natural [Cro75, Bas77, Coh76, Gin77b, GC79, GB75, Jok74, KMS77, Llo77, RT75, Ros78d, Van74a]. nature [Dra70, Hal73, PB77, Ros75d, Ros76b]. near [And74a]. near-equilibrium [And74a]. nearest [TS78a]. Necessary [Den75a, SM76, BH70, Mar71b]. nectar [Cal79, KD79]. Nectria [Win73]. need [Mag74b]. Negative [Jor75a, Lev74a, BW78c, BW78a, BW78b, CB75c, Fri79, GPP70a, Gil73b, LR76]. negatively [AG75, SO79a, ST78b]. Negoro [Hon73]. neighbor [TS78a]. neo [HS79b]. neo-Darwinism [HS79b]. neocortex [Sco79b, Sco79c]. Neolithic [All77]. neoplasia [WKG73]. neoplastic [FP72, Gau73]. nephron [KTS71]. Nernst [Off71, Shi79]. Nerve [Doe71b, Alb74, Bas74, Boy73, Cla76, Cla77b, CS77, Cos75, Fis73, Fit73, Gil73a, Gil73c, Gil73b, Gil76b, Gil76a, HJ70, Hol76, Lun74, LM74b, LS78, MJ71, SP71, dO79]. nervous [CM78, WSMS77]. nesting [BGD79]. net [AC77, CS75c, Koc71a, LA76]. nets [PAZ79]. Network [Bes70, Ami74, MWS77, Mik77, MT78b, MBG73, Nov76, Oon76, Seg79, Utt70]. Networks [Tre77, FD73, GK73b, Gla75, GL75, LG76, MT73, New72, OYT74, OS71, OS74, Par74b, Par75b, Par75c, Pfa73, SO74, Tay72, VL78, Yag76]. Neural [Gro70, Wri77, ABC^{+}70, AC77, Coy79, GL75, HJ70, HCBL70, LA76, LF77, LG76, Mac77b, MT73, OYT74, Oon76, PAZ79, SO74, Tay72, Tuc79, Ven76, WH73b]. neuroanatomical [TR78b]. neuromuscular [HL79, HLC75, SBB^{+}75]. neuron [CG73, GRDC72, Nol73, ST76a, SB75b]. Neuronal [Tre77, Tre78, TR78b, Bas71, CR73, NB76, Pfa73, PL77]. neuronic [RO74]. neurons [CT79, MO73, Sch76b, Tuc77]. neurophysins [SBB^{+}76]. neurophysiological [GZ76, Pfa71, PMN73, TR78b]. neurotoxin [SBB^{+}74b]. neurotransmission [Boy79a]. neurotransmitter [Boy79a, WB74a]. Neurotransmitters [Coy79, GM71]. neutral [RBP78, Sam77]. neutralization [LG76]. neutron [NB72]. neutrophil [Ros76a]. neutrophils [Ros75c]. News [Ano76l, Ano78h]. Newton [Kle72]. Newtonian [BG74, Str71]. niche [Kro77, Med72, NGP78]. Nitella [AK76, Kob77, KOR^{+}77]. nitrate [Fal77]. nitrogen [Paa78, Rei76, SP73b]. nitrogenase [Meg79]. nitrosamine [AA76]. nitrous [PPS75]. no [KH79b]. Noise [Tho73a, Ano79h, dC75, dC77b, Cla77b, CS77, Hol76, LM74b, MZ79, NB76]. nomenclature [Ain77a]. Non [Han72, Ish73, IM75, Ish77, LW79b, Mur75, OS74, The71, Abe73, Ada76, Ano75b, BP79a, BW78a, BY74, BJM70, dC77b, Che79, CB75a, CB76, Chu74a, Chu78, CC70, Cza73, Den75b, FSG^{+}72, GK73b, GN76a, HM71b, Ima75, Ima76, JS79, Koo77, Les74, LW76, Lin79b, MWM76, Mar70, Mar76b, Mik77, Mof73, OO77, Pap71, Par74a, Par78, PO79, PL77, SBP77, Sta79b, SB75b, STU78, Vai71b, Wes73, Wes74c, Wic76, Wil70a, dH76, Bun78a]. non-bonded [OO77]. non-carcinogenic [BJM70]. non-chromatographic [MWM76]. non-chromosomal [Mar70]. non-conservative [BY74]. non-constant [SB75b]. non-electrolytes [Abe73, Vai71b]. non-empirical [Mof73]. non-enzymatic [SBP77]. Non-equilibrium [Ish73, IM75, Ish77, LW79b, The71, Che79, Chu74a, Chu78]. non-equivalence [LW76]. non-excitable [Wil70a]. Non-existence [Mur75]. non-helical [Den75b]. non-hyperbolic [CB75a]. Non-linear [Han72, OS74, BW78a, dC77b, CB76, Cza73, GK73b, GN76a, JS79, Koo77, Lin79b, Mar76b, Mik77, PL77, Bun78a]. non-linearity [Ada76]. non-local [STU78]. non-polar [CC70]. non-random [Ima75, Ima76, Par78]. non-randomness [Sta79b]. non-refractive [Pap71]. non-satellite [BP79a]. non-self [Wes73, Wes74c]. non-self-complementary [FSG^{+}72]. non-stationary [Wic76]. non-steady [Ano75b, Par74a]. non-stoichiometric [dH76]. non-uniform [HM71b]. non-watson-crick [Les74]. noncleavage [Doe76]. nonequilibrium [Chu74b]. Nonlinear [LP75, Gro78, Kit75, Lin76, SO74, WSF76]. Nonmonotonic [Ber76a]. noradrenaline [KT72b]. normal [BK71, Con71b, KTS71, RMF78, Rom76, Whe75]. Note [BP70a, DB71, New70, Rig79a, Shi79, Blu74c, GH76, GN76a, HSS79, Kno74, Kou73, LST74a, Lym75a, Lym75b, Nov76, Ros76c, So79b, Vis77, Woo72]. Notes [BG74]. Notices [Ano79i]. novel [MC73, Org77, Yag77]. novelty [Bat73]. November [Ano70o, Ano71r, Ano72q, Ano73h, Ano73m, Ano74p, Ano75q, Ano76m, Ano77p, Ano77v, Ano78j, Ano78p, Ano79l, Ano79r]. N − [MSG^{+}78]. Nuclear [GS78a, GPP70b, GPP70a, GPB71, Got72, Ord76, TS78b, Vig75]. nucleation [LH78a]. Nucleic [Ril74, Ber72, BP79c, GS71a, GS71b, IFN72, Jur77, Koc71b, Les70, PB78, SG71a, Sum79b, TNG74, ZG75, BS79]. nucleoli [BW77, Has74]. nucleosides [CC73b, GPP73, GPP77a, GPP77b, Jor73b, PGPP78]. nucleotide [BP79c, FH71a, HKAT77, NDF79, Sum79a, Tor79]. nucleotide-linked [FH71a]. nucleotides [FN76, GPP74, GPP77a, GPP77b, Jor73b, PGPP78]. nullities [MRS77]. nullity [MHB79]. number [BBT70, Cli77, CZ76, HM79, MYF77, Nin79, PW79]. numbers [Ber78, KM75, Ros75c]. Numerical [Koo77, Off71, Pra72, Pra74, SH72c, ZS72, Hil73, MR74b, NSS74, Rus73, SSR74]. nutrient [VS71]. nutrients [Bha77]. nutritional [BD75b].

obeying [Dia76, May75, Pop79]. object [Gil72]. objects [FH71b, Jic72]. observability [KGN76]. observations [Jör75b]. observed [CM78, MHY77]. Obtaining [Mur74b, Whi79]. occupancy [Tre75]. occupation [Coo78a]. occurrence [Wal70]. Octave [Bar72b]. October [Ano70l, Ano71o, Ano72m, Ano73s, Ano74u, Ano76q, Ano76-27, Ano77w, Ano77-29, Ano78x, Ano78-29, Ano79y, Ano79-31]. Odor [Wri77, KDH77]. O⋯ [MSG^{+}78]. odour [vDHD78]. odours [LM75a, LM75b]. Offspring [Cha78c, Man75d, MSK79]. Ohki [Lak75]. Ohmic [EM78]. oil [OO76]. oil/water [OO76]. old [Mac73]. olfaction [Pol73]. olfactory [LD73, Wri77, vDHD78]. oligomer [MTM77]. oligomeric [Blu75, KDKY76a, KDKY76b, KDKY76c]. oligomers [Dou72a]. oligopeptide [SUE76]. olo [DLK75]. oncogene [Man75d]. oncogenesis [Con71b, WK73b]. oncogenic [Ama74]. oncotic [Rei79]. One [Dar75, Dar77a, Pan78, SG75a, AM76, BP70b, Dar73a, Dar73b, Dar77b, Fuj77, HH71, KH79b, MLM77, SH79]. One-prey-and-two-predator [Pan78]. one-substrate [SH79]. One-substrate-one-product [Dar75]. ones [Rou79]. onto [NKB72]. Opalina [Bla75a]. open [BD70, Fox71, Ish73, Per76b, See71c, ST76b]. opening [SS73a]. operant [Gro71]. operation [TT75]. operational [Mac73]. operationally [Ros72a]. operative [GK72, KK78, KO78, LR76, Lav71, NKB72, Rak77, Rak78a, Sch72c, ST78b, TT75, VFKL74]. operativity [BW78c, BW78a, BW78b, CB75c, KK78, KKDY74, LR76, Lev74a, Mar72a, MN73, Rak78b]. operon [BB77b, Ege79, Man79a]. opposed [WK75]. optical [Boh77, Ciu75, CMV76, JI78, Lat75, See71a]. optical-conformational [Lat75]. optically [Boh77]. Optimal [Ale79, CJ77, How74, Iwa78, Nol76, RC77, Cal79, CP76, PC76]. Optimality [Zam76]. optimization [Has78d, Man75c]. Optimizing [And74b, DP74]. Orbital [dO79, Bre75a, Fer73, FIN74, FN76, GK73a, GK74, GPP73, Goo70, GGM75, IFN72, Jor73b, JI78, Kan73, KHK79, KF76, LCC72, LT72, MPD70, MPP70, MPP71, MME74, PCC72, UINH73, UIN74, WB75, YC77]. orbitals [AL72, GKYG70, KGYG70, Mof70]. order [Del71, Fle72, MT78a, Oon76, SM76]. ordered [LH78a]. orders [BJM70]. ordinate [Bla78b, Mac77b]. organ [Bra79]. organelle [MRS77]. organic [Ano75a, Bat74, DMRD78, McC77, Zem73]. organism [Cra71, Cra73, SCDC77, Sre74]. organismic [Kro74]. organisms [Age72, Bra70, Bra71, CH72, Eld73, Eld79b, Fra74b, HS75, Koz79, Lap78, Lin75, McC71, ON77, PC76, Sag73, Sin76, Sin78, Tho77b, TS71b, dKJS76]. Organization [Mar72a, AK76, Atl74, Bab77, BBZ79, Bec72, GGM75, HA72, Has79, Itz79, Kri77, Nie71, Odu77, RD77a, RD77b, Sac72]. organize [Mac73]. organs [DP74, Sam71]. Orientation [MO75a, MO75b, OMF72, Jon75, MV76, See73a, VtM79]. orientational [Hew78]. oriented [DS73a]. Origin [AO72, BH70, Mik75a, Wil70a, Alf74, Bar77d, CSIW72, CC75d, Goo73b, GD79, Hal73, Ima78, Kam74, Les70, LK77b, Mey73, Org77, PBS72, Rea74, Rei76, Rob78, Ros76d, Rus72, SS79, Sti77, WK75, YS71a]. original [DBS^{+}70]. origins [How72, MM79, Rea77]. ornamental [Wil78]. orotic [GKYG70, KGYG70]. orotidine [KGYG70]. orotidine-5’-phosphate [KGYG70]. orthomolecular [Hue77]. oscillating [IM75, Ish77, Jor77]. Oscillations [RB77, Tor79, Blu75, Cha79, Gro78, HOR73, Hyv72, Jor74, LNHK71, LN71, MGNR78, PMC78, SO74, See70, TE76]. oscillator [KS72, KW75, Kou73, Lin79c, Mac77a, See71b, SG71b, TS78b]. oscillators [BS76, BM79, BC73b, DB78, Lin79b, Pav71]. Oscillatory [LM74a, AG78b, DS74, GK72, Naz77, Ros75d, SIS78, SSZ^{+}72, Tag78, Tys79, Woo72]. osmosis [MW78]. Osmotic [Hil72a, Gin71b, Gor78, Hil72b, Mar77, PF78, Rei70, Rei79, Sch79a, Sch77a, WP71, You73]. other [Bar77c, CB76, GR78, HKAT77, Hod71a, Med72, RB77, Tho77b, War74]. Out-of-plane [Miz77]. outer [Ciu75]. Outline [Sch77b]. output [HS77b, MTL76]. over-dispersion [CP79b]. Overall [Koc70, Cam77]. overcompensation [Ale79]. overlapping [Yoc79]. ovulation [LST74a]. oxidase [Che73b]. oxidase-oxygen [Che73b]. oxidation [DeV76, HWD76]. oxidative [Ji76, Str74, Sun77]. oxides [MKLKV76]. oxidizing [HZ77]. oxy [KGH79]. Oxygen [KS71c, Lin76, TkC74, Alb71, Alb73, BLP76, Che73b, CG75, Goc78, Gol76, Jon71, Lav76, MG72, MC73, McE78, Sto76, Sto75, Str77, SHC71, Thi78]. Oxyhemoglobin [Eas79, CMSC77a, CMSC77b]. oxytocin [SBB^{+}76].

P [Ano70d, GPP74]. pace [Tor76]. pace-maker [Tor76]. pacemaker [Bes70, DB78, RB77]. packaging [SB71]. packing [Adl77a, Bel77b, BCDK77]. Pages [Ano70t, Ano70m, Ano70n, Ano70s, Ano70q, Ano70v, Ano70k, Ano70u, Ano70p, Ano70o, Ano70l, Ano70r, Ano71l, Ano71q, Ano71u, Ano71s, Ano71n, Ano71t, Ano71m, Ano71w, Ano71p, Ano71r, Ano71o, Ano71v, Ano72j, Ano72p, Ano72r, Ano72o, Ano72k, Ano72s, Ano72l, Ano72t, Ano72n, Ano72q, Ano72m, Ano72u, Ano73l, Ano73p, Ano73t, Ano73r, Ano73o, Ano73j, Ano73v, Ano73k, Ano73u, Ano73q, Ano73h, Ano73m, Ano73s, Ano73i, Ano73n, Ano74v, Ano74s, Ano74w, Ano74t, Ano74n, Ano74x, Ano74r, Ano74m, Ano74q, Ano74p, Ano74u, Ano74o, Ano75w, Ano75x, Ano75y, Ano75v, Ano75t, Ano75p, Ano75o, Ano75m, Ano75q, Ano75n, Ano75r, Ano75u, Ano75l, Ano75s, Ano76w, Ano76v, Ano76u, Ano76-28, Ano76y, Ano76r, Ano76t, Ano76z, Ano76n, Ano76p, Ano76s, Ano76m]. Pages [Ano76q, Ano76-27, Ano76o, Ano76x, Ano77-27, Ano77-34, Ano77x, Ano77-33, Ano77y, Ano77-30, Ano77-28, Ano77-32, Ano77m, Ano77s, Ano77z, Ano77-31, Ano77l, Ano77r, Ano77o, Ano77u, Ano77n, Ano77t, Ano77p, Ano77v, Ano77w, Ano77-29, Ano77k, Ano77q, Ano78w, Ano78-30, Ano78z, Ano78-31, Ano78u, Ano78-28, Ano78v, Ano78y, Ano78i, Ano78o, Ano78-27, Ano78-32, Ano78m, Ano78s, Ano78l, Ano78r, Ano78n, Ano78t, Ano78j, Ano78p, Ano78x, Ano78-29, Ano78k, Ano78q, Ano79w, Ano79-29, Ano79x, Ano79-30, Ano79-28, Ano79-33, Ano79v, Ano79-27, Ano79j, Ano79p, Ano79z, Ano79-32, Ano79k, Ano79q, Ano79o, Ano79u, Ano79n, Ano79t, Ano79l, Ano79r, Ano79y, Ano79-31, Ano79m, Ano79s]. pair [CP79a, MSG^{+}78, SKS77, YC77]. pairing [Con79, Doy78, Hol72, Jor72, McG79]. pairings [Les74]. pairs [DZOP71, McG79]. palindrome [MTM77]. palindromic [CE77, Doy78, Gal78]. pancreatic [BS75]. panting [Sey72]. paradigm [Eas79, Sch78]. paradox [BMS77]. parallel [MWS77, Mik77, MT78b]. paralysis [Bel71b]. Paramecium [Rob70, Bla74]. Parameter [MWH73, AG75, Koi73, MGNR78, Van77]. Parameters [Bre75b, CBPT78, Cou78, Dar73b, Dar75, Dug79, EH73, GT77, Hon71, KS74, LS72a, MAM79, PGPP78, Sch79b, Sha77, Sim73a, SBR77b, SH72c, SM71, TD77a, TR78b]. Parametric [Tri75, Sha78]. Parametrization [TW77]. parasite [Dia74]. parasitism [Cha78c]. Parent [MSK79, Cza73, Man75d]. parent-mutant [Cza73]. Parent-offspring [MSK79, Man75d]. parental [Cha78c, Shi78b]. Parity [Bar74c]. Park [Ano75b]. parsimony [MBG73, Moo77]. part [Blu73, Bun78b]. partial [Ana73, BT71b, Har74]. partially [RF73]. Participation [Ost77]. particles [Boh77, NB72, ON77]. particular [Mac78, Wis77]. particulate [Fis77]. partitioning [How72]. passage [GCS72]. passive [AHH^{+}76, LT77]. patch [DBS^{+}70, DP76, Wes75, Wie79]. patchiness [MM78a, Zei77]. Patching [Ros79b]. Path [Civ70, Moo77]. pathogenesis [Fri74, MHB79]. pathogenic [Med72]. pathogenicity [Ker75]. pathology [MRS77]. pathway [BS75, FPB72, Har78, SJ79, Tor79, WSF76]. pathways [Hor74, MO76, Mol76, Sav79, SJ79]. patients [Ste78]. Pattern [DDTG73, Den75c, Hon73, AG79, AMR^{+}75a, AMR^{+}75b, CS73, Ede72, FD73, Gro70, GN76b, Hol79, Jor74, MT73, OS71, OS74, SC75a, Utt70, WE75, Win73, WG70]. pattern-reversal [AMR^{+}75a, AMR^{+}75b]. Patterns [WB73, BRS79, BHD74, FK79, Gro78, HS77a, Hon78, Leo71, Pap73d, Ran75, Sla78, Van72]. paucity [GS78c]. Pauling [CMSC77b]. PC [GPP78]. peak [WB77b]. pectin [RL77]. Pediastrum [Hon73]. pelletability [Yu77]. penalty [NTR77]. penetrating [RP73]. pentane [MRTG79]. pentene [KBH76]. penultimate [Sav79]. peptide [LCC72, Mof70, Zem73]. peptides [HR79, Hod71a, Mof73, VPBD78]. perception [BL71, Mik75b, Nel77, Tre77]. percolation [AK78]. performance [Lin79a]. performing [Bri78]. pericentric [IM78]. period [BD75b, RP76]. Periodic [Bra77, Tys79, BS77, Fra70, FG77, Has78a, MGNR78, TS77]. periodically [NG76]. periodicity [Kir79b, Sea76]. periodogram [SB78a]. periods [Lap78]. peripheral [HA72]. perm [Tag78]. perm-selective [Tag78]. permeability [Gin73, HOR73, KM75, LL70, LHC76, MJ71, SH72c, dO79]. permeable [JWS76, PST74, Wie79]. Permease [YSK^{+}72]. peroxidases [JD77]. peroxide [Poo75]. peroxisomes [Poo75]. perpersistent [CPC78]. Perron [Jea79a]. Persistence [Zei77, HL79, Ros77a]. personal [Orl79b]. Perspectives [Smi78a]. Perturbation [TLJ76, BGK77, Nol73]. Perturbations [Des77b, And74b, TR78a, TS77]. pest [LBGK76]. pesticide [Com77b, Com79]. petal [KJ72]. PGE [Smy79]. PH [GPP78, MB78, Cha79, CD72, DCMR74, Dun74, FC70, Wig72, Wig73, WP71]. pH-dependent [CD72, DCMR74, Dun74]. phage [BT71b]. Phagocytosis [CBBD74, BCBD76, Dol75]. pharmacokinetic [Kib79]. pharmacokinetics [Cut78]. Phase [NS76b, Dar77b, Den75a, KRL71, Mak76a, Par75d, Rot73, Sch74, Sco74, ST79, Smo76, Spe78, WD72, WL76, BK77a]. phase-sensitive [WD72]. Phase-shift [NS76b]. phases [CC70]. phenethylamine [HMS75, WH73a]. phenobarbital [VAA71]. phenomena [AG78b, Bab77, CB75a, DP76, FL74, Lin79c, Man75b]. Phenomenological [Che75, HM72, ZZ72]. phenomenology [AR74, Com75]. phenomenon [BFM75, Ciu77, DK76, Olo73, PBS72, SSZ^{+}72, So79b]. phenotype [Nov75]. phenotypic [Llo77]. phenylalanine [KOH70, LR76]. pheromone [KT72a]. phorbol [Smy79]. phosphatase [Wel77b]. phosphate [GS71a, KGYG70, McC72a, Nyh78, Wel77b]. phosphates [AL72, GPP78]. phospho [DLK75]. phospholipase [Smy79]. phospholipid [GW70, Mar79a, Rot73]. phospholipids [BCDK77, GGM75, KFH74, Smy79, Spe76]. phosphonate [Jor73a]. phosphopropionic [DLK75]. phosphoribosyl [GKYG70]. phosphoribosyl-pyrophosphate [GKYG70]. phosphorus [NS78]. phosphorylation [CC73a, GR73, Ji76, Str74, Sun77]. Photochemical [Les70, RNS71, Fon74]. photography [GBH70]. Photoinduced [CP79a, CMV76]. photometer [Lat75]. Photomonomerization [CHH^{+}76]. photophosphorylation [Sti77]. photoreceptor [MZ79]. photosensitivity [Win72]. photosensor [Bri79]. photosynthesis [BP79b, Cri77, Fon74, Fon77, Lav76, MG72, MR74a, MC73, Men76, Met75, Ott74, Thi78]. photosynthesis-light [Cri77]. photosynthesizing [Sam71]. photosynthetic [CC73a, CMV76, Mar78a, Mar78b, MV76, Pai72, Pai76a, Pai76b, See73a, See73b]. photosystem [Pai76a, Pai76b]. phototaxis [BJ72]. phototropism [MCO77]. Phycomyces [Coh76, MCO77, OG74]. phyllotaxis [Adl74, Adl75, Adl77a, Adl77b, Jea78, Jea79a, Jea79b, Rob77, Rob78, You78]. phylogenetic [HPF78, KS74]. Phylogenism [Jea78]. phylogeny [CCR72, Tod70]. Physarum [KW75, TS78b]. Physical [STD77, Spe70, Blu76, Cra73, Fri74, HPMH76, Mak76b, PP79, Str76, Sug75]. physico [LD73]. physico-chemical [LD73]. Physicochemical [GM71]. physics [Mik78]. physiologic [NP71]. physiological [CB75c, Fri79, Kle75a, LR76, Shu79, Van77]. physiologically [RAH75]. phytochrome [Yu77]. phytoplankton [Fal77, OW73]. picogram [NB72]. picture [HR71b]. Pielou [Sta79b]. piezoelectric [BJ72, CC73a]. Pigment [MV76]. pigments [BY75]. pinocytosis [Dol75]. pinocytotic [Rub77]. pipe [RT77]. pipettes [Pur77a]. pits [CRC77]. pivotal [SR76]. Placental [BLP76]. Planck [Off71, Shi79]. plane [Eld79a, Hua73b, Miz77, Par75d]. Plankton [RN76]. Plant [PD74, Sac75, BRS79, Den75a, Fon74, Gat78, Gin71b, Gin72, Hea74, HO74, HO78, Hol79, Jon72, KGL75, Mil78b, MHC79, Mol76, Pre73, SE79]. plant-pollinator [KGL75]. plants [Gra77, GL73, GS78c, HG78, Ker75, LR76, MR74a, Sac72, Sat74, Sch77a, See73b]. plaque [DG75, Del75a, Del75b]. plaques [DG75, Del75a]. plasma [Atk71]. plasmid [Meg79]. plasmodesmata [Bla78a, Jun77]. plasticity [Nel77, Sej77]. plate [Ada76, GKP75, Kem79, KT72b, Lan75, Sou71, WB74b, WB77b]. platelet [RMG77]. PLM [Sha74, BT70, KV74]. PLM-function [BT70]. plot [BW78b, Rak78b, ST78b]. plots [BNW78, Bar77a, BW78a, Han72, KRL71, NWW79, TR79, Zer75]. poikilotherm [SD77]. poikilotherms [Mro73]. point [Blu71, CR75b, Kub76, Ram77, Sim79]. points [BNW78, May75]. Poisson [Bar79b, MYF77, Mar79b, Off71, Shi78a, WB74a]. Pol [Lin79c]. polar [CC70, DUK73a, DUK73b, Mac77b]. polarizability [PB79]. polarization [LCH78, MV76]. Polarons [KG73]. policy [CP76]. pollination [BH78]. pollinator [KGL75, Wis77]. poly [Car78, Lew72]. poly-l-aspartic [Lew72]. poly-l-glutamic [Lew72]. polyacrylamide [Mad76]. polycephalum [KW75]. polycyclic [BJM70, VAA71]. Polyelectrolyte [Bar79b, KS77a]. polygenic [Sla78]. polyion [BM73]. polymer [GR74, RD77b]. Polymerase [Man79a, Bre75b]. polymerization [Dav71]. polymers [FIN74, NKB72, Oos70]. Polymorphic [Win73]. polymorphism [Dea72, DT74a, Har74, TD76a]. polynomial [Bar77b, BW78b]. polynucleotides [Les74, Olo73]. polyoma [LB72a]. Polypeptide [BPF74, Dav71, Den75c, DJ78, FP73, JW76, Lan77, Nov73, RCW73, Sim79]. polypeptides [Cas74, CSVR78, MPD70, MPP70, MPP71]. Polyribosomes [SG71c, Sin73]. polysome [Pap73a]. polytene [Ger75]. polyvalent [Hew77]. Pomme [MCB77]. pool [Mac73, Maz76]. pools [HC78, Koc71a]. Population [FY74, FY76, Mou71, NG76, Bed74, Ber78, BS76, Bra75, Cas72, Cha78a, CS73, Coo78b, De 75, Gar74, GN76a, GN76b, Gur73, HB70, Hel78, HH77a, JS79, Lay78, Maz73, MWH73, Mur75, NGP77, Rom76, Ros72b, Ros73c, RS79, Rot72, SM79, TC79, TF74, Van77, VB78, Whi78, WO71, Zei77, vdD74]. Populations [Pav71, Blu74a, BBT70, Bol73, BP70a, BP70b, Bre79, CO75, Con71a, DG77, Dem76, Dia76, Fra70, FJThH73, Fri72, FG77, GB75, Goo73a, GN75, Heu78, Iwa78, JR77, KOWG70, Lap78, Leó75a, Ler78, May75, MHC79, PMC78, RS77, Rot75, SP75, Sha73, Sha74, Sha77, Sza77, Ven76, VAC76, Wil75a, WH73b]. populous [Moo77]. pores [DM78, EM78, FL74, KH79b, LH78a, MPH72, KH79b]. porous [Bla77b, Mar77, MW78]. porphyrins [GPB71, HS76, Rus72, Rus74]. posed [MGB73]. position [GK74, VSW^{+}79]. Positional [Fra74b, HGR79, Hol79, Kam74, Mac77b, Mad77a]. positive [BW78c, BW78a, BW78b, Fri79]. possibilities [DP74]. possibility [AG75, CMSC77b, MS71, Mag71, McG79]. Possible [CSIW72, CS75b, Fis76, GS78c, Hod71a, JW76, Kau76, LR76, LT77, Lav77, Lew70a, Mit76, NG77, SUE76, TS71b, Wal73a, Wes78, YS71a, Arc71, AA76, Bar77a, Can70, CC70, Coo73, DK76, Fri76b, Gau73, GD79, GJ76, HS70, HO76, HWD76, Jun77, Lev74a, LCC72, New79, OV78, Paa78, Sac72, See71b, Smy79, Sun77, Tri75, Wes74b, WTR75, Yor79, Zem73]. possibly [BH70, RS78]. post [Bon78a]. post-transcriptional [Bon78a]. Postulates [How72, Ste75]. Potassium [BHK78, Gra77, HG78, Web70, Alb74, Gri77, Gri79, WJ78, dO79]. potencies [FCSS78]. potency [HL79]. potent [WP71]. potential [AO72, Bar77d, BFM75, BM73, Fis73, Fle77, Goo72a, Gou76, GL73, Hai79, Hit76, HZ77, Jor73b, KT72b, Lak75, Les74, MFS77, MSG^{+}78, MR74b, MKLKV76, NP71, Ohk73, OV78, PR76, Pur76a, Rei79, RBP78, SO79a, She73a, Shi77a, Shi77b, Shi78a, Sou71, Str76]. potentialities [Bri74]. potentials [Ada76, KT72b, Mar76b, Nel78, RAJM76, SP71, TSP78, Wil70a, WB74b, WB77b, YK71]. power [HR71b, Sav70]. power-law [Sav70]. PR [Cal73]. Practical [SR72]. Pratt [Hil73]. pre [PMN73]. pre-processing [PMN73]. prebiotic [And76b, FAT77, LCC72, Sco79a]. precipitation [Del74]. precipitin [Kub76]. precursor [LCC72, SUE76]. precursors [MKLKV76]. Predation [CH76, EH70, Van74b, CM72, PS70, TKS79]. Predator [Cro79, FJThH73, RT75, AE79, BGK77, CB74, ER74, Fuj77, GL79, Hel78, How74, Kil72, Lap79, LK76, LK77a, MM78a, Pan78, RT70, RS79, SG73, Tan78, Van73, Vin71, Zei77]. Predator-mediated [Cro79]. predators [Koc74b, Vin73]. predict [Bea79, CB79b, CB79a]. predictability [Fri76a]. Predicted [SP71, FC70, Utt77b]. predicting [GL73, Lin79a, NMK78, SM79]. Prediction [GK73a, Gil73b, DDTG73, Den75c, Fin72, IPH^{+}79, Lan77, MYF77]. Predictions [HS77a, GG77b, Nel78, Sco79c]. predictive [Cla77a, OSLM79, WT77]. predicts [Yoc77c]. Preferences [Bat73, RT70]. preferential [PMC74]. Preferred [PM76, KH75a, KH75b, GGM75]. pregenetic [TS71a]. Preliminary [BH70, BBG70, MPD70]. preparations [Pfa71]. prephenate [AH79a]. prescription [Yoc77c]. Prescriptive [Rös76f]. presence [Com77a, Gor78, Mad76, NP79, TLJ76, Wel77b]. present [Dar77b]. presented [Lei77]. Preservation [UI73, DP74]. pressure [Adl74, Adl77a, Adl77b, Doe71a, Gor78, Gra77, HG78, Hvi75, Les73, MTL76, Rob77, Rob78, RWD76, Sag73, Sch79a, WP71]. pressure-volume [Hvi75]. presynaptic [Dow70]. Previc [Ano70d]. Prey [CB74, AE79, BGK77, CH76, ER74, FJThH73, Fuj77, GL79, Hel78, Kil72, Koc74b, Lap79, LK76, LK77a, MM78a, Pan78, RT75, RS79, SG73, Tan78, TKS79, Van73, Vin73, Zei77]. Prey-predator [CB74, AE79, BGK77, GL79, Kil72, MM78a, SG73, Tan78]. pricing [Han76]. Prigogine [dS75]. primacy [All77]. primarily [Ber77b]. Primary [BW76, CS73, Fon74, Fon77, GL73, HT73, HA72, MR74a]. primate [Ber76c]. primed [MTM77]. primitive [Bar74b, Goo73b, VLB78, Nyb75]. Principle [Con77, Zah77, Ano72v, ARW73, Bas71, Esh78, Goe74, Gre76a, Has78d, LW76, Mat78, MK76, Nie71, NLC79, RNS71, Smi76b, Smi78b, Zuc73]. principles [HS78a, KSR74, Zam76]. prior [IPH^{+}79]. Probabilities [Lew72, Thi78]. probability [Has74, Oon76, Vol72b, WC76, Yoc77a]. probe [SHC71]. probed [WB73]. probes [SFT75]. problem [Ale72, CC75d, DT74b, Edw78, Har76b, LBGK76, MHMS75, McC78, MGB73, Off71, OG74, Ros75a, Ros75b]. problems [Jic72, Kur71, SP75, Ses71]. procaryotes [Hor74]. procedure [BH79]. Procedures [KNJ74, BNT71, TF74]. process [Blo77, Chu78, Com75, DP74, EWRA76, GH75, Got75, IE76, IM78, KS74, KSR74, Koz79, Lym74b, MT74a, PM77, Rub77, SLK76, SBC79]. processes [Ber77b, Blu74b, BC73a, Cil75b, Coo74, DLV77a, DLV77b, Han72, Har70, HKK76, Koc71a, Koc71b, Koc72, Kur77, Lad77, Naz77, PSM70, RS78, Ros70, Sib78, SR77, SB75b, TD77a, Tre77, Tyr70, VG79, Wal73a, Wel77a, Wic76, ZZ73]. processing [BNT71, BT71a, BT71b, PCT76, PMN73, Tor70]. processt [Cha79]. processus [MP79]. produced [Coh70, Doe71a]. produces [Pow75]. producing [MC73, Paa78]. product [Ain75, Cho72, Dar73a, Dar73b, Dar75, Dar77a, Han76, HH71, Mac77c, Sav79, WB73]. Production [Bat74, Bel70, Bel71a, Bel71b, CSIW72, Cos75, Kre71, Moo71]. Products [PBV74]. profile [BC78b, BK71, Pol73]. profile-multiple [Pol73]. profiles [Gra73, MFS77]. prognosis [BMS77]. Program [Ano78-33, MA78]. programming [Mac77d]. programs [KR76]. progress [Goc78]. progression [Pre73, WK73a]. progressive [HKK76, Ste78]. projection [BH79, Whi78]. projections [Bar74a, Gil72]. prokaryotes [BPF74, FP73, Hal71, Rea74]. proliferating [Fri72]. Proliferation [Kem79, AA78, Bar74c, JR77, KV77, PT75, PR77, Rot79, SS79, YZI77, vW76]. proliferon [Zaj77]. prolyl [MPP70]. promoter [CE77, Eld73]. promoting [Smy79]. Proof [Moo77, Che75]. propagation [CR71b, Dav71, Doe71b, LB71, Lun74, Nan76, Nol73, RAJM76, SSR74]. Propagule [Shi78b]. proper [Kac76, OO76]. Properties [MC72, Ohk70, Cho76, Chu74a, Chu74b, Chu78, CMV76, Del71, GL78, HA72, JK72, Jon75, Kri77, NM75, Pap73b, PP79, ST76a, SD75, Sto71, STU78, Wal74b, WSF76, WP71, WMHG77, Jac77]. property [CR75b, KF76, UI73]. prophage [Luc79]. propionic [DLK75]. Proposal [AAQ71, Fri72, BPP70, Hai79, Mil78b, Sam77, War70b]. Proposed [Abr71, Tor70, Bro74, CDS77, HZ71, LR74, LM74b, Orl79b, Par70, Wal73b]. proposition [PSK76]. propres [Hyv73]. Propulsion [CH72, Bla77b, WJ74]. propulsion-gravity [WJ74]. prostatic [Got72]. proteases [Pog71, PA74]. protection [BPF74]. Protein [Bro75, KP72, APA79, Bel77a, Bel77b, Ber78, BSH75, BH77, BD75b, BP79c, CB77, CB78, Dav77, Des77a, DJ78, DM78, Ehr79, GY75, GI77, Hvi75, JO77, Jör75b, KO78, MSF71, MS71, MO75a, MO75b, McG70, Ols75a, Ols75b, Ond79, Oos70, OMF72, OV78, Pap73c, RS77, Rig79a, Rig79b, RS78, Sch78, Sim79, SFT75, SG71c, SBR77b, SOF72, Ste74, TD72, Via71, Wel77b, Wes74b, Woo78, Yoc77c, vHB79]. protein-ligand [APA79, Via71]. Protein-subunit [KP72]. proteins [CA79, Cas74, CHH^{+}76, CB79b, CB79a, Cri74, Cri75, CK77, Dav73, DDTG73, Den75b, Den75c, Eps71a, Eps71b, FAT77, GPB71, GY79, GC73, HT73, Haw76, HR79, Jon75, Jör75b, Jur77, Kel71, KPU78, Mad76, Mag71, MPD70, MPP70, MPP71, Man79a, Nic73, NO74, OO77, Org77, Par70, PP79, ST78b, Sie72, SC77, SBR77b, VLB78, War70a, War74, Wes73, Wes74c, WP78, YGJ78]. Proton [BJM70, RSS^{+}73, CP79a, GPP73, GPP77a, GPP77b, Ji76, Lan78, Pog71, PGPP78, GPP77a, GPP77b, PGPP78]. proton-Lactose [Lan78]. proton-proton [GPP73]. protonated [Mag73]. protonation [Mof73]. protonmotive [Mit76]. protons [GPP70b, GPP70a]. protoplasm [KOR^{+}77]. protoplasmic [AK76, Don72, Kob77]. prototype [MMT73]. Protozoa [WJ74, McK73]. proximal [RLB74, TD77b, WSL79]. pseudo [Miz77]. pseudo-Jahn [Miz77]. pseudomonads [Meg79]. pseudoobscura [Win72]. pseudopodium [OF75]. psychological [IP71]. Psychoneural [GH75]. puckering [GS71b]. pull [PB77]. pulsatile [Nor75b, PS78]. Pulse [JvS77, RP73]. Pulse-labeling [JvS77]. pump [Cut72]. pumps [Jon71]. purine [EF76, GPP77b, Jor73b, Tor79]. Purinergic [Bur76, Rib79]. purines [GPP70b, GPP70a]. purpose [MO73, MA78]. pursuit [Vin71]. push [PB77]. push-pull [PB77]. Putting [Orl79a]. pylorus [SB79]. pyrene [KGH79]. pyridine [FH71a]. pyrimidine [CHH^{+}76, GPP77a, TS78a]. pyrophosphate [GKYG70]. pyrophosphorylase [KGYG70].

Q [Bre75a, LRH77]. quadrat [Pre73]. Qualitative [LK77a, PMC78, AE79, Gla75, Seg79]. quantal [Eld79a, WB74a, WB74b]. Quantification [RT77]. Quantitative [BCH73, CT73a, CT73b, DW71, DW75b, DW75a, PT79, RRS72, SB71, Adl70, ASS79, Bar78, Bar79a, CAA79, MCO77, MRS77, PBV74, PTB74, ST78b, SM78, WMHG77, dKJS76, vDHD78]. quantities [vHB79]. Quantum [GPP74, GS71a, GS71b, JY72, KG74, KD76, LST74b, SG71a, TNG74, GG77b, HR79, Mof70, GG77a]. Quantum-mechanical [JY72, Mof70]. quasi [SH79, Wal74a]. quasi-steady [Wal74a]. quasi-steady-state [SH79]. quasicontinuous [DUK73a, DUK73b]. quasithreshold [Gil73c]. question [TD77b]. queue [Gil76b]. Quine [MRS77]. r [BD75b]. radar [Lei77]. radiation [BT71b, Cal73, Cra71, Cra73, DBS^{+}70, DT74b, Mil70, Øri74, Res78, SP73a]. radiation-induced [SP73a]. radiations [Cra77]. radical [DMRD78]. radicals [HWD76]. radiographic [Fis77, NB72]. radioisotope [Orr79]. radionuclides [Tho77a]. radiosensitivity [WC75]. radiotracer [LS72a]. Rall [ST76a]. ramp [Ano79h]. Random [Blu70, FPB72, AK78, BW77, BCS78, CT79, Ima75, Ima76, Lad77, LK79, Les74, Par78, RF73, Vas74, WSF76, ZV77]. randomly [McM75, OYT74]. randomness [Sta79b, Yod78]. range [Coo78a, Coo78b, HKM79, KS77a, Nir75, Wev72]. Rapid [Sto76, Chu74b, Gol76, MWM76, MGNR78]. rat [Bar78, Bar79a, DMRD78, FY76, KTS71, Poo75, RZMF75, Sho74a]. Rate [AB71, Wat72, Abe79, Ain74b, Ain74a, Bar77a, Bar77c, BHD74, Bre75b, BC76, CB76, CC70, Coo70, Dav71, Eco79, FC70, GC79, HH77b, KP77b, Kir79a, Kle75a, Kra78, KDKY76b, KDKY76c, Lym74b, Pai76c, Pap73a, Ros76d, RP77, RMG76, Sin76, SKTC78, Van76, VB70, WD72, WH72, Wun75a, Wun75b]. rate-limiting [CC70, Van76]. rate-sensitive [Ros76d]. rates [Bor74, CT79, Fle72, HC75a, HC75b, Koc71a, Koc71b, NS78, Nor75b, PSK76, SB79, WB73]. rather [Ege79]. ratio [Bar72b, Ben78a, CO75, Mac78]. rationale [Pra74]. ratios [BP79a, BW78b]. rats [Sho74b]. Rattlesnake [WL71]. ray [GBH70, Ran76, Ros78b, SR72, TW76]. Rayleigh [PWW76]. re [Jak73, Mar73, McE78]. re-entered [Jak73, Mar73]. re-examination [McE78]. Reaction [SD77, AG78b, AG79, BD75a, Che73b, Dar77b, EFH77, Fon74, GKYG70, Kan73, KDKY76b, KDKY76c, Mac77c, Man78, MR74a, MB78, Mur75, Mur76, Ott74, Per76b, PA71, SH79, Sch79c, See70, See71c, SG71b, SIS78, Ses74, She73a, Str77, TIO77, UINH73, Van76, Yag76]. reaction-diffusion [Mur75, Ses74]. reactions [Ain74b, Ain74a, Ain75, BW76, Bun78b, Bun78a, Cho72, Cze71a, Dar73b, Dar75, Dar77a, Dar77b, Del74, DUK73a, DUK73b, DKU77, Dug79, Dun74, EH79, Fon77, GL78, GK71, Han72, HSK72, HS78b, IFN72, IM75, Ish77, Kei75a, Kei75b, KNJ74, Lew70b, Mal76, Ses72, SB76, SBP77, Vai71b, Vai71a, Zer75]. reactive [BS77, Bye78, MT76, Naz77]. reactivity [Cas71, KF76, Ros76c]. reagents [CB75b]. real [HLR73, HLR74, Ste75]. realistic [Kil72]. reality [CB79a]. realization [Ros75e, See71b]. reasons [GS78c]. Reassessment [BV70]. recall [Tre79]. recently [Sho74b, Sho74a]. recently-castrated [Sho74b, Sho74a]. receptor [CT73a, CT73b, CS71, GC76, HK76, KA74, KH75a, KH75b, LHC76, Pol73, Seg76, SBR77a, SBB^{+}75, Sza77, TLJ76, Tsu74, WFW75, dH76]. receptor-ionophore [SBB^{+}75]. receptors [Bur76, Hew78, Lev74a, Ros79b, Smy72]. receptotopic [Sch77b]. Reciprocal [Fre70, VPBD78, BW78a, Che75]. recirculates [LS72a]. recognition [BB79, BPP70, Blo77, BP79c, DDTG73, Den75c, ER78, FD73, FC76, HK76, KH75a, KH75b, MT73, SC75a, Utt70, Wes78, Woe70]. recolonization [Heu78]. Recombination [Chi76, Mar70, Bec78c, Ber77b, DT79, Moo74, RE76, Res76, Smi76a, Sub77, Tea76, TD77b]. reconciliation [Orl79b]. Reconsideration [Tys79, McG79, Met75]. reconstructing [FH71b]. Reconstruction [GBH70, BBGR71, BBG70, CK71, Gil72, HR71b, HLR73, HLR74, LL79, MT78a, San74]. reconstructions [Bar74a]. recorded [SP71]. recovery [CM78]. recruitment [Zuc73]. Rectification [Kir79a, Kir79b]. recurrent [GL75, Heu78]. red [Bec78b, BB73b, BKB76, Can70, LS77, Mak76b, Mil73, Pin72, ZV77]. Redox [Sto71]. reductase [SBR77a]. Reduction [VL78, Bar77c, DeV76, Fox71, Whi79]. redundancy [Eng78]. reference [Cho76, DST79, GY75, KMS77, Mac78, Niv77]. Refined [MPP71]. refinement [BCDK77, HR79]. reflection [Hil72a]. reflectivity [GP72]. Refractive [NS76a, Pap71]. regard [Win70b]. regarding [DP74]. regeneration [BSM77, Bar78, Bar79a, Mad77a]. regenerative [Eld79b]. regimens [KMSP74]. Regioelectronic [TBP78]. region [CE77, DDTG73, FM72, Man79a]. regional [Van73]. Regions [GPP70a, BJM70, OYT74]. registration [NB72]. Regulation [AP70, Bha77, KL79, Alb75, CBS75, Coy79, DW75b, DW75a, Des77a, EFH77, FGP^{+}75, Goo73a, GN75, GN76b, JAS76, JA75a, JA75b, MLM77, SB71, TT75, TF73, Win76, YM78]. regulators [OV78]. Regulatory [GA77, Ain77b, AG78a, Ain79, Ala72, KKDY74, LH78b, Man79a, Ols75a, Ols75b, Rea75a, Rea75b, Sch77a, Smi77]. reinvestigation [GPP78]. Rejection [LS71]. related [BHK78, CB75a, HT73, LK76, Rap76, Rob77, vHB79]. relatedness [OW78, Woo78]. relating [Atk71, GR73]. Relation [BC78b, BCY79, Che73b, SYNY76, BJ72, Cas74, Che75, Con72, Gre76a, Har70, Jur77, Lew70a, Mil78a, Ott74, Pas76, RE76, Shi79, Vig75, Wes75]. relational [HB70, Rös76f]. Relations [Bec78b, Bas74, BB77b, Bri78, Goo78, Jaf74, Sma76]. Relationship [SD75, WD73a, Abr71, Blu77, Dar73b, Dar75, DW71, Der78, Fuj77, Gau73, GT77, GD79, Man75d, Mor78, Orl75a, Orl75b, OW78, Osb74, Pat70, Sin73, Vog77]. relationships [BH77, Doe76, Van73]. Relative [FCSS78, Sre74, How74, LS72a, Web74]. relatives [HS79a]. Relaxation [Luc79, Per73, Ano71f, Cze71a, Cze71b, LNHK71, MJ71, Sch72c, Sta73b]. relay [Ami74, KF76, Nan76]. relaying [Nan73]. release [Ano78h, Bac74, Cou78, Dow70, Pur77a, Rib79, Sch79b, WB74a]. releasing [AFB76, GK73a]. relevance [CSIW72]. relevant [HH71]. Reliability [AFS^{+}79, PO79]. reliably [CB79b]. relicts [BY75]. Remarks [Per76b, Zah77, BT71b]. remote [HK76, IPH^{+}79, KH75a, KH75b]. renal [BDS73, Cam77, DS70, HS77b, PSM70]. reorientation [KP77a]. repair [Ber77b, DBS^{+}70, Luc79, Res76]. repeated [CZ76]. repertoire [PO79]. Repetition [Pie74]. repetitions [WW77]. repetitious [Kir79a, Kir79b]. Repetitive [SS70, WT78]. replacement [DVM79, Moy70, Osb74]. replication [Ama74, Ble71, BC77, BCY79, DH79, Den72, Din74, Dou72a, Dou72b, Jur77, KK73, LK70, LK77b, MTM77, Sha78, Ste72, Vig75]. replicon [RP77]. Replisome [Ble71]. Reply [Ant74, Lon76, Vin74, CS71, Whi71]. report [BBG70, HLR73, HLR74]. représentables [Hyv73]. representation [Kat72, Kib79, LBM70]. representations [CB76]. represser [JO77]. repression [BM78, BK77b, FT74, Hed71, TF73, TFB74, Tys79, VAA71]. repressor [SB71]. reproduce [WM73]. reproducing [Eps79, Sug75]. reproduction [Ber76c, Bli77, Lai77, Man75b, Man75d, Nag76, Ost77, Tre76]. reproductive [Cra71, Cra73, Lai75, PD74]. reptiles [Osb74]. required [PW79]. requirements [Cal76a, PBV74]. requiring [Les70]. reserve [CJ77, PC76]. resetting [Win70a]. residue [LS72a, MPP71]. residues [AFB76, Chu78, MPP70, PM76, PP79, VSW^{+}79, Yoc77c]. resilience [Dia76]. Resistance [MFS77, Com77a, Com77b, Com79, Gor78, GS78c, Hay74, Koi73, Pow75]. resistive [Shi71]. resistivity [Nel78]. Resolution [FH71b, HR71b, Via71, CSL75, Ciu75]. resolving [HR71b]. resonance [BJM70]. resonant [Kau76, Kou73]. resorption [SK72]. resource [AM76]. resources [Coh71, LT75, WO71]. respiration [BP79b, MAM79, Mur74a]. Respiratory [WT77, Arc71, HC78, Sey72, Sto70, Sto71, WK78]. respiring [Str77]. Response [Fin73, Jak73, Cal73, CFV75, Cla76, Coh70, DW71, Fis73, KK78, LK79, LF77, Mil78a, Nos77, NR71, Par73b, Par75e, Ros76d, SS70, Sre74, TR79, TL74, WY75, Win76]. responses [GL73, HJ70, JAS76, Pur77b, Res78]. responsible [Tho74a]. restarts [BPF74]. resting [AO72, Dra70, Ler78, RT77, Wun75a, Wun75b]. Restricted [CCY^{+}75, Eld79b]. restriction [ABCT75, SK79, SHG^{+}71]. restrictions [CAA79, Gar76, Sch73b]. result [AC74, PN73, Tea76]. resulting [Bab77, Koc74a, Sac75]. results [BNW78, BT71b, Gup75, Jor73b, KJ72, NWW79]. retention [Tho77a]. retina [CR75a, CR75b, Mac77b]. retinal [BFM75, Lei70]. retinal-cortical [BFM75]. retrieval [GZ76]. retrospective [OSLM79]. returning [LTP78]. Reversal [KS72, AMR^{+}75a, AMR^{+}75b, OG74]. reverse [Coo77, Wen77]. Reversible [GGS78, Cot73, Sta70, Wal74a]. review [RS78]. revision [Ran77]. revisited [OK76a]. rewards [BCS78]. rheological [GGSW72]. Rhizobium [Paa78, Rei76]. RHO [OW78]. rhythm [Bro75, KJ72]. Rhythmic [EF76]. rhythms [Ber76a, DB78, JK72, KJ72, KV77, RB77, SB78a, TE76]. ribonuclease [BS75]. ribonucleic [KR76]. ribose [GPP77a, GPP77b, GS71a, GS71b]. Ribosomal [BHD74, CE77, LR70]. ribosome [Bar74b, BBG70, Blo77, Cas74, Pap73a, Sch78, Woe70, Woe73, Zem73]. ribosome-bound [Cas74]. ribosome-catalyzed [Zem73]. ribosomes [BBG70, Bre75b, Koc70]. riding [Hew79]. Ring [Dan76, GPP77a, GPP77b, ST76c]. Risk [Vin71]. RNA [ABCT75, Bal72, Bal73, BPP70, Blu74b, BHD74, Bre75b, BD75b, CFV75, CCR72, CE77, EF76, FSG^{+}72, Fit76, FN76, Jor72, KK73, Klä75, LR70, LRH77, Man79a, Man75d, NDF79, Rea75a, Rea75b, Rig79a, SG71c, Sin73, Teo73, VNB77, WTR75, YM78]. RNA-derived [NDF79]. RNA-linked [KK73]. RNAs [CC73b, Haw76, ST77, BPP70]. Robinson [Vin74]. rod [GWZR77, RGZW78, WTR75, Zar75]. rod-shaped [GWZR77, RGZW78, WTR75, Zar75]. rods [Isr73]. Rogers [Ant74]. Role [Bra71, Eps71a, Eps71b, Her70, Kom71, OK76b, Tho71, BV70, Bra70, CT73b, CDS77, Che73b, Cil75a, FY74, Goo73b, GG75a, GG75b, GJ76, Hai79, HS70, Har74, Hea74, HWD76, Jon72, Lan75, Lew76, Loe71, Mur74a, NB76, Rea75a, Rea75b, RT77, Ril74, Sha78, Spe76, SR76, Teo73, TS71a, Wat72, Wer76, Zem73, And76b]. roles [Hod71a, Jun77]. root [Bar76, Gin72, Lav77, RL77]. roots [Gin71b, Gre77a, GL73, LSGFdlT^{+}75]. rotates [Mor77]. rotating [Woe73]. Rotation [Ada77, Jor73a, Jor73b, JI78, RHA75, Sil76]. rotation-induced [Sil76]. rotational [Don72, McG70]. rotatory [Boh77, YSK^{+}72]. rule [Ain74b, Ain74a, LH79b]. rules [Dun74, Lüc75]. ruminants [Adl70]. run [EF76].

S [Ano71x, Sch74, LR70, MHY77]. S-phase [Sch74]. Saccharomyces [KD72]. safe [Shi78b]. saline [NP71]. salt [Fle77, RLB74, SK72, WSL79]. saltatory [Mad77b]. same [BP79a, Koc74b, Sza77]. samples [Orl70, Per76a]. sampling [SP75, SZ71]. sarcomere [Mit75]. Satellite [Ano78-33, BP79a]. saturable [MA78, Nea72]. saturated [KH79b]. saturation [Bar77b, Chu74a, Pap73a]. scale [Kau74]. scatchard [BNW78, NWW79, TR79, CB76]. scattering [WP72]. SCF [WB75]. scheme [Via70, WW77]. schemes [BD75a, Cze71b, Jor72, Via71]. schizophrenia [BMS77, Tho73a]. Schizophyllum [KP72]. school [SATT79]. school-attendance [SATT79]. science [BBGR71, Blu73, CK71]. sclerosis [Wes74a]. scope [Bod78]. scopolamine [WSMS77]. scoring [Eng78]. scorpion [SBB^{+}74b]. SCRF [TSP78]. second [Rud79, WSF76]. second-site [WSF76]. Secondary [LR70, LRH77, ST77, Bal72, Bal73, Den75c, Klä75, Mag71, PPS75, VNB77]. secretion [Dea75]. secretory [DM75]. sedimentation [Kum75a, Mag70]. seed [BY75, GB75, HJ75, Hag76]. seeming [Wil78]. segment [Car78]. segments [Den75b]. segregation [Hit76, SKT79]. selected [Jor73b]. Selection [Bli77, Par78, And76b, AF77, Bel70, Bel71a, Bel71b, Bry73, Cha77, Cro75, DO76, EWRA76, Gin77a, Gin77b, GC79, Joh79, Llo77, Mic79, Orl75a, Orl75b, Orl75c, Orl77, OW78, PO79, Sag73, Smi76b, Tea76, Tre76, Wil75a, Wil78, Zah75]. Selective [Rot77b, CGM77, Les73, MMRE77, Tag78, Vin71, dO79]. Self [And79, Bab77, GL70, GCG^{+}70, Wes74c, AK76, AM75, ABCT75, BM73, Eps79, FSG^{+}72, KP72, Lai77, LG71, MF79, MHC79, Nag76, PO79, Sha78, Sug75, Wes73, Wil75a]. self-complementarity [ABCT75]. self-consistent [BM73]. Self-electrophoresis [And79]. self-excitation [Sha78]. self-fertilization [Nag76]. self-incompatibility [KP72]. self-inspection [Lai77]. self-limiting [Wil75a]. self-non-self [PO79]. Self-organization [Bab77, AK76]. self-replication [Sha78]. self-reproducing [Eps79, Sug75]. Self-sorting [GL70, GCG^{+}70, LG71, MF79]. self-thinning [MHC79]. self-tolerance [AM75]. selfish [Ham71]. semi [KHK79, PN73, Rob77]. semi-decussate [Rob77]. semi-empirical [KHK79]. semi-infinite [PN73]. Semiempirical [MSG^{+}78, Jor73b]. semilogarithmic [Zer75]. senescence [Duc77, DMRD78, KH75c, Ros78c]. senescing [Den75a]. sense [WY75, Win72]. sensing [IPH^{+}79]. sensitive [Lei77, Ros76d, SC75b, WD72, dH76]. Sensitivity [Mil74, And74b, KC78, MWH73, Nos77, Pfa71, UF71, vDHD78]. sensitization [DW71]. sensitized [CS71]. sensor [Sil76]. sensory [Ate73, Bar72b]. separate [Leó75a, Man75b]. Separation [Tho74a]. separations [Spe78]. September [Ano70r, Ano71v, Ano72u, Ano73i, Ano73n, Ano74o, Ano75n, Ano76o, Ano76x, Ano77k, Ano77q, Ano78k, Ano78q, Ano79m, Ano79s]. septoria [SBC79]. Sequence [JO77, Sum79b, CE77, CB77, CB78, CB79b, CB79a, Edw78, Eng78, FSG^{+}72, HT73, Les74, Mac77c, Mal79, RE76, War74, Yoc74]. sequence-independent [Edw78]. Sequence-specific [JO77, Sum79b]. sequences [BP79a, BSH75, CS75a, CL70, Fit76, Gal78, Itz79, Klä75, Les77, MBG73, NDF79, RCW73, Ril73, Sum79a, Tem75, Yoc77c]. sequencing [Fri76b]. sequential [Tre79]. series [Bon78b, Cut72, MT74b, MWS77, Mik77, MT78b]. series-parallel [MWS77, Mik77, MT78b]. serine [Pog71, PA74]. serotonin [Kie73, Nic76, Wes74b]. Set [RC77, Ain77a, GK71, LF77, Vol72b]. Set-back [RC77]. sets [Jic72, MGB73]. several [MLM77]. sex [Ben78a, CO75, KT72a, LW79a, Mac78, Nyb75, Smi71, Smi76a, Tay79]. Sexual [DO76, Man75d, Smi76b, Wil78, Man75b, Tre76]. Sexuality [Bre79]. sexually [WM73]. SHAM [CSL75]. Shannon [Hut70]. Shape [Pin72, Bec78a, Bec78b, Blu73, BKB76, Can70, Hon71, Pal73]. shaped [Boh77, GWZR77, RGZW78, WTR75, Zar75]. shapes [Bar77a]. sharper [Chr76]. shear [Lub73]. shedding [DK76]. sheep [CS73]. shell [BKB76, IHI79, Swa74]. shells [Pap71]. shielding [GPP70b, GPP70a, GPB71]. shift [BD75b, CMSC77a, Eld79a, NS76b, Sin78]. shift-down [Sin78]. shift-up [BD75b, Sin78]. shifts [DVM79, GPP77a, GPP77b, Nel77, PGPP78]. shocks [Nyh78, RC77]. shoot [Sat74]. Short [HKM79, GL75, Orr79, Smi76a, Tay79]. short-lived [Orr79]. short-term [Smi76a, Tay79]. Shunting [Hay74]. shuttles [Coy79]. sialo [RRB76]. sialo-glyco-macromolecules [RRB76]. sib [Smi76a]. sib-competition [Smi76a]. siblings [Cha78b]. sibships [Joh79]. sickle [Pow75, VP77]. sickle-cell [Pow75]. side [Bel77a, GC73, Jon75]. side-chain [Jon75]. sieve [Lan75]. sigmoid [BW78a, CB75a, Eas79]. sigmoidal [Wal74b]. sigmoidality [NWW79]. Signal [Nan76, BPF74, FP73, Nan73, Tre75]. signals [VD78]. Significance [Leb73, SB71, Ano79h, Bar77d, CB75c, DM75, DK76, Fab76, Fri79, Gil73c, Ima76, LR76, Olo73, Oos73, Ros71, Wal70, Wic76]. significant [Coh76]. signs [Den75c, Sch73a]. Similar [Res78]. Similarities [Via70, CB79b, CB79a, FH71a, Kie73]. Similarity [NDF79, Utt77b, Che76, Der77, GB77a, NLC79, WS78]. Simple [Kel71, Mac79, Nos77, Sch79c, SB79, Ain77b, ABR72, AR75, BD75a, Bed74, Cri77, Fis73, Mac77a, MPD70, MPH72, MWS77, Mik77, MT78b, MG77a, NM78, Sch76a, Shi71, SM79, TS77, Via71]. Simplified [TIO77]. Simpson [LH79b]. simulate [Ran77]. Simulated [Mar72b]. simulating [BB73b, Odu77, Utt77a]. Simulation [DS70, Gar70, HJ70, HC75a, HC75b, HS75, Hul71, LG71, Nyh78, AC74, BRS79, BR75, Cam77, CS75a, Cas72, Cla77b, Con74b, GKP75, GS78a, GR78, HMT79, HPCH79, Koc71b, KTS71, Mil74, MA71, NKB72, OMF72, PS78, PCT76, RG78, RD77b, Sal74, SK72, SD76, TV75, TR78a, ZW70]. Simulations [DB78, AA78, Gre78, MFS77, Ste75]. simulators [Par74b]. Simultaneous [Hea77, Bat73, HPMH76, LR74, AHH^{+}76]. single [BP73, Gil76b, KH79b, Kra78, KNJ74, LD73, MT74a, MHY77, Mor78, Nan76, RS79, Tuc77, WP72, CC75a, Han72, TFB74]. single-file [KH79b]. single-ion [Gil76b]. Single-locus [CC75a]. singularities [Bar79b]. sinusoidal [See70]. sinusoids [BR79]. SIRT [LL79]. Site [Lan78, AH79a, BP73, BP74, BP76, Bro74, CB75b, CS71, HC78, KGYG70, Kan73, KA74, LSB78, LS77, MLM77, Ove77, Pol73, RSL79, Ril73, SBR77a, Smi77, Sto71, WSF76]. site-directed [Smi77]. Site-exposure [Lan78]. site-single [BP73]. sited [Har71, Har73a]. Sites [Ger75, And76a, And76b, Cho77, Gar76, Har73a, ST78b, Smy79, Sto71, WZ72, Yoc77c]. situ [Wes74c]. six [Ain74a]. Size [Oos70, Cas74, Coo78b, FGP^{+}75, Kle72, PO79, RP77, Shi78b, Wes75, Zuc73]. sizes [Cas74, Pre73, Ril74]. skeletal [HPCH79, Mit75, RSL79, SRML79, WE75, WD78]. skin [CS73, HMT79, HH77b, HH77c]. slave [Kau76]. sleep [Fre70, Hun74, KMS77, SSZ^{+}72]. sleep-electroencephalogram [SSZ^{+}72]. slices [CG75, Gol76, Pop79, Sto76]. slide [Ain74b, Ain74a]. Sliding [Dou75, Hil70, Che75, GG77a, GG77b, LB77, Mit75, Sch73b]. slime [CR71a, CR71b, KS70, LH78b, Nan76]. slope [BW78b, CSL75]. slopes [Ano79h]. slow [KDKY76a, WJ78]. slowly [HH77a]. small [AG75, BR79, BBT70, GCS72, KDKY76b, Lei78, Smi78a]. Smoes [Ros75d]. smoke [AA76]. smooth [CT73a, Hay74, Koi73, RAJM76]. snail [JS79]. snubbed [Meg79]. Sobell [Doy78]. social [Ben78a, Mac78, Sey77]. societies [Ber76c]. Society [Ano78h]. sodium [Abe73, Cha72, CA77, MJ71, SFN77, SBB^{+}74a, dO79]. sodium-transporting [SFN77]. sof [Des77b]. sokal [NSS74]. sol [Doe71b, PB77]. sol-gel [Doe71b]. solar [Øri74, TS71a]. solid [Gre76b, MCM79, Str74]. Solitary [AG79]. Solitons [Dav77]. solubility [Ste70]. soluble [dLA76]. Solute [CC70, Hil72a, KHP77, KH79a, Mik77, MT78b, RWD76, Seg70, Win70b]. solutes [RBP78]. Solution [Off74, BD70, BGK77, LCH78, Mag71, Mar79b, NSS74, NP71, Nov75, Off71, PB79, Ros75a, Ros75b, Shi78a, Shi79]. solutions [ABR72, BS77, Bra77, Fin76a, Fra74a, Has78a, Hvi75, Jor75a, Kac76, KH79a, Mar71a, Mur75, Mur76, Naz77, Oos73, Ros75d, Ros76b, Sav70, She73a, UF71]. solvable [Pan78]. solve [Edw78]. solvent [APA79, PM76]. somatic [Con71a]. Some [Cha78b, CC75d, Chu74a, Chu74b, Chu78, Din74, Doe76, EH70, FP73, Gau73, GL75, Gup75, Jea79b, Jic72, Jok74, Jun77, Lai75, Orl75c, ST76a, SP75, Wal74b, WH72, Atk71, APK75, Bec78b, Ber76a, Bra77, GKYG70, Gou76, Haw76, Jon72, KGYG70, KC78, Kil72, KSR74, LS71, Măr76a, PMN73, SC77, SBR77b, TSP78, WH73a, vHB79, SG71c]. sommation [Ada76]. songs [Wic76]. sorting [GL70, GCG^{+}70, Goo72b, GGSW72, Har76a, LG71, MF79, RG78, Ste75]. sorting-out [Goo72b]. SOS [Luc79]. SOS-repair [Luc79]. sounds [Hjo70]. source [Bas77, Coo73, New72]. sources [Fle75, Pal70, Ram77]. Space [Mei78, Adl75, BL71, Bol73, CAA79, Ciu77, FK79, Lau72, LBM70, Par73b, Wal73a]. Space-dependent [Mei78]. spaces [KH79a]. spacing [Hil70, RC77, Tay76]. spacing-out [Tay76]. span [Duc77, Lei78]. spasmolytic [CT73b]. Spatial [BD75a, GN76b, SKT79, Sla78, GK72, LF77, PP79, Wil75b, Woo72]. spatially [AG79, CR75a, CR75b, CB74]. special [Cil75a, DC77c, DST79, KMS77, Niv77, Ste75]. Specialization [BCBD76, Wis77]. species [Con72, CM72, Dar77b, Dem77, FSFD70, Fra74a, Fuj77, Lap79, LT75, Pan78, PS70, Pre73, RS79, SKT79, Sma76, SBP77, TKS79, Van74b, Wis77]. species-abundance [Sma76]. specific [Coh70, Ede71b, Har73b, Hod71a, JO77, Lap79, McG79, Ond79, Rot75, Sch77b, Sum79a, Sum79b, TR78a]. specificity [Ehr79, Gla79, HA72, Hol72, Les78, Lew70a, Rea77]. spectra [Boh77, DD77, DUK73a]. spectral [BH79, Gri77, Jea79a, See73a]. spectroscopy [CMSC77b]. spectrum [DUK73b]. speculations [Jun77, WH72]. speed [How74]. sperm [AW75, Dou75, NR70]. spermatozoa [KP77a]. sphere [BB73b]. spheres [Adl77a, NS76a]. spherical [Lin76, MH74, McE78, Pap71, Pur76b, WL72]. spherocyte [BKB76]. sphingomyelin [KBH76, Pat70]. spike [KS72]. Spin [TD72, GPP74]. Spin-lattice [TD72]. spin-spin [GPP74]. spinal [Odu77]. spindle [Doe71a]. spiral [Kel77, OG74]. spirochetes [Ber76b, WJ72]. spleen [SR76]. split [RLB74]. Spontaneous [Tay72, Bre79, Goo72b, RO74, See71a, See72, Yoc77a]. Spore [Cot73, CGM79, Lef78]. Spores [Hit76]. sporulation [Hod71a]. spot [YS71a]. Spread [SB73, TV72, Blu70, BB73a, Lau72, Nor75a, SATT79]. spreading [Mar75a, Mar75b]. squamous [AWD77]. square [Lav77, SB78a]. square-root [Lav77]. squared [PMN73]. squared-error [PMN73]. squares [BH79, LL79, SB77]. squatting [MLM77]. squid [AR74, Got75, Gri79, Has78a, SS70, Sta73b, Sta73c, Sta76]. stabffity [TM74]. Stability [BM78, Bel77b, BP70b, Con72, Has79, KGN76, Lad76, Maz76, NM78, Rot79, vdD74, AC74, Bed74, Coo78a, De 75, Dia76, Fuj77, FIN74, Gin77a, GC73, Goo70, Gre76b, McM75, Mil78a, New72, Nov76, Par75b, Par75c, PT79, Ros76e, Rot75, RBM76, SB75a, So79b, STU78, TD76a, Tha73, Wu77, dS75, Lad77]. stabilization [ABD^{+}72, HS76, Van73]. Stable [BS77, BD75b, Kum75b, May75, NTR77, Sch76a, SG71b, AF77, Bre79, BGD79, Has78c, Hin77, Hin78, HSS79, See71b]. stage [Maz73]. stages [CR71a, CR71b, Fed77, Hod71a, Lap78]. Standing [Seg70, WSL79]. Standing-gradient [Seg70]. standpoint [MGB73]. start [Dar77b]. starvation [Bra75, Hel78]. Starzak [BHK74]. State [TW77, WT79, WRT79, AG78b, Ano75b, ABR72, AR75, BN72, Bar77a, BK76, Bol73, dC75, Che79, Cho77, CD72, Dar73b, Dar75, DC73, Dra70, DCMR74, Eva79, FPB72, Hil78a, HC78, Koo77, KNJ74, Lap79, LB72a, Lym75a, Lym75b, MSG^{+}78, MG77a, ON77, OD73, Par74a, Par75b, Par75c, Per73, PR77, RO74, Rap77, SH79, SFN77, Ses71, Shi77a, Shi77b, Smi77, Str74, Thi78, VB70, Wal74a, WSF76, Whi79, Wig75, WD73a]. state-space [Bol73]. State-vector [TW77]. states [Bas71, DZOP71, Ede70a, GL78, HOR73, Mal76, Ses74, TE76, WH73b, Yca74]. static [BY75, Rob70]. statics [TP79]. Stationary [GL78, WH73b, DC73, Shi77a, Shi77b, Wic76]. Statistical [Ano78-33, Fag78, Goo70, KS74, LD75, MYF77, Pfa71, Sej77, Sma76, Sue78, Ven76, BW78b, Ber78, Sou71]. statistics [Van74a, Web74, WB74a]. status [SATT79]. Steady [HC78, KNJ74, Ano75b, ABR72, BN72, Bar77a, Bas71, BP70b, dC75, Che79, Cho77, Dar73b, Dar75, Don72, Ede70a, FPB72, HOR73, Hil78a, Koo77, Lap79, Lym75a, Lym75b, OD73, Par74a, Par75b, Par75c, Per73, PR77, Ram77, SH79, SFN77, Ses71, Ses74, VB70, Wal74a, WSF76, Whi79]. Steady-state [HC78, ABR72, Bar77a, dC75, Cho77, Dar73b, Dar75, FPB72, Hil78a, Koo77, Lym75a, Lym75b, OD73, PR77, SFN77, VB70, Whi79]. Steinberg [Har76a, Ove77]. step [CC70, Lym74b]. stepwise [Lew71]. stereochemical [BCDK77]. stereochemistry [PA74]. stereoscopic [Nel75, Tre78]. steric [GJ76, SBR77a]. sterilization [Bra70, Bra71]. Steroid [MKLKV76, Got72]. steroids [JA75a, JA75b]. stick [Sma76]. stiffen [McC75b]. stiffness [HC76, Huk78, MJ76]. stimulated [DG77]. stimulating [RP73]. stimulation [EGF77, LD73, Leb73]. stimuli [WY75]. stimulus [DW71, Dea75, GT77, Ger78]. stimulus-secretion [Dea75]. Stochastic [Blu74b, HB70, Han78, HH71, RS77, Rig79b, She71, SG75a, SG75b, Cla76, CG75, CFS78, Dow73, DT79, GRDC72, Goo77, Gup75, Has78b, HKK76, IM78, KH75a, KH75b, KSR74, Par73a, Ran73, SS73b, Sta70, Tuc77, Tuc79, AK78, SM79]. stochastics [Rig79a]. stoichiometric [HL79, dH76]. Stokes [Gol76, Lon76]. stomach [RBSS70]. stomatal [APK75, DC77a, SS73a]. stopping [LH79b]. storage [CP76, Tor70]. strain [Doe76]. strait [HC76]. strait-jacket [HC76]. straitjacketing [McC75b]. strand [Chi76, Res76, Res78]. stranded [Edw78, FIN74, NDF79, Ril73, SK79]. strategies [CJ77, Has78c, Hin78, HSS79, How74, Lai75, Leó76, NTR77, Ran73, Ril79]. strategy [Bre79, BGD79, Hin77, Iwa78, Par74c, PC76, Pin70]. stratified [AWD77, IHI79]. streaming [Ada77, Don72, MPHG76]. streams [Leo71].

Stress [KOWG70, Doe76, LM74a, Oli73]. stress-strain [Doe76]. striated [Coo73, GG75a, GG75b, Hat73]. Strictly [Hil78a]. Striping [WG70]. stroma [Fri71]. stromal [San74]. strongly [See71c, See71a]. structaral [SBR77b]. structon [CE77]. Structural [Che76, FH71a, HO78, HC76, Nic76, Ros76e, Abr71, ABCT75, CK77, Eco79, Hol76, Jör75b, Mag73, Mag74a, Mur74b, Van74a]. Structure [DHK78, HK74, KDH77, Sir70, WP72, Adl77a, Bal72, Bal73, BCH73, BBG70, Dem77, Den75c, DCT73, Ede70b, FP73, Fin76b, Fis76, FK75, GY79, GC73, GS71a, GS71b, Gre76a, HMS75, Han73, HZ77, JY72, JW76, Jor73a, KOH70, Kau74, KS77a, Klä75, Kro77, Kur71, LR70, LRH77, Lev77, LM75a, LM75b, Mag71, Mag73, MSG^{+}78, MG77b, MMT73, New79, OSLM79, Par74b, PPS75, SG71a, SG72, Sch77b, SJ72, ST77, SBB^{+}74a, SBB^{+}75, Smy79, SK79, TNG74, VBCL76, VNB77, Whi79, WO71, Wig71, Wig72, WMHG77, YGJ78, DW75b]. Structure-activity [DHK78, HK74, KDH77, WMHG77]. structured [Dem76, SG75a, SG75b]. structures [And77, ABC^{+}70, Bar74a, Bas73, BJ72, CR78, CZ76, Mar72b, McG70, MK76, NO74, Nov73, PN73, Tri75, Zgi75]. studied [RD77b]. Studies [AFB76, DHK78, GK74, GS71a, GS71b, GGM75, Han72, Haw76, Hit76, HK74, HH77b, JvS77, Jor73b, KC78, Kem79, KDH77, PO79, SG71a, Sha74, Sha77, TNG74, TW76, VtM79, Yu77, MRTG79]. Study [HR71a, Mar72b, YK71, AHH^{+}76, AW70, BS79, BW76, BD74, Bre75a, BCLV76, Cas72, CP79a, DVM79, Dun74, Fin76a, FIN74, FN76, GKYG70, GH76, HR75, HTSH72, IFN72, JI78, JFR76, KGYG70, Kan73, KA74, KOWG70, KF76, KV77, KD76, LT77, Lau72, Les74, LCC72, MLM77, MME74, MRTG79, Mof70, Mof73, MT76, MT73, Ned75, NS76b, PCC72, RWD76, SSR74, Seg79, SFT75, SV75, TSP78, Thi78, UINH73, UIN74, VFK75, WSMS77, WW77, YC77, vdSJ71]. studying [KR76, Pre73]. Sub [NB72, Fin76a]. Sub-picogram [NB72]. sub-zero [Fin76a]. subassembly [Ros70]. subcellular [DC77c]. subject [STA73a, Sil76]. subnetworks [Ano75b, Par74a, Par75b, Par75c]. substance [Boy73]. substances [Boy79a]. substituents [Cil75a]. substitutable [LT75]. substitutional [KS74]. substitutions [VPBD78]. substrata [Mar79a]. Substrate [MH74, BM79, Cho72, Chu74b, Dar73a, Dar73b, Dar75, Dar77a, ER78, EH73, EH79, FC76, Har71, HH71, Kat72, LR76, MB78, Rap77, RF73, Sam71, SH79, STA73a, SG75a, Smi77, SD75, WB73]. substrate-analogues [LR76]. substrate-multiple [Cho72]. substrate-one [Dar73a, Dar77a]. substrate/one [Dar73b]. substrates [ABD^{+}72, BKWT76, BB77a, BRB78, HK74, KNJ74]. subsystem [Odu77]. subthreshold [WT78]. subtilis [EGF77]. Subunit [Ros70, BHK78, Fis76, KP72]. subunits [KO78, New79]. succinoxydase [MCB77]. such [WP71]. suctorians [McN79]. sufficiency [Gil76b]. sufficient [BH70, GGSW72, SM76]. sugar [LS77]. Suggested [Nie71, KK73]. sulfhydryl [WP78]. sulfur [DeV76]. sulphate [CV76, Mad76]. summation [Mar76b, WB74b]. Supercoiled [Kum74, Kum75a]. supercritical [LM74a]. superficial [Wig73]. supergenes [CC75b]. superposition [Alb74, Con77]. superposition-free [Con77]. supply [Sto75]. supports [JWS76]. supposed [Rus73]. suppression [Moy70]. suprafusion [BFM75]. supramolecular [Gla79]. Surf [Hew79]. Surf-riding [Hew79]. Surface [CV76, RGZW78, SA70, CE78, Des77b, DK76, HR79, MM78b, NCM75, OV78, Pas76, PWW76, PR76, RL77, Ros79a, Seb79, TBP76]. surface-chemical [Seb79]. Surface-volume [SA70]. surfaces [CW74, Mar75a, Mar75b, May74, NP71, NW74, Par73b, PR76, RP75, SO79a, WB75, WZ72]. surrface [RBP78]. surveillance [Fre70, Wal73a]. survey [Lin75]. survival [Bra75, Cra71, Cra73, Cra77, Kie71, Mil70, Mou71, RNS71, Rom76, WC76, Wit78, ZV77]. surviving [MYF77]. suspension [AHH^{+}76, HTSH72]. sustained [LN71]. SV40 [LB72a, MM79]. sweat [RT77]. swelling [Fri71, Kom71, You73]. swells [Hod71b]. swim [Ber76b]. swimming [Wei74]. switch [PD74]. Switching [Tan78, TKS79, CE77, Dou72b]. Sylvester [Bar77c]. symbolic [MHB79]. symmetric [Sim73b]. symmetries [Gal78]. Symmetry [DC73, Fer73, LT77, McG70]. sympathetic [Odu77]. symplasm [Mol76]. symplast [Tyr70]. symplastic [Tyr70]. symport [Lan78]. symposium [Ano75a]. synapses [Bri74, CM78, Sch79b, SC75b]. Synaptic [Tuc79, LP75, Mar76b, PL77, RRB76, Sej77, Tre79, WT78, Wil75b]. synchronization [RC77, Tor76, Wev72]. synchronized [Fra70]. synchronous [SP75]. synchrony [Rot77b]. syncytium [Pur76a]. synergetic [SY75, SYNY76]. synergism [AA76, HM71a]. syntheses [Sin73]. synthesis [BY74, Bre75b, BD75b, Bro75, BP79c, CE77, Coo74, HO76, Hea74, KP77b, Koc71a, Lew70a, Lew71, Olo73, PB78, RS77, Rig79a, Sac75, Sch78, See72, Sim73a, SG71c, Tys79, YM78, vHB79]. synthesizing [GY75, GI77]. synthetase [BPP70, Bro74, Sav79, SJ79]. synthetases [Van76]. synthetic [ST77]. syphilis [Med72]. System [SLK76, ZW70, AG75, Ala72, Ami74, AG79, AW70, BS77, BGK77, Coh77, Coh78, CM78, DW75b, DW75a, Ede71a, EJB^{+}78, FC71, Fuj77, Goo70, Gur73, Has78b, HH71, Höf71, IMMY79, Jac77, Jor77, Kat72, Kem79, KF76, Kit75, Lav76, Lei77, Mac77b, Man78, Man75c, Mat78, MTL76, Oos75, Pan78, Pap73b, Pap73c, Pap73d, RT77, Ros72a, Ros78c, See71c, Ses74, She71, SM76, ST76b, Ste74, Tan78, Thi78, TM74, VB70, WB73, WSMS77, Woo70, Yag76, Yca74, ZZ73, vDHD78]. systematic [Mob73]. systèmes [Hyv73]. systemic [FP72, Med72]. Systems [AHH^{+}76, BDS73, HTSH72, See71a, Abe73, And74b, Ano71f, AE79, BM78, BV75, Bat72, CNS79, CCY^{+}75, dC75, dC77b, Cho77, Com75, CS75c, Cze71a, Cze71b, Del71, Del74, DWS76, DS74, DJ78, DP77b, EH73, Fox71, Fra74a, FT74, GP72, GA77, Gro78, HK76, HB70, Har73b, Har77, HR71a, HS75, HK77a, HZ77, HS78a, Hul71, Hyv72, Ish73, Jea79a, Joh76, Jor74, KVB77, KGL75, KWG79, KDKY76b, KDKY76c, KL79, Kur71, Lad77, Lap79, LK76, Lin71, Lin75, Lin79c, McM75, MRTG79, Mit76, New70, OB73, OD73, Per76b, RD77a, RLS73, Ros75d, Ros76d, Ros77a, Rot73, RWD76, RF73, Rya72, Sav70, Sch73a, Sch79c, See70, SIS78, Sey72, Sim73b, SG75a, SG75b, Sme76, Sug75, TSP78, Tas70, Tha73, The71, Tho78, Tho77a, Wal70, Wal74b, Woh78]. systems [WA78, Zei77]. Systems-theoretic [See71a].

T [SKS77, MM79, Cal73, CCR72]. t-RNA [CCR72]. T4 [BT71a, BT71b]. tables [Pie72]. tail [BSM77]. tandem [Dan76]. tandemly [Kir79a, Kir79b]. tape [SG71c]. Target [Kie71, GL78]. tautomeric [MSG^{+}78]. Tautomerism [FIN70]. taxonomies [Fab76]. taxonomy [Hil73, Pra72, Pra74, Rus73]. Taylor [MT74b, PWW76]. technique [BBG70, HR71b, Les77, SB77, Yag76]. Techniques [GBH70, AAQ71, BBGR71, HLR73, HLR74, Mac77d, NB72, RP77]. Tellegen [OD71]. Teller [Miz77, Wei72]. telocentric [Ima78]. Telomere [DH79]. Temperature [KRL71, Mro73, Cal79, Cal76b, FC70, GS78b, Gün77, Han72, MAM79, Mit75, Syn79]. temperature-dependency [Cal79]. temperature-dependent [Han72]. temperatures [Fin76a, KP79, Ros75c, WL71]. template [Olo73]. templates [HZ71]. temporal [Coo78a, LF77, Ran73, Sin73]. temporally [IM75, Ish77, MPH72]. tenacity [Yod78]. tension [DH78, Sue78]. tensions [Alb73]. tentacle [McN79]. term [GL75, Orl79b, Smi76a, Tay79]. terminal [BSH75, Jör75b, NDF79, SC77]. terminating [Cla72]. termination [BPF74, Bro74, FP73]. terminology [Boy79b]. terms [KM75, MA78, MJ71]. ternary [AK78]. Terre [MCB77]. terrestrial [WO71]. territorial [Har73b]. tertiary [GY79, GC73, NO74, Nov73, YGJ78]. test [Bec78a, FCSS78, Hut70, Kub76, Mad77b]. tested [BNT71]. testing [BNT71, BT71a, BT71b, Măr76a, Sum79b]. tests [Sma76]. tetracycline [Kre71]. tetrahydrobenzo [KGH79]. tetrahymena [WJ74]. tetrameric [Bat72]. thalli [AH79b]. their [BJ72, Cas74, Dav73, FPB72, Gla75, Hul78, Jea79a, JFR76, Kur71, Lin71, Nov73, RD77a, Wat72]. them [MS71]. theor [Ano70d, Ano71f, Ano71x, Ano75b, CB79a]. Theorem [LRB71, OD71, SBP77]. theoretic [See71a, Vol72b, ZW70]. Theoretical [Adl70, Ami74, BL71, CC75a, CC75b, CC75c, CP72, DLK75, DP74, Elt74, FP72, Joh76, Jon71, Kem79, KKJ77, Lin79c, Mag73, Mag74a, MZ79, Mil70, NB72, PO79, Ros71, RWD76, Tay78, Tho77a, VtM79, VAA71, VS71, VFK75, WL71, Yos77, Zuc73, Atk71, BW76, Bea79, Bor74, CC75d, CB75b, DVM79, Fal77, Fle75, FM72, Gou76, HR75, JFR76, KGYG70, KS71b, KA74, KF76, KDKY76a, KDKY76b, KDKY76c, Kur77, Mat78, Mil73, MTL76, Mof73, MT76, Nel78, Odu77, Ran76, Ros73a, Rub77, ST76a, SH72a, Spe76, Ste76a, SR72, Vog77, WSMS77, Wel77b, Wes75, Whi79, Wic79, ZV77, CP79a, HTSH72, KD72, Rei70]. theories [Eco79, Fri73, RRS72]. Theory [AR75, Bas73, Bec72, Bre79, Cas71, CCY^{+}75, Cra71, Cra73, DeV76, Goo72b, Mad77b, MG77b, Rot72, VFKL74, AK76, Ala72, And76a, Ash72, Ate73, Bac74, BL74, Bar79a, BNT71, BT71a, BT71b, Ben78a, Ber77a, BFM75, BJ72, BC77, CSVR78, Cha77, CS77, CR75a, Con71b, Dan76, Dav73, Del74, Den72, Der77, Doe72, DUK73a, DUK73b, Doy78, DBU76, Duc77, Ebr75b, Ebr75a, EM78, Fon74, For75, Fri71, GK73a, GT77, Gil73a, Gil73c, Gil73b, Gil76b, Gil76a, Gla79, Goo77, GZ76, Han78, Hat73, Hil72b, HS77b, HKK76, Hua71a, Hua71b, Hua73a, Hue77, HC76, IHI79, Jea79a, Kie71, KK78, KMS77, Lew73, Mac77a, Mik77, MJ71, NRW72, NRW73, NGP78, NW74, Nos77, OF75, Ode76, Olo73, Ols79, Par75e, PA71, RT77, RCW73, Rei78, Ros76a, Rud79, RBM76]. theory [SG71a, Sch77b, SS79, SY75, SYNY76, SG71c, SM78, Smi74, SB76, SBP77, Str74, SHG^{+}71, SB71, Sub77, Sug75, Swa74, Tho73a, Tor76, Tsu79, Tyr70, Tys79, Utt77a, Utt77b, Vai71a, VAC76, Vol72b, WJ72, Yoc74, Yoc77a, Yoc78, Yoc79, You78, ZZ72, dS75, CR75b, GDA75]. theory-testing [BT71b]. therapy [FP72, Har70, Vog77]. there [LS77, Ove77, Sch77a, TD76b]. thermal [Kit75, RC77, MO75a, MO75b]. thermally [BS75]. thermally-induced [BS75]. Thermodynamic [Gar76, Kei75a, Kei75b, Kre71, Man75a, Ned75, Oli73, SFT75, BW78c, KRL71, MWS77, Mik77, MT78b, OD71, PB77, Wic79, dS75]. Thermodynamics [Cal76b, CM70, Gün77, Hua77, Syn79, Whi71, Cot73, Gla78, HJ75, Ish73, IM75, Ish77, LW79b, The71, Tyr70, Ano71f, Cze71a]. thermokinetic [MT79]. thermoregulation [RT77]. thesis [Hul78]. thickness [Ohk70]. thin [Isr73, MNR72, NRW72, NRW73, SMN73, Swa74]. thinning [MHC79]. thioguanine [CS75b]. thiol [Rea77]. third [Bar77b, Bar77a]. thoracic [MJ76]. thoraco [Odu77]. thoraco-lumbar [Odu77]. those [BP79a]. Thoughts [Bra74, Gau73]. three [AR75, Bar74a, BBGR71, BBG70, CH72, CK71, Gil72, GBH70, Jac77, JW76, Lym74b, Pan78, Par77, Pur76a, RF73, San74, SMN73, WT79]. three-dimensional [BBGR71, BBG70, CH72, CK71, Gil72, GBH70, JW76, Par77, Pur76a, San74]. three-state [AR75]. three-step [Lym74b]. Threshold [AG78b, BM79, CG73, HC77, WB77a]. Thresholds [LSW77]. thrombi [RMG77]. thylakoid [You74]. thylakoids [DST79]. thymidine [Sha74, Sha77, SD76]. thymine [MSG^{+}78]. thymocyte [PT75]. thymus [Tem75]. Thyroglobulin [CD75]. thyroid [CD75, Cil75a, Hul78]. thyrotropin [AFB76, GK73a]. thyrotropin-releasing [AFB76, GK73a]. thyroxin [Dra74, GW70]. tight [KHP77, KH79a]. tilt [LG76]. Time [BV75, BB77a, CB75b, Lap78, Mac77c, Man78, Bed74, BC76, Cas72, Coh71, Eva79, Fri72, GKS78, Lad76, LS72a, Lau72, Lef78, MGNR78, NR71, Off74, Pur77b, RR79, RLS73, SP73a, Sta76, TR78b, WT79, WRT79, Win72, Wu77]. time-constant [NR71]. time-course [GKS78]. time-delay [Lad76]. Time-dependent [BB77a, CB75b, Off74]. time-independent [Eva79]. time-sense [Win72]. times [Hag76, LK70, Per73, Res73, Rot77b, SH72a, SCDC77, WY75]. Tissue [Per76a, Tar72, Alb71, Alb73, CG75, Fri71, Gol76, Hay74, HC73, JWS76, LK79, Lev72, MMS74, MAM79, Mol76, NR71, Pop79, RLS73, Rud79, SG76, Sto76, Str77, SHC71, SB73, TkC74, TV72, WR77, Wes75, Wig71]. tissues [Ana73, DP74, GR78, HC76, Hut73, Măr76a, PMC74, WR77]. titratable [NCM75]. toad [Civ70]. tobacco [AA76]. tolerance [AM75, Coh70, Rap71]. Tonic [Sch73a]. tooth [Osb74]. Topological [Del71, Dou72a, Dou72b, Din74, Hit76, MF79, SBR77a]. Topology [Cri74, Cri75, PAZ79]. torsion [GPP77a, GPP77b]. Total [Han79, Fit76, KM75]. toxic [Ord76]. Toxicity [HL79]. toxin [CBS75, HL79]. tracer [Fle72, Koc71a, Koc71b, MMS74, Mil73, RM73, Res73, RZMF75, She71, SM76, SM71]. tracer-kinetic [RZMF75]. tracers [HL70, Nor71, Orr79]. tracheids [CRC77]. track [Dou72b, NB72]. tracts [TS78a]. traits [Bli77, Wil78]. transcarbamylase [Smi77]. transcriptase [Wen77]. transcription [Bla77a, PCT76, STD77]. transcriptional [Bon78a, Eld73]. transcriptionally [New79]. transducer [CC73a, GG77b, Lak74, Yor79]. transduction [Ate73, LRS79, MV75]. transfection [EJB^{+}78]. Transfer [KVB77, Ain77a, Bak76, BPP70, BLP76, CP79a, CC73b, CC70, DS73a, Dav77, DUK73a, DUK73b, DKU77, Erh76, Kem74, KPU78, MW78, MV76, NS78, Pai76a, Pog71, Res73, Sco79b, Sco79c, See73b, Sey72, Sun77, Tre75]. transfilter [WN79]. transform [Ciu75, Orr79, SH72a]. transformation [Bha77, BB73b, Del71, Gau73, Hyv73]. transformations [Bec78b, BK77a, Wic78]. transformed [KH75c, LB72a, RMF78]. transforms [KV74]. transhydrogenase [Fox78]. Transient [Dar77b, FL74, Sin78, Mil78a, NR71]. transients [Ber76a, BFM75, ST76a, WH73b]. transit [LS72a, RLS73, SH72a, WT79]. transit-time [LS72a]. Transition [BD75b, AG78b, CD72, DCMR74, Han72, Mak76a, Mal76, Man75a, Rap77, Smi77, Thi78, Wel77a, vHB76]. transitions [BN72, BK76, Bla79, Cla72, Hua77, Kil70, Rot73, Sch72c, Sco74]. Transitivity [AMR^{+}75a, AMR^{+}75b]. transitory [MT79]. translates [Mor77]. Translation [VNB77, Coo77, Woe73]. translational [BPF74]. translocation [Pap73a, TBP76]. transmembrane [Bla78b, JvS77, MM78b, Wil70a]. transmissibility [MHY77]. transmission [RRB76, Tuc79, VD78]. transmitter [Boy73, Cou78, Dow70, Rib79, Sch79b, WT78]. transplantation [CM75]. Transport [NP79, Pai72, PSM70, Sam77, UF71, AW75, AHH^{+}76, Arc71, BL75, Blu74c, Blu75, BP73, BP74, BP76, Bun78b, Bun78a, CKJ79, Cha79, CA77, Chr76, CS75c, CG75, DWS76, Dra71, DM78, EFH77, FL74, GR73, GN76a, HC75a, HC75b, HTSH72, HPMH76, HS78a, Jac77, JvS77, KHP77, Kle75b, Koc72, LS72a, LS72b, LS77, Mar78a, Mol76, Nea72, Ode76, RBSS70, RLB74, RBP78, Ros76a, Ros76b, Rot72, Rub77, Seg70, Ses71, Sin75, SS73b, Sta76, ST76b, Ste76b, Str77, TkC74, TD77a, Tyr70, Vai71a, WSL79, WD78, dLA76]. transport-model [LS72a]. transporting [Cut72, Fle77, Höf71, SFN77]. transports [LT77]. transversions [Cla72]. trapping [See73a]. traps [Ott74]. Traveling [KS71b, OK76a, Ros75a, Ros75b, SSR74]. travelling [Mur76]. Treatment [RF73, BW76, BC76, Cha77, CB75b, Fri74, JY72, KKJ77, Sum79a, Wel77b]. treatments [Ord76]. Tree [MK76, Hon71, KS74, SB76, SBP77]. tree-like [Hon71]. Trees [Leo71, BCH73, HPF78, Hon71, Moo77, Mur74b, Pal73, Tho77b, Wal79, WSSB77]. Tribolium [Niv77]. trichromatic [Bri79]. tricycloquinazoline [HR75]. triggered [Bla79]. trimethylammonium [KA74]. triphosphatase [Wal73b]. triphosphate [LSL77]. triplets [Cla72].

Tritiated [Sha77, Sha74]. Tritonia [HLC75]. tRNA [Abr71, Blo77, FC76, Gar74, Kam74, Les78, Sav79, SJ79]. tRNA-tRNA [Les78]. Trophic [Dem77, WO71]. tropomyosin [TT75]. troponin [VSW^{+}79]. troponin-C [VSW^{+}79]. true [EH73]. tryptic [Hol71]. tsou [Rak78b]. tube [RK72]. tubercule [MCB77]. Tuberculin [CS71]. tubularization [You73]. tubule [PST74, RLB74, WSL79]. tubules [Blu74c, WN79]. tumor [JR77, KBW75, SLK76, Smy79, WZ72]. tumor-promoting [Smy79]. tumorigenesis [AA76]. tumors [Arc71, Gre76b, Sac75, Sha77]. Tumour [Man75d, Har76b, Ker75, WNHS75, WK73a]. tumours [MCM79]. Tunneling [MSG^{+}78]. Tunnelling [BC73a]. turbidimetric [ST79]. turbulent [JP79]. turgor [Gra77, HG78]. Turing [BL74, CR78, LH78b, LH79a]. turn [VSW^{+}79]. turnover [AW70, Blu74b, Kle75a, Sie72, Wel77b, FY76]. turns [Den75c]. twenty [RS78]. Twin [Vol72b]. Two [Ano71x, BP73, DB78, Hel78, Wei71, AG78a, AM76, BD70, BP74, BP76, BP70a, Dar77a, Den75a, EH79, Fra74a, Fuj77, GK71, Har71, Har73a, Har78, Hon78, Kam74, KVB77, Koc74b, LT75, MMS74, MFS77, MO75a, MO75b, MHY77, MA78, MLM77, MG77a, MB78, Nea72, OMF72, Pan78, PN73, Ros73b, Rot79, SY75, SYNY76, SG75b, Ste70, SOF72, Sza77, TE76, Via71, WB73, WD73a, ZR73, BK76]. two-carrier [BP74]. two-compartment [BD70, MMS74]. two-dimensional [Hon78, MO75a, MO75b, OMF72, ZR73]. two-enzyme [GK71, TE76]. two-headed [SY75, SYNY76]. two-phase [Den75a]. two-prey-one-predator [Fuj77]. two-site [BP76]. two-sited [Har71, Har73a]. two-state [MG77a, WD73a, BK76]. two-substrate [EH79]. type [AMR^{+}75a, AMR^{+}75b, BN72, CSL75, CR78, Cho76, KHK79, Kre71, MCB77, ST79]. types [Vai71a]. Typical [BY75]. typist [Mac73]. tyrosine [Dra74].

u.v [Les73]. UHF [Leb73]. UHF-microwave [Leb73]. ultra [MT74a]. ultra-filtration [MT74a]. Ultrastructural [Arc71]. Ultraviolet [Sag73, IE76, VFKL74]. unbranched [Ain74b, Ain75]. Uncertainty [Blu74a]. Undamped [See70, Hyv72]. underlying [BC77]. understanding [Bar76, HR79]. undulations [HM71b]. Uneven [Sim79]. unfolding [BS75]. unicellular [CP76, Fra74b, Shu79]. Unidirectional [KH79b]. Unified [CS77, Con71b, Doe72, Hea74, HS77b, dLA76, KK78]. Uniform [Dol75, BKB76, CR75a, CR75b, HM71b, JRP74]. Unifying [And79, Bak76, Bla78b, Duc77, Tho73a]. unit [Bar72b, Kle75a, Mar78a, See73a, See73b, WB74a]. units [Gin73, HR79, Mar78b, Pai76a, WB74b]. universal [Ala72, Tri75]. unknown [Bar74a]. unmyelinated [SP71]. unperfused [Gol76, Sto76]. unregulated [Sac75]. unsaturated [Hai79]. Unstable [Con79, Rak78a, SG71c, Sin73, Van73]. Unstirred [NR71, LS72b, PF78, TD77a]. Unstirred-layer [NR71]. unwinding [Edw78, Moo74, VFK75]. upconversion [Fon74, Men76]. uphill [Höf71, RBSS70]. upon [FC70, MAM79, NB72, SP73a, Ste75]. Upper [Ada76]. upside [HP77]. uptake [Bha77, Che73b, Fal77, HH77b, Jon71, Lin76, MH74, McE78, Par77, RBP78]. uracil [SKS77]. urban [All77]. Urea [VP77]. ureteral [SK72]. urinary [Civ70]. urine [SK72]. usage [Sie72]. Use [Alb71, BB73a, GT77, HM71a, Rom76, SH72a, Ale79, BSM77, CD72, Fin76a, FK79, Fri76b, Gra73, KV74, LR74, NB72, Orl70, SFT75, Smi71, Whi78]. used [PMN73]. useful [Smi78a, WB77b, Woo78]. using [AHH^{+}76, Bar77c, CSL75, Coo74, Cza73, DHK78, GP72, HTSH72, HK74, KS74, KDH77, KRH77, MYF77, Mof70, Pre73, RF73, SB77, Sav70, SV75, Sta76, SD76, Sug75, Wal74a, WB75, Zer75]. uterine [BLP76]. utility [SB78a]. utilization [Ege79, FC70, Men76]. utilizing [KP77b, Koc74b].

V [AHH^{+}76, Cze71b, Hil70, Sug75]. vacancy [KH79b]. vagal [WD72]. valent [DST79]. Valeurs [Hyv73]. validation [Mil74]. validity [LR74, Shi79, SF79, Wal74a, Zer75]. value [Off71, Ros75a, Ros75b, Vol79, You76]. values [GPP70b, GPP70a, Kau76]. valve [Tsu74]. vapor [WL76]. variability [BR79, Dem76, Fri72, Teo73]. variable [Bry73, HOR73, KZ79, Mou71, SP73a]. variables [EH70, ON77, OB73]. variance [How72]. Variation [Sin76]. variations [EF76]. variegation [Hut73]. Varieties [VB78]. variety [See73a]. variola [SATT79]. various [FSFD70, KD76, Rig79a, Ros75c, Rya72, Sre74]. varying [CG73, FC70, NG76, Rei79]. vascular [KKJ77, MCM79]. vasogenic [Rap78]. vasomotor [Kit75]. vasopressin [SBB^{+}76]. vector [BH79, TW77, WT79, WRT79]. Vegetation [Pie74, Orl70]. vegetative [Nag76, PD74]. velocities [Zer75]. velocity [Ain75, And74a, CS75c, Dar77a, Mor78]. ventricular [Nol76]. verify [MG77a]. Vernon [Whi71]. versus [Wes73, Wes74c]. vertebrate [GG75a, GG75b, Mac77b, Man75d, Pol73, Res78]. vertebrates [LD73]. Very [SA70]. vesicles [Bon78b, GCS72]. vesicular [Rub77]. vestibular [Leb73]. via [PSK76]. vibration [Wri77]. view [Cha72, Kub76, Win70a]. viewed [KS70, KSR74]. viewpoint [RS79, RBM76]. VII. [MPP71]. Viking [LSB78]. Villefranche [RN76]. villous [Win75]. vine [ER74]. violate [Yoc79]. Viral [Ros77b, Bec72, Rea75a, Rea75b, Sum79a, Wes74a]. virally [Sum79a]. virally-mediated [Sum79a]. virial [Mag70]. viridis [CM75]. Virtual [Wev72]. Virus [BNT71, BT71a, BT71b, Ama74, FSST78, Har76b, MYF77, Man75d, Ost77, WTR75]. Virus-genetic [BNT71, BT71a, BT71b]. virus/host [Man75d]. viruses [Car78, WL72, WTR75]. visceral [Hay74]. viscosity [Lew76, SKTC78]. viscous [RK72]. vision [BFM75, Nel75, Ric75, Tre78]. vision-suprafusion [BFM75]. Visual [LG76, Blu73, Ciu75, ER74, HC77, Lei70, Nin79, Sch77b, Vin71, WY75]. visualization [Ciu77]. vitamin [JI78]. vitro [Goo77, RMF78, ST79]. vivo [BV70, BB77b, Che73b, SB71, Sug75]. vocal [Bra79]. vocalizations [SW78]. voltage [AR74, Bas74, Jaf74, NR71]. voltage-clamp [AR74]. voltage-concentration [Jaf74]. voltages [Ano79h, Sta76]. Volterra [BS77, BGK77, Fra74a, Gro78, HB70, Jor77, Mur75, So79b]. volume [Ano70j, Ano74k, Ano74l, Ano75k, Ano76i, Ano76j, Ano76k, Hvi75, Mik77, MT78b, Nel78, SA70, SB79]. volumes [HS77b]. vs [CGM77, WO71, Bar77a].

Waals [DM75, GP72, Isr73, MNR72, PN73, SMN73]. walk [Blu70]. wall [Hea74, LH78a, MJ76, NP79]. walls [Tho71]. war [BC78a, BCS78]. Warfare [Pit78]. warning [MT74b]. wasp [BGD79]. wastewater [BC76]. Water [Goo78, Lew70b, Măr76a, Met75, Mol76, Wig71, Bra71, Civ70, DMRD78, FK75, Gin71b, Goo73b, GL73, KHP77, KH79a, NS78, OO76, RBSS70, RLB74, SK72, TD77a, TD72, WSL79, WT77, Wig72, Win70b, YC77]. watson [Les74, Edw78, SKS77]. Watson-Crick [Edw78, SKS77]. Wave [CR71b, AG79, BM79, Lun74, Mur75, Mur76, Nan76]. waveforms [SH72b]. wavefront [CZ76]. wavelike [Mar71a]. waves [Blu75, CH72, LB72b, SZ71, Smo76]. way [MG77a]. ways [Lew71]. weak [Jac77, WY75]. weakly [Lin79c, See71c]. wearing [Abe79]. wearing-out [Abe79]. webs [JDTF73]. Weibull [Sta79a]. weight [Kle75a, Mag70, Par75e, Ste78]. weighting [HE73]. Weir [Ano71x]. well [BL74, GDA75, Nic76, WP71]. wet [Bra70]. wheat [IPH^{+}79, SBC79]. where [Cra71, Cra73, Dar77b, SR76]. Which [Rou79, Bli77, Har78, Jon71, KDKY76b, KDKY76c, MM78a, Tre76, Yoc77c]. Whittaker [GB77a]. whose [Ain75, Rot75]. Wigner [BGK77]. will [Bro77]. Willard [Ant74]. wind [Ale79]. windchill [Øri74]. winter [WL71]. within [AW75, FH71a, MH74]. without [CJ76a, KV74, Lin71]. Wobbling [Bar77d]. women [Gup75]. words [BPP70]. work [BL74, Har76a]. working [Ama74, CV76, LB72a]. wound [Swa74]. writing [Ste76b]. Wyman [KDKY76a].

X [GBH70, Ran76, Ros78b, SR72, TW76]. X- [Ros78b]. X-ray [GBH70, Ran76, SR72, TW76]. X174 [FSST78].

Yeast [Bro74, BPP70, Höf71, Oli77]. yield [Coh71, Den75a, NB72, Ott74, PD74, SP73a, See71a]. yield-time [SP73a]. yielding [Hor74]. yields [IPH^{+}79, KMSP74]. young [NMK78, Bro77].

Zahavi [DO76]. Zea [GL73, SS73a]. zeatin [Kam74]. zero [Fin76a]. Zhabotinskii [Mur76]. Zhabotinsky [TIO77]. Zone [Dra71]. zones [Hil72a, Pie74].

References

Argus:1976:HNS

[AA76] Mary F. Argus and Joseph C. Arcos. Hydrocarbon-nitrosamine synergism as a possible amplifying factor in lung tumorigenesis by tobacco smoke. Journal of Theoretical Biology, 56(2):491– 498, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800884.

Absher:1978:MMC

[AA78] Richard G. Absher and P. Marlene Absher. Mathematical models and computer simulations of proliferation of human diploid fibroblast clones. Journal of Theoretical Biology, 72(4):627–638, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937890276X.

Allen:1971:PIT

[AAQ71] Li-Ming C. Allen, Henry W. Allen, and J. J. Quilligan. Proposal for improved techniques in electron microscopy. Journal of Theoretical Biology, 30(1):195–202, January 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900440.

Anderson:1971:RGC

[AB71] E. C. Anderson and G. I. Bell. Rate of growth of cells. Journal of Theoretical Biology, 31(1):161–164, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901299.

Anninos:1970:DNS

[ABC^{+}70] P. A. Anninos, B. Beek, T. J. Csermely, E. M. Harth, and G. Pertile. Dynamics of neural structures. Journal of Theoretical Biology, 26(1):121–148, January 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800364.

Arroyo:1975:LSC

[ABCT75] P. Arroyo, R. Benguria, E. Collados, and J. C. Tohá. Low self-complementarity of cytochrome c messenger RNA, a structural and functional restriction for evolution at the molecular level. Journal of Theoretical Biology, 54(2):153–166, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801202.

Ajo:1972:ESC

[ABD^{+}72] D. Ajò, M. Bossa, A. Damiani, R. Fidenzi, and … A. Lapiccirella. Electrostatic stabilization of cholinergic substrates. Journal of Theoretical Biology, 34(1):15–20, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900501.

Abdulnur:1976:WFM

[Abd76] S. Abdulnur. Why is 5-fluorouracil a mutagen? Journal of Theoretical Biology, 58(1):165–175, May 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901454.

Abernethy:1973:IBS

[Abe73] John Abernethy. Interaction between sodium ion and nonelectrolytes in the countercurrent systems of the kidney. Journal of Theoretical Biology, 39(3):589–600, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900702.

Abernethy:1979:EIM

[Abe79] John D. Abernethy. The exponential increase in mortality rate with age attributed to wearing-out of biological components. Journal of Theoretical Biology, 80(3):333–354, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900985.

Abraham:1971:PDS

[Abr71] Donald J. Abraham. Proposed detailed structural model for tRNA and its geometric relationship to a messenger. Journal of Theoretical Biology, 30(1):83–86, IN1, 87–91, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900385.

Arndt:1972:EAS

[ABR72] Richard A. Arndt, James D. Bond, and L. David Roper. Electroneutral approximate solutions of steady-state electrodiffusion equations for a simple membrane. Journal of Theoretical Biology, 34(2):265–276, February 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901609.

Austin:1974:ESR

[AC74] M. P. Austin and B. G. Cook. Ecosystem stability: A result from an abstract simulation. Journal of Theoretical Biology, 45(2): 435–458, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901234.

Anninos:1977:NNM

[AC77] Photios A. Anninos and Ronald Cyrulnik. A neural net model for epilepsy. Journal of Theoretical Biology, 66(4):695–709, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902399.

Adams:1976:ULB

[Ada76] William B. Adams. Upper and lower bounds on the nonlinearity of sommation of end-plate potentials. Journal of Theoretical Biology, 63(1):217–224, November 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690093X.

Adam:1977:RBF

[Ada77] G. Adam. Rotation of bacterial flagella as driven by cytomembrane streaming. Journal of Theoretical Biology, 65(4):713–726, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900170.

Adler:1970:TQA

[Adl70] J. H. Adler. Theoretical quantitative approach to the mechanism of hypoglycemic ketosis in ruminants. Journal of Theoretical Biology, 28(1):101–109, July 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900664.

Adler:1974:MCP

[Adl74] Irving Adler. A model of contact pressure in phyllotaxis. Journal of Theoretical Biology, 45(1):1–79, May 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900435.

Adler:1975:MSF

[Adl75] Irving Adler. A model of space filling in phyllotaxis. Journal of Theoretical Biology, 53(2):435–444, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800142.

Adler:1977:ACP

[Adl77a]  Irving Adler.  An application of the contact pressure

model of phyllotaxis to the close packing of spheres around a cylinder in biological fine structure. Journal of Theoretical Biology, 67(3):447–458, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900480.

Adler:1977:CCP

[Adl77b] Irving Adler. The consequences of contact pressure in phyllotaxis. Journal of Theoretical Biology, 65(1):29–77, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900777.

Assimacopoulos:1979:QMA

[AE79] D. Assimacopoulos and F. J. Evans. A qualitative method for analysis of prey-predator systems under enrichment. Journal of Theoretical Biology, 80(4):467–484, October 21, 1979.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900894.

Asmussen:1977:DDS

[AF77] Marjorie A. Asmussen and Marcus W. Feldman. Density dependent selection 1: A stable feasible equilibrium may not be attainable. Journal of Theoretical Biology, 64(4):603–618, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902636.

Abdulnur:1976:STR

[AFB76] Suheil F. Abdulnur, R. L. Flurry, and J. M. Bopp. Studies of the thyrotropin-releasing hormone I. CNDO/2 studies of the conformations of the residues. Journal of Theoretical Biology, 59(1): 189–203, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937680032X.

Ajo:1974:CCC

[AFDL74] D. Ajò, R. Fidenzi, A. Damiani, and A. Lapiccirella. A comment on CNDO/2 conformational energy of acetylcholine. Journal of Theoretical Biology, 44(2):403–404, April 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901714.

Alvarado:1979:RGC

[AFS^{+}79] S. Alvarado, R. Figueroa, A. Sepúlveda, M. A. Soto, and J. Tohá. Reliability in the genetic code. Journal of Theoretical Biology, 77(4):505–512, April 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900237.

Aarons:1975:ESP

[AG75] Leon J. Aarons and Brian F. Gray. On the effect of a small parameter and the possibility of limit cycle behaviour in a negatively inductive control system. Journal of Theoretical Biology, 50(2):501–505, April 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590096X.

Ainsworth:1978:EMR

[AG78a] Stanley Ainsworth and Roger B. Gregory. The exponential model for a regulatory enzyme: Its extension to describe the binding of two ligands. Journal of Theoretical Biology, 75(1):97–114, November 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902047.

Anan:1978:BCR

[AG78b] YoichirO Anan and Nobuhiro Go. Behavior of chemical reaction catalyzed by allosteric enzyme: Threshold phenomena and discontinuous transition to oscillatory state. Journal of Theoretical Biology, 72(2):283–296, May 22, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900942.

Anan:1979:SWS

[AG79] Yoichiro Anan and Nobuhiro Go. Solitary wave and spatially locked solitary pattern in a chemical reaction system. Journal of Theoretical Biology, 80(2):171–183, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902030.

Ageno:1972:MAL

[Age72] M. Ageno. On molecular asymmetry in living organisms. Journal of Theoretical Biology, 37(1):187–192, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901257.

Andrews:1979:CAS

[AH79a] P. R. Andrews and Elizabeth Heyde. A common active site model for catalysis by chorismate mutase-prephenate dehydrogenase. Journal of Theoretical Biology, 78(3):393–403, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903382.

Aplin:1979:GAC

[AH79b] P. S. Aplin and D. J. Hill. Growth analysis of circular lichen thalli. Journal of Theoretical Biology, 78(3):347–363, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903357.

Ando:1976:SAS

[AHH^{+}76] H. Y. Ando, N. F. H. Ho, W. I. Higuchi, J. Turi, and C. Shipman. Systems approach to the study of drug transport across membranes using suspension cultures of mammalian cells V. Simultaneous passive transport and biosynthesis. Journal of Theoretical Biology, 62(1):211–225, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900576.

Ainsworth:1974:SRDb

[Ain74a] S. Ainsworth. A slide rule for deriving the rate equations of enzyme catalysed reactions with mechanisms involving up to six enzyme-containing intermediates. Journal of Theoretical Biology, 48(1):141–147, November 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901854.

Ainsworth:1974:SRDa

[Ain74b] S. Ainsworth. A slide rule for deriving the rate equations of enzyme catalysed reactions with unbranched mechanisms. Journal of Theoretical Biology, 44(1):161–165, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800366.

Ainsworth:1975:MII

[Ain75] Stanley Ainsworth. The mechanistic interpretation of initial velocity and product inhibition data arising from enzyme catalysed reactions whose catalytic cycle is unbranched. Journal of Theoretical Biology, 50(1):129–151, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900284.

Ainsworth:1977:ITE

[Ain77a] Stanley Ainsworth. Isotope transfer at equilibrium. An application of the set nomenclature for mechanisms of enzymic catalysis. Journal of Theoretical Biology, 64(2):381–389, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903654.

Ainsworth:1977:SMR

[Ain77b] Stanley Ainsworth. A simple model for a regulatory enzyme. Journal of Theoretical Biology, 68(3):391–413, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900686.

Ainsworth:1979:EMR

[Ain79] Stanley Ainsworth. The exponential model for a regulatory enzyme: its extension to describe catastrophic changes in function. Journal of Theoretical Biology, 77(1):27–35, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901346.

Andrews:1979:CAM

[AJ79] P. R. Andrews and G. A. R. Johnston. Conformational analysis of muscimol, a GABA agonist. Journal of Theoretical Biology, 79(3):263–273, August 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990345X.

Aizawa:1976:TSO

[AK76] Yoji Aizawa and Yonosuke Kobatake. A theory of selforganization of an excitable membrane formed on the protoplasmic droplet isolated from nitella. Journal of Theoretical Biology, 60(2):393–399, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900655.

Argyrakis:1978:EPI

[AK78] P. Argyrakis and R. Kopelman. Exciton percolation III. Stochastic and coherent migration in binary and ternary random lattices. Journal of Theoretical Biology, 73(2):205–236, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890187X.

Alving:1972:EEO

[AL72] R. E. Alving and K. Laki. Energy and electron orbitals in adenosine phosphates. Journal of Theoretical Biology, 34(2):199–214, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901555.

Alarcon:1972:ACU

[Ala72] R. A. Alarcon. Acrolein, a component of a universal cellgrowth regulatory system: A theory. Journal of Theoretical Biology, 37(1):159–167, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901221.

Albanese:1971:UMC

[Alb71] Richard A. Albanese. Use of membrane-covered oxygen cathodes in tissue. Journal of Theoretical Biology, 33(1):91–103, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902189.

Albanese:1973:MDT

[Alb73] Richard A. Albanese. On microelectrode distortion of tissue oxygen tensions. Journal of Theoretical Biology, 38(1):143– 154, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902300.

Albano:1974:CMS

[Alb74] A. M. Albano. Cole-Moore superposition and kinetic models for the potassium conductance of nerve fiber membranes. Journal of Theoretical Biology, 45(1):275–282, May 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900551.

Alberghina:1975:MRG

[Alb75] F. A. M. Alberghina. A model for the regulation of growth in mammalian cells. Journal of Theoretical Biology, 55(2):533– 545, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801007.

Alberghina:1977:DCC

[Alb77] Lilia Alberghina. Dynamics of the cell cycle in mammalian cells. Journal of Theoretical Biology, 69(4):633–643, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903721.

Alexandrov:1972:PBC

[Ale72] V. Ya. Alexandrov. The problem of behaviour at the cellular level (cytoethology). Journal of Theoretical Biology, 35 (1):1–8, IN1–IN4, 9–26, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901890.

Alerstam:1979:OUW

[Ale79] Thomas Alerstam. Optimal use of wind by migrating birds: Combined drift and overcompensation. Journal of Theoretical Biology, 79(3):341–353, August 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903515.

Alfano:1974:OGF

[Alf74] Michael C. Alfano. The origin of gingival fluid. Journal of Theoretical Biology, 47(1):127–136, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901039.

Allen:1977:NUP

[All77] T. F. H. Allen. Neolithic urban primacy: The case against the invention of agriculture. Journal of Theoretical Biology, 66(1):169–180, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903186.

Allison:1975:MGS

[AM75] David Allison and Thomas McKearn. A mechanism for the generation of self-tolerance. Journal of Theoretical Biology, 54(2): 201–206, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801251.

Armstrong:1976:CTC

[AM76] R. A. Armstrong and R. McGehee. Coexistence of two competitors on one resource. Journal of Theoretical Biology, 56(2):499– 502, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800896.

Amati:1974:WMO

[Ama74] P. Amati. A working model for oncogenic DNA virus replication. Journal of Theoretical Biology, 46(1):221–227, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901489.

Amidon:1974:TCA

[Ami74] Gordon L. Amidon. Theoretical calculations on an acetic acid, 5-methylimidazole, methanol hydrogen bond network: A model charge relay system. Journal of Theoretical Biology, 46(1):101– 109, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901416.

Antonelli:1975:TPRa

[AMR^{+}75a] P. Antonelli, D. I. McLaren, T. D. Rogers, M. Lathrop, and M. A. Willard. Transitivity, pattern-reversal, engulfment and duality in exchange-type cell aggregation kinetics. Journal of Theoretical Biology, 49(1):385–400, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800427.

Antonelli:1975:TPRb

[AMR^{+}75b] P. Antonelli, D. I. McLaren, T. D. Rogers, M. Lathrop, and M. A. Willard. Transitivity, pattern-reversal, engulfment and duality in exchange-type cell aggregation kinetics. Journal of Theoretical Biology, 49(2):385–400, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901800.

Anagnostopoulos:1973:PCL

[Ana73] Takis Anagnostopoulos. The partial conductances of limiting membranes in epithelial tissues. Journal of Theoretical Biology, 42(1):177–179, November 5, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901562.

Anderson:1974:CEV

[And74a] Julius H. Anderson. Control of enzymatic velocity under near-equilibrium conditions. Journal of Theoretical Biology, 47(1):153–161, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901064.

Anderson:1974:OSE

[And74b] Julius H. Anderson. Optimizing the sensitivity of enzyme systems to perturbations from equilibrium. Journal of Theoretical Biology, 47(1):163–176, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901076.

Anderson:1976:IMSa

[And76a] Norman G. Anderson. Interactive macromolecular sites I. Basic theory. Journal of Theoretical Biology, 60(2):401–412, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900667.

Anderson:1976:IMSb

[And76b] Norman G. Anderson. Interactive macromolecular sites II. Role in prebiotic macromolecular selection and early cellular evolution. Journal of Theoretical Biology, 60(2):413–419, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900679.

Anderson:1977:BCS

[And77] Richard G. W. Anderson. The biogenesis of cell structures and the expression of assembly information. Journal of Theoretical Biology, 67(3):535–548, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900558.

Andersson:1979:SEM

[And79] G. Andersson. Self-electrophoresis of membrane molecules and chromosomes: Unifying hypothesis of cell cycle controls. Journal of Theoretical Biology, 77(1):1–18, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901322.

Anonymous:1970:Ac

[Ano70a] Anonymous. Announcemen. Journal of Theoretical Biology, 29(3):493, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901128.

Anonymous:1970:Ab

[Ano70b] Anonymous. Announcement. Journal of Theoretical Biology, 28(2):301, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900585.

Anonymous:1970:Aa

[Ano70c] Anonymous. Announcements. Journal of Theoretical Biology, 27(3):499–500, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937080011X.

Anonymous:1970:BDB

[Ano70d] Anonymous. Biochemical determination of bacterial morphology and the geometry of cell divison: E. P. Previc, J. theor. Biol., 27, 471–497. Journal of Theoretical Biology, 29(3):i, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901013.

Anonymous:1970:EBa

[Ano70e] Anonymous. Editorial Board. Journal of Theoretical Biology, 26 (1):i, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800248.

Anonymous:1970:EBb

[Ano70f] Anonymous. Editorial Board. Journal of Theoretical Biology, 27 (1):ifc, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901256.

Anonymous:1970:EBc

[Ano70g] Anonymous. Editorial Board. Journal of Theoretical Biology, 28(1):ifc, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900597.

Anonymous:1970:EBd

[Ano70h] Anonymous. Editorial Board. Journal of Theoretical Biology, 29(1):CO2, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937090113X.

Anonymous:1970:E

[Ano70i] Anonymous. Erratum. Journal of Theoretical Biology, 27(2):i, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901463.

Anonymous:1970:IV

[Ano70j] Anonymous. Index to volume 26. Journal of Theoretical Biology, 26(1):509–514, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937080025X.

Anonymous:1970:PJb

[Ano70k] Anonymous. Pages 1–154 (July 1970). Journal of Theoretical Biology, 28(1):??, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PO

[Ano70l] Anonymous. Pages 1–154 (October 1970). Journal of Theoretical Biology, 29(1):??, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PAa

[Ano70m] Anonymous. Pages 1–173 (April 1970). Journal of Theoretical Biology, 27(1):??, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PAb

[Ano70n] Anonymous. Pages 155–301 (August 1970). Journal of Theoretical Biology, 28(2):??, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PN

[Ano70o] Anonymous. Pages 155–328 (November 1970). Journal of Theoretical Biology, 29(2):??, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PMb

[Ano70p] Anonymous. Pages 175–342 (May 1970). Journal of Theoretical Biology, 27(2):??, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PF

[Ano70q] Anonymous. Pages 181–342 (February 1970). Journal of Theoretical Biology, 26(2):??, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PS

[Ano70r] Anonymous. Pages 303–546 (September 1970). Journal of Theoretical Biology, 28(3):??, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PD

[Ano70s] Anonymous. Pages 329–493 (December 1970). Journal of Theoretical Biology, 29(3):??, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:P

[Ano70t] Anonymous. Pages 343–500 (1970). Journal of Theoretical Biology, 27(3):??, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PMa

[Ano70u] Anonymous. Pages 343–508 (March 1970). Journal of Theoretical Biology, 26(3):??, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1970:PJa

[Ano70v] Anonymous. Pages 509–514, 1-179 (January 1970). Journal of Theoretical Biology, 26(1):??, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:Aa

[Ano71a] Anonymous. Announcement. Journal of Theoretical Biology, 30(1):223, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937190049X.

Anonymous:1971:Ab

[Ano71b] Anonymous. Announcement. Journal of Theoretical Biology, 30(2):427, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900634.

Anonymous:1971:Ad

[Ano71c] Anonymous. Announcement. Journal of Theoretical Biology, 32(2):427, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901792.

Anonymous:1971:Ae

[Ano71d] Anonymous. Announcement. Journal of Theoretical Biology, 33(3):605, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901007.

Anonymous:1971:Ac

[Ano71e] Anonymous. Announcements. Journal of Theoretical Biology, 32(1):217, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901494.

Anonymous:1971:CRC

[Ano71f] Anonymous. Chemical relaxation of cyclic enzyme systems IV. Thermodynamics of isomerase systems: G. H. Czerlinski, J. theor. Biol. (1971) 31, 493–499. Journal of Theoretical Biology, 33(1):i, October 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902268.

Anonymous:1971:EBa

[Ano71g] Anonymous. Editorial Board. Journal of Theoretical Biology, 30(1):ifc, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900324.

Anonymous:1971:EBb

[Ano71h] Anonymous. Editorial Board. Journal of Theoretical Biology, 31 (1):ifc, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901160.

Anonymous:1971:EBc

[Ano71i] Anonymous. Editorial Board. Journal of Theoretical Biology, 32(1):ifc, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901317.

Anonymous:1971:EBd

[Ano71j] Anonymous. Editorial Board. Journal of Theoretical Biology, 33(1):ifc, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371902128.

Anonymous:1971:E

[Ano71k] Anonymous. Errata. Journal of Theoretical Biology, 32(3):i, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901615.

Anonymous:1971:PAa

[Ano71l] Anonymous. Pages 1–169 (April 1971). Journal of Theoretical Biology, 31(1):??, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PJc

[Ano71m] Anonymous. Pages 1–217 (July 1971). Journal of Theoretical Biology, 32(1):??, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PJa

[Ano71n] Anonymous. Pages 1–223 (January 1971). Journal of Theoretical Biology, 30(1):??, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PO

[Ano71o] Anonymous. Pages 1–223 (October 1971). Journal of Theoretical Biology, 33(1):??, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PMb

[Ano71p] Anonymous. Pages 171–375 (May 1971). Journal of Theoretical Biology, 31(2):??, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PAb

[Ano71q] Anonymous. Pages 219–427 (August 1971). Journal of Theoretical Biology, 32(2):??, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PN

[Ano71r] Anonymous. Pages 225–418 (November 1971). Journal of Theoretical Biology, 33(2):??, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PF

[Ano71s] Anonymous. Pages 225–427 (February 1971). Journal of Theoretical Biology, 30(2):??, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PJb

[Ano71t] Anonymous. Pages 377–559 (June 1971). Journal of Theoretical Biology, 31(3):??, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PD

[Ano71u] Anonymous. Pages 419–605 (December 1971). Journal of Theoretical Biology, 33(3):??, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PS

[Ano71v] Anonymous. Pages 429–632 (September 1971). Journal of Theoretical Biology, 32(3):??, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:PMa

[Ano71w] Anonymous. Pages 429–645 (March 1971). Journal of Theoretical Biology, 30(3):??, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1971:TLI

[Ano71x] Anonymous. Two locus inbreeding functions: B. S. Weir, J. theor. Biol. (1971) 30, 423–425. Journal of Theoretical Biology, 32(2):i, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901809.

Anonymous:1972:Aa

[Ano72a] Anonymous. Announcement. Journal of Theoretical Biology, 34(2):405, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901701.

Anonymous:1972:Ab

[Ano72b] Anonymous. Announcement. Journal of Theoretical Biology, 35 (3):605, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901543.

Anonymous:1972:Ac

[Ano72c] Anonymous. Announcement. Journal of Theoretical Biology, 36(2):425, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901105.

Anonymous:1972:Ad

[Ano72d] Anonymous. Announcement. Journal of Theoretical Biology, 37(1):207, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901300.

Anonymous:1972:Ae

[Ano72e] Anonymous. Announcement. Journal of Theoretical Biology, 37(2):387, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937290032X.

Anonymous:1972:EBa

[Ano72f] Anonymous. Editorial Board. Journal of Theoretical Biology, 34(1):ifc, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900483.

Anonymous:1972:EBb

[Ano72g] Anonymous. Editorial Board. Journal of Theoretical Biology, 35 (1):ifc, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901889.

Anonymous:1972:EBc

[Ano72h] Anonymous. Editorial Board. Journal of Theoretical Biology, 36(1):ifc, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901713.

Anonymous:1972:EBd

[Ano72i] Anonymous. Editorial Board. Journal of Theoretical Biology, 37(1):ifc, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901117.

Anonymous:1972:PAa

[Ano72j] Anonymous. Pages 1–189 (April 1972). Journal of Theoretical Biology, 35(1):??, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PJa

[Ano72k] Anonymous. Pages 1–198 (January 1972). Journal of Theoretical Biology, 34(1):??, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PJc

[Ano72l] Anonymous. Pages 1–202 (July 1972). Journal of Theoretical Biology, 36(1):??, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PO

[Ano72m] Anonymous. Pages 1–207 (October 1972). Journal of Theoretical Biology, 37(1):??, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PMb

[Ano72n] Anonymous. Pages 191–397 (May 1972). Journal of Theoretical Biology, 35(2):??, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PF

[Ano72o] Anonymous. Pages 199–405 (February 1972). Journal of Theoretical Biology, 34(2):??, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PAb

[Ano72p] Anonymous. Pages 203–425 (August 1972). Journal of Theoretical Biology, 36(2):??, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PN

[Ano72q] Anonymous. Pages 209–387 (November 1972). Journal of Theoretical Biology, 37(2):??, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PD

[Ano72r] Anonymous. Pages 389–595 (December 1972). Journal of Theoretical Biology, 37(3):??, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PJb

[Ano72s] Anonymous. Pages 399–605 (June 1972). Journal of Theoretical Biology, 35(3):??, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PMa

[Ano72t] Anonymous. Pages 407–590 (March 1972). Journal of Theoretical Biology, 34(3):??, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PS

[Ano72u] Anonymous. Pages 427–645 (September 1972). Journal of Theoretical Biology, 36(3):??, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1972:PCM

[Ano72v] Anonymous. A principle for chemical multivibration. Journal of Theoretical Biology, 36(2):413–417, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901087.

Anonymous:1973:A

[Ano73a] Anonymous. Announcement. Journal of Theoretical Biology, 40 (1):205, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901744.

Anonymous:1973:EBa

[Ano73b] Anonymous. Editorial Board. Journal of Theoretical Biology, 38(1):ifc, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902208.

Anonymous:1973:EBb

[Ano73c] Anonymous. Editorial Board. Journal of Theoretical Biology, 39 (1):ifc, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902002.

Anonymous:1973:EBc

[Ano73d] Anonymous. Editorial Board. Journal of Theoretical Biology, 40(1):ifc, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901598.

Anonymous:1973:EBd

[Ano73e] Anonymous. Editorial Board. Journal of Theoretical Biology, 41(1):ifc, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901859.

Anonymous:1973:EBe

[Ano73f] Anonymous. Editorial Board. Journal of Theoretical Biology, 42(1):ifc, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901446.

Anonymous:1973:EC

[Ano73g] Anonymous. Erratum and comments. Journal of Theoretical Biology, 39(3):i, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900842.

Anonymous:1973:PNa

[Ano73h] Anonymous. Pages 1–187 (5 November 1973). Journal of Theoretical Biology, 42(1):??, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PSa

[Ano73i] Anonymous. Pages 1–200 (14 September 1973). Journal of Theoretical Biology, 41(1):??, September 14, 1973. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PJa

[Ano73j] Anonymous. Pages 1–204 (January 1973). Journal of Theoretical Biology, 38(1):??, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PJc

[Ano73k] Anonymous. Pages 1–205 (July 1973). Journal of Theoretical Biology, 40(1):??, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PAa

[Ano73l] Anonymous. Pages 1–248 (April 1973). Journal of Theoretical Biology, 39(1):??, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PNb

[Ano73m] Anonymous. Pages 189–395 (15 November 1973). Journal of Theoretical Biology, 42(2):??, November 15, 1973. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PSb

[Ano73n] Anonymous. Pages 201–399 (21 September 1973). Journal of Theoretical Biology, 41(2):??, September 21, 1973. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PF

[Ano73o] Anonymous. Pages 205–422 (February 1973). Journal of Theoretical Biology, 38(2):??, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PAb

[Ano73p] Anonymous. Pages 207–407 (15 August 1973). Journal of Theoretical Biology, 40(2):??, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PMb

[Ano73q] Anonymous. Pages 249–486 (May 1973). Journal of Theoretical Biology, 39(2):??, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PD

[Ano73r] Anonymous. Pages 397–596 (December 1973). Journal of Theoretical Biology, 42(3):??, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PO

[Ano73s] Anonymous. Pages 401–569 (October 1973). Journal of Theoretical Biology, 41(3):??, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PAc

[Ano73t] Anonymous. Pages 409–609 (22 August 1973). Journal of Theoretical Biology, 40(3):??, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PMa

[Ano73u] Anonymous. Pages 423–679 (March 1973). Journal of Theoretical Biology, 38(3):??, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1973:PJb

[Ano73v] Anonymous. Pages 487–683 (June 1973). Journal of Theoretical Biology, 39(3):??, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:Aa

[Ano74a] Anonymous. Announcement. Journal of Theoretical Biology, 43(1):227, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937480055X.

Anonymous:1974:Ab

[Ano74b] Anonymous. Announcement. Journal of Theoretical Biology, 47(1):251, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901143.

Anonymous:1974:EBa

[Ano74c] Anonymous. Editorial Board. Journal of Theoretical Biology, 43 (1):i, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937480038X.

Anonymous:1974:EBb

[Ano74d] Anonymous. Editorial Board. Journal of Theoretical Biology, 44(1):i, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800226.

Anonymous:1974:EBc

[Ano74e] Anonymous. Editorial Board. Journal of Theoretical Biology, 45(1):ifc, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900423.

Anonymous:1974:EBd

[Ano74f] Anonymous. Editorial Board. Journal of Theoretical Biology, 46(1):ifc, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901362.

Anonymous:1974:EBe

[Ano74g] Anonymous. Editorial Board. Journal of Theoretical Biology, 47(1):ifc, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374900952.

Anonymous:1974:EBf

[Ano74h] Anonymous. Editorial Board. Journal of Theoretical Biology, 48(1):ifc, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901738.

Anonymous:1974:Ea

[Ano74i] Anonymous. Erratum. Journal of Theoretical Biology, 45(1): 293, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900575.

Anonymous:1974:Ec

[Ano74j] Anonymous. Erratum. Journal of Theoretical Biology, 48(2): 501, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800214.

Anonymous:1974:IVa

[Ano74k] Anonymous. Index to volume 43. Journal of Theoretical Biology, 43(1):vii–xii, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800391.

Anonymous:1974:IVb

[Ano74l] Anonymous. Index to volume 44. Journal of Theoretical Biology, 44(1):vii–xi, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800238.

Anonymous:1974:PMa

[Ano74m] Anonymous. Pages 1–190 (March 1974). Journal of Theoretical Biology, 44(1):??, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PJa

[Ano74n] Anonymous. Pages 1–227 (January 1974). Journal of Theoretical Biology, 43(1):??, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PS

[Ano74o] Anonymous. Pages 1–251 (September 1974). Journal of Theoretical Biology, 47(1):??, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PN

[Ano74p] Anonymous. Pages 1–256 (November 1974). Journal of Theoretical Biology, 48(1):??, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PMb

[Ano74q] Anonymous. Pages 1–293 (May 1974). Journal of Theoretical Biology, 45(1):??, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PJc

[Ano74r] Anonymous. Pages 1–294 (July 1974). Journal of Theoretical Biology, 46(1):??, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PAa

[Ano74s] Anonymous. Pages 191–404 (April 1974). Journal of Theoretical Biology, 44(2):??, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PF

[Ano74t] Anonymous. Pages 229–396 (February 1974). Journal of Theoretical Biology, 43(2):??, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PO

[Ano74u] Anonymous. Pages 253–489 (October 1974). Journal of Theoretical Biology, 47(2):??, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:P

[Ano74v] Anonymous. Pages 257–501 (1974). Journal of Theoretical Biology, 48(2):??, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PAb

[Ano74w] Anonymous. Pages 295–589 (August 1974). Journal of Theoretical Biology, 46(2):??, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1974:PJb

[Ano74x] Anonymous. Pages 295–597 (June 1974). Journal of Theoretical Biology, 45(2):??, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:JSI

[Ano75a] Anonymous. The 9th Jerusalem symposium an international symposium on metal-ligand interactions in organic and biochemistry. Journal of Theoretical Biology, 52(2):500, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590017X.

Anonymous:1975:ADN

[Ano75b] Anonymous. An algorithm for detecting non-steady state moieties in steady state subnetworks: David J. M. Park, J. theor. Biol. (1974) 48 125–131. Journal of Theoretical Biology, 50 (2):507, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900971.

Anonymous:1975:Aa

[Ano75c] Anonymous. Announcemen. Journal of Theoretical Biology, 50 (2):509, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900983.

Anonymous:1975:Ab

[Ano75d] Anonymous. Announcement. Journal of Theoretical Biology, 52(2):499, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900168.

Anonymous:1975:EBa

[Ano75e] Anonymous. Editorial Board. Journal of Theoretical Biology, 50 (1):ifc, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900181.

Anonymous:1975:EBb

[Ano75f] Anonymous. Editorial Board. Journal of Theoretical Biology, 51(1):ifc, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901344.

Anonymous:1975:EBc

[Ano75g] Anonymous. Editorial Board. Journal of Theoretical Biology, 52(1):ifc, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900363.

Anonymous:1975:EBd

[Ano75h] Anonymous. Editorial Board. Journal of Theoretical Biology, 53(1):ifc, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900995.

Anonymous:1975:EBe

[Ano75i] Anonymous. Editorial Board. Journal of Theoretical Biology, 55(1):i, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801032.

Anonymous:1975:E

[Ano75j] Anonymous. Errata. Journal of Theoretical Biology, 51(1):241– 242, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901514.

Anonymous:1975:IV

[Ano75k] Anonymous. Index to volume 55. Journal of Theoretical Biology, 55(1):vii–xi, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801044.

Anonymous:1975:Pd

[Ano75l] Anonymous. Pages 1–128 (1975). Journal of Theoretical Biology, 54(1):??, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PMb

[Ano75m] Anonymous. Pages 1–242 (May 1975). Journal of Theoretical Biology, 51(1):??, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PS

[Ano75n] Anonymous. Pages 1–245 (September 1975). Journal of Theoretical Biology, 53(1):??, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PMa

[Ano75o] Anonymous. Pages 1–252 (March 1975). Journal of Theoretical Biology, 50(1):??, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PJb

[Ano75p] Anonymous. Pages 1–257 (July 1975). Journal of Theoretical Biology, 52(1):??, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PN

[Ano75q] Anonymous. Pages 1–282 (November 1975). Journal of Theoretical Biology, 55(1):??, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:Pb

[Ano75r] Anonymous. Pages 1–497 (1975). Journal of Theoretical Biology, 49(1):??, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:Pe

[Ano75s] Anonymous. Pages 129–401 (1975). Journal of Theoretical Biology, 54(2):??, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PJa

[Ano75t] Anonymous. Pages 243–524 (June 1975). Journal of Theoretical Biology, 51(2):??, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:Pc

[Ano75u] Anonymous. Pages 247–501 (1975). Journal of Theoretical Biology, 53(2):??, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PF

[Ano75v] Anonymous. Pages 249–497 (February 1975). Journal of Theoretical Biology, 49(2):??, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PAa

[Ano75w] Anonymous. Pages 253–509 (April 1975). Journal of Theoretical Biology, 50(2):??, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PAb

[Ano75x] Anonymous. Pages 259–500 (August 1975). Journal of Theoretical Biology, 52(2):??, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1975:PDa

[Ano75y] Anonymous. Pages 283–557 (December 1975). Journal of Theoretical Biology, 55(2):??, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:EBa

[Ano76a] Anonymous. Editorial Board. Journal of Theoretical Biology, 56 (1):i, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800410.

Anonymous:1976:EBb

[Ano76b] Anonymous. Editorial Board. Journal of Theoretical Biology, 57(1):i, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937680001X.

Anonymous:1976:EBc

[Ano76c] Anonymous. Editorial Board. Journal of Theoretical Biology, 58 (1):ifc, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937690134X.

Anonymous:1976:EBd

[Ano76d] Anonymous. Editorial Board. Journal of Theoretical Biology, 59(1):i, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800197.

Anonymous:1976:EBe

[Ano76e] Anonymous. Editorial Board. Journal of Theoretical Biology, 60 (1):ifc, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937690151X.

Anonymous:1976:EBf

[Ano76f] Anonymous. Editorial Board. Journal of Theoretical Biology, 61(1):ifc, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900989.

Anonymous:1976:EBg

[Ano76g] Anonymous. Editorial Board. Journal of Theoretical Biology, 62(1):ifc, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900485.

Anonymous:1976:EBh

[Ano76h] Anonymous. Editorial Board. Journal of Theoretical Biology, 63(1):ifc, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900795.

Anonymous:1976:IVa

[Ano76i] Anonymous. Index to volume 56. Journal of Theoretical Biology, 56(1):v–vii, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800422.

Anonymous:1976:IVb

[Ano76j] Anonymous. Index to volume 57. Journal of Theoretical Biology, 57(1):vii–xi, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800021.

Anonymous:1976:IVc

[Ano76k] Anonymous. Index to volume 59. Journal of Theoretical Biology, 59(1):vii–xi, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800203.

Anonymous:1976:NI

[Ano76l] Anonymous. News item. Journal of Theoretical Biology, 63(2): 243, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900321.

Anonymous:1976:PN

[Ano76m] Anonymous. Pages 1–241 (November 1976). Journal of Theoretical Biology, 63(1):??, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PJd

[Ano76n] Anonymous. Pages 1–245 (21 July 1976). Journal of Theoretical Biology, 60(1):??, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PSa

[Ano76o] Anonymous. Pages 1–248 (7 September 1976). Journal of Theoretical Biology, 61(1):??, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PMa

[Ano76p] Anonymous. Pages 1–248 (March 1976). Journal of Theoretical Biology, 57(1):??, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:POa

[Ano76q] Anonymous. Pages 1–251 (7 October 1976). Journal of Theoretical Biology, 62(1):??, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PJa

[Ano76r] Anonymous. Pages 1–251 (January 1976). Journal of Theoretical Biology, 56(1):??, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PMb

[Ano76s] Anonymous. Pages 1–252 (7 May 1976). Journal of Theoretical Biology, 58(1):??, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PJb

[Ano76t] Anonymous. Pages 1–252 (June 1976). Journal of Theoretical Biology, 59(1):??, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PD

[Ano76u] Anonymous. Pages 243–484 (December 1976). Journal of Theoretical Biology, 63(2):??, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PAb

[Ano76v] Anonymous. Pages 247–497 (7 August 1976). Journal of Theoretical Biology, 60(2):??, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PAa

[Ano76w] Anonymous. Pages 249–500 (April 1976). Journal of Theoretical Biology, 57(2):??, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PSb

[Ano76x] Anonymous. Pages 249–504 (21 September 1976). Journal of Theoretical Biology, 61(2):??, September 21, 1976. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PI

[Ano76y] Anonymous. Pages 253–370, IN1, 371-498 (1976). Journal of Theoretical Biology, 58(2):??, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PJc

[Ano76z] Anonymous. Pages 253–501 (7 July 1976). Journal of Theoretical Biology, 59(2):??, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:POb

[Ano76-27] Anonymous. Pages 253–503 (21 October 1976). Journal of Theoretical Biology, 62(2):??, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1976:PF

[Ano76-28] Anonymous. Pages 253–506 (February 1976). Journal of Theoretical Biology, 56(2):??, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:Aa

[Ano77a] Anonymous. Announcement. Journal of Theoretical Biology, 66 (2):i, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901692.

Anonymous:1977:Ab

[Ano77b] Anonymous. Announcement. Journal of Theoretical Biology, 67(3):607, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900625.

Anonymous:1977:EBa

[Ano77c] Anonymous. Editorial Board. Journal of Theoretical Biology, 64(1):ifc, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377901102.

Anonymous:1977:EBb

[Ano77d] Anonymous. Editorial Board. Journal of Theoretical Biology, 65(1):ifc, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900741.

Anonymous:1977:EBc

[Ano77e] Anonymous. Editorial Board. Journal of Theoretical Biology, 66 (1):ifc, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937790306X.

Anonymous:1977:EBd

[Ano77f] Anonymous. Editorial Board. Journal of Theoretical Biology, 67 (1):ifc, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901801.

Anonymous:1977:EBe

[Ano77g] Anonymous. Editorial Board. Journal of Theoretical Biology, 68(1):ifc, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902223.

Anonymous:1977:EBf

[Ano77h] Anonymous. Editorial Board. Journal of Theoretical Biology, 69(1):ifc, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903836.

Anonymous:1977:Ea

[Ano77i] Anonymous. Errata. Journal of Theoretical Biology, 65(3): i, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902211.

Anonymous:1977:Ec

[Ano77j] Anonymous. Erratum. Journal of Theoretical Biology, 69(1):i, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903988.

Anonymous:1977:PSa

[Ano77k] Anonymous. Pages 1–165 (7 September 1977). Journal of Theoretical Biology, 68(1):??, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJe

[Ano77l] Anonymous. Pages 1–180 (7 July 1977). Journal of Theoretical Biology, 67(1):??, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJa

[Ano77m] Anonymous. Pages 1–197 (7 January 1977). Journal of Theoretical Biology, 64(1):??, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PMc

[Ano77n] Anonymous. Pages 1–202 (7 May 1977). Journal of Theoretical Biology, 66(1):??, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PMa

[Ano77o] Anonymous. Pages 1–213 (7 March 1977). Journal of Theoretical Biology, 65(1):??, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PNa

[Ano77p] Anonymous. Pages 1–237 (7 November 1977). Journal of Theoretical Biology, 69(1):??, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PSb

[Ano77q] Anonymous. Pages 167–320 (21 September 1977). Journal of Theoretical Biology, 68(2):??, September 21, 1977. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJf

[Ano77r] Anonymous. Pages 181–336 (21 July 1977). Journal of Theoretical Biology, 67(2):??, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJb

[Ano77s] Anonymous. Pages 199–399 (21 January 1977). Journal of Theoretical Biology, 64(2):??, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PMd

[Ano77t] Anonymous. Pages 203–398 (21 May 1977). Journal of Theoretical Biology, 66(2):??, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PMb

[Ano77u] Anonymous. Pages 215–398 (21 March 1977). Journal of Theoretical Biology, 65(2):??, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PNb

[Ano77v] Anonymous. Pages 239–389 (21 November 1977). Journal of Theoretical Biology, 69(2):??, November 21, 1977. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:POa

[Ano77w] Anonymous. Pages 321–458 (7 October 1977). Journal of Theoretical Biology, 68(3):??, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PAc

[Ano77x] Anonymous. Pages 337–607 (7 August 1977). Journal of Theoretical Biology, 67(3):??, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PDa

[Ano77y] Anonymous. Pages 391–562 (7 December 1977). Journal of Theoretical Biology, 69(3):??, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJc

[Ano77z] Anonymous. Pages 399–594 (7 June 1977). Journal of Theoretical Biology, 66(3):??, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PAa

[Ano77-27] Anonymous. Pages 399–607 (7 April 1977). Journal of Theoretical Biology, 65(3):??, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PFa

[Ano77-28] Anonymous. Pages 401–601 (7 February 1977). Journal of Theoretical Biology, 64(3):??, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:POb

[Ano77-29] Anonymous. Pages 459–612 (21 October 1977). Journal of Theoretical Biology, 68(4):??, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PDb

[Ano77-30] Anonymous. Pages 563–769 (21 December 1977). Journal of Theoretical Biology, 69(4):??, December 21, 1977. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PJd

[Ano77-31] Anonymous. Pages 595–814 (21 June 1977). Journal of Theoretical Biology, 66(4):??, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PFb

[Ano77-32] Anonymous. Pages 603–803 (21 February 1977). Journal of Theoretical Biology, 64(4):??, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PAd

[Ano77-33] Anonymous. Pages 609–773 (22 August 1977). Journal of Theoretical Biology, 67(4):??, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1977:PAb

[Ano77-34] Anonymous. Pages 609–799 (21 April 1977). Journal of Theoretical Biology, 65(4):??, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:EBa

[Ano78a] Anonymous. Editorial Board. Journal of Theoretical Biology, 70(1):ifc, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902977.

Anonymous:1978:EBb

[Ano78b] Anonymous. Editorial Board. Journal of Theoretical Biology, 71(1):ifc, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902084.

Anonymous:1978:EBc

[Ano78c] Anonymous. Editorial Board. Journal of Theoretical Biology, 72 (1):ifc, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378900139.

Anonymous:1978:EBd

[Ano78d] Anonymous. Editorial Board. Journal of Theoretical Biology, 73 (1):ifc, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901765.

Anonymous:1978:EBe

[Ano78e] Anonymous. Editorial Board. Journal of Theoretical Biology, 74(1):ifc, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902849.

Anonymous:1978:EBf

[Ano78f] Anonymous. Editorial Board. Journal of Theoretical Biology, 75(1):ifc, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901984.

Anonymous:1978:E

[Ano78g] Anonymous. Erratum. Journal of Theoretical Biology, 70(1): i, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903065.

Anonymous:1978:NRC

[Ano78h] Anonymous. News release creation of the Latin American society of biomathematics. Journal of Theoretical Biology, 74(3):335, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902175.

Anonymous:1978:PJa

[Ano78i] Anonymous. Pages 1–134 (7 January 1978). Journal of Theoretical Biology, 70(1):??, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PNa

[Ano78j] Anonymous. Pages 1–140 (7 November 1978). Journal of Theoretical Biology, 75(1):??, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PSa

[Ano78k] Anonymous. Pages 1–154 (20 September 1978). Journal of Theoretical Biology, 74(1):??, September 20, 1978. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PMa

[Ano78l] Anonymous. Pages 1–166 (7 March 1978). Journal of Theoretical Biology, 71(1):??, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PJe

[Ano78m] Anonymous. Pages 1–204 (6 July 1978). Journal of Theoretical Biology, 73(1):??, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PMc

[Ano78n] Anonymous. Pages 1–204 (8 May 1978). Journal of Theoretical Biology, 72(1):??, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PJb

[Ano78o] Anonymous. Pages 135–252 (20 January 1978). Journal of Theoretical Biology, 70(2):??, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PNb

[Ano78p] Anonymous. Pages 141–247 (21 November 1978). Journal of Theoretical Biology, 75(2):??, November 21, 1978. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PSb

[Ano78q] Anonymous. Pages 155–334 (21 September 1978). Journal of Theoretical Biology, 74(2):??, September 21, 1978. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PMb

[Ano78r] Anonymous. Pages 167–278 (20 March 1978). Journal of Theoretical Biology, 71(2):??, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PJf

[Ano78s] Anonymous. Pages 205–366 (20 July 1978). Journal of Theoretical Biology, 73(2):??, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PMd

[Ano78t] Anonymous. Pages 205–376 (22 May 1978). Journal of Theoretical Biology, 72(2):??, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PDa

[Ano78u] Anonymous. Pages 249–390 (7 December 1978). Journal of Theoretical Biology, 75(3):??, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PFa

[Ano78v] Anonymous. Pages 253–337 (6 February 1978). Journal of Theoretical Biology, 70(3):??, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PAa

[Ano78w] Anonymous. Pages 279–478 (6 April 1978). Journal of Theoretical Biology, 71(3):??, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:POa

[Ano78x] Anonymous. Pages 335–473 (7 October 1978). Journal of Theoretical Biology, 74(3):??, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PFb

[Ano78y] Anonymous. Pages 339–469 (20 February 1978). Journal of Theoretical Biology, 70(4):??, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PAc

[Ano78z] Anonymous. Pages 367–574 (8 August 1978). Journal of Theoretical Biology, 73(3):??, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PJc

[Ano78-27] Anonymous. Pages 377–588 (6 June 1978). Journal of Theoretical Biology, 72(3):??, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PDb

[Ano78-28] Anonymous. Pages 391–536 (21 December 1978). Journal of Theoretical Biology, 75(4):??, December 21, 1978. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:POb

[Ano78-29] Anonymous. Pages 475–606 (21 October 1978). Journal of Theoretical Biology, 74(4):??, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PAb

[Ano78-30] Anonymous. Pages 479–667 (20 April 1978). Journal of Theoretical Biology, 71(4):??, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PAd

[Ano78-31] Anonymous. Pages 575–800 (21 August 1978). Journal of Theoretical Biology, 73(4):??, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:PJd

[Ano78-32] Anonymous. Pages 589–759 (20 June 1978). Journal of Theoretical Biology, 72(4):??, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1978:SPS

[Ano78-33] Anonymous. Satellite program in statistical ecolog. Journal of Theoretical Biology, 71(4):669, April 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903314.

Anonymous:1979:EBa

[Ano79a] Anonymous. Editorial Board. Journal of Theoretical Biology, 76(1):ifc, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379903722.

Anonymous:1979:EBb

[Ano79b] Anonymous. Editorial Board. Journal of Theoretical Biology, 77(1):ifc, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901310.

Anonymous:1979:EBc

[Ano79c] Anonymous. Editorial Board. Journal of Theoretical Biology, 78 (1):ifc, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903199.

Anonymous:1979:EBd

[Ano79d] Anonymous. Editorial Board. Journal of Theoretical Biology, 79 (1):ifc, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902522.

Anonymous:1979:EBe

[Ano79e] Anonymous. Editorial Board. Journal of Theoretical Biology, 80(1):ifc, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901735.

Anonymous:1979:EBf

[Ano79f] Anonymous. Editorial Board. Journal of Theoretical Biology, 81(1):ifc, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379900754.

Anonymous:1979:E

[Ano79g] Anonymous. Erratum. Journal of Theoretical Biology, 78(2): 308, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902728.

Anonymous:1979:MNM

[Ano79h] Anonymous. Membrane noise and microscopical significance of critical slopes of ramp voltages. Journal of Theoretical Biology, 80(4):589–593, October 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990095X.

Anonymous:1979:N

[Ano79i] Anonymous. Notices. Journal of Theoretical Biology, 80(2):305, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902145.

Anonymous:1979:PJa

[Ano79j] Anonymous. Pages 1–107 (7 January 1979). Journal of Theoretical Biology, 76(1):??, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PJe

[Ano79k] Anonymous. Pages 1–136 (7 July 1979). Journal of Theoretical Biology, 79(1):??, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PNa

[Ano79l] Anonymous. Pages 1–149 (7 November 1979). Journal of Theoretical Biology, 81(1):??, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PSa

[Ano79m] Anonymous. Pages 1–153 (7 September 1979). Journal of Theoretical Biology, 80(1):??, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PMc

[Ano79n] Anonymous. Pages 1–165 (7 May 1979). Journal of Theoretical Biology, 78(1):??, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PMa

[Ano79o] Anonymous. Pages 1–170 (7 March 1979). Journal of Theoretical Biology, 77(1):??, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PJb

[Ano79p] Anonymous. Pages 109–231 (21 January 1979). Journal of Theoretical Biology, 76(2):??, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PJf

[Ano79q] Anonymous. Pages 137–262 (21 July 1979). Journal of Theoretical Biology, 79(2):??, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PNb

[Ano79r] Anonymous. Pages 151–375 (21 November 1979). Journal of Theoretical Biology, 81(2):??, November 21, 1979. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PSb

[Ano79s] Anonymous. Pages 155–305 (21 September 1979). Journal of Theoretical Biology, 80(2):??, September 21, 1979. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PMd

[Ano79t] Anonymous. Pages 167–308 (21 May 1979). Journal of Theoretical Biology, 78(2):??, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PMb

[Ano79u] Anonymous. Pages 171–234 (21 March 1979). Journal of Theoretical Biology, 77(2):??, March 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PFa

[Ano79v] Anonymous. Pages 233–367 (7 February 1979). Journal of Theoretical Biology, 76(3):??, February 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PAa

[Ano79w] Anonymous. Pages 235–384 (7 April 1979). Journal of Theoretical Biology, 77(3):??, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PAc

[Ano79x] Anonymous. Pages 263–414 (7 August 1979). Journal of Theoretical Biology, 79(3):??, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:POa

[Ano79y] Anonymous. Pages 307–443 (7 October 1979). Journal of Theoretical Biology, 80(3):??, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PJc

[Ano79z] Anonymous. Pages 309–450 (7 June 1979). Journal of Theoretical Biology, 78(3):??, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PFb

[Ano79-27] Anonymous. Pages 369–515 (21 February 1979). Journal of Theoretical Biology, 76(4):??, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PDa

[Ano79-28] Anonymous. Pages 377–620 (7 December 1979). Journal of Theoretical Biology, 81(3):??, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PAb

[Ano79-29] Anonymous. Pages 385–538 (21 April 1979). Journal of Theoretical Biology, 77(4):??, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PAd

[Ano79-30] Anonymous. Pages 415–553 (21 August 1979). Journal of Theoretical Biology, 79(4):??, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:POb

[Ano79-31] Anonymous. Pages 445–593 (21 October 1979). Journal of Theoretical Biology, 80(4):??, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PJd

[Ano79-32] Anonymous. Pages 451–610 (21 June 1979). Journal of Theoretical Biology, 78(4):??, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1979:PDb

[Ano79-33] Anonymous. Pages 621–822 (21 December 1979). Journal of Theoretical Biology, 81(4):??, December 21, 1979. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Antonelli:1974:RCG

[Ant74] Peter L. Antonelli. Reply to comments by N. Goel on the aggregation kinetics of Antonelli, Rogers and Willard. Journal of Theoretical Biology, 48(2):493–495, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800196.

AoNo:1972:ORP

[AO72] Osamu AoNo and Shinpei Ohki. Origin of resting potential of axon membrane. Journal of Theoretical Biology, 37(2):273–282, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900227.

Apffel:1970:RAE

[AP70] Charles A. Apffel and John H. Peters. Regulation of antigenic expression. Journal of Theoretical Biology, 26(1):47–59, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800315.

Alerstam:1977:WDM

[AP77] Thomas Alerstam and Sven-Göran Pettersson. Why do migrating birds fly along coastlines? Journal of Theoretical Biology, 65(4):699–712, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900169.

Amidon:1979:SCF

[APA79] G. L. Amidon, R. S. Pearlman, and S. T. Anik. The solvent contribution to the free energy of protein-ligand interactions. Journal of Theoretical Biology, 77(1):161–170, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901450.

Aylor:1975:CSR

[APK75] Donald E. Aylor, Jean-Yves Parlange, and Abraham D. Krikorian. Comment on some recent models of stomatal mechanics. Journal of Theoretical Biology, 54(2):395–397, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801421.

Applewhite:1972:FIL

[App72] Philip B. Applewhite. The flow of ions in learning and memory. Journal of Theoretical Biology, 36(2):419–423, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901099.

Arndt:1974:NPS

[AR74] R. A. Arndt and L. D. Roper. A new phenomenology for squid axon voltage-clamp currents. Journal of Theoretical Biology, 48(2):373–387, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937480007X.

Arndt:1975:TEM

[AR75] R. A. Arndt and L. D. Roper. Theory of excitable membranes. I. A simple model for a three-state artificial membrane. Journal of Theoretical Biology, 54(2):249–283, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801305.

Arcos:1971:UAM

[Arc71] Joseph C. Arcos. Ultrastructural alteration of the mitochondrial electron transport chain involving electron leak possible basis of “respiratory impairment” in certain tumors. Journal of Theoretical Biology, 30(3):533–543, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900063.

Antonelli:1973:GEP

[ARW73] P. L. Antonelli, T. D. Rogers, and M. A. Willard. Geometry and the exchange principle in cell aggregation kinetics. Journal of Theoretical Biology, 41(1):1–21, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901860.

Ashley:1972:HTM

[Ash72] Richard Ashley. A hybrid theory of muscle contraction. Journal of Theoretical Biology, 36(2):339–354, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901038.

Awerbuch:1979:QMD

[ASS79] T. E. Awerbuch, R. Samson, and A. J. Sinskey. A quantitative model of diffusion bioassays. Journal of Theoretical Biology, 79(3):333–336, IN2, 337–340, August 7, 1979.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903503.

Atema:1973:MTS

[Ate73] Jelle Atema. Microtubule theory of sensory transduction. Journal of Theoretical Biology, 38(1):181–190, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902336.

Atkins:1971:IST

[Atk71] Gordon L. Atkins. Investigation of some theoretical models relating the concentrations of glucose and insulin in plasma. Journal of Theoretical Biology, 32(3):471–494, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901524.

Atlan:1974:FDO

[Atl74] H. Atlan. On a formal definition of organization. Journal of Theoretical Biology, 45(2):295–304, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901155.

Aub:1970:AFC

[AW70] M. R. Aub and J. C. Waterlow. Analysis of a fivecompartment system with continuous infusion and its application to the study of amino acid turnover. Journal of Theoretical Biology, 26(2):243–250, February 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800157.

Adam:1975:MTA

[AW75] Donald E. Adam and James Wei. Mass transport of ATP within the motile sperm. Journal of Theoretical Biology, 49(1):125– 145, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800233.

Appleton:1977:ADC

[AWD77] David R. Appleton, Nicholas A. Wright, and Paul Dyson. The age distribution of cells in stratified squamous epithelium. Journal of Theoretical Biology, 65(4):769–779, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900224.

Babloyantz:1977:SOP

[Bab77] Agnessa Babloyantz. Self-organization phenomena resulting from cell-cell contact. Journal of Theoretical Biology, 68 (4):551–561, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377901059.

Bach:1974:MTE

[Bac74] Michael K. Bach. A molecular theory to explain the mechanisms of allergic histamine release. Journal of Theoretical Biology, 45(1):131–151, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374900472.

Bakken:1976:HTA

[Bak76] George S. Bakken. A heat transfer analysis of animals: Unifying concepts and the application of metabolism chamber data to field ecology. Journal of Theoretical Biology, 60(2):337–384, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900631.

Ball:1972:ISS

[Bal72] L. A. Ball. Implications of secondary structure in messenger RNA. Journal of Theoretical Biology, 36(2):313–320, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901002.

Ball:1973:MIS

[Bal73] L. Andrew Ball. Mutual influence of the secondary structure and information content of a messenger RNA. Journal of Theoretical Biology, 41(2):243–247, September 21, 1973.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901161.

Bard:1972:CHM

[Bar72a] Jonathan Bard. A comment on homeostatic mechanisms. Journal of Theoretical Biology, 37(2):385–386, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900318.

Barrett:1972:ORA

[Bar72b] Terence W. Barrett. Octave ratio as the absolute measurement unit of sensory communication in the cerebral cortex. Journal of Theoretical Biology, 34(1):87–97, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900562.

Barbieri:1974:CET

[Bar74a] Marcello Barbieri. A criterion to evaluate three dimensional reconstructions from projections of unknown structures. Journal of Theoretical Biology, 48(2):451–467, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800123.

Barbieri:1974:PRM

[Bar74b] Marcello Barbieri. The primitive ribosome model. Journal of Theoretical Biology, 47(2):269–280, October 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901970.

Barrett:1974:PMC

[Bar74c] J. C. Barrett. Parity models of cell proliferation. Journal of Theoretical Biology, 44(2):319–336, April 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901635.

Barlow:1976:TUB

[Bar76] Peter W. Barlow. Towards an understanding of the behaviour of root meristems. Journal of Theoretical Biology, 57(2):433– 451, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937690014X.

Bardsley:1977:FEK

[Bar77a] William G. Bardsley. The 3:3 function in enzyme kinetics possible shapes of vS and (1v)(1s) plots for third degree steady-state rate equations. Journal of Theoretical Biology, 65(2):281–316, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903265.

Bardsley:1977:FAB

[Bar77b] William G. Bardsley. Factorability of the allosteric binding polynomial and graphical manifestations of cooperativity in third degree saturation functions. Journal of Theoretical Biology, 67(3):407–431, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900467.

Bardsley:1977:RDA

[Bar77c] William G. Bardsley. The reduction in degree of allosteric and other complex rate equations using Sylvester’s dialytic method of elimination. Journal of Theoretical Biology, 67(1):121–139, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901886.

Barricelli:1977:OEG

[Bar77d] Nils Aall Barricelli. On the origin and evolution of the genetic code I. Wobbling and its potential significance. Journal of Theoretical Biology, 67(1):85–109, July 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901862.

Bard:1978:QML

[Bar78] J. B. L. Bard. A quantitative model of liver regeneration in the rat. Journal of Theoretical Biology, 73(3):509–530, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901558.

Bard:1979:QTL

[Bar79a] J. B. L. Bard. A quantitative theory of liver regeneration in the rat II: Matching an improved mitotic inhibitor model to the data. Journal of Theoretical Biology, 79(1):121–136, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902613.

Barrett:1979:PHI

[Bar79b] T. W. Barrett. Polyelectrolyte hysteresis interpreted as singularities in the Poisson–Boltzmann description. Journal of Theoretical Biology, 77(4):523–536, April 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900250.

Bass:1971:DSS

[Bas71] Ludvik Bass. Discrete steady states of neuronal membranes consistent with an evolutionary principle. Journal of Theoretical Biology, 33(3):577–597, December 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900981.

Basu:1973:TLD

[Bas73] S. Basu. Theory of linear dichroism of DNA and DNAcontaining structures. Journal of Theoretical Biology, 42(3):419– 441, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902397.

Bass:1974:CVR

[Bas74] Ludvik Bass. Current-voltage relations in nerve membranes. Journal of Theoretical Biology, 48(1):133–140, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901842.

Basu:1977:NSF

[Bas77] Samarendra Basu. A natural source of fluorescence depolarization in dye-DNA complexes. Journal of Theoretical Biology, 67(3):459–470, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900492.

Batke:1972:MAI

[Bat72] J. Batke. Modelling of allosteric interactions in dissociable tetrameric isoenzyme systems. Journal of Theoretical Biology, 34(2):313–324, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901646.

Bateson:1973:PFN

[Bat73] P. P. G. Bateson. Preferences for familiarity and novelty: A model for the simultaneous development of both. Journal of Theoretical Biology, 41(2):249–259, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901173.

Batzli:1974:PAA

[Bat74] George O. Batzli. Production, assimilation and accumulation of organic matter in ecosystems. Journal of Theoretical Biology, 45(1):205–217, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374900514.

Blumenson:1973:UMM

[BB73a] Leslie E. Blumenson and Irwin D. J. Bross. Use of a mathematical model to bridge the clinic-laboratory gap: Local spread of endometrial cancer. Journal of Theoretical Biology, 38(2):397– 411, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901823.

Brailsford:1973:RCM

[BB73b] J. Douglas Brailsford and Brian S. Bull. The red cell — a macromodel simulating the hypotonic-sphere isotonic-disc transformation. Journal of Theoretical Biology, 39(2):325–326, IN1, 327– 332, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901021.

Bass:1977:TDE

[BB77a] L. Bass and A. J. Bracken. Time-dependent elimination of substrates flowing through the liver or kidney. Journal of Theoretical Biology, 67(4):637–652, August 22, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790251X.

Berg:1977:MAR

[BB77b] Otto G. Berg and Clas Blomberg. Mass action relations in vivo with application to the Lac operon. Journal of Theoretical Biology, 67(3):523–533, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900546.

Barnard:1979:DHE

[BB79] C. J. Barnard and T. Burk. Dominance hierarchies and the evolution of “individual recognition”. Journal of Theoretical Biology, 81(1):65–73, November 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990081X.

Bender:1970:ARP

[BBG70] Robert Bender, S. H. Bellman, and Richard Gordon. ART and the ribosome: A preliminary report on the threedimensional structure of individual ribosomes determined by an algebraic reconstruction technique. Journal of Theoretical Biology, 29(3):483–484, IN5–IN10, 485–487, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901104.

Bellman:1971:ASB

[BBGR71] S. H. Bellman, Robert Bender, Richard Gordon, and James E. Rowe. ART is science being a defense of algebraic reconstruction techniques for three-dimensional electron microscopy. Journal of Theoretical Biology, 32(1):205–206, IN1–IN2, 207–216, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901482.

Bogyo:1970:IFS

[BBT70] Thomas P. Bogyo, Donald G. Burnell, and Shirley Wu Ting. Inbreeding function in small populations with an arbitrary number of linked loci. Journal of Theoretical Biology, 26(1):33– 45, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800303.

Bak:1979:HLO

[BBZ79] A. Leth Bak, P. Bak, and J. Zeuthen. Higher levels of organization in chromosomes. Journal of Theoretical Biology, 76(2):205–217, January 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903709.

Blumenfeld:1973:TEB

[BC73a] L. A. Blumenfeld and D. S. Chernavskii. Tunnelling of electrons in biological processes. Journal of Theoretical Biology, 39(1): 1–7, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902014.

Burton:1973:BCB

[BC73b] A. C. Burton and P. B. Canham. The behaviour of coupled biochemical oscillators as a model of contact inhibition of cellular division. Journal of Theoretical Biology, 39(3):555–580, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900684.

Bush:1976:ETD

[BC76] A. W. Bush and A. E. Cool. The effect of time delay and growth rate inhibition in the bacterial treatment of wastewater. Journal of Theoretical Biology, 63(2):385–395, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900412.

Bremer:1977:ECH

[BC77] H. Bremer and G. Churchward. An examination of the Cooper-helmstetter theory of DNA replication in bacteria and its underlying assumptions. Journal of Theoretical Biology, 69(4):645–654, December 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903733.

Bishop:1978:GWA

[BC78a] D. T. Bishop and C. Cannings. A generalized war of attrition. Journal of Theoretical Biology, 70(1):85–124, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903041.

Bremer:1978:AFB

[BC78b] Hans Bremer and Gordon Churchward. Age fractionation in bacteria by membrane elution: Relation between age distribution and elution profile. Journal of Theoretical Biology, 74(1):69–81, September 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902904.

Bongrand:1976:SP

[BCBD76] P. Bongrand, C. Capo, A. M. Benoliel, and R. Depieds. Specialization in phagocytosis. Journal of Theoretical Biology, 58(1):197–205, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901478.

Brosio:1977:SMP

[BCDK77] E. Brosio, F. Conti, A. Di Nola, and A. L. Kovacs. A stereochemical model for phospholipids: I. Conformational energy refinement and molecular packing of l-α-dipalmitoyl lecithin (DPL). Journal of Theoretical Biology, 67(2):319–334, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902041.

Barker:1973:QMB

[BCH73] S. B. Barker, G. Cumming, and K. Horsfield. Quantitative morphometry of the branching structure of trees. Journal of Theoretical Biology, 40(1):33–43, July 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390163X.

Broch:1976:CAA

[BCLV76] H. Broch, R. Cornillon, J. N. Lespinasse, and D. Vasilescu. Conformational analysis of aminoacyladenylates II. EHT study of glycyladenylate. Journal of Theoretical Biology, 57(2):407– 418, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900126.

Bishop:1978:WAR

[BCS78] D. T. Bishop, C. Cannings, and J. Maynard Smith. The war of attrition with random rewards. Journal of Theoretical Biology, 74(3):377–388, October 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902205.

Bremer:1979:RBG

[BCY79] Hans Bremer, Gordon Churchward, and Ry Young. Relation between growth and replication in bacteria. Journal of Theoretical Biology, 81(3):533–545, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900511.

Bischoff:1970:GSL

[BD70] Kenneth B. Bischoff and Robert L. Dedrick. Generalized solution to linear, two-compartment, open model for drug distribution. Journal of Theoretical Biology, 29(1):63–83, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901190.

Bhargava:1974:GSD

[BD74] T. N. Bhargava and P. H. Doyle. A geometric study of diversity. Journal of Theoretical Biology, 43(2):241–251, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800573.

Balslev:1975:SIS

[BD75a] I. Balslev and H. Degn. Spatial instability in simple reaction schemes. Journal of Theoretical Biology, 49(1):173–177, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800257.

Bremer:1975:TPF

[BD75b] Hans Bremer and Patrick P. Dennis. Transition period following a nutritional shift-up in the bacterium Escherichia coli B/r: Stable RNA and protein synthesis. Journal of Theoretical Biology, 52(2):365–382, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900077.

Bigelow:1973:SAR

[BDS73] J. H. Bigelow, J. C. DeHaven, and M. L. Shapley. Systems analysis of the renal function. Journal of Theoretical Biology, 41(2):287–322, September 21, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901203.

Beach:1979:TMP

[Bea79] Kirk W. Beach. A theoretical model to predict the behavior of glycosylated hemoglobin levels. Journal of Theoretical Biology, 81(3):547–561, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900523.

Becker:1972:TOM

[Bec72] Yechieil Becker. Theory of the organization of the mammalian cell and viral DNA genomes and the mechanism of carcinogenesis. Journal of Theoretical Biology, 34(1):47–59, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900537.

Beck:1978:EFT

[Bec78a] James S. Beck. Echinocyte formation: A test case for mechanisms of cell shape changes. Journal of Theoretical Biology, 71(4):515–524, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378903223.

Beck:1978:RBM

[Bec78b] James S. Beck. Relations between membrane monolayers in some red cell shape transformations. Journal of Theoretical Biology, 75(4):487–488, IN1–IN3, 489–501, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903582.

Becker:1978:MIR

[Bec78c] Yechiel Becker. A model for intramolecular recombination of herpesvirus DNA leading to the formation of circular, circular-linear and fragmented genomes. Journal of Theoretical Biology, 75(3):339–342, IN1–IN2, 343–347, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903399.

Beddington:1974:ADS

[Bed74] J. R. Beddington. Age distribution and the stability of simple discrete time population models. Journal of Theoretical Biology, 47(1):65–74, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490099X.

Bell:1970:MMC

[Bel70] George I. Bell. Mathematical model of clonal selection and antibody production. Journal of Theoretical Biology, 29(2):191– 232, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900196.

Bell:1971:MMCa

[Bel71a] George I. Bell. Mathematical model of clonal selection and antibody production. II. Journal of Theoretical Biology, 33(2):339– 378, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900713.

Bell:1971:MMCb

[Bel71b] George I. Bell. Mathematical model of clonal selection and antibody production III. The cellular basis of immunological paralysis. Journal of Theoretical Biology, 33(2):379–398, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900725.

Bello:1977:FEM

[Bel77a] Jake Bello. On the free energy of mixing of hydrophobic side chains in protein. Journal of Theoretical Biology, 67(2):335– 336, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902053.

Bello:1977:SPC

[Bel77b] Jake Bello. Stability of protein conformation: Internal packing and enthalpy of fusion of model compounds. Journal of Theoretical Biology, 68(1):139–142, September 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902326.

Bell:1978:EA

[Bel78] Graham Bell. The evolution of anisogamy. Journal of Theoretical Biology, 73(2):247–270, July 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901893.

Benford:1978:FTS

[Ben78a] Frank A. Benford. Fisher’s theory of the sex ratio applied to the social hymenoptera. Journal of Theoretical Biology, 72(4):701– 727, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902795.

Bengtsson:1978:AIW

[Ben78b] B. O. Bengtsson. Avoiding inbreeding: at what cost? Journal of Theoretical Biology, 73(3):439–444, August 8, 1978.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901510.

Bertran:1972:IBN

[Ber72] J. Bertran. Interactions between nucleic acid bases. Journal of Theoretical Biology, 34(2):353–361, February 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901671.

Berde:1976:NTS

[Ber76a] Charles Berde. Nonmonotonic transients and some mathematical models of circadian rhythms. Journal of Theoretical Biology, 56(2):435–436, IN1, 437–441, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800847.

Berg:1976:HSM

[Ber76b] Howard C. Berg. How spirochetes may swim. Journal of Theoretical Biology, 56(2):269–273, February 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800744.

Bernstein:1976:DAR

[Ber76c] Irwin S. Bernstein. Dominance, aggression and reproduction in primate societies. Journal of Theoretical Biology, 60(2):459– 472, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900722.

Berger:1977:CIT

[Ber77a] J. Berger. A comment on information theory and central dogma of molecular biology. Journal of Theoretical Biology, 65(2):393– 395, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903332.

Bernstein:1977:GLR

[Ber77b] Harris Bernstein. Germ line recombination may be primarily a manifestation of DNA repair processes. Journal of Theoretical Biology, 69(2):371–380, November 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901448.

Berg:1978:MSF

[Ber78] Otto G. Berg. A model for the statistical fluctuations of protein numbers in a microbial population. Journal of Theoretical Biology, 71(4):587–603, April 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903260.

Bess:1970:NMB

[Bes70] L. Bess. Network model as a biological pacemaker. Journal of Theoretical Biology, 28(1):59–80, July 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900640.

Bird:1975:EPV

[BFM75] J. F. Bird, R. W. Flower, and G. H. Mowbray. An elemental phenomenon of vision-suprafusion transients: General theory, retinal-cortical manifestations, potential application. Journal of Theoretical Biology, 55(2):553–557, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801020.

Bakken:1974:NHL

[BG74] G. S. Bakken and D. M. Gates. Notes on “Heat loss from a Newtonian animal”. Journal of Theoretical Biology, 45(1):283– 292, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900563.

Bisignani:1975:AAA

[BG75] W. T. Bisignani and S. C. Greenhouse. Automated analysis of autoradiographic imagery. Journal of Theoretical Biology, 54(1): 121–128, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800580.

Brockmann:1979:ESN

[BGD79] H. Jane Brockmann, Alan Grafen, and Richard Dawkins. Evolutionarily stable nesting strategy in a digger wasp. Journal of Theoretical Biology, 77(4):473–496, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900213.

Biswas:1977:BWP

[BGK77] S. N. Biswas, K. C. Gupta, and B. B. Karmakar. Brillouin– Wigner perturbation solution of Volterra’s prey-predator system. Journal of Theoretical Biology, 64(2):253–260, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903551.

Bruni:1976:DAD

[BGKS76] Carlo Bruni, Alfredo Germani, Giorgio Koch, and Roberto Strom. Derivation of antibody distribution from experimental binding data. Journal of Theoretical Biology, 61(1):143–170, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901107.

Bianchi:1970:OLP

[BH70] Luigi M. Bianchi and Joh R. Hamann. The origin of life: Preliminary considerations on necessary and (possibly) sufficient conditions. Journal of Theoretical Biology, 28(3):489–496, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900834.

Black:1977:AAC

[BH77] John A. Black and Richard N. Harkins. Amino acid compositions and evolutionary relationships with protein families. Journal of Theoretical Biology, 66(2):281–295, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901722.

Buchmann:1978:BMB

[BH78] Stephen L. Buchmann and James P. Hurley. A biophysical model for buzz pollination in angiosperms. Journal of Theoretical Biology, 72(4):639–657, June 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902771.

Berman:1979:CLL

[BH79] Stephen A. Berman and Sue Hanlon. Comparison of the linear least mean squares procedure and the vector projection method of spectral analysis. Journal of Theoretical Biology, 78(2):167– 183, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902625.

Bhargava:1977:RCD

[Bha77] Pushpa M. Bhargava. Regulation of cell division and malignant transformation: A new model for control by uptake of nutrients. Journal of Theoretical Biology, 68(1):101–137, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902314.

Bremer:1974:RRC

[BHD74] Hans Bremer, Jeffrey Hymes, and Patrick P. Dennis. Ribosomal RNA chain growth rate and RNA labeling patterns in Escherichia coli Br. Journal of Theoretical Biology, 45(2):379– 403, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901209.

Bretag:1974:SMC

[BHK74] A. H. Bretag, C. A. Hurst, and D. I. B. Kerr. The starzak membrane conductance model: A comment. Journal of Theoretical Biology, 48(2):477–479, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800159.

Bretag:1978:PCM

[BHK78] A. H. Bretag, C. A. Hurst, and D. I. B. Kerr. Potassium conductance models related to an interactive subunit membrane II: Monte Carlo analysis and subunit migration. Journal of Theoretical Biology, 73(3):367–382, August 8, 1978.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901480.

Bovee:1972:TPA

[BJ72] Eugene C. Bovee and Theodore L. Jahn. A theory of piezoelectric activity and ion-movements in the relation of flagellar structures and their movements to the phototaxis of euglena. Journal of Theoretical Biology, 35(2):259–276, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900380.

Bartle:1970:PMR

[BJM70] K. D. Bartle, D. W. Jones, and R. S. Matthews. Proton magnetic resonance coupling constants and π-bond orders in the active regions of carcinogenic and non-carcinogenic polycyclic hydrocarbons. Journal of Theoretical Biology, 27(1):117– 126, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901335.

Buysman:1971:ICP

[BK71] John R. Buysman and Frank T. Koide. Ion concentration profile normal to cell membrane. Journal of Theoretical Biology, 32(1): 1–23, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901329.

Benham:1976:DEC

[BK76] Craig J. Benham and John J. Kozak. Denaturation: an example of a catastrophe II. Two-state transitions. Journal of Theoretical Biology, 63(1):125–149, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900898.

Benham:1977:DEC

[BK77a] Craig J. Benham and John J. Kozak. Denaturation: An example of a catastrophe III. Phase diagrams for multistate transformations. Journal of Theoretical Biology, 66(4):679–693, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902387.

Burns:1977:ARA

[BK77b] J. A. Burns and H. Kacser. Allosteric repression: An analysis. Journal of Theoretical Biology, 68(2):199–213, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790159X.

Brailsford:1976:RCS

[BKB76] J. Douglas Brailsford, Ralph A. Korpman, and Brian S. Bull. The red cell shape from discocyte to hypotonic spherocyte — a mathematical delineation based on a uniform shell hypothesis. Journal of Theoretical Biology, 60(1):131–145, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901594.

Bass:1976:EES

[BKWT76] L. Bass, S. Keiding, K. Winkler, and N. Tygstrup. Enzymatic elimination of substrates flowing through the intact liver. Journal of Theoretical Biology, 61(2):393–409, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900266.

Balslev:1971:TAB

[BL71] Erik Balslev and K. N. Leibovic. Theoretical analysis of binocular space perception. Journal of Theoretical Biology, 31(1): 77–100, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901238.

Bard:1974:HWD

[BL74] Jonathan Bard and Ian Lauder. How well does Turing’s theory of morphogenesis work? Journal of Theoretical Biology, 45(2): 501–531, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901283.

Bailey:1975:AMT

[BL75] J. E. Bailey and C. K. Lee. Active membrane transport driven by dynamic asymmetry. Journal of Theoretical Biology, 50(2): 309–316, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900831.

Blake:1974:HCM

[Bla74] John Blake. Hydrodynamic calculations on the movements of cilia and flagella I. Paramecium. Journal of Theoretical Biology, 45(1):183–203, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374900502.

Blake:1975:HCM

[Bla75a] John Blake. Hydrodynamic calculations on the movements of cilia and flagella. II. Opalina. Journal of Theoretical Biology, 52 (1):67–82, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900405.

Blank:1975:MCB

[Bla75b] Martin Blank. A model for calculating the Bohr effect in hemoglobin equilibria. Journal of Theoretical Biology, 51(1): 127–134, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901435.

Blake:1977:MMT

[Bla77a] Iván Ortega Blake. A molecular mechanism for transcription of a closed DNA molecule. Journal of Theoretical Biology, 64(3):557–569, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902880.

Blake:1977:APM

[Bla77b] J. R. Blake. An active porous medium model for ciliary propulsion. Journal of Theoretical Biology, 64(4):697–701, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902685.

Blake:1978:HP

[Bla78a] J. R. Blake. On the hydrodynamics of plasmodesmata. Journal of Theoretical Biology, 74(1):33–47, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902886.

Blanquet:1978:UHM

[Bla78b] Pierre R. Blanquet. Unifying heuristic model of transmembrane co-ordinate control for cell growth and cell movement. Journal of Theoretical Biology, 70(4):345–399, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902485.

Blanquet:1979:ECC

[Bla79] Pierre R. Blanquet. Effect of charge changes in ligandtriggered membrane cooperative transitions. Journal of Theoretical Biology, 81(3):587–598, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900560.

Bleecken:1971:RCI

[Ble71] Stefan Bleecken. “replisome”-controlled initiation of DNA replication. Journal of Theoretical Biology, 32(1):81–92, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901378.

Blick:1977:STW

[Bli77] J. Blick. Selection for traits which lower individual reproduction. Journal of Theoretical Biology, 67(3):597–601, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900601.

Blomberg:1977:KRP

[Blo77] Clas Blomberg. A kinetic recognition process for tRNA at the ribosome. Journal of Theoretical Biology, 66(2):307–325, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901746.

Butler:1976:PBF

[BLP76] L. Allan Butler, Lawrence D. Longo, and Gordon G. Power. Placental blood flows and oxygen transfer during uterine contractions: A mathematical model. Journal of Theoretical Biology, 61(1):81–95, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901065.

Blumenson:1970:RWS

[Blu70] Leslie E. Blumenson. Random walk and the spread of cancer. Journal of Theoretical Biology, 27(2):273–290, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901426.

Blum:1971:EBP

[Blu71] J. J. Blum. Existence of a breaking point in cilia and flagella. Journal of Theoretical Biology, 33(2):257–263, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900658.

Blum:1973:BSV

[Blu73] Harry Blum. Biological shape and visual science (part i). Journal of Theoretical Biology, 38(2):205–287, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901756.

Blum:1974:UGC

[Blu74a] Harold F. Blum. Uncertainty of growth of cell populations in cancer. Journal of Theoretical Biology, 46(1):143–166, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901453.

Blum:1974:SPM

[Blu74b] Haywood Blum. Stochastic processes in messenger RNA turnover. Journal of Theoretical Biology, 48(1):161–171, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901878.

Blum:1974:NFT

[Blu74c] J. J. Blum. A note on fluid transport in ciliated tubules. Journal of Theoretical Biology, 46(1):287–290, July 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901544.

Blumenthal:1975:IOC

[Blu75] Robert Blumenthal. Instabilities, oscillations and chemical waves in an oligomeric model for membrane transport. Journal of Theoretical Biology, 49(1):219–239, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800300.

Blumenfeld:1976:PAE

[Blu76] L. A. Blumenfeld. The physical aspects of enzyme functioning. Journal of Theoretical Biology, 58(2):269–284, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801208.

Blum:1977:GFD

[Blu77] J. J. Blum. On the geometry of four-dimensions and the relationship between metabolism and body mass. Journal of Theoretical Biology, 64(3):599–601, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902922.

Brenner:1973:SCC

[BM73] Stephen L. Brenner and Donald A. McQuarrie. A self-consistent calculation of the free energy and electrostatic potential for a cylindrical polyion. Journal of Theoretical Biology, 39(2):343– 361, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901045.

Banks:1978:SCG

[BM78] H. T. Banks and J. M. Mahaffy. Stability of cyclic gene models for systems involving repression. Journal of Theoretical Biology, 74(2):323–334, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900796.

Britton:1979:TWC

[BM79] N. F. Britton and J. D. Murray. Threshold, wave and cell-cell avalanche behaviour in a class of substrate inhibition oscillators. Journal of Theoretical Biology, 77(3):317–332, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990359X.

Branchey:1977:PPS

[BMS77] Marl Branchey, Morris Meisner, and George M. Simpson. A paradox in the prognosis of schizophrenia. Explanation by a mathematical model. Journal of Theoretical Biology, 66(2):267– 280, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901710.

Babloyantz:1972:CIM

[BN72] Agnessa Babloyantz and G. Nicolis. Chemical instabilities and multiple steady state transitions in Monod-Jacob type models. Journal of Theoretical Biology, 34(1):185–192, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900628.

Barricelli:1971:VGTa

[BNT71] Nils Aall Barricelli, Louis Nelson, and Robert Toombs. Virus-genetic theory testing by data processing machines: I. Basic procedures and models tested. Journal of Theoretical Biology, 32(3):559–580, September 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901561.

Baghurst:1978:CEC

[BNW78] P. A. Baghurst, L. W. Nichol, and D. J. Winzor. Conditions for the existence of critical points in scatchard plots of binding results. Journal of Theoretical Biology, 74(4):523–534, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902394.

Boddington:1978:AMS

[Bod78] M. J. Boddington. An absolute metabolic scope for activity. Journal of Theoretical Biology, 75(4):443–449, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903557.

Bohren:1977:CDO

[Boh77] Craig F. Bohren. Circular dichroism and optical rotatory dispersion spectra of arbitrarily shaped optically active particles. Journal of Theoretical Biology, 65(4):755–767, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900212.

Boling:1973:TSS

[Bol73] Robert H. Boling. Toward state-space models for biological populations. Journal of Theoretical Biology, 40(3):485–506, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900064.

Boncinelli:1978:MPT

[Bon78a] E. Boncinelli. A model of post-transcriptional control in eukaryotic cells. Journal of Theoretical Biology, 72(1):75–79, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900188.

Bont:1978:DMV

[Bon78b] W. S. Bont. The diameters of membrane vesicles fit in geometric series. Journal of Theoretical Biology, 74(3):361–375, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902199.

Borgmann:1974:TDB

[Bor74] Uwe Borgmann. A theoretical description of biological rates. Journal of Theoretical Biology, 45(1):171–182, May 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900496.

Boyd:1973:ANT

[Boy73] Helen Boyd. ATP as a nerve transmitter substance. Journal of Theoretical Biology, 42(1):49–52, IN3, 53, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901471.

Boyd:1979:CNH

[Boy79a] C. A. R. Boyd. Chemical neurotransmission: an hypothesis concerning the evolution of neurotransmitter substances. Journal of Theoretical Biology, 76(4):415–417, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900109.

Boyer:1979:TMC

[Boy79b] John F. Boyer. The terminology of Markov chains. Journal of Theoretical Biology, 78(3):443–447, June 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903436.

Bradford:1970:NAP

[BP70a] E. Bradford and J. R. Philip. Note on asocial populations dispersing in two dimensions. Journal of Theoretical Biology, 29(1):27–33, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901165.

Bradford:1970:SSD

[BP70b] E. Bradford and J. R. Philip. Stability of steady distributions of asocial populations dispersing in one dimension. Journal of Theoretical Biology, 29(1):13–26, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901153.

Borst-Pauwels:1973:TSS

[BP73] G. W. F. H. Borst-Pauwels. Two site-single carrier transport kinetics. Journal of Theoretical Biology, 40(1):19–31, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901628.

Borst-Pauwels:1974:MST

[BP74] G. W. F. H. Borst-Pauwels. Multi-site two-carrier transport comparison with enzyme kinetics. Journal of Theoretical Biology, 48(1):183–195, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901891.

Borst-Pauwels:1976:DCA

[BP76] G. W. F. H. Borst-Pauwels. Discrimination criteria for apparent two-site transport models. Journal of Theoretical Biology, 56(1):191–204, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800525.

Bard:1979:CED

[BP79a] Jonathan B. L. Bard and Marshall E. Presser. The construction and evolution of DNA sequences having the same doublet ratios as those of non-satellite, mammalian DNA. Journal of Theoretical Biology, 76(2):181–204, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903692.

Broda:1979:DRP

[BP79b] E. Broda and G. A. Peschek. Did respiration or photosynthesis come first? Journal of Theoretical Biology, 81(2):201–212, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901607.

Bruskov:1979:MMN

[BP79c] V. I. Bruskov and V. I. Poltev. On molecular mechanisms of nucleic acid synthesis. Fidelity aspects: 2. Contribution of protein–nucleotide recognition. Journal of Theoretical Biology, 78(1):29–32, IN1, 33–41, May 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903230.

Benassi:1974:PCT

[BPF74] M. Benassi, R. Paoluzi, and A. Ferracin. Polypeptide chain termination signal in prokaryotes: Further considerations about protection against the appearance of translational restarts out of frame. Journal of Theoretical Biology, 43(2):385–387, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937480069X.

Bhargava:1970:ATR

[BPP70] P. M. Bhargava, T. Pallaiah, and E. Premkumar. Aminoacyltransfer RNA synthetase recognition code-words in yeast transfer RNA’s: A proposal. Journal of Theoretical Biology, 29(3):447–469, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901086.

Bezem:1975:CSE

[BR75] J. J. Bezem and Chr. P. Raven. Computer simulation of early embryonic development. Journal of Theoretical Biology, 54(1): 47–61, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800543.

Bass:1979:HSF

[BR79] L. Bass and P. Robinson. How small is the functional variability of liver sinusoids? Journal of Theoretical Biology, 81 (4):761–769, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902807.

Brannen:1970:RDW

[Bra70] J. P. Brannen. On the role of DNA in wet heat sterilization of micro-organisms. Journal of Theoretical Biology, 27(3):425– 432, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800078.

Brannen:1971:RWA

[Bra71] J. P. Brannen. Role of water activity in the dry heat sterilization of micro-organisms. Journal of Theoretical Biology, 32(2):331– 334, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937190169X.

Braitenberg:1974:TCC

[Bra74] V. Braitenberg. Thoughts on the cerebral cortex. Journal of Theoretical Biology, 46(2):421–447, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900071.

Brannen:1975:PMD

[Bra75] J. P. Brannen. A population model for dermestid beetle survival under starvation and cannibalism. Journal of Theoretical Biology, 49(1):179–189, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800269.

Brauer:1977:PSS

[Bra77] Fred Brauer. Periodic solutions of some ecological models. Journal of Theoretical Biology, 69(1):143–152, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903927.

Brackenbury:1979:AVO

[Bra79] J. H. Brackenbury. Aeroacoustics of the vocal organ of birds. Journal of Theoretical Biology, 81(2):341–349, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901711.

Bass:1978:HEF

[BRB78] L. Bass, P. Robinson, and A. J. Bracken. Hepatic elimination of flowing substrates: The distributed model. Journal of Theoretical Biology, 72(1):161–184, May 8, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900231.

Breen:1975:CQM

[Bre75a] Donna L. Breen. Coenzyme q: A molecular orbital study. Journal of Theoretical Biology, 53(1):101–113, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901058.

Bremer:1975:PAR

[Bre75b] Hans Bremer. Parameters affecting the rate of synthesis of ribosomes and RNA polymerase in bacteria. Journal of Theoretical Biology, 53(1):115–124, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590106X.

Bremermann:1979:TSC

[Bre79] Hans J. Bremermann. Theory of spontaneous cell fusion. Sexuality in cell populations as an evolutionary stable strategy. Applications to immunology and cancer. Journal of Theoretical Biology, 76(3):311–334, February 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903151.

Brindley:1974:PDK

[Bri74] G. S. Brindley. The potentialities of different kinds of modifiable synapses. Journal of Theoretical Biology, 43(2):393–396, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800718.

Brill:1978:DPI

[Bri78] Michael H. Brill. A device performing illuminant-invariant assessment of chromatic relations. Journal of Theoretical Biology, 71(3):473–478, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901753.

Brill:1979:FFI

[Bri79] Michael H. Brill. Further features of the illuminant-invariant trichromatic photosensor. Journal of Theoretical Biology, 78(2):305–308, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902716.

Brown:1974:YFA

[Bro74] Olen R. Brown. Yeast fatty acid synthetase: a proposed binding site for coenzyme A with consequences for chain initiation and chain termination. Journal of Theoretical Biology, 47(1):137–144, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901040.

Brodsky:1975:PSR

[Bro75] W. Ya. Brodsky. Protein synthesis rhythm. Journal of Theoretical Biology, 55(1):167–200, November 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801159.

Broda:1977:FWE

[Bro77] E. Broda. Free will and evolution (comments on Young’s article). Journal of Theoretical Biology, 69(4):755–760, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903800.

Bell:1979:BPS

[BRS79] A. D. Bell, D. Roberts, and A. Smith. Branching patterns: the simulation of plant architecture. Journal of Theoretical Biology, 81(2):351–375, November 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901723.

Bryant:1973:HSV

[Bry73] Edwin H. Bryant. Habitat selection in a variable environment. Journal of Theoretical Biology, 41(3):421–429, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900520.

Burgess:1975:HPT

[BS75] A. W. Burgess and H. A. Scheraga. A hypothesis for the pathway of the thermally-induced unfolding of bovine pancreatic ribonuclease. Journal of Theoretical Biology, 53(2):403–420, ????

1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800129.

Boon:1976:PCO

[BS76] P. N. M. Boon and J. Strackee. A population of coupled oscillators. Journal of Theoretical Biology, 57(2):491–500, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900175.

Bhargava:1977:SPS

[BS77] S. C. Bhargava and R. P. Saxena. Stable periodic solutions of the reactive–diffusive Volterra system of equations. Journal of Theoretical Biology, 67(3):399–406, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900455.

Board:1978:ECM

[BS78] P. G. Board and J. E. Smith. Energy cost of the modified γ glutamyl cycle. Journal of Theoretical Biology, 75(1):139–140, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902072.

Bandekar:1979:CIN

[BS79] Jagdeesh Bandekar and B. K. Sathyanarayana. Copper(II)- Nucleic acid interactions — a conformational study. Journal of Theoretical Biology, 79(4):443–454, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902364.

Black:1975:TAA

[BSH75] John A. Black, Peter Stenzel, and Richard N. Harkins. The terminal amino acids of protein sequences and protein maturation. Journal of Theoretical Biology, 50(1):161–166, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900302.

Baranowitz:1977:UGC

[BSM77] S. A. Baranowitz, S. N. Salthe, and P. F. A. Maderson. The use of the Gompertz curve in an analysis of the dynamics of lizard tail regeneration. Journal of Theoretical Biology, 65(2):267–279, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903253.

Brockwell:1970:DGP

[BT70] P. J. Brockwell and E. Trucco. On the decomposition by generations of the PLM-function. Journal of Theoretical Biology, 26(1):149–179, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519370800376.

Barricelli:1971:VGTb

[BT71a] Nils Aall Barricelli and Robert Toombs. Virus-genetic theory testing by data processing machines: II. Fit of classical genetic T4 data. Journal of Theoretical Biology, 32(3):581–601, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901573.

Barricelli:1971:VGTc

[BT71b] Nils Aall Barricelli and Robert Toombs. Virus-genetic theory testing by data processing machines: III. The interpretation of radiation genetic and partial phage experiments in T4. General remarks on the theory-testing results. Journal of Theoretical Biology, 32(3):603–623, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901585.

Bunow:1978:CRMb

[Bun78a] Barry Bunow. Chemical reactions and membranes: A macroscopic basis for active transport II. Non-linear aspects. Journal of Theoretical Biology, 75(1):79–96, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902035.

Bunow:1978:CRMa

[Bun78b] Barry Bunow. Chemical reactions and membranes: A macroscopic basis for facilitated transport, chemiosmosis and active transport part i: Linear analysis. Journal of Theoretical Biology, 75(1):51–78, November 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902023.

Burnstock:1976:PR

[Bur76] G. Burnstock. Purinergic receptors. Journal of Theoretical Biology, 62(2):491–503, October 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901338.

Buysman:1973:ICC

[Buy73] John R. Buysman. Ion concentration and charge adjacent to electrically charged cell membrane. Journal of Theoretical Biology, 38(1):51–59, January 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902245.

Banks:1970:RRA

[BV70] Barbara E. C. Banks and C. A. Vernon. Reassessment of the role of ATP in vivo. Journal of Theoretical Biology, 29(2):301– 326, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937090024X.

Barclay:1975:TLE

[BV75] H. Barclay and P. Van Den Driessche. Time lags in ecological systems. Journal of Theoretical Biology, 51(2):347–356, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590065X.

Bardsley:1976:PKI

[BW76] William G. Bardsley and Roy D. Waight. Primary kinetic isotope effects in enzyme-catalysed reactions. A theoretical treatment and new experimental methods of study. Journal of Theoretical Biology, 63(2):325–345, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900370.

Beckman:1977:RDN

[BW77] R. J. Beckman and M. S. Waterman. On the random distribution of nucleoli in metabolic cells. Journal of Theoretical Biology, 69(3):561–562, December 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901564.

Bardsley:1978:DPN

[BW78a] William G. Bardsley and Roy D. Waight. The determination of positive and negative co-operativity with allosteric enzymes and the interpretation of sigmoid curves and nonlinear double reciprocal plots for the MWC and KNF models. Journal of Theoretical Biology, 70(2):135–156, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903430.

Bardsley:1978:FHB

[BW78b] William G. Bardsley and Roy D. Waight. Factorability of the Hessian of the binding polynomial. The central issues concerning statistical ratios between binding constants, hill plot slope and positive and negative co-operativity. Journal of Theoretical Biology, 72(2):321–372, May 22, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900966.

Bardsley:1978:CTD

[BW78c] William G. Bardsley and Jeffries Wyman. Concerning the thermodynamic definition and graphical manifestations of positive and negative co-operativity. Journal of Theoretical Biology, 72(2):373–376, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378900978.

Barnhill:1974:MAD

[BY74] C. W. Barnhill and K. Lemone Yielding. A model for antibody diversity based on non-conservative DNA synthesis errors. Journal of Theoretical Biology, 43(1):197–209, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800536.

Bauman:1975:MSC

[BY75] A. J. Bauman and Henry Yokoyama. Magnolia seed carotenoid pigments: Typical evolutionarily-static relicts? Journal of Theoretical Biology, 53(2):277–284, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580004X.

Byers:1978:BRI

[Bye78] Larry D. Byers. Binding of reactive intermediate analogs to enzymes. Journal of Theoretical Biology, 74(4):501–512, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902370.

Chizmadjev:1977:ITA

[CA77] Yu. A. Chizmadjev and S. Kh. Aityan. Ion transport across sodium channels in biological membranes. Journal of Theoretical Biology, 64(3):429–453, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902818.

Calhoon:1979:DAA

[CA79] Robert E. Calhoon and Sheldon Aaronson. A discriminant analysis of amino acid frequency in collagen-like proteins. Journal of Theoretical Biology, 78(2):225–239, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902662.

Chapman:1979:CSR

[CAA79] Robert W. Chapman, John C. Avise, and Marjorie A. Asmussen. Character space restrictions and boundary conditions in the evolution of quantitative multistate characters.

Journal of Theoretical Biology, 80(1):51–64, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901796.

Calkins:1973:PMR

[Cal73] John Calkins. The T–N–PR model of radiation response. Journal of Theoretical Biology, 39(3):609–622, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900726.

Calladine:1976:DRC

[Cal76a] C. R. Calladine. Design requirements for the construction of bacterial flagella. Journal of Theoretical Biology, 57(2):469– 489, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900163.

Calloway:1976:BTT

[Cal76b] Nathaniel O. Calloway. Body temperature: Thermodynamics of homeothermism. Journal of Theoretical Biology, 57(2):331– 344, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900059.

Calder:1979:TDO

[Cal79] William A. Calder. On the temperature-dependency of optimal nectar concentrations for birds. Journal of Theoretical Biology, 78(2):185–196, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902637.

Cameron:1977:MFC

[Cam77] W. Hugh Cameron. A model framework for computer simulation of overall renal function. Journal of Theoretical Biology, 66(3):551–572, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903022.

Canham:1970:MEB

[Can70] P. B. Canham. The minimum energy of bending as a possible explanation of the biconcave shape of the human red blood cell. Journal of Theoretical Biology, 26(1):61–76, IN7–IN8, 77– 81, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800327.

Carlin:1978:PSM

[Car78] Richard K. Carlin. The poly(a) segment of mRNA: (1) evolution and function and (2) the evolution of viruses. Journal of Theoretical Biology, 71(3):323–338, April 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901637.

Castro:1971:THR

[Cas71] Charles E. Castro. Theory of hemeprotein reactivity. Journal of Theoretical Biology, 33(3):475–490, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900920.

Caswell:1972:SST

[Cas72] Hal Caswell. A simulation study of a time lag population model. Journal of Theoretical Biology, 34(3):419–439, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901336.

Cassada:1974:SRB

[Cas74] Randall C. Cassada. The size relation between finished proteins of mixed sizes and their ribosome-bound nascent polypeptides. Journal of Theoretical Biology, 46(2):581–584, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900186.

Castrodeza:1978:ECF

[Cas78] Carlos Castrodeza. Evolution, complexity, and fitness. Journal of Theoretical Biology, 71(3):469–471, April 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901741.

Comins:1974:PPM

[CB74] Hugh N. Comins and David W. E. Blatt. Prey-predator models in spatially heterogeneous environments. Journal of Theoretical Biology, 48(1):75–83, November 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901805.

Childs:1975:ARP

[CB75a] Robert E. Childs and William G. Bardsley. Allosteric and related phenomena: An analysis of sigmoid and non-hyperbolic functions. Journal of Theoretical Biology, 50(1):45–58, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900235.

Childs:1975:TDI

[CB75b] Robert E. Childs and William G. Bardsley. Time-dependent inhibition of enzymes by active-site-directed reagents a theoretical treatment of the kinetics of affinity labelling. Journal of Theoretical Biology, 53(2):381–394, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800105.

Cornish-Bowden:1975:PSN

[CB75c] Athel Cornish-Bowden. The physiological significance of negative co-operativity. Journal of Theoretical Biology, 51(1):233– 235, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901496.

Childs:1976:ANL

[CB76] Robert E. Childs and William G. Bardsley. An analysis of non-linear Eadie–Hofstee–Scatchard representations of ligandbinding and initial-rate data for allosteric and other complex enzyme mechanisms. Journal of Theoretical Biology, 63(1):1– 18, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900801.

Cornish-Bowden:1977:APS

[CB77] Athel Cornish-Bowden. Assessment of protein sequence identity from amino acid composition data. Journal of Theoretical Biology, 65(4):735–742, April 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900194.

Cornish-Bowden:1978:IDI

[CB78] Athel Cornish-Bowden. Interpretation of the difference index as a guide to protein sequence identity. Journal of Theoretical Biology, 74(2):155–161, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900681.

Cornish-Bowden:1979:HRDb

[CB79a] A. Cornish-Bowden. How reality do amino acid composition comparisons predict sequence similarities between proteins?: J. theor. Biol. (1979) 76, 369–386. Journal of Theoretical Biology, 80(2):i, September 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902157.

Cornish-Bowden:1979:HRDa

[CB79b] Athel Cornish-Bowden. How reliably do amino acid composition comparisons predict sequence similarities between proteins? Journal of Theoretical Biology, 76(4):369–386, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900079.

Capo:1974:P

[CBBD74] C. Capo, P. Bongrand, A. M. Benoliel, and R. Depieds. Phagocytosis. Journal of Theoretical Biology, 47(1):177–188, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901088.

Cornish-Bowden:1978:EDF

[CBPT78] Athel Cornish-Bowden, William R. Porter, and William F. Trager. Evaluation of distribution-free confidence limits for enzyme kinetic parameters. Journal of Theoretical Biology, 74(2):163–175, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900693.

Coleman:1975:MRB

[CBS75] G. Coleman, S. Brown, and D. A. Stormonm. A model for the regulation of bacterial extracellular enzyme and toxin biosynthesis. Journal of Theoretical Biology, 52(1):143–148, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900454.

Coe:1970:STB

[CC70] Elmon L. Coe and Mary H. Coe. Solute transfer between intramolecular polar and non-polar phases as a possible rate-limiting step in chymotrypsin deacylation. Journal of Theoretical Biology, 29(3):411–425, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901062.

Caserta:1973:PTM

[CC73a] Giuseppe Caserta and Tommaso Cervigni. A piezoelectric transducer model for phosphorylation in photosynthetic membranes. Journal of Theoretical Biology, 41(1):127–142, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901938.

Cedergren:1973:DMN

[CC73b] R. J. Cedergren and J. R. Cordeau. The distribution of modified nucleosides in transfer RNAs. Journal of Theoretical Biology, 39(2):477–486, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901124.

Charlesworth:1975:TGBa

[CC75a] D. Charlesworth and B. Charlesworth. Theoretical genetics of Batesian mimicry I. Single-locus models. Journal of Theoretical Biology, 55(2):283–303, December 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800816.

Charlesworth:1975:TGBb

[CC75b] D. Charlesworth and B. Charlesworth. Theoretical genetics of Batesian mimicry II. Evolution of supergenes. Journal of Theoretical Biology, 55(2):305–324, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800828.

Charlesworth:1975:TGBc

[CC75c] D. Charlesworth and B. Charlesworth. Theoretical genetics of Batesian mimicry III. Evolution of dominance. Journal of Theoretical Biology, 55(2):325–337, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580083X.

Chernavskii:1975:STA

[CC75d] D. S. Chernavskii and N. M. Chernavskaya. Some theoretical aspects of the problem of life origin. Journal of Theoretical Biology, 50(1):13–23, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590020X.

Cedergren:1972:PR

[CCR72] R. J. Cedergren, Jeff R. Cordeau, and Pierre Robillard. On the phylogeny of t-RNA’s. Journal of Theoretical Biology, 37(2):209–220, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900173.

Chang:1975:RDB

[CCY^{+}75] David B. Chang, Robert L. Cooper, Allan C. Young, Carroll J. Martin, and Betsy Ancker-Johnson. Restricted diffusion in biophysical systems: Theory. Journal of Theoretical Biology, 50(2): 285–308, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937590082X.

Critchlow:1972:UTS

[CD72] J. E. Critchlow and H. B. Dunford. The use of transition state acid dissociation constants in pH-dependent enzyme kinetics. Journal of Theoretical Biology, 37(2):307–320, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900252.

Cantraine:1975:TDT

[CD75] F. R. L. Cantraine and B. Dewandre. Thyroglobulin diffusion in the thyroid follicular lumen. Journal of Theoretical Biology, 54(2):229–242, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801287.

Chatteriee:1977:PFR

[CDS77] A. N. Chatteriee, R. J. Doyle, and U. N. Streips. A proposed functional role for bacterial N-acetylmuramyl-l-alanine amidases. Journal of Theoretical Biology, 68(3):385–390, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900674.

Collins:1977:MSR

[CE77] John Collins and Jan Engberg. A model for switching on ribosomal RNA synthesis by creating a palindromic DNA sequence in the promoter region of the ribosomal RNA cistron: the “structon”. Journal of Theoretical Biology, 66(3):573–582, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903034.

Catalano:1978:MME

[CE78] Giorgio Catalano and J. Chris Eilbeck. A mathematical model for embryonic cell division based on a surface “cleavage field”. Journal of Theoretical Biology, 75(1):123–137, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902060.

Curry:1978:FSD

[CFS78] Guy L. Curry, Richard M. Feldman, and Peter J. H. Sharpe. Foundations of stochastic development. Journal of Theoretical Biology, 74(3):397–410, October 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902229.

Cavallo:1975:RAI

[CFV75] G. Cavallo, G. Form, and L. Varesio. RNA in amplification of the immunological response. Journal of Theoretical Biology, 51(2):383–392, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900673.

Clay:1973:DMF

[CG73] John R. Clay and Narendra S. Goel. Diffusion models for firing of a neuron with varying threshold. Journal of Theoretical Biology, 39(3):633–644, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390074X.

Curran:1975:SMF

[CG75] Thomas C. Curran and Harvey J. Gold. A stochastic model for facilitated transport of oxygen through tissue slices. Journal of Theoretical Biology, 50(2):397–411, April 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900892.

Costantino:1977:ECF

[CGM77] R. F. Costantino, L. R. Ginzburg, and A. M. Moffa. An experimental check of fitness entropy vs. selective delay. Journal of Theoretical Biology, 68(2):317–320, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901680.

Clouston:1979:AMB

[CGM79] J. G. Clouston, C. P. Gilbert, and P. Miskelly. An autocatalytic model for bacterial spore germination. Journal of Theoretical Biology, 76(4):387–402, February 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900080.

Coakley:1972:PMO

[CH72] C. J. Coakley and M. E. J. Holwill. Propulsion of microorganisms by three-dimensional flagellar waves. Journal of Theoretical Biology, 35(3):525–542, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290149X.

Carrano:1973:FCA

[CH73] Anthony V. Carrano and John A. Heddle. The fate of chromosome aberrations. Journal of Theoretical Biology, 38(2):289– 304, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901768.

Comins:1976:PMP

[CH76] H. N. Comins and M. P. Hassell. Predation in multi-prey communities. Journal of Theoretical Biology, 62(1):93–114, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900539.

Chapman:1972:MVS

[Cha72] J. B. Chapman. A macroscopic view of sodium exchange diffusion. Journal of Theoretical Biology, 37(1):201–204, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901282.

Charnov:1977:ETG

[Cha77] Eric L. Charnov. An elementary treatment of the genetical theory of kin-selection. Journal of Theoretical Biology, 66(3):541– 550, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903010.

Charlesworth:1978:PGA

[Cha78a] Brian Charlesworth. The population genetics of anisogamy. Journal of Theoretical Biology, 73(2):347–357, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901959.

Charlesworth:1978:SME

[Cha78b] Brian Charlesworth. Some models of the evolution of altruistic behaviour between siblings. Journal of Theoretical Biology, 72(2):297–319, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378900954.

Charnov:1978:EEB

[Cha78c] Eric L. Charnov. Evolution of eusocial behavior: Offspring choice or parental parasitism? Journal of Theoretical Biology, 75(4):451–465, December 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903569.

Chay:1979:POT

[Cha79] Teresa Ree Chay. pH oscillations in transport processt. Journal of Theoretical Biology, 80(1):83–99, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901814.

Chen:1973:CCA

[Che73a] Augustine C. Chen. Comment on comparative approach. Journal of Theoretical Biology, 42(3):587–589, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902488.

Chevillotte:1973:RBR

[Che73b] Paule Chevillotte. Relation between the reaction cytochrome oxidase-oxygen and oxygen uptake in cells in vivo: The role of diffusion. Journal of Theoretical Biology, 39(2):277–295, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900982.

Chen:1975:PCG

[Che75] Yi-Der Chen. Phenomenological coefficients and general proof of reciprocal relation for a certain class of sliding filament models of muscle contraction. Journal of Theoretical Biology, 51(2): 419–428, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900715.

Cheng:1976:SSB

[Che76] C. C. Cheng. Structural similarity between febrifugine and chloroquine. Journal of Theoretical Biology, 59(2):497–501, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901867.

Chen:1979:AMA

[Che79] Yi-Der Chen. On the admittance of membranes associated with channel conduction. Application to channels at non-equilibrium steady state. Journal of Theoretical Biology, 81(4):633–644, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902741.

Chen:1976:PPD

[CHH^{+}76] J. Chen, C. W. Huang, L. Hinman, M. P. Gordon, and D. A. Deranleau. Photomonomerization of pyrimidine dimers by indoles and proteins. Journal of Theoretical Biology, 62(1):53–67, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900515.

Chipchase:1976:RSA

[Chi76] Margaret Chipchase. Recombination by strand assimilation and strand crossover. Journal of Theoretical Biology, 57(2):249– 279, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900011.

Chou:1972:CAM

[Cho72] Ting-Chao Chou. Combinatorial analysis of multiple substratemultiple product enzyme reactions. Journal of Theoretical Biology, 35(2):285–297, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900409.

Chou:1976:DPM

[Cho76] Ting-Chao Chou. Derivation and properties of Michaelis– Menten type and Hill type equations for reference ligands. Journal of Theoretical Biology, 59(2):253–254, IN1, 255–276, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901697.

Chou:1977:DAL

[Cho77] Ting-Chao Chou. On the determination of availability of ligand binding sites in steady-state systems. Journal of Theoretical Biology, 65(2):345–356, March 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903290.

Christensen:1976:TSD

[Chr76] Halvor N. Christensen. Towards a sharper definition of energetic coupling through integration of membrane transport into bioenergetics. Journal of Theoretical Biology, 57(2):419–431, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900138.

Chuknyiski:1974:SENa

[Chu74a] P. Chuknyiski. Some equilibrium and non-equilibrium properties of the allosteric interactions in enzymes I. Ligand binding and saturation equations. Journal of Theoretical Biology, 44(1):131– 140, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800330.

Chuknyiski:1974:SENb

[Chu74b] P. Chuknyiski. Some equilibrium and nonequilibrium properties of the allosteric interactions in enzymes: II. Allosteric effect in the absence of a rapid equilibrium between substrate and enzyme. Journal of Theoretical Biology, 48(1):173–181, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490188X.

Chuknyiski:1978:SEN

[Chu78] P. Chuknyiski. Some equilibrium and non-equilibrium properties of the allosteric interactions in enzymes III. Dynamics of the amino acid residues and ligand binding process. Journal of Theoretical Biology, 74(3):389–395, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902217.

Cilento:1975:DSR

[Cil75a] G. Cilento. In defence of a special role for the iodo substituents in thyroid hormones. Journal of Theoretical Biology, 52(1):255– 257, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900570.

Cilento:1975:DIB

[Cil75b] Giuseppe Cilento. Dioxetanes as intermediates in biological processes. Journal of Theoretical Biology, 55(2):471–479, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800944.

Ciures:1975:VIT

[Ciu75] Alexander Ciures. The visual inner transform of the outer optical medium and the visual resolution. Journal of Theoretical Biology, 55(1):145–151, November 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801135.

Ciures:1977:MBV

[Ciu77] Alexandru Ciures. The Mach bands: The visualization of Gibbs phenomenon in the space domain. Journal of Theoretical Biology, 66(1):195–196, IN5, 197, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903204.

Civan:1970:PBW

[Civ70] Mortimer M. Civan. Path of bulk water movement through the urinary bladder of the toad. Journal of Theoretical Biology, 27(3):387–391, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519370800042.

Colding-Jorgensen:1976:CME

[CJ76a] Morten Colding-Jorgensen. Can membrane excitation be described without a membrane capacity? Journal of Theoretical Biology, 57(2):373–383, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900084.

Colding-Jorgensen:1976:DAE

[CJ76b] Morten Colding-Jørgensen. A description of adaptation in excitable membranes. Journal of Theoretical Biology, 63(1):61– 87, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900849.

Calow:1977:OSM

[CJ77] P. Calow and J. B. Jennings. Optimal strategies for the metabolism of reserve materials in microbes and metazoa. Journal of Theoretical Biology, 65(3):601–603, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902193.

Crowther:1971:ASC

[CK71] R. A. Crowther and A. Klug. ART and science or conditions for three-dimensional reconstruction from electron microscope images. Journal of Theoretical Biology, 32(1):199–203, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901470.

Crippen:1977:DSF

[CK77] G. M. Crippen and I. D. Kuntz. Directional structural features of globular proteins. Journal of Theoretical Biology, 66(1):47– 61, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903113.

Chapman:1979:KMD

[CKJ79] J. Brian Chapman, J. Mailen Kootsey, and Edward A. Johnson. A kinetic model for determining the consequences of electrogenic active transport in cardiac muscle. Journal of Theoretical Biology, 80(3):405–424, October 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901012.

Chatfield:1970:ASB

[CL70] Christopher Chatfield and Robert E. Lemon. Analysing sequences of behavioural events. Journal of Theoretical Biology, 29(3):427–445, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901074.

Clarke:1972:ILC

[Cla72] C. H. Clarke. Intragenic localization of chain-terminating triplets arising by transitions and transversions. Journal of Theoretical Biology, 34(3):411–418, March 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901324.

Clay:1976:SAG

[Cla76] John R. Clay. A stochastic analysis of the graded excitatory response of nerve membrane. Journal of Theoretical Biology, 59(1):141–158, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800288.

Clarke:1977:FIC

[Cla77a] John R. Clarke. Fluidity-induced changes in diffusion through membranes: a predictive model. Journal of Theoretical Biology, 68(2):293–298, September 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901667.

Clay:1977:MCS

[Cla77b] John R. Clay. Monte Carlo simulation of membrane noise: An analysis of fluctuations in graded excitation of nerve membrane. Journal of Theoretical Biology, 64(4):671–680, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902661.

Clifford:1977:EAN

[Cli77] Peter Clifford. The effect of arm number on chromosome exchange distortion. Journal of Theoretical Biology, 64(4): 723–728, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902703.

Chidichimo:1977:CCA

[CLR77] G. Chidichimo, F. Lelj, and N. Russo. CNDO/2 conformational analysis of acetylselenocholine. Journal of Theoretical Biology, 66(4):811–814, June 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790248X.

Ciures:1970:TBI

[CM70] A. Ciures and D. Mărgineanu. Thermodynamics in biology: An intruder? Journal of Theoretical Biology, 28(1):147–150, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900706.

Cramer:1972:ICP

[CM72] N. F. Cramer and R. M. May. Interspecific competition, predation and species diversity: A comment. Journal of Theoretical Biology, 34(2):289–293, February 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901622.

Cohen:1975:CFF

[CM75] Joel E. Cohen and Harry K. MacWilliams. The control of foot formation in transplantation experiments with hydra viridis. Journal of Theoretical Biology, 50(1):87–105, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900259.

Cragg:1978:MOR

[CM78] Brian Cragg and Elspeth McLachlan. A mechanism for the observed recovery from ineffectiveness of synapses in the central nervous system. Journal of Theoretical Biology, 71(3):433– 440, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901704.

Cianchi:1977:MIS

[CMSC77a] L. Cianchi, M. Mancini, G. Spina, and R. Cappelletti. On the Mössbauer isomer shift in oxyhemoglobin. Journal of Theoretical Biology, 67(4):765–768, August 22, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902600.

Cianchi:1977:PDB

[CMSC77b] L. Cianchi, M. Mancini, G. Spina, and R. Cappelletti. On the possibility of deciding between Pauling’s and Griffith’s models of oxyhemoglobin by means of Mössbauer spectroscopy. Journal of Theoretical Biology, 67(4):757–763, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902594.

Conjeaud:1976:PEF

[CMV76] Hélène Conjeaud and M. Michel-Villaz. Photoinduced electric field effect on the optical properties of photosynthetic membranes. Journal of Theoretical Biology, 62(1):1–16, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900497.

Calvert:1979:BEI

[CNS79] P. D. Calvert, L. W. Nichol, and W. H. Sawyer. Binding equations for interacting systems comprising multivalent acceptor and bivalent ligand: Application to antigen-antibody systems. Journal of Theoretical Biology, 80(2):233–247, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990208X.

Cannings:1975:ASR

[CO75] C. Cannings and Luis M. Cruz Orive. On the adjustment of the sex ratio and the gregarious behaviour of animal populations. Journal of Theoretical Biology, 55(1):115–136, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801111.

Cohen:1970:MMA

[Coh70] Stanley Cohen. A model for the mechanism of antibody induction and tolerance, with specific attention to the affinity characteristics of antibodies produced during the immune response. Journal of Theoretical Biology, 27(1):19–29, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090127X.

Cohen:1971:MFY

[Coh71] Dan Cohen. Maximizing final yield when growth is limited by time or by limiting resources. Journal of Theoretical Biology, 33(2):299–307, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900683.

Cohen:1976:ABP

[Coh76] Robert J. Cohen. The avoidance behavior of phycomyces blakesleeanus: Is mediation by natural convection significant? Journal of Theoretical Biology, 58(1):207–217, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690148X.

Cohen:1977:CAC

[Coh77] Marcus S. Cohen. The cyclic AMP control system in the development of Dictyostelium discoideum: I. Cellular dynamics. Journal of Theoretical Biology, 69(1):57–85, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903885.

Cohen:1978:CAC

[Coh78] Marcus S. Cohen. The cyclic AMP control system in the development of dictyostelium discoideum II. An allosteric model. Journal of Theoretical Biology, 72(2):231–255, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900917.

Comorosan:1975:MPB

[Com75] Sorin Comorosan. The measurement process in biological systems: A new phenomenology. Journal of Theoretical Biology, 51 (1):35–49, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937590137X.

Comins:1977:DIR

[Com77a] Hugh N. Comins. The development of insecticide resistance in the presence of migration. Journal of Theoretical Biology, 64(1):177–197, January 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901199.

Comins:1977:MPR

[Com77b] Hugh N. Comins. The management of pesticide resistance. Journal of Theoretical Biology, 65(3):399–420, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902065.

Comins:1979:AMM

[Com79] Hugh N. Comins. Analytic methods for the management of pesticide resistance. Journal of Theoretical Biology, 77(2):171– 188, March 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903059.

Cone:1971:MMH

[Con71a] Clarence D. Cone. Maintenance of mitotic homeostasis in somatic cell populations. Journal of Theoretical Biology, 30 (1):183–194, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900439.

Cone:1971:UTB

[Con71b] Clarence D. Cone. Unified theory on the basic mechanism of normal mitotic control and oncogenesis. Journal of Theoretical Biology, 30(1):151–181, January 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900427.

Conrad:1972:SFR

[Con72] Michael Conrad. Stability of foodwebs and its relation to species diversity. Journal of Theoretical Biology, 34(2):325–335, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901658.

Conrad:1974:ELC

[Con74a] Michael Conrad. Evolutionary learning circuits. Journal of Theoretical Biology, 46(1):167–188, July 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901465.

Conrad:1974:LBS

[Con74b] Michael Conrad. The limits of biological simulation. Journal of Theoretical Biology, 45(2):585–590, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901325.

Conrad:1977:PSF

[Con77] Michael Conrad. Principle of superposition-free memory. Journal of Theoretical Biology, 67(2):213–219, July 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901953.

Conrad:1979:UEP

[Con79] M. Conrad. Unstable electron pairing and the energy loan model of enzyme catalysis. Journal of Theoretical Biology, 79(2):137– 156, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902431.

Cooper:1970:MDG

[Coo70] S. Cooper. Model for the determination of growth rate in Escherichia coli. Journal of Theoretical Biology, 28(1):151–154, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900718.

Cooper:1973:APS

[Coo73] Alan Cooper. An analysis of the possible source of contractile forces in striated muscle. Journal of Theoretical Biology, 42(3):545–562, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902464.

Cooper:1974:CUC

[Coo74] Stephen Cooper. A criterion for using chloramphenicol to define different processes in the initiation of DNA synthesis in bacteria. Journal of Theoretical Biology, 46(1):117–127, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490143X.

Cook:1977:CRT

[Coo77] Norman D. Cook. The case for reverse translation. Journal of Theoretical Biology, 64(1):113–135, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901163.

Cooper:1978:HRC

[Coo78a] William E. Cooper. Home range criteria based on temporal stability of areal occupation. Journal of Theoretical Biology, 73(4):687–695, August 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901303.

Cooper:1978:HRS

[Coo78b] William E. Cooper. Home range size and population dynamic. Journal of Theoretical Biology, 75(3):327–337, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903387.

Coster:1975:HPA

[Cos75] H. G. L. Coster. Heat production and absorption during excitation in nerve. Journal of Theoretical Biology, 54(2):225– 227, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801275.

Cotter:1973:SGD

[Cot73] David A. Cotter. Spore germination in dictyostelium discoideum: I. The thermodynamics of reversible activation. Journal of Theoretical Biology, 41(1):41–51, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901884.

Courtney:1978:EMA

[Cou78]  Kenneth R. Courtney.  Extended moment analysis for

binomial parameters of transmitter release. Journal of Theoretical Biology, 73(2):285–292, July 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901911.

Coyne:1979:NIS

[Coy79] Lelia M. Coyne. Neurotransmitters as ion shuttles in regulation of neural activity. Journal of Theoretical Biology, 79(4):455–472, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902376.

Conrad:1970:EEA

[CP70]  Michael Conrad and H. H. Pattee.  Evolution exper-

iments with an artificial ecosystem. Journal of Theoretical Biology, 28(3):393–409, September 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900779.

Crane:1972:TMC

[CP72] Joe S. Crane and Herbert A. Pohl. Theoretical models of cellular dielectrophoresis. Journal of Theoretical Biology, 37(1):15– 41, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901130.

Cohen:1976:OPM

[CP76] Dan Cohen and Hanna Parnas. An optimal policy for the metabolism of storage materials in unicellular algae. Journal of Theoretical Biology, 56(1):1–18, January 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800434.

Catalan:1979:PDP

[CP79a] J. Catalán and P. Pérez. Photoinduced double proton transfer in a model hydrogen bonded base pair. Theoretical study. Journal of Theoretical Biology, 81(2):213–221, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901619.

Couzin:1979:DBC

[CP79b] D. Couzin and D. G. Papworth. The over-dispersion between cells of chromosomal aberrations. Journal of Theoretical Biology, 80(2):249–258, September 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902091.

Coste:1978:DCF

[CPC78] J. Coste, J. Peyraud, and P. Coullet. Does complexity favor the existence perpersistent ecosystems? Journal of Theoretical Biology, 73(2):359–362, July 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901960.

Cohen:1971:CES

[CR71a] Morrel H. Cohen and Anthony Robertson. Chemotaxis and the early stages of aggregation in cellular slime molds. Journal of Theoretical Biology, 31(1):119–130, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901251.

Cohen:1971:WPE

[CR71b] Morrel H. Cohen and Anthony Robertson. Wave propagation in the early stages of aggregation of cellular slime molds. Journal of Theoretical Biology, 31(1):101–118, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190124X.

Capocelli:1973:CMM

[CR73]  R. M. Capocelli and L. M. Ricciardi.  A continuous

Markovian model for neuronal activity. Journal of Theoretical Biology, 40(2):369–387, August 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901380.

Coleman:1975:CDLa

[CR75a] Bernard D. Coleman and George H. Renninger. Consequences of delayed lateral inhibition in the retina of Limulus I. Elementary theory of spatially uniform fields. Journal of Theoretical Biology, 51(2):243–265, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900582.

Coleman:1975:CDLb

[CR75b] Bernard D. Coleman and George H. Renninger. Consequences of delayed lateral inhibition in the retina of Limulus II. Theory of spatially uniform fields, assuming the four-point property. Journal of Theoretical Biology, 51(2):267–291, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900594.

Chernavskii:1978:DSM

[CR78] D. S. Chernavskii and Th. W. Ruijgrok. Dissipative structures in morphogenetic models of the Turing type. Journal of Theoretical Biology, 73(4):585–607, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901248.

Craig:1971:TRD

[Cra71] A. G. Craig. Theory of radiation dose-survival curves for the case where organism lethality is defined as loss of reproductive capacity. Journal of Theoretical Biology, 32(2):259–267, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901640.

Craig:1973:TRD

[Cra73] A. G. Craig. Theory of radiation dose-survival curves for the case where organism lethality is defined as loss of reproductive capacity: II. Application to cells and physical interpretation. Journal of Theoretical Biology, 38(1):81–99, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902269.

Craig:1977:MLM

[Cra77] A. G. Craig. Multi-lesion models for cell survival following exposure to ionising radiations. Journal of Theoretical Biology, 64(4):765–770, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902752.

Chapman:1977:HMB

[CRC77] D. C. Chapman, R. H. Rand, and J. R. Cooks. A hydrodynamical model of bordered pits in conifer tracheids. Journal of Theoretical Biology, 67(1):11–24, July 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901825.

Crippen:1974:TGP

[Cri74] Gordon M. Crippen. Topology of globular proteins. Journal of Theoretical Biology, 45(2):327–338, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901180.

Crippen:1975:TGP

[Cri75] Gordon M. Crippen. Topology of globular proteins. II. Journal of Theoretical Biology, 51(2):495–500, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900764.

Crill:1977:PLC

[Cri77] Peter A. Crill. The photosynthesis-light curve: A simple analog model. Journal of Theoretical Biology, 64(3):503–516, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902843.

Crowley:1975:NSM

[Cro75] Philip H. Crowley. Natural selection and the Michaelis constant. Journal of Theoretical Biology, 50(2):461–475, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900934.

Crowley:1979:PMC

[Cro79] Philip H. Crowley. Predator-mediated coexistence: an equilibrium interpretation. Journal of Theoretical Biology, 80 (1):129–144, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937990184X.

Coulson:1971:TRS

[CS71] A. S. Coulson and N. V. Smith. Tuberculin receptor site on sensitized human lymphocytes: A reply to Dr Fang’s criticism. Journal of Theoretical Biology, 31(3):555–556, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900294.

Claxton:1973:MPF

[CS73] J. H. Claxton and C. A. Sholl. A model of pattern formation in the primary skin follicle population of sheep. Journal of Theoretical Biology, 40(2):353–367, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901379.

Carlson:1975:CSF

[CS75a] A. D. Carlson and B. Soucek. Computer simulation of firefly flash sequences. Journal of Theoretical Biology, 55(2):353–370, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800853.

Chojnacki:1975:IIB

[CS75b] H. Chojnacki and W. A. Sokalski. Interactions of 6-thioguanine in b-DNA: Possible mechanism of its mutagenic action. Journal of Theoretical Biology, 54(2):167–174, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801214.

Cuppoletti:1975:KAA

[CS75c] John Cuppoletti and Irwin H. Segel. Kinetic analysis of active membrane transport systems: Equations for net velocity and isotope exchange. Journal of Theoretical Biology, 53(1):125–144, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901071.

Clay:1977:UTC

[CS77] John R. Clay and Michael F. Shlesingert. Unified theory of 1f and conductance noise in nerve membrane. Journal of Theoretical Biology, 66(4):763–773, June 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902442.

Cairns-Smith:1972:FPM

[CSIW72] A. G. Cairns-Smith, P. Ingram, and G. L. Walker. Formose production by minerals: Possible relevance to the origin of life. Journal of Theoretical Biology, 35(3):601–604, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901531.

Caprani:1975:SMB

[CSL75] Ole Caprani, Edda Sveinsdottir, and Niels Lassen. SHAM, a method for biexpone ial carve resolution using initial slope, height, area and moment of the experimental decay type curve. Journal of Theoretical Biology, 52(2):299–315, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590003X.

Chandra:1978:TME

[CSVR78] A. K. Chandra, B. S. Sudhindra, M. Vijayan, and T. Ramasarma. A theory on the migration of an extraneous electron across hydrogen bonds in polypeptides. Journal of Theoretical Biology, 74(1):1–9, September 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902850.

Chang:1973:QADa

[CT73a] K.-J. Chang and D. J. Triggle. Quantitative aspects of drugreceptor interactions: I. Ca^{2+} and cholinergic receptor activation in smooth muscle: A basic model for drug-receptor interactions. Journal of Theoretical Biology, 40(1):125–154, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901689.

Chang:1973:QADb

[CT73b] K.-J. Chang and D. J. Triggle. Quantitative aspects of drugreceptor interactions: II. The role of Ca_{mem}^{2+} in desensitization and spasmolytic activity. Journal of Theoretical Biology, 40(1): 155–172, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901690.

Cope:1979:FRN

[CT79] Davis K. Cope and Henry C. Tuckwell. Firing rates of neurons with random excitation and inhibition. Journal of Theoretical Biology, 80(1):1–14, September 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901747.

Cummings:1975:BMC

[Cum75] F. W. Cummings. A biochemical model of the circadian clock. Journal of Theoretical Biology, 55(2):455–470, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800932.

Cuthbert:1972:DSP

[Cut72] A. W. Cuthbert. A double (series) pump model for transporting epithelia. Journal of Theoretical Biology, 36(3):555–568, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900082.

Cutler:1978:DCC

[Cut78] D. J. Cutler. On the definition of the compartment concept in pharmacokinetics. Journal of Theoretical Biology, 73(2):329– 345, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901947.

Chiarugi:1976:SHS

[CV76] Vincenzo P. Chiarugi and Simonetta Vannucchi. Surface heparan sulphate as a control element in eukariotic cells: A working model. Journal of Theoretical Biology, 61(2):459–475, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900308.

Chan:1974:EIB

[CW74] D. Chan and L. R. White. The electrostatic interaction between surfaces bearing amphoteric ionizable groups across a 1:1 electrolyte. Journal of Theoretical Biology, 48(1):253–256, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901957.

Cooke:1976:CWM

[CZ76] J. Cooke and E. C. Zeeman. A clock and wavefront model for control of the number of repeated structures during animal morphogenesis. Journal of Theoretical Biology, 58(2):455–476, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801312.

Czaplewski:1973:MEP

[Cza73] Raymond L. Czaplewski. A methodology for evaluation of parent-mutant competition using a generalized non-linear ecosystem model. Journal of Theoretical Biology, 40(3):429–439, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900027.

Czerlinski:1971:CRCa

[Cze71a] George H. Czerlinski. Chemical relaxation of cyclic enzyme reactions IV. Thermodynamics of isomerase systems. Journal of Theoretical Biology, 31(3):493–499, June 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900245.

Czerlinski:1971:CRCb

[Cze71b] George H. Czerlinski. Chemical relaxation of cyclic enzyme systems: V. The Michaelis and Henri schemes. Journal of Theoretical Biology, 32(2):373–380, IN1–IN2, 381–382, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901743.

Daly:1979:WDM

[Dal79] Martin Daly. Why don’t male mammals lactate? Journal of Theoretical Biology, 78(3):325–345, June 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903345.

Danielli:1970:E

[Dan70] J. F. Danielli. Editorial. Journal of Theoretical Biology, 28(3):303, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937090072X.

Danielli:1975:E

[Dan75] James F. Danielli. Editorial. Journal of Theoretical Biology, 49 (2):i, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901708.

Dancis:1976:RTI

[Dan76] Barry Martin Dancis. Ring theory II. Fractional tandem model. Journal of Theoretical Biology, 61(2):329–351, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900229.

Darvey:1973:KIE

[Dar73a] I. G. Darvey. Kinetics of isotope exchange at equilibrium for a one substrate-one product enzyme mechanism. Journal of Theoretical Biology, 42(1):55–62, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901483.

Darvey:1973:RBS

[Dar73b] I. G. Darvey. The relationship between steady-state kinetic parameters and equilibrium parameters for one substrate/ one product enzymic reactions. Journal of Theoretical Biology, 41(3):441–450, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900544.

Darvey:1975:OSO

[Dar75] I. G. Darvey. One-substrate-one-product enzymic reactions: The relationship between isotope-exchange kinetic, steady-state kinetic and equilibrium parameters. Journal of Theoretical Biology, 49(1):201–211, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800282.

Darvey:1977:OSO

[Dar77a] I. G. Darvey. One substrate-one product enzymic reactions the determination of the velocity constants for an enzyme mechanism involving two conformational forms of enzyme and one enzyme intermediate. Journal of Theoretical Biology, 66(4):653– 663, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902363.

Darvey:1977:TPK

[Dar77b] Ivan G. Darvey. Transient phase kinetics of enzyme reactions where more than one species of enzyme is present at the start of the reaction. Journal of Theoretical Biology, 65(3):465–478, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902089.

Davis:1971:CPP

[Dav71] B. K. Davis. Chain propagation and polypeptide polymerization rate. Journal of Theoretical Biology, 30(1):203–210, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900452.

Davydov:1973:TCP

[Dav73] A. S. Davydov. The theory of contraction of proteins under their excitation. Journal of Theoretical Biology, 38(3):559– 569, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902567.

Davydov:1977:SET

[Dav77] A. S. Davydov. Solitons and energy transfer along protein molecules. Journal of Theoretical Biology, 66(2):379–387, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901783.

Davidson:1971:NCG

[DB71] Eric H. Davidson and Roy J. Britten. Note on the control of gene expression during development. Journal of Theoretical Biology, 32(1):123–130, July 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901408.

Daan:1978:TCO

[DB78] Serge Daan and Charles Berde. Two coupled oscillators: Simulations of the circadian pacemaker in mammalian activity rhythms. Journal of Theoretical Biology, 70(3):297–313, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903788.

Dalrymple:1970:DMC

[DBS^{+}70] G. V. Dalrymple, M. L. Baker, J. L. Sanders, A. J. Moss, and … K. P. Wilkinson. Do mammalian cells repair radiation injury by the “cut-and-patch” mechanism? II. An extension of the original model. Journal of Theoretical Biology, 28(1):121–142, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900688.

Dragomir:1976:NHT

[DBU76] C. T. Dragomir, Anca Barbier, and Daria Ungureanu. A new hybrid theory of muscle contraction. Journal of Theoretical Biology, 61(1):221–244, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901168.

Desimone:1973:SSS

[DC73] John A. Desimone and S. Roy Caplan. Symmetry and the stationary state behavior of enzyme membranes. Journal of Theoretical Biology, 39(3):523–544, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900660.

Chen:1975:FNK

[dC75] Yi der Chen. Fluctuations and noise in kinetic systems III. Cycling steady-state models. Journal of Theoretical Biology, 55(1):229–243, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801172.

Delwiche:1977:AMH

[DC77a] Michael J. Delwiche and J. Robert Cooke. An analytical model of the hydraulic aspects of stomatal dynamics. Journal of Theoretical Biology, 69(1):113–141, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903915.

Chen:1977:FNN

[dC77b] Yi der Chen. Fluctuations and noise in non-linear kinetic systems. Journal of Theoretical Biology, 65(2):357–367, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903307.

Dragomir:1977:DSC

[DC77c] C. T. Dragomir and S. A. Cananau. Distribution of subcellular components under gravity changes. The special condition of the dystrophic cell. Journal of Theoretical Biology, 65(1):209–213, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900856.

Dunford:1974:ATS

[DCMR74] H. B. Dunford, J. E. Critchlow, R. J. Maguire, and R. Roman. The advantages of transition state and group acid dissociation constants for pH-dependent enzyme kinetics. Journal of Theoretical Biology, 48(2):283–298, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800020.

Dounce:1973:SHE

[DCT73] Alexander L. Dounce, Subir K. Chanda, and Philip L. Townes. The structure of higher eukaryotic chromosomes. Journal of Theoretical Biology, 42(2):275–280, IN3–IN6, 281–285, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900908.

Denis:1977:MSA

[DD77] Michll Denis and Richard Deyrieux. Is multicomponent spectra analysis coming to a deadlock? Journal of Theoretical Biology, 69(2):301–309, November 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901394.

Denisov:1973:PRM

[DDTG73] D. A. Denisov, L. N. Drozdov-Tichomirov, and D. N. Grigoryeva. Pattern recognition methods for prediction of helical region in proteins. Journal of Theoretical Biology, 41(3):431– 439, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900532.

DeAngelis:1975:GAS

[De 75] D. L. De Angelis. Global asymptotic stability criteria for models of density-dependent population growth. Journal of Theoretical Biology, 50(1):35–43, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900223.

Deakin:1972:MIP

[Dea72] Michael A. B. Deakin. A model for inversion polymorphism. Journal of Theoretical Biology, 35(2):191–212, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900331.

Dean:1975:EME

[Dea75] P. M. Dean. Exocytosis modelling: An electrostatic function for calcium in stimulus-secretion coupling. Journal of Theoretical Biology, 54(2):289–308, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801329.

Delattre:1971:TOP

[Del71] P. Delattre. Topological and order properties in transformation systems. Journal of Theoretical Biology, 32(2):269–282, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901652.

Delisi:1974:TPA

[Del74] Charles Delisi. A theory of precipitation and agglutination reactions in immunological systems. Journal of Theoretical Biology, 45(2):555–575, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901301.

Delisi:1975:KHPb

[Del75a] Charles Delisi. The kinetics of hemolytic plaque formation III. Inhibition of plaques by antigen. Journal of Theoretical Biology, 51(2):337–345, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900648.

Delisi:1975:KHPc

[Del75b] Charles Delisi. The kinetics of hemolytic plaque formation IV. IgM plaque inhibition. Journal of Theoretical Biology, 52(2):419–440, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900107.

Demetrius:1976:MVA

[Dem76] L. Demetrius. Measures of variability in age-structured populations. Journal of Theoretical Biology, 63(2):397–404, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900424.

Demetrius:1977:TSS

[Dem77] Lloyd Demetrius. Trophic structure and species diversity. Journal of Theoretical Biology, 69(3):447–455, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901515.

Denhardt:1972:TDR

[Den72] David T. Denhardt. A theory of DNA replication. Journal of Theoretical Biology, 34(3):487–508, March 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901373.

Denholm:1975:NCM

[Den75a] J. V. Denholm. Necessary condition for maximum yield in a senescing two-phase plant. Journal of Theoretical Biology, 52(1): 251–254, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900569.

Denisov:1975:CLH

[Den75b] D. A. Denisov. On characteristic lengths of helical and nonhelical segments in proteins. Journal of Theoretical Biology, 54(2):383–387, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801391.

Denisov:1975:PRM

[Den75c] D. A. Denisov. Pattern recognition methods for prediction of secondary structure of proteins: II. Elimination of “contradicting” signs and prediction of turns of polypeptide chains in proteins. Journal of Theoretical Biology, 55(1):107–114, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580110X.

Derome:1977:BSG

[Der77] Jean-Robert Derome. Biological similarity and group theory. Journal of Theoretical Biology, 65(2):369–378, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903319.

Derome:1978:RBM

[Der78] Jean-Robert Derome. On the relationship between metabolism and body mass. Journal of Theoretical Biology, 72(4):757–759, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902837.

Deshcherevsky:1977:KMR

[Des77a] V. I. Deshcherevsky. Kinetic model of regulation of muscle protein activity. Journal of Theoretical Biology, 64(3):517–534, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902855.

Desimone:1977:PSD

[Des77b] John A. Desimone. Perturbations in the sof the double layer at an enzymic surface. Journal of Theoretical Biology, 68(2):225–240, September 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901618.

DeVault:1976:TIS

[DeV76] Don DeVault. Theory of iron-sulfur center n-2 oxidation and reduction by ATP. Journal of Theoretical Biology, 62(1):115–139, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900540.

Delisi:1975:KHPa

[DG75] Charles Delisi and Buron Goldstein. The kinetics of hemolytic plaque formation II. Inhibition of plaques by Hapten. Journal of Theoretical Biology, 51(2):313–335, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900636.

DeMaertelaer:1977:CKM

[DG77] V. De Maertelaer and P. Galand. Cell kinetics: A model of stimulated populations. Journal of Theoretical Biology, 69 (3):429–445, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377901503.

deHaen:1976:NSF

[dH76] Christoph de Haën. The non-stoichiometric floating receptor model for hormone sensitive adenylyl cyclase. Journal of Theoretical Biology, 58(2):383–400, ???? 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801269.

Dolowy:1978:CIT

[DH78] Krzysztof Dolowy and Frank J. Holly. Contribution of interfacial tension changes during cellular interaction to the energy balance. Journal of Theoretical Biology, 75(3):373–380, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903417.

Dancis:1979:TRF

[DH79] Barry M. Dancis and Gerald P. Holmquist. Telomere replication and fusion in eukaryotes. Journal of Theoretical Biology, 78(2):211–224, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902650.

Dipaolo:1978:SAS

[DHK78] Therese Dipaolo, Lowell H. Hall, and Lemont B. Kier. Structureactivity studies on hallucinogenic amphetamines using model interaction calculations. Journal of Theoretical Biology, 71 (3):295–309, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901613.

Diamond:1974:ADI

[Dia74] Phil Diamond. Area of discovery of an insect parasite. Journal of Theoretical Biology, 45(2):467–471, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901258.

Diamond:1976:DSR

[Dia76] Phil Diamond. Domains of stability and resilience for biological populations obeying difference equations. Journal of Theoretical Biology, 61(2):287–306, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900205.

Dingman:1974:BCR

[Din74] C. Wesley Dingman. Bidirectional chromosome replication: Some topological considerations. Journal of Theoretical Biology, 43(1):187–195, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800524.

Duerden:1978:AAP

[DJ78] M. Duerden and M. N. Jones. Amphipathicity and aggregation in polypeptide and protein systems. Journal of Theoretical Biology, 75(1):1–19, November 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901996.

Doljanski:1976:CSS

[DK76] F. Doljanski and M. Kapeller. Cell surface shedding — the phenomenon and its possible significance. Journal of Theoretical Biology, 62(2):253–270, October 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901193.

deKwaadsteniet:1976:QDH

[dKJS76] J. W. de Kwaadsteniet, J. C. Jager, and A. H. Stouthamer. A quantitative description of heterotrophic growth in microorganisms. Journal of Theoretical Biology, 57(1):103–120, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800070.

Dogonadze:1977:CDB

[DKU77] R. R. Dogonadze, A. M. Kuznetsov, and J. Ulstrup. Conformational dynamics in biological electron and atom transfer reactions. Journal of Theoretical Biology, 69(2):239–263, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901345.

deLong:1977:BD

[dL77] Richard de Long. On bacterial dissociation. Journal of Theoretical Biology, 64(4):761–764, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902740.

deLevie:1976:UAI

[dLA76] Robert de Levie and Kathleen M. Abbey. A unified approach to ion transport through membranes: I. Membrane-soluble ions. Journal of Theoretical Biology, 56(1):151–173, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800501.

Damiani:1975:TCA

[DLK75] A. Damiani, A. Lapiccirella, and A. L. Kovacs. Theoretical conformational analysis of 2-phosphopropionic acid, 2-phospho,3-olo-propionic acid and d-2,3-diphosphoglycerate. Journal of Theoretical Biology, 52(2):383–397, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900089.

Dibrov:1977:MMIa

[DLV77a] B. F. Dibrov, M. A. Livshits, and M. V. Volkenstein. Mathematical model of immune processes. Journal of Theoretical Biology, 65(4):609–631, April 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900121.

Dibrov:1977:MMIb

[DLV77b] B. F. Dibrov, M. A. Livshits, and M. V. Volkenstein. Mathematical model of immune processes: II. Kinetic features of antigen-antibody interrelations. Journal of Theoretical Biology, 69(1):23–39, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903861.

Dean:1975:LVW

[DM75] P. M. Dean and E. K. Matthews. The London–van der Waals attraction constant of secretory granules and its significance. Journal of Theoretical Biology, 54(2):309–321, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801330.

Dunker:1978:MMT

[DM78] A. Keith Dunker and D. A. Marvin. A model for membrane transport through α-helical protein pores. Journal of Theoretical Biology, 72(1):9–16, May 8, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900152.

Duchesne:1978:EHW

[DMRD78] J. Duchesne, F. Mosora, M. Rinné, and A. Dresse. The effect of heavy water on the organic free radical content in rat liver and kidney: Consequences for senescence and cancer. Journal of Theoretical Biology, 71(2):275–278, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902722.

Davis:1976:SSH

[DO76] J. W. F. Davis and P. O’Donald. Sexual selection for a handicap: A critical analysis of Zahavi’s model. Journal of Theoretical Biology, 57(2):345–354, April 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900060.

deOrtiz:1979:SPS

[dO79] Conchita Zuazaga de Ortiz. The selective permeability of sodium and potassium channels in nerve: Orbital control hypothesis. Journal of Theoretical Biology, 78(1):1–10, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903205.

Doerner:1971:FMS

[Doe71a] Karl F. Doerner. Forces on the mitotic spindle produced by an internal pressure gradient. Journal of Theoretical Biology, 30(3):573–577, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900105.

Doerner:1971:NIP

[Doe71b] Karl F. Doerner. Nerve impulse propagation affected by solgel changes in axoplasm. Journal of Theoretical Biology, 32(1): 159–164, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901433.

Doetsch:1972:UTB

[Doe72] R. N. Doetsch. A unified theory of bacterial motile behavior. Journal of Theoretical Biology, 35(1):55–66, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901919.

Doerner:1976:SSS

[Doe76] Karl Doerner. Some stress-strain relationships of cellular cleavage and noncleavage. Journal of Theoretical Biology, 61 (1):205–209, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901144.

Dolowy:1975:UHC

[Dol75] Kryysztof Dolowy. Uniform hypothesis of cell behaviourmovement, contact inhibition of movement, adhesion, chemotaaris, phagocytosis, pinocytosis, division, contact inhibition of division, fusion. Journal of Theoretical Biology, 52(1):83–97, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900417.

Donaldson:1972:CLF

[Don72] Ian G. Donaldson. Cyclic longitudinal fibrillar motion as a basis for steady rotational protoplasmic streaming. Journal of Theoretical Biology, 37(1):75–91, October 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901154.

Douthart:1972:TCDa

[Dou72a] Richard J. Douthart. Topological constraints daring replication I. Generation of catenanes, circular dimers and higher oligomers. Journal of Theoretical Biology, 35(2):315–335, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900422.

Douthart:1972:TCDb

[Dou72b] Richard J. Douthart. Topological constraints during replication II. Complex forms generated by multiple replication forks or track switching. Journal of Theoretical Biology, 35(2):337–358, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900434.

Douglas:1975:SFS

[Dou75] G. J. Douglas. Sliding filaments in sperm flagella. Journal of Theoretical Biology, 53(2):247–252, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800014.

Downie:1970:NMT

[Dow70] D. E. Downie. New model for transmitter release at the presynaptic membrane. Journal of Theoretical Biology, 28(2):297– 300, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900573.

Dowman:1973:ISI

[Dow73] J. E. Dowman. Implications of stochastic inheritance. Journal of Theoretical Biology, 39(1):55–72, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902051.

Doyle:1978:GTC

[Doy78] Gregory G. Doyle. A general theory of chromosome pairing based on the palindromic DNA model of Sobell with modifications and amplifications. Journal of Theoretical Biology, 70(2):171–184, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378903454.

Dragomir:1974:TCR

[DP74] C. T. Dragomir and E. Pausescu. Theoretical considerations regarding the cell membrane function during cold-induced preservation of tissues and organs: New possibilities for optimizing the process. Journal of Theoretical Biology, 47(2):281–293, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901982.

DeLisi:1976:KAP

[DP76] Charles DeLisi and Alan Perelson. The kinetics of aggregation phenomena I. Minimal models for patch formation on lymphocyte membranes. Journal of Theoretical Biology, 62(1):159–210, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900564.

Delong:1977:EA

[DP77a] Richard Delong and Loyd Poplin. On the etiology of aging. Journal of Theoretical Biology, 67(1):111–120, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901874.

Duncan:1977:MEE

[DP77b] Ruth Duncan and Margaret K. Pratten. Membrane economics in endocytic systems. Journal of Theoretical Biology, 66(4):727– 735, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902417.

Dragomir:1970:NFA

[Dra70] Caius Traian Dragomir. On the nature of the forces acting between myofilaments in resting state and under contraction. Journal of Theoretical Biology, 27(3):343–356, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800017.

Dragomir:1971:ZMM

[Dra71] Caius Traian Dragomir. Zone melting as a model for active transport across the cell membrane. Journal of Theoretical Biology, 31(3):453–468, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937190021X.

Dratman:1974:MAT

[Dra74] Mary B. Dratman. On the mechanism of action of thyroxin, an amino acid analog of tyrosine. Journal of Theoretical Biology, 46(1):255–270, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901519.

Dehaven:1970:SRE

[DS70] J. C. Dehaven and N. Z. Shapiro. Simulation of the renal effects of antidiuretic hormone (ADH) in man. Journal of Theoretical Biology, 28(2):261–286, August 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090055X.

Damianovich:1973:MEM

[DS73a] Sándor Damianovich and Bela Somogyi. A molecular enzyme model based on oriented energy transfer. Journal of Theoretical Biology, 41(3):567–569, October 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900635.

DeMichele:1973:AMG

[DS73b] D. W. DeMichele and Peter J. H. Sharpe. An analysis of the mechanics of guard cell motion. Journal of Theoretical Biology, 41(1):77–78, IN1–IN2, 79–96, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901902.

Dreitlein:1974:MOK

[DS74] Joseph Dreitlein and Marie-Luce Smoes. A model for oscillatory kinetic systems. Journal of Theoretical Biology, 46(2):559– 572, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900150.

deSobrino:1975:GPT

[dS75] Luis de Sobrino. The Glansdorff–Prigogine thermodynamic stability criterion in the light of Lyapunov’s theory. Journal of Theoretical Biology, 54(2):323–333, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801342.

Duniec:1979:AEM

[DST79] J. T. Duniec, M. J. Sculley, and S. W. Thorne. An analysis of the effect of mono- and di-valent cations on the forces between charged lipid membranes with special reference to the grana thylakoids of chloroplasts. Journal of Theoretical Biology, 79(4):473–484, August 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902388.

Deakin:1974:GMI

[DT74a] Michael A. B. Deakin and Rodney B. Teague. A generalized model of inversion polymorphism. Journal of Theoretical Biology, 48(1):105–123, November 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901829.

Dugan:1974:FPR

[DT74b] Virgil Dugan and Ralph Trujillo. On a fundamental problem in radiation biology. Journal of Theoretical Biology, 44(2):397– 401, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901702.

Dromashko:1979:SMR

[DT79] S. E. Dromashko and N. A. Troitsky. A stochastic model of recombination during conjugation in Escherichia coli k-12. Journal of Theoretical Biology, 77(1):37–45, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901358.

Duchesne:1977:UBT

[Duc77] Jules Duchesne. A unifying biochemical theory of cancer, senescence and maximal life span. Journal of Theoretical Biology, 66(1):137–145, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903162.

Duggleby:1979:EDE

[Dug79] Ronald G. Duggleby. Experimental designs for estimating the kinetic parameters for enzyme-catalysed reactions. Journal of Theoretical Biology, 81(4):671–684, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902765.

Dogonadze:1973:TBA

[DUK73a] R. R. Dogonadze, J. Ulstrup, and Yu. I. Kharkats. A theory of bridge-assisted polar medium electron transfer reactions between molecules having quasicontinuous electronic energy spectra. Journal of Theoretical Biology, 40(2):279–283, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390132X.

Dogonadze:1973:TPM

[DUK73b] R. R. Dogonadze, J. Ulstrup, and Yu. I. Kharkats. A theory of polar medium electron transfer reactions through bridge groups with a quasicontinuous energy spectrum. Journal of Theoretical Biology, 40(2):259–277, August 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901318.

Dunford:1974:KER

[Dun74] H. B. Dunford. Kinetic exclusion rules applied to the study of pH-dependent enzyme reactions: Application of the diffusioncontrolled limit. Journal of Theoretical Biology, 46(2):467– 479, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900095.

Durston:1973:DDA

[Dur73] A. J. Durston. Dictyostelium discoideum aggregation fields as excitable media. Journal of Theoretical Biology, 42(3):483–504, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902427.

Deland:1979:TES

[DVM79] E. C. Deland, M. F. Villamil, and J. V. Maloney. A theoretical and experimental study of ionic shifts induced by k depletion and replacement. Journal of Theoretical Biology, 76(1):31–51, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903746.

Davies:1971:QRB

[DW71] J. I. Davies and P. A. Williams. Quantitative relationship between stimulus and response in hormone action: Amplification and sensitization. Journal of Theoretical Biology, 30(1):41– 57, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937190035X.

Davies:1975:QARb

[DW75a] J. I. Davies and P. A. Williams. Quantitative aspects of regulation of cellular cyclic AMP levels: II. Kinetics of drug action in a model cyclic AMP generating system. Journal of Theoretical Biology, 53(1):31–50, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901010.

Davies:1975:QARa

[DW75b] J. I. Davies and P. A. Williams. Quantitative aspects of the regulation of cellular cyclic AMP levels: I. Structure and kinetics of a model system. Journal of Theoretical Biology, 53(1):1–30, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901009.

Denizeau:1976:KMT

[DWS76] F. Denizeau, J. Wyse, and T. L. Sourkes. Kinetics of multiple transport systems for amino acids in the brain. Journal of Theoretical Biology, 63(1):99–110, November 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900862.

Danilov:1971:ESD

[DZOP71] V. I. Danilov, N. V. Zheltovsky, V. V. Ogloblin, and V. I. Pechenaya. Excited states of the DNA base pairs. Journal of Theoretical Biology, 30(3):559–571, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900099.

Easton:1979:ODC

[Eas79] Dexter M. Easton. Oxyhemoglobin dissociation curve as expo-exponential paradigm of asymmetric sigmoid function. Journal of Theoretical Biology, 76(3):335–349, February 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903163.

Ebringer:1975:ITL

[Ebr75a] A. Ebringer. Information theory and limitations in antibody diversity. Journal of Theoretical Biology, 51(2):293–302, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900600.

Ebringer:1975:CCD

[Ebr75b] Alan Ebringer. A comment on code degeneracy, “information theory and the central dogma”. Journal of Theoretical Biology, 53(1):243–245, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901150.

Economos:1979:STB

[Eco79] Angelos C. Economos. On structural theories of basal metabolic rate. Journal of Theoretical Biology, 80(4):445–450, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900870.

Edelstein:1970:BMM

[Ede70a] Barry B. Edelstein. Biochemical model with multiple steady states and hysteresis. Journal of Theoretical Biology, 29(1):57– 62, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901189.

Edelstein:1970:IAD

[Ede70b] Barry B. Edelstein. Instabilities associated with dissipative structure. Journal of Theoretical Biology, 26(2):227–241, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800145.

Edelstein:1971:ABS

[Ede71a] Barry B. Edelstein. Autocatalysis in a biological system. Journal of Theoretical Biology, 32(1):191–197, July 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901469.

Edelstein:1971:CSD

[Ede71b] Barry B. Edelstein. Cell specific diffusion model of morphogenesis. Journal of Theoretical Biology, 30(3):515–532, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900051.

Edelstein:1972:DCD

[Ede72] Barry B. Edelstein. The dynamics of cellular differentiation and associated pattern formation. Journal of Theoretical Biology, 37(2):221–243, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900185.

Edstrom:1975:EEB

[Eds75] J.-E. Edström. Eukaryotic evolution based on information in chromosomes on allele frequencies. Journal of Theoretical Biology, 52(1):163–174, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937590048X.

Edwards:1978:SIF

[Edw78]  Paul A. W. Edwards.  A sequence-independent, four-

stranded, double Watson-Crick DNA helix that could solve the unwinding problem of double helices. Journal of Theoretical Biology, 70(3):323–334, February 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903806.

Elton:1976:RVP

[EF76] R. A. Elton and W. Fiers. Rhythmic variations in purine run frequencies in bacteriophage RNA. Journal of Theoretical Biology, 63(1):49–59, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900837.

Engasser:1977:HRC

[EFH77] Jean-Marc Engasser, Michael Flamm, and Csaba Horvath. Hormone regulation of cellular metabolism: Interplay of membrane transport and consecutive enzymic reaction. Journal of Theoretical Biology, 67(3):433–445, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900479.

Egel:1979:LOL

[Ege79] Richard Egel. The lac-operon for lactose degradation, or rather for the utilization of galactosylglycerols from galactolipids? Journal of Theoretical Biology, 79(1):117–119, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902601.

Espinosa:1977:MAS

[EGF77] M. Espinosa, E. Garcia, and J. M. Fernaud. Mathematical approach to the stimulation of the competence development in bacillus subtilis. Journal of Theoretical Biology, 67(1):155– 174, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901904.

Engstrom-Heg:1970:PCE

[EH70] Verna Louise Engstrom-Heg. Predation, competition and environmental variables: Some mathematical models. Journal of Theoretical Biology, 27(2):175–195, May 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901360.

Engasser:1973:EID

[EH73] Jean-Marc Engasser and Csaba Horvath. Effect of internal diffusion in heterogeneous enzyme systems: Evaluation of true kinetic parameters and substrate diffusivity. Journal of Theoretical Biology, 42(1):137–155, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901537.

Engasser:1979:DEH

[EH79] Jean-Marl Engasser and Patrick Hisland. Diffusional effects on the heterogeneous kinetics of two-substrate enzymic reactions. Journal of Theoretical Biology, 77(4):427–440, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900183.

Ehrlich:1979:EMS

[Ehr79] Paul H. Ehrlich. The effect of multivalency on the specificity of protein and cell interactions. Journal of Theoretical Biology, 81(1):123–127, November 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900857.

Eijkman:1970:LFB

[Eij70] E. G. Eijkman. Learning and fluctuating behaviour experiments with goldfish. Journal of Theoretical Biology, 26(2):251–264, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800169.

Elseth:1978:KMB

[EJB^{+}78] Gerald D. Elseth, George Juetersonke, Mary Ann Bierman, Lawrence P. Karcher, and Kandy D. Baumgardner. A kinetic model for bacterial transfection: The lambda DNA system. Journal of Theoretical Biology, 72(3):429–441, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903107.

Elder:1973:MPM

[Eld73] D. Elder. A multiple promoter model for transcriptional control in differentiated organisms. Journal of Theoretical Biology, 39(3):673–675, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900805.

Elder:1979:BQM

[Eld79a] D. Elder. BUdR and the quantal mitosis: A cleavage plane shift model. Journal of Theoretical Biology, 78(2):299–304, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902704.

Elder:1979:WRC

[Eld79b] D. Elder. Why is regenerative capacity restricted in higher organisms? Journal of Theoretical Biology, 81(3):563–568, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900535.

Elsasser:1972:MBI

[Els72] Walter M. Elsasser. A model of biological indeterminacy. Journal of Theoretical Biology, 36(3):627–633, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900136.

Elton:1974:TMH

[Elt74] R. A. Elton. Theoretical models for heterogeneity of base composition in DNA. Journal of Theoretical Biology, 45(2):533– 553, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901295.

Eldridge:1978:HTI

[EM78] Charles A. Eldridge and Harold J. Morowitz. A hydrodynamic theory of ion conductance through Ohmic pores. Journal of Theoretical Biology, 73(3):539–548, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901571.

Eisen:1979:MAI

[EMM79] Martin Eisen, Nicholas Macri, and Jatinder Mehta. Microfluorometry analysis II. A Bayesian approach. Journal of Theoretical Biology, 78(1):43–49, May 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903242.

Engel:1978:DAA

[Eng78] Paul C. Engel. The detection of amino acid sequence homology: A new scoring method based on the genetic code with allowance for redundancy. Journal of Theoretical Biology, 72(1):1–8, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900140.

Epstein:1971:REIa

[Eps71a] Henry F. Epstein. Role of electrostatic interactions in allosteric proteins. Journal of Theoretical Biology, 31(1):69–76, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901226.

Epstein:1971:REIb

[Eps71b] Henry F. Epstein. Role of electrostatic interactions in allosteric proteins. Journal of Theoretical Biology, 31(3):i, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900312.

Epstein:1979:CCS

[Eps79] Irving R. Epstein. Competitive coexistence of self-reproducing macromolecules. Journal of Theoretical Biology, 78(2):271–298, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902698.

Estabrook:1974:CVP

[ER74] G. F. Estabrook and H. H. Robinson. A comment on vine’s predator–prey visual detection model. Journal of Theoretical Biology, 47(1):245–247, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490112X.

Edelstein:1978:ESR

[ER78] L. Edelstein and R. Rosen. Enzyme-substrate recognition. Journal of Theoretical Biology, 73(1):181–204, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901856.

Erhan:1976:IDC

[Erh76] Semih Erhan. Information density, cellular immunity and transfer factor. Journal of Theoretical Biology, 63(1):111–115, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900874.

Eisen:1977:MA

[ES77] Martin Eisen and John Schiller. Microfluorometry analysis. Journal of Theoretical Biology, 66(4):799–809, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902478.

Eshel:1978:HPC

[Esh78] Ilan Eshel. On the handicap principle — a critical defence. Journal of Theoretical Biology, 70(2):245–250, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903508.

Evans:1979:STI

[Eva79] Christopher H. Evans. Is the A(G_{0}) state time-independent? Journal of Theoretical Biology, 79(2):259–262, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902510.

Ellis:1976:CMD

[EWRA76] James E. Ellis, John A. Wiens, Charles F. Rodell, and Jerry C. Anway. A conceptual model of diet selection as an ecosystem process. Journal of Theoretical Biology, 60(1):93–108, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901570.

Fabrega:1976:BST

[Fab76] Horacio Fabrega. The biological significance of taxonomies of disease. Journal of Theoretical Biology, 63(1):191–216, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900928.

Fagen:1978:IMS

[Fag78] Robert M. Fagen. Information measures: Statistical confidence limits and inference. Journal of Theoretical Biology, 73(1):61– 79, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901807.

Falkowski:1977:TDN

[Fal77] Paul G. Falkowski. A theoretical description of nitrate uptake kinetics in marine phytoplankton based on bisubstrate kinetics. Journal of Theoretical Biology, 64(2):375–379, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903642.

Ferrara:1977:CBP

[FAT77] L. Ferrara, S. Andini, and P. A. Temussi. Correlation between prebiotic amino acid compositions and contemporary proteins. Journal of Theoretical Biology, 67(2):241–254, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901977.

Filmer:1970:EVT

[FC70] D. L. Filmer and T. G. Cooper. Effect of varying temperature and pH upon the predicted rate of “CO_{2}” utilization by carbylases. Journal of Theoretical Biology, 29(1):131–145, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901220.

Fraidenraich:1971:KAI

[FC71] Naum Fraidenraich and Marcelino Cereijido. Kinetic analysis of isotope exchange in a system of concentric cylinders. Journal of Theoretical Biology, 32(1):107–122, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901391.

Freist:1976:BCM

[FC76] Wolfgang Freist and Friedrich Cramer. A binary code model for substrate recognition of aminoacyl-tRNA ligases. Journal of Theoretical Biology, 58(2):401–416, ???? 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801270.

Friedman:1978:RMP

[FCSS78] Marvin A. Friedman, Walter H. Carter, Jack Staub, and Anthony Segreti. Relative mutagenic potencies determined in the micronucleus test. Journal of Theoretical Biology, 73 (4):575–584, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901236.

Findlay:1973:MPR

[FD73] J. M. Findlay and G. J. Daniell. A model for pattern recognition by cell networks. Journal of Theoretical Biology, 38(3):641– 645, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902634.

Feduccia:1977:HSE

[Fed77] Alan Feduccia. Hypothetical stages in the evolution of modern ducks and flamingos. Journal of Theoretical Biology, 67(4):715–721, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902569.

Ferreira:1973:EOS

[Fer73] Ricardo Ferreira. Enzymes as orbital symmetry adapters. Journal of Theoretical Biology, 39(3):665–668, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900787.

Frisch:1977:PAC

[FG77] H. L. Frisch and I. J. Gotham. On periodic algal cyclostat populations. Journal of Theoretical Biology, 66(4):665–678, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902375.

Fantes:1975:RCS

[FGP^{+}75] P. A. Fantes, W. D. Grant, R. H. Pritchard, P. E. Sudbery, and A. E. Wheals. The regulation of cell size and the control of mitosis. Journal of Theoretical Biology, 50(1):213–244, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590034X.

Fondy:1971:SSW

[FH71a] Thomas P. Fondy and Peter D. Holohan. Structural similarities within groups of pyridine nucleotide-linked dehydrogenases. Journal of Theoretical Biology, 31(2):229–244, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901846.

Frieder:1971:RRO

[FH71b] Gideon Frieder and Gabor T. Herman. Resolution in reconstructing objects from electron micrographs. Journal of Theoretical Biology, 33(1):189–196, IN2, 197–211, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902244.

Fujita:1970:TAB

[FIN70] Hirokg Fujita, Akira Imamura, and Chikayoshi Nagata. Tautomerism of the anticodon base and the code degeneracy. Journal of Theoretical Biology, 28(1):143–145, July 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090069X.

Finkel:1972:BPC

[Fin72] Robert Warren Finkel. Biological prediction: A comparative approach. Journal of Theoretical Biology, 37(2):245–250, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900197.

Finkel:1973:RCC

[Fin73] Robert W. Finkel. Response to Chen’s comment. Journal of Theoretical Biology, 42(3):591–592, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390249X.

Fujita:1974:MOS

[FIN74] Hiroko Fujita, Akira Imamura, and Chikayoshi Nagata. A molecular orbital study of stability and the conformation of double-stranded DNA-like polymers. Journal of Theoretical Biology, 45(2):411–433, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901222.

Fink:1976:CUS

[Fin76a] Anthony L. Fink. Cryoenzymology: The use of sub-zero temperatures and fluid solutions in the study of enzyme mechanisms. Journal of Theoretical Biology, 61(2):419–445, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690028X.

Finn:1976:MES

[Fin76b] John T. Finn. Measures of ecosystem structure and function derived from analysis of flows. Journal of Theoretical Biology, 56(2):363–380, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937680080X.

Fishman:1973:CAL

[Fis73] S. N. Fishman. Conditions of appearance of local response and action potential in a simple mathematical model of nerve. Journal of Theoretical Biology, 41(3):409–420, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900519.

Fish:1976:FSP

[Fis76] Wayne W. Fish. Ferritin structure: Possible models for apoferritin subunit arrangement. Journal of Theoretical Biology, 60(2):385–386, IN1–IN4, 387–392, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900643.

Fischer:1977:IRC

[Fis77] Harry W. Fischer. Improvement in radiographic contrast media through the development of colloidal or particulate media an analysis. Journal of Theoretical Biology, 67(4):653–670, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902521.

FitzHugh:1973:DAN

[Fit73] Richard FitzHugh. Dimensional analysis of nerve models. Journal of Theoretical Biology, 40(3):517–541, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900088.

Fitch:1976:DMR

[Fit76] W. M. Fitch. The deduction of messenger RNA sequences from amino acid sequences is, to date, a total failure. Journal of Theoretical Biology, 59(1):251–252, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800409.

Fredrickson:1973:PPI

[FJThH73] A. G. Fredrickson, J. L. Jost, H. M. Tsuchiya, and Ping hwa Hsu. Predator–prey interactions between Malthusian populations. Journal of Theoretical Biology, 38(3):487–526, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902531.

Forslind:1975:SML

[FK75] Erik Forslind and Roland Kjellander. A structure model for the lecithin-cholesterol-water membrane. Journal of Theoretical Biology, 51(1):97–102, IN1–IN6, 103–109, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901411.

Ford:1979:ASU

[FK79] R. Glenn Ford and David W. Krumme. The analysis of space use patterns. Journal of Theoretical Biology, 76(2):125–155, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903667.

Frehland:1974:ITT

[FL74] E. Frehland and P. Läuger. Ion transport through pores: Transient phenomena. Journal of Theoretical Biology, 47(1):189–207, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937490109X.

Fleck:1972:GFO

[Fle72] George M. Fleck. On the generality of first-order rates in isotopic tracer kinetics. Journal of Theoretical Biology, 34(3):509– 514, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901385.

Flesch:1975:DHS

[Fle75] U. Flesch. The distribution of heat sources in the human head: A theoretical consideration. Journal of Theoretical Biology, 54(2):285–287, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801317.

Fletcher:1977:EPD

[Fle77] C. R. Fletcher. Electrical potential differences across salt transporting membranes. Journal of Theoretical Biology, 67(2):255– 268, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901989.

Frontali:1972:TMI

[FM72] L. Frontali and G. Morpurgo. A theoretical model for the intercistronic region. Journal of Theoretical Biology, 36(1):147–151, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901841.

Fujita:1976:MOS

[FN76] Hiroko Fujita and Chikayoshi Nagata. A molecular orbital study on the incorporation of nucleotides into RNA. Journal of Theoretical Biology, 57(1):187–202, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800124.

Fong:1974:EUT

[Fon74] Francis K. Fong. Energy upconversion theory of the primary photochemical reaction in plant photosynthesis. Journal of Theoretical Biology, 46(2):407–420, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490006X.

Fong:1977:CDI

[Fon77] F. K. Fong. Current developments in the investigation of the primary light reactions in photosynthesis. Journal of Theoretical Biology, 66(1):199–202, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903216.

Forsdyke:1975:FIT

[For75] D. R. Forsdyke. Further implications of a theory of immunity. Journal of Theoretical Biology, 52(1):187–198, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900508.

Fox:1971:ERO

[Fox71] Ronald Forrest Fox. Entropy reduction in open systems. Journal of Theoretical Biology, 31(1):43–46, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901196.

Fox:1978:MT

[Fox78] J. Lawrence Fox. A model for transhydrogenase. Journal of Theoretical Biology, 70(3):335–337, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903818.

Fischer:1972:TCC

[FP72] James J. Fischer and Rose J. Papac. Theoretical considerations in combinations of localized and systemic therapy for neoplastic disease. Journal of Theoretical Biology, 37(1):105–114, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901178.

Ferracin:1973:SHA

[FP73] Alessandro Ferracin and Renato Paoluzi. Some hypotheses about the structure of polypeptide chain termination signal in prokaryotes. Journal of Theoretical Biology, 39(1):179– 182, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902130.

Fisher:1972:RPM

[FPB72] James R. Fisher, David G. Priest, and Janice S. Barton. Random pathway models and their application to steady-state enzyme kinetics. Journal of Theoretical Biology, 37(2):335–352, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900276.

Franke:1970:MMS

[Fra70] E. K. Franke. A mathematical model of synchronized periodic growth of cell populations. Journal of Theoretical Biology, 26(3):373–382, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900901.

Frame:1974:EST

[Fra74a] J. S. Frame. Explicit solutions in two species Volterra systems. Journal of Theoretical Biology, 43(1):73–81, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800469.

Frankel:1974:PIU

[Fra74b] Joseph Frankel. Positional information in unicellular organisms. Journal of Theoretical Biology, 47(2):439–481, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902094.

Freemon:1970:RES

[Fre70] Frank R. Freemon. Reciprocal environmental surveillance model of sleep. Journal of Theoretical Biology, 27(2):339–340, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090144X.

Frech:1973:BEE

[Fre73] H. E. Frech. Biological externalities and evolution: A comment. Journal of Theoretical Biology, 39(3):669–672, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900799.

Farnell:1974:CCF

[FRG74] L. Farnell, W. G. Richards, and C. R. Ganellin. Calculation of conformational free energy of histamine. Journal of Theoretical Biology, 43(2):389–392, February 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800706.

Friedman:1971:GTT

[Fri71] M. H. Friedman. General theory of tissue swelling with application to the corneal stroma. Journal of Theoretical Biology, 30(1):93–109, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900397.

Fried:1972:PDG

[Fri72] Jerrold Fried. Proposal for the determination of generation time variability and dormancy of proliferating cell populations by a modification of the grain-count halving method. Journal of Theoretical Biology, 34(3):535–555, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901403.

Friedman:1973:CCT

[Fri73] M. H. Friedman. Critique of current theories of corneal hydration control. Journal of Theoretical Biology, 42(2):287–306, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937390091X.

Friedman:1974:PDP

[Fri74] M. H. Friedman. A physical description of the pathogenesis, histopathology and treatment of corneal epithelial edema. Journal of Theoretical Biology, 45(1):153–169, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900484.

Friedman:1976:EMH

[Fri76a] M. H. Friedman. The effect of membrane heterogeneity on the predictability of fluxes, with application to the cornea. Journal of Theoretical Biology, 61(2):307–328, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900217.

Frizell:1976:PED

[Fri76b] John Frizell. A possible electron dense label for the use in the sequencing of DNA by electron microscopy. Journal of Theoretical Biology, 59(1):241–242, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800379.

Friedrich:1979:PSP

[Fri79] Peter Friedrich. On the physiological significance of positive and negative cooperativity in enzymes. Journal of Theoretical Biology, 81(3):527–532, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990050X.

Forbes:1970:MAM

[FSFD70] W. F. Forbes, D. A. Sprott, M. Feldstein, and A. L. Dounce. A model to account for mortality curves of various species. Journal of Theoretical Biology, 29(2):293–299, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900238.

Figueroa:1972:BSN

[FSG^{+}72] Raúl Figueroa, A. Soto, G. González, M. Pieber, and … J. C. Tohá. Base sequence of a non-self-complementary messenger RNA of human heart cytochrome c. Journal of Theoretical Biology, 36(2):321–326, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901014.

Figueroa:1978:IAM

[FSST78] R. Figueroa, A. Sepúlveda, M. A. Soto, and J. Tohá. Informational analysis of MS_{2} and θx174 virus genomes. Journal of Theoretical Biology, 74(2):203–207, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900711.

Fraser:1974:GFR

[FT74] Alex Fraser and Jawahar Tiwari. Genetical feedback-repression: II. Cyclic genetic systems. Journal of Theoretical Biology, 47(2):397–412, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374902069.

Fujii:1977:CSR

[Fuj77] Koichi Fujii. Complexity-stability relationship of two-prey-one-predator species system model: Local and global stability. Journal of Theoretical Biology, 69(4):613–623, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903708.

Fukuda:1974:PDM

[FY74] Nobuo Fukuda and Nagasumi Yago. Population dynamics of mitochondria I. A model for the role of ACTH in the degradation of adrenocortical mitochondria. Journal of Theoretical Biology, 46 (1):21–30, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901386.

Fukuda:1976:PDM

[FY76] Nobuo Fukuda and Nagasumi Yago. Population dynamics of mitochondria II. Turnover and ageing of rat liver mitochondria. Journal of Theoretical Biology, 58(1):131–142, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901430.

Go:1977:RFA

[GA77] Nobuhiro Go and Yoichiro Anan. Regulatory functions of allosteric enzymes in far-from-equilibrium systems. Journal of Theoretical Biology, 66(3):475–483, June 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902971.

Galas:1978:SMP

[Gal78] David J. Galas. On the symmetries of multi-palindromic DNA sequences. Journal of Theoretical Biology, 72(1):57–73, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900176.

Gamble:1975:MAH

[Gam75] W. Gamble. Mechanism of action of hypolipidemic and herbicidal aryloxy acids. Journal of Theoretical Biology, 54(2):181– 190, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801238.

Garfinkel:1970:SKC

[Gar70] David Garfinkel. Simulation of the kinetics of ^{14}C in the Krebs cycle in liver and its application to the design of experiments. Journal of Theoretical Biology, 29(1):113–130, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901219.

Garel:1974:FAT

[Gar74] Jean-Pierre Garel. Functional adaptation of tRNA population. Journal of Theoretical Biology, 43(1):211–225, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800548.

Garay:1976:TRA

[Gar76] Ricardo P. Garay. Thermodynamic restrictions on the allosteric models through an analysis of the free energy of interaction between sites. Journal of Theoretical Biology, 63(2):421–441, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900448.

Gates:1978:BEA

[Gat78] D. J. Gates. Bimodality in even-aged plant monocultures. Journal of Theoretical Biology, 71(4):525–540, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903235.

Gaudin:1973:STP

[Gau73] David Gaudin. Some thoughts on a possible relationship between known gene dosage effects and neoplastic transformation. Journal of Theoretical Biology, 41(1):191–200, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901999.

Goloff:1975:GMN

[GB75] Alexander A. Goloff and F. A. Bazzaz. A germination model for natural seed populations. Journal of Theoretical Biology, 52(2):259–283, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900016.

Gallaher:1977:EFW

[GB77a] Edward E. Gallaher and Norman J. Blake. On equivalent forms of Whittaker’s similarity index. Journal of Theoretical Biology, 68(2):259–265, September 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901643.

Green:1977:ACC

[GB77b] Paul B. Green and Keith Bauer. Analysing the changing cell cycle. Journal of Theoretical Biology, 68(2):299–315, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901679.

Gordon:1970:ART

[GBH70] Richard Gordon, Robert Bender, and Gabor T. Herman. Algebraic reconstruction techniques (ART) for three-dimensional electron microscopy and X-ray photography. Journal of Theoretical Biology, 29(3):471–476, IN1–IN2, 477–481, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901098.

Goldsack:1973:CFE

[GC73] D. E. Goldsack and R. C. Chalifoux. Contribution of the free energy of mixing of hydrophobic side chains to the stability of the tertiary structure of proteins. Journal of Theoretical Biology, 39(3):645–651, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900751.

Gero:1976:EDA

[GC76] Alexander Gero and Robert J. Capetola. Exploration of drug action on a morphine receptor by methods of enzyme kinetics. Journal of Theoretical Biology, 61(1):129–142, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901090.

Ginzburg:1979:RGA

[GC79] L. R. Ginzburg and R. F. Costantino. On the rate of genetic adaptation under natural selection. Journal of Theoretical Biology, 77(3):307–316, April 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903588.

Goel:1970:SSI

[GCG^{+}70] Narendra Goel, Richard D. Campbell, Richard Gordon, Robert Rosen, and … Martynas Ycas. Self-sorting of isotropic cells. Journal of Theoretical Biology, 28(3):423–468, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900809.

Green:1972:CPS

[GCS72] H. S. Green and J. R. Casley-Smith. Calculations on the passage of small vesicles across endothelial cells by Brownian motion. Journal of Theoretical Biology, 35(1):103–111, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901956.

Gould:1979:PRB

[GD79] G. W. Gould and G. J. Dring. On a possible relationship between bacterial endospore formation and the origin of eukaryotic cells. Journal of Theoretical Biology, 81(1):47–53, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900791.

Goldstein:1975:IGC

[GDA75] Byron Goldstein, Charles Delisi, and Joseph Abate. Immunodiffusion in gels containing erythrocyte antigen I. Theory for diffusion of antiserum from a circular well. Journal of Theoretical Biology, 52(2):317–334, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900041.

Gersh:1975:SGA

[Ger75] Eileen Sutton Gersh. Sites of gene activity and of inactive genes in polytene chromosomes of diptera. Journal of Theoretical Biology, 50(2):413–428, April 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900909.

Gero:1978:BSD

[Ger78] Alexander Gero. Biological stimulus and drug effect. Journal of Theoretical Biology, 74(3):469–473, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902254.

Greenspan:1977:HCA

[GF77] Harvey P. Greenspan and Judah Folkman. Hypotheses on cell adhesion and actin cables. Journal of Theoretical Biology, 65(2):397–398, IN1, March 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903344.

Gray:1975:MERa

[GG75a] Brian F. Gray and Igor Gonda. A molecular explanation of the role of actin in acceleration of the myosin-catalysed ATP hydrolysis in vertebrate striated muscles. Journal of Theoretical Biology, 49(1):493–497, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580049X.

Gray:1975:MERb

[GG75b] Brian F. Gray and Igor Gonda. A molecular explanation of the role of actin in acceleration of the myosin-catalysed ATP hydrolysis in vertebrate striated muscles. Journal of Theoretical Biology, 49(2):493–497, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901873.

Gray:1977:SFMa

[GG77a] B. F. Gray and I. Gonda. The sliding filament model of muscle contraction: I. Quantum mechanical formalism. Journal of Theoretical Biology, 69(1):167–185, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903940.

Gray:1977:SFMb

[GG77b] B. F. Gray and I. Gonda. The sliding filament model of muscle contraction: II. The energetic and dynamical predictions of a quantum mechanical transducer model. Journal of Theoretical Biology, 69(1):187–230, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903952.

Gupta:1975:MOS

[GGM75] S. P. Gupta, G. Govil, and R. K. Mishra. Molecular orbital studies on the conformation of phospholipids. II. Preferred conformations of hydrocarbon chains and molecular organization in biomembranes. Journal of Theoretical Biology, 51(1):13–34, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901368.

Gandolfi:1978:RBM

[GGS78] Alberto Gandolfi, Maria A. Giovenco, and Roberto Strom. Reversible binding of multivalent antigen in the control of b lymphocyte activation. Journal of Theoretical Biology, 74 (4):513–521, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902382.

Gordon:1972:RMS

[GGSW72] Richard Gordon, Narendra S. Goel, Malcolm S. Steinberg, and Lawrence L. Wiseman. A rheological mechanism sufficient to explain the kinetics of cell sorting. Journal of Theoretical Biology, 37(1):43–73, October 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901142.

Gibson:1975:PMA

[GH75] John M. Gibson and Henry R. Hirsch. Psychoneural models of the auditory masking process. Journal of Theoretical Biology, 51(1):135–147, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375901447.

Gupta:1976:NEI

[GH76] Prithwis Das Gupta and Lynn Hickman. A note on the effect of ignoring dropout cases on fecundability as estimated from a follow-up study. Journal of Theoretical Biology, 61(2):411–417, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900278.

Goel:1977:ECE

[GI77] Narendra S. Goel and Sirajul Islam. Error catastrophe in and the evolution of the protein synthesizing machinery. Journal of Theoretical Biology, 68(2):167–182, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901576.

Gilbert:1972:IMT

[Gil72] Peter Gilbert. Iterative methods for the three-dimensional reconstruction of an object from projections. Journal of Theoretical Biology, 36(1):105–117, July 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901804.

Gillespie:1973:TMTa

[Gil73a] C. J. Gillespie. Towards a molecular theory of the nerve membrane. Journal of Theoretical Biology, 40(3):409–428, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900015.

Gillespie:1973:TMTc

[Gil73b] C. J. Gillespie. Towards a molecular theory of the nerve membrane: Prediction of the maximum negative conductance. Journal of Theoretical Biology, 42(3):533–543, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902452.

Gillespie:1973:TMTb

[Gil73c] C. J. Gillespie. Towards a molecular theory of the nerve membrane: The significance of the quasithreshold behaviour. Journal of Theoretical Biology, 42(3):519–531, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902440.

Gilpin:1974:LFC

[Gil74] Michael E. Gilpin. A Liapunov function for competition communities. Journal of Theoretical Biology, 44(1):35–48, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800287.

Gillespie:1976:TMTb

[Gil76a] C. J. Gillespie. Towards a molecular theory of the nerve membrane: Inactivation. Journal of Theoretical Biology, 60(1):19– 35, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901533.

Gillespie:1976:TMTa

[Gil76b] C. J. Gillespie. Towards a molecular theory of the nerve membrane: The sufficiency of a single-ion queue. Journal of Theoretical Biology, 58(2):477–498, ???? 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801324.

Gingell:1971:CFE

[Gin71a] D. Gingell. Computed force and energy of membrane interaction. Journal of Theoretical Biology, 30(1):121–149, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900415.

Ginsburg:1971:MIO

[Gin71b] Hagai Ginsburg. Model for iso-osmotic water flow in plant roots. Journal of Theoretical Biology, 32(1):147–158, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901421.

Ginsburg:1972:APR

[Gin72] Hagai Ginsburg. Analysis of plant root electropotentials. Journal of Theoretical Biology, 37(3):389–412, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900823.

Gingell:1973:MPC

[Gin73] D. Gingell. Membrane permeability change by aggregation of mobile glycoprotein units. Journal of Theoretical Biology, 38(3):677–679, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937390266X.

Ginzburg:1977:ESA

[Gin77a] Lev R. Ginzburg. The equilibrium and stability for n alleles under the density-dependent selection. Journal of Theoretical Biology, 68(4):545–550, October 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901047.

Ginzburg:1977:MEN

[Gin77b] Lev R. Ginzburg. A macro-equation of natural selection. Journal of Theoretical Biology, 67(4):677–686, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902545.

Greig:1976:PRS

[GJ76] R. G. Greig and M. N. Jones. The possible role of steric forces in cellular cohesion. Journal of Theoretical Biology, 63(2):405– 419, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900436.

Goldman:1971:KBT

[GK71] Rachel Goldman and Ephraim Katchalski. Kinetic behavior of a two-enzyme membrane carrying out a consecutive set of reactions. Journal of Theoretical Biology, 32(2):243–257, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901639.

Glass:1972:COC

[GK72] Leon Glass and Stuart A. Kauffman. Co-operative components, spatial localization and oscillatory cellular dynamics. Journal of Theoretical Biology, 34(2):219–237, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901579.

George:1973:CTR

[GK73a] Jack M. George and Lemont B. Kier. Conformation of thyrotropin-releasing hormone: Prediction from molecular orbital theory. Journal of Theoretical Biology, 40(2):393–394, IN1, 395–396, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901409.

Glass:1973:LAC

[GK73b] Leon Glass and Stuart A. Kauffman. The logical analysis of continuous, non-linear biochemical control networks. Journal of Theoretical Biology, 39(1):103–129, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902087.

George:1974:MOS

[GK74] Jack M. George and Lemant B. Kier. Molecular orbital studies of the conformation around the 5 position in angiotensin II. Journal of Theoretical Biology, 46(1):111–115, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901428.

Gabrovec:1975:ASE

[GKP75] I. Gabrovec, M. Kordas, and B. Popović. An attempt at a simulation of the end-plate current. Journal of Theoretical Biology, 55(1):29–45, November 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580107X.

Gradisek:1978:MAF

[GKS78] A. Gradisek, M. Kordas, and S. Svetina. A mathematical analysis of factors determining the time-course of membrane conductance change evoked by electrophoretic drug application. Journal of Theoretical Biology, 71(3):311–322, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901625.

Golovinsky:1970:SSO

[GKYG70] Evgeny Golovinsky, Jose Kaneti, Ivan Yukhnovsky, and Dobri Genchev. A study of some orotic acid derivatives and analogues by Hückel’s method of molecular orbitals: II. Analogues of the orotic acid in the reaction with phosphoribosylpyrophosphate. Journal of Theoretical Biology, 26(1):29–32, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800297.

Goel:1970:SSA

[GL70] Narendra S. Goel and Ardean G. Leith. Self-sorting of anisotropic cells. Journal of Theoretical Biology, 28(3):469–482, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900810.

Grenetz:1973:MPG

[GL73] Philip S. Grenetz and Albert List. A model for predicting growth responses in plants to changes in external water potential: Zea mays primary roots. Journal of Theoretical Biology, 39(1):29– 45, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902038.

Grossberg:1975:SDA

[GL75] Stephen Grossberg and Daniel Levine. Some developmental and attentional biases in the contrast enhancement and short term memory of recurrent neural networks. Journal of Theoretical Biology, 53(2):341–380, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800099.

Garay:1978:KAI

[GL78] R. P. Garay and R. Lefever. A kinetic approach to the immunology of cancer: Stationary states properties of efffector-target cell reactions. Journal of Theoretical Biology, 73(3):417–438, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901509.

Goswami:1979:GPP

[GL79] Dilip Goswami and Avijit Lahiri. A generalized prey-predator model. Journal of Theoretical Biology, 79(2):243–246, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902492.

Glass:1975:CBN

[Gla75] Leon Glass. Classification of biological networks by their qualitative dynamics. Journal of Theoretical Biology, 54(1):85– 107, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800567.

Gladyshev:1978:TBE

[Gla78] G. P. Gladyshev. On the thermodynamics of biological evolution. Journal of Theoretical Biology, 75(4):425–441, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903545.

Glabe:1979:STS

[Gla79] Charles G. Glabe. A supramolecular theory for specificity in intracellular adhesion. Journal of Theoretical Biology, 78(3):417– 423, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903400.

Galzigna:1971:PCC

[GM71] Lauro Galzigna and Patrick Maurel. Physicochemical characteristics of central neurotransmitters. Journal of Theoretical Biology, 31(3):553–554, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900282.

Glick:1972:LMI

[GM72] J. Leslie Glick and Alaknanda Majumdar. A “loop” model for integration of donor DNA into host DNA of eukaryote cells. Journal of Theoretical Biology, 36(3):503–512, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900057.

Gainey:1975:CCM

[GMR75] Ralph Gainey, Marvin Malone, and Manfred Rimpler. Chitin as a causative mechanism for inflammatory disease. Journal of Theoretical Biology, 52(1):249–250, July 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900557.

Gurney:1975:RIP

[GN75] W. S. C. Gurney and R. M. Nisbet. The regulation of inhomogeneous populations. Journal of Theoretical Biology, 52(2):441– 457, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900119.

Gurney:1976:NNL

[GN76a] W. S. C. Gurney and R. M. Nisbet. A note on non-linear population transport. Journal of Theoretical Biology, 56(1):249– 251, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800562.

Gurney:1976:SPM

[GN76b] W. S. C. Gurney and R. M. Nisbet. Spatial pattern and the mechanism of population regulation. Journal of Theoretical Biology, 59(2):361–370, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901764.

Gochlerner:1978:FOE

[Goc78] Galina B. Gochlerner. Free oxygen and evolutionary progress. Journal of Theoretical Biology, 75(4):467–486, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903570.

Goel:1974:CGE

[Goe74] Narendra S. Goel. Comments on “geometry and the exchange principle in cell aggregation kinetics”. Journal of Theoretical Biology, 46(2):585–586, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900198.

Gold:1976:CLS

[Gol76] Harvey J. Gold. Comment on the letter of A. N. Stokes on rapid oxygen entry into unperfused tissue slices. Journal of Theoretical Biology, 60(2):477, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900746.

Goodwin:1970:MBG

[Goo70] Brian C. Goodwin. Model of the bacterial growth cycle: Statistical dynamics of a system with asymptotic orbital stability. Journal of Theoretical Biology, 28(3):375–391, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900767.

Good:1972:HML

[Goo72a] Phillip Good. Homeostatic mechanisms and the loss of mitotic potential. Journal of Theoretical Biology, 34(1):99–102, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900574.

Good:1972:TAC

[Goo72b] R. J. Good. Theory of the adhesion of cells and the spontaneous sorting-out of mixed cell aggregates. Journal of Theoretical Biology, 37(3):413–434, December 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900835.

Good:1973:AMH

[Goo73a] Phillip I. Good. Alternative mechanisms for homeostatic regulation of mammalian cell populations. Journal of Theoretical Biology, 40(3):543–552, August 22, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390009X.

Good:1973:RWO

[Goo73b] W. Good. The role of water in the origin of life and its function in the primitive gene. Journal of Theoretical Biology, 39(2): 249–276, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900970.

Good:1977:SMV

[Goo77] Phillip I. Good. A stochastic model for in vitro ageing II. A theory of marginotomy. Journal of Theoretical Biology, 64(2):261–275, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903563.

Goon:1978:WRM

[Goo78] W. Goon. Water relations in the malonamide-induced haemolysis of mammalian erythrocytes. Journal of Theoretical Biology, 74(2):279–296, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900760.

Gordon:1978:ERE

[Gor78] Lyndsay G. M. Gordon. The electrical resistance of epithelia in the presence of osmotic and hydrostatic pressure gradients. Journal of Theoretical Biology, 72(3):545–550, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903168.

Gottfried:1972:PCD

[Got72] H. Gottfried. The prostatic cell DNA-steroid association: A hypothesis of nuclear events and cytoplasmic consequences. Journal of Theoretical Biology, 37(3):447–454, IN1, 455–475, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900859.

Gotoh:1975:MAP

[Got75] Hideki Gotoh. A model of the activation process of Na + conductance in the squid axon: An approach with interactive desorption kinetics of divalent cations. Journal of Theoretical Biology, 53(2):309–325, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800063.

Goulden:1976:BMP

[Gou76] Peter T. Goulden. The biological membrane potential some theoretical considerations. Journal of Theoretical Biology, 58(2):

425–438, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376801294.

Gingell:1972:CVW

[GP72] David Gingell and V. Adrian Parsegian. Computation of van der Waals interactions in aqueous systems using reflectivity data. Journal of Theoretical Biology, 36(1):41–52, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901750.

Giessner-Prettre:1971:INS

[GPB71] Claude Giessner-Prettre and Pullman Bernard. Intermolecular nuclear shielding due to the aromatic amino acids of proteins and to porphyrins. Journal of Theoretical Biology, 31(2):287– 294, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901883.

Giessner-Prettre:1970:RNV

[GPP70a] C. Giessner-Prettre and B. Pullman. Regions of negative values of intermolecular nuclear shielding for protons of purines. Journal of Theoretical Biology, 27(2):341–342, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901451.

Giessner-Prettre:1970:INS

[GPP70b] Claude Giessner-Prettre and Bernard Pullman. Intermolecular nuclear shielding values for protons of purines and flavins. Journal of Theoretical Biology, 27(1):87–95, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090130X.

Giessner-Prettre:1973:MOC

[GPP73] Claude Giessner-Prettre and Bernard Pullman. Molecular orbital calculations of the proton-proton coupling constants in nucleosides. Journal of Theoretical Biology, 40(3):441–444, IN1, 445–454, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900039.

Giessner-Prettre:1974:QMC

[GPP74] Claude Giessner-Prettre and Bernard Pullman. Quantum mechanical calculations of P–C and P–H spin-spin coupling constants in model compounds and in nucleotides. Journal of Theoretical Biology, 48(2):425–443, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937480010X.

Giessner-Prettre:1977:CDPa

[GPP77a] Claude Giessner-Prettre and Bernard Pullman. On the conformational dependence of the proton chemical shifts in nucleosides and nucleotides. I. Proton shifts in the ribose ring of pyrimidine nucleosides as a function of the torsion angle about the glycosyl bond. Journal of Theoretical Biology, 65(1):171–188, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900820.

Giessner-Prettre:1977:CDPb

[GPP77b] Claude Giessner-Prettre and Bernard Pullman. On the conformational dependence of the proton chemical shifts in nucleosides and nucleotides. II. Proton shifts in the ribose ring of purine nucleosides as a function of the torsion angle about the glycosyl bond. Journal of Theoretical Biology, 65(1):189–201, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900832.

Giessner-Prettre:1978:IRP

[GPP78] C. Giessner-Prettre and B. Pullman. An ab initio reinvestigation of PC and PH coupling constants in phosphates. Journal of Theoretical Biology, 72(4):751–756, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902825.

Gray:1973:HRL

[GR73] Gary D. Gray and Harold E. Renis. A hypothesis relating lymphocyte phosphorylation and transport of ara-c (cytarabine) to its antiviral activity. Journal of Theoretical Biology, 39(3):623– 632, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900738.

Grest:1974:IEE

[GR74] G. S. Grest and A. K. Rajagopal. Impurity and end effects on finite polymer chains — Ising model approach. Journal of Theoretical Biology, 47(1):43–56, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900976.

Green:1975:CAM

[GR75] J. R. Green and P. E. B. Reilly. Choosing an appropriate multiexponential model. Journal of Theoretical Biology, 55(1):137– 143, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801123.

Goel:1978:CSE

[GR78] Narendra S. Goel and Gary Rogers. Computer simulation of engulfment and other movements of embryonic tissues. Journal of Theoretical Biology, 71(1):103–140, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902151.

Granlund:1973:UDF

[Gra73] Goesta H. Granlund. The use of distribution functions to describe integrated density profiles of human chromosomes. Journal of Theoretical Biology, 40(3):573–589, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900118.

Gradmann:1977:PTP

[Gra77] Dietrich Gradmann. Potassium and turgor pressure in plants. Journal of Theoretical Biology, 65(3):597–599, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902181.

Goel:1972:DSM

[GRDC72] Narendra S. Goel, Nira Richter-Dyn, and John R. Clay. Discrete stochastic models for firing of a neuron. Journal of Theoretical Biology, 34(1):155–184, January 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900616.

Green:1976:SBM

[Gre76a] D. E. Green. The structure of biological membranes in relation to the principle of energy coupling. Journal of Theoretical Biology, 62(2):271–285, October 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690120X.

Greenspan:1976:GSC

[Gre76b] H. P. Greenspan. On the growth and stability of cell cultures and solid tumors. Journal of Theoretical Biology, 56(1):229– 242, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800549.

Green:1977:GCC

[Gre77a] Paul B. Green. On growth and the cell cycle in roots. Journal of Theoretical Biology, 64(3):597–598, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902910.

Greenspan:1977:DCC

[Gre77b] H. P. Greenspan. On the dynamics of cell cleavage. Journal of Theoretical Biology, 65(1):79–99, March 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900789.

Greenspan:1978:FMS

[Gre78] H. P. Greenspan. On fluid-mechanical simulations of cell division and movement. Journal of Theoretical Biology, 70 (1):125–128, IN1–IN2, 129–130, IN3, 131–134, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903053.

Grisell:1977:MPC

[Gri77] Ronald D. Grisell. Models of potassium conduction in excitable membrane examined by spectral analysis. Journal of Theoretical Biology, 66(3):399–436, June 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902934.

Grisell:1979:TMM

[Gri79] R. D. S. Grisell. Toward a multi-membrane model for potassium conduction in squid giant axon. Journal of Theoretical Biology, 76(3):233–266, February 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903126.

Grossberg:1970:NPD

[Gro70] Stephen Grossberg. Neural pattern discrimination. Journal of Theoretical Biology, 27(2):291–337, May 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901438.

Grossberg:1971:DOC

[Gro71] Stephen Grossberg. On the dynamics of operant conditioning. Journal of Theoretical Biology, 33(2):225–255, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900646.

Grossberg:1978:DPO

[Gro78] Stephen Grossberg. Decisions, patterns, and oscillations in nonlinear competitive systems with applications to Volterra–Lotka systems. Journal of Theoretical Biology, 73(1):101–130, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901820.

Goel:1972:MCC

[GRY^{+}72] Narendra S. Goel, Gita Subba Rao, Martynas Ycas, Hans J. Bremermann, and Lucy King. A method for calculating codon frequencies in DNA. Journal of Theoretical Biology, 35(3):399–420, IN1–IN2, 421–457, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901439.

Govil:1971:QCSa

[GS71a] Girjesh Govil and Anil Saran. Quantum chemical studies on the conformational structure of nucleic acids I. Extended Hückel calculations on d-ribose phosphate. Journal of Theoretical Biology, 30(3):621–630, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900142.

Govil:1971:QCSb

[GS71b] Girjesh Govil and Anil Saran. Quantum chemical studies on the conformational structure of nucleic acids II. EHT and CNDO calculations on the puckering of d-ribose. Journal of Theoretical Biology, 33(2):399–406, November 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900737.

Georgopapadakou:1977:BBE

[GS77] N. H. Georgopapadakou and A. I. Scott. On B_{12} biosynthesis and evolution. Journal of Theoretical Biology, 69(2):381–384, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937790145X.

Gani:1978:NAM

[GS78a] J. Gani and I. Saunders. Nuclear aggregation: Modelling and simulation. Journal of Theoretical Biology, 72(1):81–90, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890019X.

Ghosh:1978:MTD

[GS78b] P. K. Ghosh and D. Sengupta. A model for the temperature dependence of membrane excitability. Journal of Theoretical Biology, 73(4):609–614, August 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890125X.

Gressel:1978:PPE

[GS78c] J. Gressel and L. A. Segel. The paucity of plants evolving genetic resistance to herbicides: Possible reasons and implications. Journal of Theoretical Biology, 75(3):349–371, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903405.

Gero:1977:UBS

[GT77] Alexander Gero and Ronald J. Tallarida. Use of the biological stimulus in determining parameters of drug action, and its relationship to the drug effect a contribution to the theory of drug action. Journal of Theoretical Biology, 69(2):265–274, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901357.

Gunther:1977:BTT

[Gün77] B. Günther. Body temperature: Thermodynamics of homeothermism. Conjectural effects of high altitudes. Journal of Theoretical Biology, 66(3):593–594, June 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903058.

Gupta:1975:SFR

[Gup75] Prithwis Das Gupta. Some further results from a stochastic fertility model for women. Journal of Theoretical Biology, 50(2): 363–385, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900879.

Gurtin:1973:SEA

[Gur73] Morton E. Gurtin. A system of equations for age-dependent population diffusion. Journal of Theoretical Biology, 40(2):

389–392, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901392.

Gruenstein:1970:MMA

[GW70] Eric Gruenstein and James Wynn. A molecular mechanism of action of thyroxin: Modification of membrane phospholipid by iodine. Journal of Theoretical Biology, 26(3):343–363, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900883.

Grover:1977:ERS

[GWZR77] N. B. Grover, C. L. Woldringh, A. Zaritsky, and R. F. Rosenberger. Elongation of rod-shaped bacteria. Journal of Theoretical Biology, 67(2):181–193, July 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901928.

Goel:1975:ECH

[GY75] Narendra S. Goel and Martynas Ycas. The error catastrophe hypothesis with reference to aging and the evolution of the protein synthesizing machinery. Journal of Theoretical Biology, 55(1):245–282, November 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801184.

Goel:1979:CTS

[GY79] Narendra S. Goel and Martynas Ycas. On the computation of the tertiary structure of globular proteins II. Journal of Theoretical Biology, 77(3):253–305, April 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903576.

Grinberg-Zylberbaum:1976:RLI

[GZ76] J. Grinberg-Zylberbaum. The retrieval of learned information — a neurophysiological convergence-divergence theory. Journal of Theoretical Biology, 56(1):95–110, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800471.

Hartmann:1972:APO

[HA72] Å. O. Hartmann and H. T. Andersen. Analysis of the peripheral organization and the specificity properties of primary gustatory afferents. Journal of Theoretical Biology, 34(3):577–590, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901427.

Herman:1975:EF

[HA75] Gabor Herman and Michael A. Arbib. Editors’ foreword. Journal of Theoretical Biology, 54(1):1, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800506.

Haefner:1978:EAG

[Hae78] James W. Haefner. Ecosystem assembly grammars: Generative capacity and empirical adequacy. Journal of Theoretical Biology, 73(2):293–318, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901923.

Hageseth:1976:AEG

[Hag76] G. T. Hageseth. Activation energy and germination times for isothermal seed germination. Journal of Theoretical Biology, 59(1):247–249, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800392.

Haines:1979:PFU

[Hai79] Thomas H. Haines. A proposal on the function of unsaturated fatty acids and ionic lipids: the role of potential compaction in biological membranes. Journal of Theoretical Biology, 80(3):307–323, October 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900961.

Hall:1971:EP

[Hal71] John B. Hall. Evolution of the prokaryotes. Journal of Theoretical Biology, 30(3):429–454, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900014.

Hall:1973:NHO

[Hal73] John B. Hall. The nature of the host in the origin of the eukaryote cell. Journal of Theoretical Biology, 38(2):413–418, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901835.

Hamilton:1971:GSH

[Ham71] W. D. Hamilton. Geometry for the selfish herd. Journal of Theoretical Biology, 31(2):295–311, May 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901895.

Han:1972:NLA

[Han72] Moon H. Han. Non-linear Arrhenius plots in temperaturedependent kinetic studies of enzyme reactions: I. Single transition processes. Journal of Theoretical Biology, 35(3):543– 568, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901506.

Hannon:1973:SE

[Han73] Bruce Hannon. The structure of ecosystems. Journal of Theoretical Biology, 41(3):535–546, October 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390060X.

Hannon:1976:MPP

[Han76] Bruce Hannon. Marginal product pricing in the ecosystem. Journal of Theoretical Biology, 56(2):253–267, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800732.

Hanggi:1978:STK

[Han78] P. Hanggi. Stochastic theory for the kinetics of migration of ligands in biomolecules. Journal of Theoretical Biology, 74(3):337–359, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902187.

Hannon:1979:TEC

[Han79] Bruce Hannon. Total energy costs in ecosystems. Journal of Theoretical Biology, 80(2):271–293, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990211X.

Harington:1970:ACC

[Har70] J. S. Harington. Adaptation of cellular control processes in relation to cancer and its therapy. Journal of Theoretical Biology, 28 (1):31–45, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900627.

Harper:1971:KTS

[Har71]  Edwin T. Harper.  Kinetics of the two-sited enzyme:

I. Activation and inhibition by substrate. Journal of Theoretical Biology, 32(2):405–414, August 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901767.

Harper:1973:KTS

[Har73a] Edwin T. Harper. Kinetics of the two-sited enzyme: II. A method of distinguishing between anticooperative and independent active sites based on competitive inhibition. Journal of Theoretical Biology, 39(1):91–102, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902075.

Harrison:1973:EBS

[Har73b] L. G. Harrison. Evolution of biochemical systems with specific chiralities: A model involving territorial behaviour. Journal of Theoretical Biology, 39(2):333–341, May 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901033.

Harris:1974:BFI

[Har74] William Frith Harris. The bacterial flagellum as an imperfect cylindrical crystal: Flagellar geometry, movement and polymorphism, and the role of partial dislocations. Journal of Theoretical Biology, 47(2):295–296, IN1, 297–308, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901994.

Hartman:1975:CC

[Har75] Hyman Hartman. The centriole and the cell. Journal of Theoretical Biology, 51(2):501–509, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900776.

Harris:1976:CSC

[Har76a] Albert K. Harris. Is cell sorting caused by differences in the work of intercellular adhesion? A critique of the Steinberg hypothesis. Journal of Theoretical Biology, 61(2):267–285, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900199.

Harris:1976:LTV

[Har76b] R. J. C. Harris. Latency and the tumour virus problem. Journal of Theoretical Biology, 62(2):287–298, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901211.

Hartung:1977:IAH

[Har77] John Hartung. An implication about human mating systems. Journal of Theoretical Biology, 66(4):737–745, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902429.

Harvey:1978:ITI

[Har78] R. J. Harvey. Interaction of two inhibitors which act on different enzymes of a metabolic pathway. Journal of Theoretical Biology, 74(3):411–437, October 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902230.

Hasofer:1974:PMF

[Has74] A. M. Hasofer. A probability model for the fusion of nucleoli in metabolic cells. Journal of Theoretical Biology, 45(2):305– 310, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901167.

Hassard:1978:BPS

[Has78a] Brian Hassard. Bifurcation of periodic solutions of the Hodgkin–Huxley model for the squid giant axon. Journal of Theoretical Biology, 71(3):401–420, April 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901686.

Hasstedt:1978:SMC

[Has78b] Sandra Hasstedt. A stochastic model for a closed biochemical system at equilibrium. Journal of Theoretical Biology, 70(2):199–212, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378903478.

Hastings:1978:ESS

[Has78c] Alan Hastings. Evolutionarily stable strategies and the evolution of life history strategies: I. Density dependent models. Journal of Theoretical Biology, 75(4):527–536, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903612.

Hastings:1978:EOP

[Has78d] Alan Hastings. An evolutionary optimization principle. Journal of Theoretical Biology, 75(4):519–525, December 21, 1978.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903600.

Hastings:1979:SCC

[Has79] Harold M. Hastings. Stability considerations in community organization. Journal of Theoretical Biology, 78(1):121–127, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903308.

Hatze:1973:TCM

[Hat73] Herbert Hatze. A theory of contraction and a mathematical model of striated muscle. Journal of Theoretical Biology, 40(2):219–246, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937390129X.

Hawkins:1976:RSS

[Haw76] Shirley E. Hawkins. Recent studies of some RNAs and proteins in amoebae. Journal of Theoretical Biology, 62(2):299–313, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901223.

Hayhoe:1974:SER

[Hay74] Henry N. Hayhoe. Shunting effect on resistance measurements in visceral smooth muscle tissue. Journal of Theoretical Biology, 46(2):295–305, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900010.

Hamann:1970:SPM

[HB70] J. R. Hamann and L. M. Bianchi. Stochastic population mechanics in the relational systems formalism: Volterra–Lotka ecological dynamics. Journal of Theoretical Biology, 28(2):175– 184, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900500.

Hirsch:1973:DGM

[HC73] Henry R. Hirsch and Howard J. Curtis. Dynamics of growth in mammalian diploid tissue cultures. Journal of Theoretical Biology, 42(2):227–244, November 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900878.

Harrison:1975:SDEa

[HC75a] Lester I. Harrison and Halvor N. Christensen. Simulation of differential effects on rates in membrane transport. Journal of Theoretical Biology, 49(1):439–459, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800476.

Harrison:1975:SDEb

[HC75b] Lester I. Harrison and Halvor N. Christensen. Simulation of differential effects on rates in membrane transport. Journal of Theoretical Biology, 49(2):439–459, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590185X.

Hukins:1976:SHC

[HC76] David W. L. Hukins and John A. Chapman. Structural hierarchy in collagenous tissues and the “strait-jacket” theory of bone stiffness. Journal of Theoretical Biology, 57(1):247–248, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800185.

Hay:1977:MVT

[HC77] George A. Hay and M. Susan Chesters. A model of visual threshold detection. Journal of Theoretical Biology, 67(2):221– 240, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901965.

Hill:1978:SSK

[HC78] Terrell L. Hill and Britton Chance. Steady-state kinetics of models of respiratory chain enzymes with isopotential pools and conformational site enzymes. Journal of Theoretical Biology, 72(1):

17–56, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378900164.

Harth:1970:BFN

[HCBL70] E. M. Harth, T. J. Csermely, B. Beek, and R. D. Lindsay. Brain functions and neural dynamics. Journal of Theoretical Biology, 26(1):93–120, January 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800352.

Hearne:1978:MBA

[HDP^{+}78] E. M. Hearne, J. W. Drane, A. Pierce, G. Harris, and W. G. Johanson. A model for bacterial activity in the lung. Journal of Theoretical Biology, 72(2):205–218, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900899.

Hansell:1973:DEC

[HE73] Roger I. C. Hansell and Bland Ewing. The detection and estimation of character weighting in classifications. Journal of Theoretical Biology, 39(2):297–314, May 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900994.

Heath:1974:UHR

[Hea74] I. Brent Heath. A unified hypothesis for the role of membrane bound enzyme complexes and microtubules in plant cell wall synthesis. Journal of Theoretical Biology, 48(2):445–449, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800111.

Heath:1977:SHC

[Hea77] D. J. Heath. Simultaneous hermaphroditism; cost and benefit. Journal of Theoretical Biology, 64(2):363–373, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903630.

Heath:1979:BEH

[Hea79] D. J. Heath. Brooding and the evolution of hermaphroditism. Journal of Theoretical Biology, 81(2):151–155, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901577.

Hedges:1971:HCR

[Hed71] R. W. Hedges. Hypothesis for catabolite repression. Journal of Theoretical Biology, 30(2):297–305, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900567.

Hejnowicz:1975:MMM

[Hej75] Zygmunt Hejnowicz. A model for morphogenetic map and clock. Journal of Theoretical Biology, 54(2):345–362, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801366.

Heller:1978:TPP

[Hel78] Rolf Heller. Two predator–prey difference equations considering delayed population growth and starvation. Journal of Theoretical Biology, 70(4):401–413, February 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902497.

Herman:1970:RED

[Her70] G. T. Herman. Role of environment in developmental models. Journal of Theoretical Biology, 29(3):329–341, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901025.

Heuch:1978:MBP

[Heu78] Ivar Heuch. Maintenance of butterfly populations with allfemale broods under recurrent extinction and recolonization. Journal of Theoretical Biology, 75(1):115–122, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902059.

Hewitt:1977:IPL

[Hew77] J. A. Hewitt. On the influence of polyvalent ligands on membrane curvature. Journal of Theoretical Biology, 64(3):455–472, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937790282X.

Hewitt:1978:DGM

[Hew78] J. A. Hewitt. Diffusion gradients, membrane receptors, and the acquisition of orientational information by cells. Journal of Theoretical Biology, 74(2):297–306, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900772.

Hewitt:1979:SRM

[Hew79] J. A. Hewitt. Surf-riding model for cell capping. Journal of Theoretical Biology, 80(1):115–127, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901838.

Hastings:1978:PTP

[HG78] David F. Hastings and John Gutknecht. Potassium and turgor pressure in plants. Journal of Theoretical Biology, 73(2):363– 366, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901972.

Hayes:1979:PIB

[HGR79] Pliny H. Hayes, Jack R. Girton, and Michael A. Russell. Positional information and the bithorax complex. Journal of Theoretical Biology, 79(1):1–17, July 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902534.

Heyde:1971:SFO

[HH71] C. C. Heyde and Elizabeth Heyde. Stochastic fluctuations in a one substrate one product enzyme system: Are they ever relevant? Journal of Theoretical Biology, 30(2):395–404, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900609.

Horwitz:1973:EDC

[HH73] B. A. Horwitz and L. P. Horwitz. Electronic distribution coding in cellular differentiation. Journal of Theoretical Biology, 42(1):169–176, November 5, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901550.

Hudson:1977:LMC

[HH77a] H. M. Hudson and G. M. Hahn. The labelled mitosis curve for a population consisting of fast and slowly cycling cells. Journal of Theoretical Biology, 66(1):63–79, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903125.

Huf:1977:CMA

[HH77b] Ernst G. Huf and John R. Howell. Computer modelling. Application to studies on the initial rate of Na + uptake by frog skin epidermis. Journal of Theoretical Biology, 65(4):653–669, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900145.

Huf:1977:GDC

[HH77c] Ernst G. Huf and John R. Howell. Global Na + diffusion coefficient in a multicompartment frog skin epidermis model. Journal of Theoretical Biology, 68(1):161–165, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790234X.

Hill:1970:SFM

[Hil70] Terrell L. Hill. Sliding filament model of muscular contraction: V. Isometric force and interfilament spacing. Journal of Theoretical Biology, 29(3):395–410, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901050.

Hill:1972:OFS

[Hil72a] A. E. Hill. Osmotic flow and solute reflection zones. Journal of Theoretical Biology, 36(2):255–270, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900963.

Hill:1972:THO

[Hil72b] A. E. Hill. A theory of hydro-osmotic hormone action. Journal of Theoretical Biology, 36(2):271–281, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900975.

Hill:1973:FPC

[Hil73] L. R. Hill. A fault in Pratt’s critique of numerical taxonomy? Journal of Theoretical Biology, 40(2):397–398, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901410.

Hill:1978:EEE

[Hil78a] Terrell L. Hill. Effect of enzyme-enzyme interactions on steadystate enzyme kinetics. IV. “strictly steady-state” examples. Journal of Theoretical Biology, 75(4):391–416, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903521.

Hiller:1978:CDA

[Hil78b] Howard L. Hiller. Codon degeneracy and amino acid composition. Journal of Theoretical Biology, 74(4):599–601, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890245X.

Hines:1977:CES

[Hin77] W. G. S. Hines. Competition with an evolutionary stable strategy. Journal of Theoretical Biology, 67(1):141–153, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901898.

Hines:1978:MSS

[Hin78] W. G. S. Hines. Mutations and stable strategies. Journal of Theoretical Biology, 72(3):413–428, June 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903090.

Hitchins:1976:SPT

[Hit76]  Anthony D. Hitchins.  Spores as potential topological

markers in bacterial DNA segregation studies. Journal of Theoretical Biology, 58(1):97–111, May 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901417.

Hanna:1970:SAN

[HJ70] G. R. Hanna and R. N. Johnson. Simulation of auditory nerve responses by a neural model. Journal of Theoretical Biology, 26(3):501–505, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900998.

Hageseth:1975:KTI

[HJ75] G. T. Hageseth and R. D. Joyner. Kinetics and thermodynamics of isothermal seed germination. Journal of Theoretical Biology, 53(1):51–65, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901022.

Hjorth:1970:CGD

[Hjo70] Ingemar Hjorth. A comment on graphic displays of bird sounds and analyses with a new device, the melograph mona. Journal of Theoretical Biology, 26(1):1–4, IN1, 5–6, IN3–IN4, 7– 10, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800261.

Holtje:1974:SAS

[HK74] Hans-Dieter Höltje and Lemont B. Kier. Structure-activity studies of enzyme substrates using model interaction calculations. Journal of Theoretical Biology, 48(1):197–205, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901908.

Hall:1976:MDR

[HK76] Lowell H. Hall and Lemont B. Kier. A model for drug-receptor remote recognition applied to certain hydrogen bonding systems. Journal of Theoretical Biology, 58(1):177–195, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901466.

Ho:1977:KDM

[HK77a] S. P. Ho and M. D. Kostin. Kinetics and diffusion in multienzyme systems. Journal of Theoretical Biology, 64(3):421–427, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902806.

Hulbert:1977:CC

[HK77b] Matthew H. Hulbert and Steven Krawiec. Cometabolism: A critique. Journal of Theoretical Biology, 69(2):287–291, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901370.

Hideshima:1977:AON

[HKAT77] T. Hideshima, H. Kimizuka, L. G. Abood, and R. Tanaka. ATP and other nucleotide interaction with model compounds of amino acids. Journal of Theoretical Biology, 65(1):15–27, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900765.

Hovinen:1976:TSD

[HKK76] E. Hovinen, M. Kekki, and S. Kuikka. A theory to the stochastic dynamic model building for chronic progressive disease processes with an application to chronic gastritis. Journal of Theoretical Biology, 57(1):131–152, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800094.

Haynes:1979:SLR

[HKM79] Duncan H. Haynes, Michael A. Kolber, and Stephen J. Morris. Short and long-range forces involved in cation-induced aggregation of chromaffin granule membranes. Journal of Theoretical Biology, 81(4):713–743, December 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902789.

Hoshiko:1970:MDI

[HL70] T. Hoshiko and B. D. Lindley. Macroscopic definition of ideal behavior of tracers. Journal of Theoretical Biology, 26(2):315– 320, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800200.

Hanig:1979:TBT

[HL79] Joseph P. Hanig and Carl Lamanna. Toxicity of botulinum toxin: A stoichiometric model for the locus of its extraordinary potency and persistence at the neuromuscular junction. Journal of Theoretical Biology, 77(1):107–113, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901413.

Harth:1975:ETD

[HLC75] E. Harth, N. S. Lewis, and T. J. Csermely. The escape of tritonia: Dynamics of a neuromuscular control mechanism. Journal of Theoretical Biology, 55(1):201–228, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801160.

Herman:1973:AMA

[HLR73] Gabor T. Herman, Arnold Lent, and Stuart W. Rowland. ART: Mathematics and applications: A report on the mathematical foundations and on the applicability to real data of the algebraic reconstruction techniques. Journal of Theoretical Biology, 42(1):1–18, IN1–IN2, 19–32, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901458.

Herman:1974:AMA

[HLR74] G. T. Herman, A. Lent, and S. W. Rowland. ART: Mathematics and applications a report on the mathematical foundations and on the applicability to real data of the algebraic reconstruction techniques. Journal of Theoretical Biology, 45(2):597, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901350.

Hamilton-Miller:1971:UMM

[HM71a] J. M. T. Hamilton-Miller. Use of Michaelis–Menten kinetics in the analysis of synergism between β-lactam antibiotics. Journal of Theoretical Biology, 31(2):171–176, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901810.

Holwill:1971:HAN

[HM71b] M. E. J. Holwill and C. A. Miles. Hydrodynamic analysis of non-uniform flagellar undulations. Journal of Theoretical Biology, 31(1):25–42, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901184.

Huang:1972:MPA

[HM72] Huey-Wen Huang and Harold J. Morowitz. A method for phenomenological analysis of ecological data. Journal of Theoretical Biology, 35(3):489–503, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901476.

Horiike:1979:CBB

[HM79] Kihachiro Horiike and Donald B. McCormick. Correlations between biological activity and the number of functional groups chemically modified. Journal of Theoretical Biology, 79(3):403– 414, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903552.

Hall:1975:CEE

[HMS75] George G. Hall, Colin J. Miller, and Gary W. Schnuelle. Conformation energies and electronic structure of phenethylamine and amphetamine. Journal of Theoretical Biology, 53(2):475– 480, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800178.

Honda:1979:EEC

[HMT79] Hisao Honda, Toshiteru Morita, and Akira Tanabe. Establishment of epidermal cell columns in mammalian skin: Computer simulation. Journal of Theoretical Biology, 81(4):745–759, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902790.

Hettiaratchi:1974:MMP

[HO74] D. R. P. Hettiaratchi and J. R. O’Callaghan. A membrane model of plant cell extension. Journal of Theoretical Biology, 45(2): 459–465, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901246.

Harris:1976:PMS

[HO76] G. Harris and I. Olsen. A possible mechanism for the synthesis of antibodies. Journal of Theoretical Biology, 58(2):417–423, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801282.

Hettiaratchi:1978:SMP

[HO78] D. R. P. Hettiaratchi and J. R. O’Callaghan. Structural mechanics of plant cells. Journal of Theoretical Biology, 74(2):235–257, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378900747.

Hodgson:1971:PRA

[Hod71a] B. Hodgson. Possible roles for antibiotics and other biologically active peptides at specific stages during sporulation of bacillaceae. Journal of Theoretical Biology, 30(1):111–119, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900403.

Hodson:1971:WCS

[Hod71b] Stuart Hodson. Why the cornea swells. Journal of Theoretical Biology, 33(3):419–427, December 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900907.

Hofer:1971:MMU

[Höf71] Milan Höfer. A model of the monosaccharide uphill transporting cell membrane system in yeast. Journal of Theoretical Biology, 33(3):599–603, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900993.

Holler:1971:ETE

[Hol71] E. Holler. Enhancement of tryptic esterase activity. Journal of Theoretical Biology, 32(2):415–424, August 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901779.

Holland:1972:DSP

[Hol72] B. W. Holland. Dynamic specificity and the pairing of homologous chromosomes in meiosis. Journal of Theoretical Biology, 35(2):395–397, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900471.

Holden:1976:FNS

[Hol76] A. V. Holden. Flicker noise and structural changes in nerve membrane. Journal of Theoretical Biology, 57(1):243–246, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800173.

Holder:1979:PIP

[Hol79] Nigel Holder. Positional information and pattern formation in plant morphogenesis and a mechanism for the involvement of plant hormones. Journal of Theoretical Biology, 77(2):195–212, March 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903072.

Honda:1971:DFT

[Hon71] Hisao Honda. Description of the form of trees by the parameters of the tree-like body: Effects of the branching angle and the branch length on the shape of the tree-like body. Journal of Theoretical Biology, 31(2):331–334, IN1–IN6, 335–338, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901913.

Honda:1973:PFC

[Hon73] Hisao Honda. Pattern formation of the coenobial algae pediastrum biwae negoro. Journal of Theoretical Biology, 42(3):461–464, IN1–IN6, 465–481, December 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902415.

Honda:1978:DCP

[Hon78] Hisao Honda. Description of cellular patterns by Dirichlet domains: The two-dimensional case. Journal of Theoretical Biology, 72(3):523–530, IN4, 531–543, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903156.

Hahn:1973:COM

[HOR73] Hong-Sup Hahn, Peter J. Ortoleva, and John Ross. Chemical oscillations and multiple steady states due to variable boundary permeability. Journal of Theoretical Biology, 41(3):503– 521, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900581.

Horvath:1974:EAE

[Hor74]  Raymond S. Horvath.  Evolution of anaerobic energy-

yielding metabolic pathways of the procaryotes. Journal of Theoretical Biology, 48(2):361–371, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800068.

Howat:1972:PPV

[How72] M. G. Howat. Postulates for partitioning variance: I. In body movement among its origins. Journal of Theoretical Biology, 37(1):125–137, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901191.

Howland:1974:OSP

[How74] Howard C. Howland. Optimal strategies for predator avoidance: The relative importance of speed and manoeuvrability. Journal of Theoretical Biology, 47(2):333–350, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902021.

Howell:1977:WBH

[HP77] D. J. Howell and Joe Pylka. Why bats hang upside down: a biomechanical hypothesis. Journal of Theoretical Biology, 69(4):625–630, IN5, 631, December 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790371X.

Hunter:1979:BGS

[HPCH79] S. A. Hunter, R. A. Peura, T. C. Crusberg, and R. J. Harvey. Bond graph simulation of skeletal muscle glucose metabolism. Journal of Theoretical Biology, 78(4):499–508, IN1, 509–517, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901887.

Hendy:1978:IPT

[HPF78] M. D. Hendy, David Penny, and L. R. Foulds. Identification of phylogenetic trees of minimal length. Journal of Theoretical Biology, 71(3):441–452, April 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901716.

Ho:1976:PMS

[HPMH76] N. F. H. Ho, J. Park, W. Morozowich, and W. I. Higuchi. A physical model for the simultaneous membrane transport and metabolism of drugs. Journal of Theoretical Biology, 61 (1):185–193, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901120.

Hawkins:1971:SCF

[HR71a] R. Hawkins and Stuart A. Rice. Study of concentration fluctuations in model systems. Journal of Theoretical Biology, 30(3):579–596, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900117.

Herman:1971:RAE

[HR71b] Gabor T. Herman and Stuart Rowland. Resolution in ART: An experimental investigation of the resolving power of an algebraic picture reconstruction technique. Journal of Theoretical Biology, 33(1):213–216, IN3–IN9, 217–223, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902256.

Hall:1975:TST

[HR75] G. G. Hall and W. R. Rodwell. A theoretical study of tricycloquinazoline carcinogenesis. Journal of Theoretical Biology, 50(1):107–120, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900260.

Hillier:1979:CBA

[HR79] Ian H. Hillier and Barry Robson. Conformational behaviour of the architectural units of peptides and proteins. Assessment of current understanding by ab initio quantum mechanical methods and refinement of the dipeptide energy surface. Journal of Theoretical Biology, 76(1):83–98, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990376X.

Harris:1970:CCC

[HS70] W. F. Harris and L. E. Scriven. Cylindrical crystals, contractile mechanisms of bacteriophages and the possible role of dislocations in contraction. Journal of Theoretical Biology, 27(2):233–238, IN1, 239–257, May 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901402.

Herman:1975:SMG

[HS75] Gabor T. Herman and Gordon L. Schiff. Simulation of multigradient models of organisms in the context of l-systems. Journal of Theoretical Biology, 54(1):35–46, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800531.

Hess:1976:ASP

[HS76] B. A. Hess and L. J. Schaad. The aromatic stabilization of porphyrins. Journal of Theoretical Biology, 59(1):237–240, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800367.

Ho:1977:PCG

[HS77a]  S. V. Ho and M. L. Shuler.  Predictions of cellular

growth patterns by a feedback model. Journal of Theoretical Biology, 68(3):415–435, October 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900698.

Homer:1977:UTE

[HS77b] L. D. Homer and Alfred Small. A unified theory for estimation of cardiac output, volumes of distribution and renal clearances from indicator dilution curves. Journal of Theoretical Biology, 64(3):535–550, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902867.

Honig:1978:DPA

[HS78a] Barry Honig and Wilfred D. Stein. Design principles for active transport systems. Journal of Theoretical Biology, 75 (3):299–305, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378903363.

Hoshikawa:1978:CRC

[HS78b] Hiroshi Hoshikawa and Nobuhiko Saitô. On the chemical reactions for chemotaxis in bacteria. Journal of Theoretical Biology, 70(1):1–15, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902989.

Hines:1979:GBR

[HS79a] W. G. S. Hines and J. Maynard Smith. Games between relatives. Journal of Theoretical Biology, 79(1):19–30, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902546.

Ho:1979:BND

[HS79b] M. W. Ho and P. T. Saunders. Beyond neo-Darwinism — an epigenetic approach to evolution. Journal of Theoretical Biology, 78(4):573–591, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901917.

Hamilton:1973:CDE

[HSC73] Bruce K. Hamilton, Larry J. Stockmeyer, and Clark K. Colton. Comments on diffusive and electrostatic effects with immobilized enzymes. Journal of Theoretical Biology, 41(3):547–560, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900611.

Hlavacek:1972:MDE

[HSK72] V. Hlavácek, J. Sinkule, and M. Kubícek. A method for determining effects of diffusion on antocatalytic chemical reactions. Journal of Theoretical Biology, 36(2):283–290, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900987.

Hofbauer:1979:NES

[HSS79] J. Hofbauer, P. Schuster, and K. Sigmund. A note on evolutionary stable strategies and game dynamics. Journal of Theoretical Biology, 81(3):609–612, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900584.

Harris:1973:EPS

[HT73] Curtis E. Harris and David C. Teller. Estimation of primary sequence homology from amino acid composition of evolutionary related proteins. Journal of Theoretical Biology, 38(2):347– 362, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901793.

Ho:1972:SAS

[HTSH72] N. F. H. Ho, J. Turi, C. Shipman, and W. I. Higuchi. Systems approach to the study of drug transport across membranes using suspension cultures of mammalian cells: I. Theoretical diffusion models. Journal of Theoretical Biology, 34(3):451–452, IN3, 453–467, March 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290135X.

Huang:1971:KTAa

[Hua71a] Huey Wen Huang. Kinetic theory of antibiotic ion carrier. Journal of Theoretical Biology, 32(2):351–361, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190172X.

Huang:1971:KTAb

[Hua71b] Huey Wen Huang. Kinetic theory of antibiotic ion carrier. II. Journal of Theoretical Biology, 32(2):363–371, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901731.

Huang:1973:KTA

[Hua73a] Huey Wen Huang. Kinetic theory of antibiotic ion carrier. III. Journal of Theoretical Biology, 38(1):191–196, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902348.

Huang:1973:MDP

[Hua73b] Huey Wen Huang. Mobility and diffusion in the plane of cell membrane. Journal of Theoretical Biology, 40(1):11–17, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901616.

Huang:1977:TAT

[Hua77] Huey W. Huang. Thermodynamics of allosteric transitions. Journal of Theoretical Biology, 67(3):557–565, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900571.

Huemer:1977:TDO

[Hue77] Richard P. Huemer. A theory of diagnosis for orthomolecular medicine. Journal of Theoretical Biology, 67(4):625–635, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902508.

Hukins:1978:BSE

[Huk78] David W. L. Hukins. Bone stiffness explained by the liquid crystal model for the collagen fibril. Journal of Theoretical Biology, 71(4):661–667, April 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903302.

Hulme:1971:SBS

[Hul71] E. C. Hulme. Simulation of biochemical systems. Journal of Theoretical Biology, 31(1):131–137, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901263.

Hulbert:1978:THT

[Hul78] A. J. Hulbert. The thyroid hormones a thesis concerning their action. Journal of Theoretical Biology, 73(1):81–100, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901819.

Hunter:1974:MAS

[Hun74] John E. Hunter. A model of activity I and sleep. Journal of Theoretical Biology, 46(2):481–499, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900101.

Hutcheson:1970:TCD

[Hut70] Kermit Hutcheson. A test for comparing diversities based on the Shannon formula. Journal of Theoretical Biology, 29(1):151– 154, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901244.

Hutchison:1973:MEE

[Hut73] H. Terry Hutchison. A model for estimating the extent of variegation in mosaic tissues. Journal of Theoretical Biology, 38(1):61–79, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902257.

Hvidt:1975:DPV

[Hvi75] Aase Hvidt. A discussion of pressure-volume effects in aqueous protein solutions. Journal of Theoretical Biology, 50(1):245– 252, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900351.

Hutchinson:1976:PRF

[HWD76] D. W. Hutchinson, R. Whittenbury, and H. Dalton. A possible role of free radicals in the oxidation of methane by methylococcus capsulatus. Journal of Theoretical Biology, 58(2):325– 335, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376801233.

Hyver:1972:IEU

[Hyv72] C. Hyver. Impossibility of existence of undamped oscillations in linear chemical systems. Journal of Theoretical Biology, 36(1): 133–138, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901828.

Hyver:1973:VPS

[Hyv73] C. Hyver. Valeurs propres des systèmes de transformation représentables par des graphes en arbres. Journal of Theoretical Biology, 42(3):397–409, December 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902373.

Hartman:1971:ITN

[HZ71] Paul A. Hartman and Richard Ziprin. Immunological templates: A new model and proposed mechanism of immunogenicity. Journal of Theoretical Biology, 33(3):539–556, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900956.

Hobza:1977:ESE

[HZ77] Pavel Hobza and Rudolf Zahradník. Electronic structure of established and potential oxidizing agents in biological systems. Journal of Theoretical Biology, 66(3):461–474, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790296X.

Ichikawa:1979:ADC

[IC79] Masahiro Ichikawa and Tsu-Wei Chou. Analysis of deformation and contraction in cylindrical crystals due to dispirations or dislocations. Journal of Theoretical Biology, 78(1):129–142, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990331X.

Ijiri:1976:MMG

[IE76] K. Ijiri and N. Egami. A mathematical model for germ cell determination process and effect of ultraviolet light on the process. Journal of Theoretical Biology, 58(1):15–16, IN1–IN2, 17– 32, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901363.

Imamura:1972:MOS

[IFN72] Akira Imamura, Hiroko Fujita, and Chikayoshi Nagata. A molecular orbital study on the alkylation and depurination reactions of the base components of nucleic acids. Journal of Theoretical Biology, 36(3):447–457, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900021.

Irimajiri:1979:DTM

[IHI79] Akihiko Irimajiri, Tetsuya Hanai, and Akira Inouye. A dielectric theory of “multi-stratified shell” model with its application to a lymphoma cell. Journal of Theoretical Biology, 78(2):251– 269, May 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902686.

Ishida:1975:NET

[IM75] Kenzi Ishida and Shigeru Matsumoto. Non-equilibrium thermodynamics of temporally oscillating chemical reactions. Journal of Theoretical Biology, 52(2):343–363, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900065.

Imai:1978:KEP

[IM78] Hirotami T. Imai and Takeo Maruyama. Karyotype evolution by pericentric inversion as a stochastic process. Journal of Theoretical Biology, 70(3):253–261, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903752.

Imai:1975:ENR

[Ima75] Hirotami T. Imai. Evidence for non-random localization of the centromere on mammalian chromosomes. Journal of Theoretical Biology, 49(1):111–123, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800221.

Imai:1976:FEB

[Ima76] Hirotami T. Imai. Further evidence and biological significance for non-random localization of the centromere on mammalian chromosomes. Journal of Theoretical Biology, 61(1):195–203, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901132.

Imai:1978:OTC

[Ima78] Hirotami T. Imai. On the origin of telocentric chromosomes in mammals. Journal of Theoretical Biology, 71(4):619–637, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903284.

Imai:1979:FAB

[IMMY79] Yusuke Imai, Hirohiko Mori, Masataka Murakami, and Hideyo Yoshida. Flow analysis in a biological system adopting the buffer capacitor concept. Journal of Theoretical Biology, 77(1):115– 128, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901425.

Irving:1971:MPF

[IP71] D. Irving and P. Proctor. Model of psychological functioning. Journal of Theoretical Biology, 33(1):149–170, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902220.

Idso:1979:RSM

[IPH^{+}79] S. B. Idso, P. J. Pinter, J. L. Hatfield, R. D. Jackson, and R. J. Reginato. A remote sensing model for the prediction of wheat yields prior to harvest. Journal of Theoretical Biology, 77(2):217–228, March 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903096.

Ishida:1973:NET

[Ish73] Kenzi Ishida. Non-equilibrium thermodynamics of marginaland conditional-open chemical systems. Journal of Theoretical Biology, 40(2):301–327, August 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901343.

Ishida:1977:NET

[Ish77] Kenzi Ishida. Non-equilibrium thermodynamics of temporally oscillating chemical reactions: The case of limit cycle. Journal of Theoretical Biology, 65(1):1–13, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900753.

Israelachvili:1973:VWF

[Isr73] J. N. Israelachvili. Van der Waals forces involving thin rods. Journal of Theoretical Biology, 42(3):411–417, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902385.

Itzkowitz:1979:OCS

[Itz79] Murray Itzkowitz. On the organization of courtship sequences in fishes. Journal of Theoretical Biology, 78(1):21–28, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903229.

Iwasa:1978:ODS

[Iwa78] Yoh Iwasa. Optimal death strategy of animal populations. Journal of Theoretical Biology, 72(4):611–626, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902758.

Jost:1975:GRMa

[JA75a] Jean-Pierre Jost and Melvin Averner. Gene regulation in mammalian cells: A model for the interaction of steroids and 3’,5’- cyclic AMP. Journal of Theoretical Biology, 49(1):337–344, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800385.

Jost:1975:GRMb

[JA75b] Jean-Pierre Jost and Melvin Averner. Gene regulation in mammalian cells: A model for the interaction of steroids and 3’,5’- cyclic AMP. Journal of Theoretical Biology, 49(2):337–344, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901769.

Jackson:1977:ETW

[Jac77] Michael J. Jackson. Epithelial transport of weak electrolytes. Properties of a model of the three compartment system. Journal of Theoretical Biology, 64(4):771–788, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902764.

Jaffe:1974:IVC

[Jaf74] L. F. Jaffe. The interpretation of voltage-concentration relations. Journal of Theoretical Biology, 48(1):11–18, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901751.

Jakobsson:1972:EDG

[Jak72a] E. Jakobsson. An entry into the debate on generalized complementarity. Journal of Theoretical Biology, 37(1):193–195, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901269.

Jakobsson:1972:AAG

[Jak72b] Eric Jakobsson. An alternative approach to generalized complementarity. Journal of Theoretical Biology, 37(1):93–103, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901166.

Jakobsson:1973:RGC

[Jak73] Eric Jakobsson. Response to “generalized complementarity reentered”. Journal of Theoretical Biology, 40(2):403–407, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901434.

Jaroszewski:1976:MRI

[JAS76] Jan Jaroszewski, Aftab Ahmed, and Kenneth W. Sell. Mechanism for regulation of immune responses. Journal of Theoretical Biology, 57(1):121–129, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800082.

Jung:1978:CAM

[JBM^{+}78] A. Jung, P. Bartholdi, B. Mermillod, J. Reeve, and R. Neer. Critical analysis of methods for analysing human calcium kinetics. Journal of Theoretical Biology, 73(1):131–157, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901832.

Jones:1977:MCF

[JD77] Peter Jones and H. Brian Dunford. On the mechanism of compound I formation from peroxidases and catalases. Journal of Theoretical Biology, 69(3):457–470, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901527.

Jost:1973:MFC

[JDTF73] J. L. Jost, J. F. Drake, H. M. Tsuchiya, and A. G. Fredrickson. Microbial food chains and food webs. Journal of Theoretical Biology, 41(3):461–484, October 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900568.

Jean:1978:GEP

[Jea78] Roger V. Jean. Growth and entropy: Phylogenism in phyllotaxis. Journal of Theoretical Biology, 71(4):639–660, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903296.

Jean:1979:DAP

[Jea79a] Roger V. Jean. Developmental algorithms in phyllotaxis and their implications: a contribution to the field of growth functions of L-systems by an application of Perron–Frobenius spectral theory. Journal of Theoretical Biology, 76(1):1–30, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903734.

Jean:1979:SCH

[Jea79b] Roger V. Jean. Some consequences of the hierarchical approach to phyllotaxis. Journal of Theoretical Biology, 81(2):309–326, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901681.

Julg:1976:TCC

[JFR76] André Julg, Philippe François, and Michel Rajzmann. A theoretical and comparative CNDO study of chlorophylls a and b and their beryllium homologs. Journal of Theoretical Biology, 57(2):391–394, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900102.

Ji:1976:MOP

[Ji76] Sungchul Ji. A model of oxidative phosphorylation that accommodates the chemical intermediate, chemiosmotic, localized proton and conformational hypotheses. Journal of Theoretical Biology, 59(2):319–330, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901739.

Judd:1978:MOS

[JI78] C. Thomas Judd and Lloyd L. Ingraham. Molecular orbital study of optical rotation in a vitamin B_{6} enzyme. Journal of Theoretical Biology, 75(4):503–517, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903594.

Jicin:1972:SPD

[Jic72] Rudolf Jicín. Some problems of description of sets of objects. Journal of Theoretical Biology, 34(2):295–311, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901634.

Johnsson:1972:FMB

[JK72] A. Johnsson and H. G. Karlsson. A feedback model for biological rhythms: I. Mathematical description and basic properties of the model. Journal of Theoretical Biology, 36(1):153–174, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901853.

Jones:1977:SSD

[JO77] Chris E. Jones and Mark O. J. Olson. Sequence-specific DNAprotein interaction: the lac represser. Journal of Theoretical Biology, 64(2):323–326, IN3–IN6, 327–332, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903605.

Johnston:1976:TCA

[Joh76] Timothy D. Johnston. Theoretical considerations in the adaptation of animal communication systems. Journal of Theoretical Biology, 57(1):43–72, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800057.

Johnson:1979:KSF

[Joh79] Lawrence L. Johnson. Kin selection in finite sibships. Journal of Theoretical Biology, 77(3):379–381, April 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903631.

Jokl:1974:SNL

[Jok74] Miloslav Jokl. Some natural laws about harmful agents in the human environment. Journal of Theoretical Biology, 48(1):1– 9, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937490174X.

Jones:1971:TAF

[Jon71] David R. Jones. Theoretical analysis of factors which may limit the maximum oxygen uptake of fish: The oxygen cost of the cardiac and branchial pumps. Journal of Theoretical Biology, 32(2):341–349, August 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901718.

Jones:1972:CRA

[Jon72] P. C. T. Jones. Central role for ATP in determining some aspects of animal and plant cell behaviour. Journal of Theoretical Biology, 34(1):1–13, January 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900495.

Jones:1975:AAP

[Jon75] Daniel D. Jones. Amino acid properties and side-chain orientation in proteins: A cross correlation approach. Journal of Theoretical Biology, 50(1):167–183, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900314.

Jordan:1972:CGP

[Jor72] B. R. Jordan. Computer generation of pairing schemes for RNA molecules. Journal of Theoretical Biology, 34(2):363–378, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901683.

Jordan:1973:ESC

[Jor73a] Frank Jordan. The electronic structure of and conformational energy barrier to rotation around the C–N glycosidic linkage in adenosine-3’, 5’-cyclicmonophosphate (cyclic AMP) and its phosphonate analog. Journal of Theoretical Biology, 41(1):23–40, September 14, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901872.

Jordan:1973:LJP

[Jor73b] Frank Jordan. Lennard-Jones potential calculations of the barrier to rotation around the glycosidic C–N linkage in selected purine nucleosides and nucleotides. A direct comparison of the results of 6–12 potential calculations with results of semiempirical molecular orbital studies. Journal of Theoretical Biology, 41(2):375–395, September 21, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901252.

Jorne:1974:EIM

[Jor74] Jacob Jorné. The effects of ionic migration on oscillations and pattern formation in chemical systems. Journal of Theoretical Biology, 43(2):375–380, February 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800676.

Jorne:1975:NIC

[Jor75a] Jacob Jorné. Negative ionic cross diffusion coefficients in electrolytic solutions. Journal of Theoretical Biology, 55(2):529– 532, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800993.

Jornvall:1975:APT

[Jör75b] Hans Jörnvall. Acetylation of protein n-terminal amino groups structural observations on α-amino acetylated proteins. Journal of Theoretical Biology, 55(1):1–12, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801056.

Jorne:1977:DLV

[Jor77] Jacob Jorné. The diffusive Lotka–Volterra oscillating system. Journal of Theoretical Biology, 65(1):133–139, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900807.

Jones:1979:KCA

[JP79] M. N. Jones and R. Perry. The kinetics of cellular aggregation induced by turbulent flow. Journal of Theoretical Biology, 81(1):75–89, November 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900821.

Jansson:1977:CED

[JR77] Birger Jansson and Laszlo Révész. Cell ecology: Deductive and dynamic models for proliferation, differentiation and competition of tumor cell populations. Journal of Theoretical Biology, 68(1):43–51, September 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902260.

Jaffe:1974:UCM

[JRP74] L. F. Jaffe, K. R. Robinson, and B. F. PicowoLou. A uniform current model of the endogenous current in an egg. Journal of Theoretical Biology, 45(2):593–595, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901349.

Jorne:1979:LNL

[JS79] Jacob Jorné and Uriel N. Safriel. Linear and non-linear diffusion models applied to the behavior of a population of an intertidal snail. Journal of Theoretical Biology, 79(3):367–380, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903539.

Juniper:1977:SSP

[Jun77] B. E. Juniper. Some speculations on the possible roles of the plasmodesmata in the control of differentiation. Journal of Theoretical Biology, 66(3):583–586, IN3–IN4, 587–592, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903046.

Jurka:1977:RNA

[Jur77] Jerzy W. Jurka. On replication of nucleic acids in relation to the evolution of the genetic code and of proteins. Journal of Theoretical Biology, 68(4):515–520, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901023.

Jaspers:1977:PLE

[JvS77] H. T. A. Jaspers and J. van Steveninck. Pulse-labeling experiments in transmembrane transport studies. Journal of Theoretical Biology, 68(4):599–605, October 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901084.

Johnson:1976:PTD

[JW76] Michael E. Johnson and Tai Te Wu. Possible three-dimensional models for the polypeptide backbone structure of alamethicin. Journal of Theoretical Biology, 60(1):183–192, IN1–IN4, 193– 195, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901636.

Jensen:1976:DTC

[JWS76] Mona D. Jensen, Donald F. H. Wallach, and Peter Sherwood. Diffusion in tissue cultures on gas-permeable and impermeable supports. Journal of Theoretical Biology, 56(2):443–458, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800859.

Jameson:1972:QMT

[JY72] Cynthia J. Jameson and William Yang. Quantum-mechanical treatment of the electronic structure and geometry of hydrogen cyanide dimer. Journal of Theoretical Biology, 35(2):247–257, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900379.

Kier:1974:TSR

[KA74] Lemont B. Kier and Haven S. Aldrich. A theoretical study of receptor site models for trimethylammonium group interaction. Journal of Theoretical Biology, 46(2):529–541, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900137.

Kacser:1976:PSK

[Kac76] H. Kacser. The proper solutions of Kacser’s developmental model. Journal of Theoretical Biology, 63(1):239–241, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900977.

Kaminek:1974:ETO

[Kam74] Miroslav Kamínek. Evolution of tRNA and origin of the two positional isomers of zeatin. Journal of Theoretical Biology, 48(2): 489–492, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800184.

Kaneti:1973:MOS

[Kan73] José Kaneti. A molecular orbital study of the reaction catalysed by the enzyme adenosine kinase and its active site. Journal of Theoretical Biology, 38(1):169–179, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902324.

Kanungo:1975:MA

[Kan75] M. S. Kanungo. A model for ageing. Journal of Theoretical Biology, 53(2):253–261, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800026.

Katsumata:1972:GRB

[Kat72] Masuyuki Katsumata. Graphic representation of Botts– Morales equation for enzyme–substrate–modifier system. Journal of Theoretical Biology, 36(2):327–338, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901026.

Katz:1979:EFC

[Kat79] R. J. Katz. Exploration as a functional correlate of endorphins. Journal of Theoretical Biology, 77(4):537–538, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900262.

Kauffman:1971:DMB

[Kau71] Stuart Kauffman. Differentiation of malignant to benign cells. Journal of Theoretical Biology, 31(3):429–451, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900208.

Kauffman:1974:LSS

[Kau74] Stuart Kauffman. The large scale structure and dynamics of gene control circuits: An ensemble approach. Journal of Theoretical Biology, 44(1):167–190, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800378.

Kaus:1976:PAV

[Kau76] Peter Kaus. Possible adaptive values of the resonant slave in biological clocks. Journal of Theoretical Biology, 61(2):249–265, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900187.

Kang:1976:CFD

[KBH76] Sungzong Kang, Michael V. W. Bergamini, and Michael Hasen. Conformation of N-formyl-1,3-dihydroxy-Δ^{31}-pentene, a model compound of sphingomyelin. Journal of Theoretical Biology, 63(1):117–124, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900886.

Kim:1975:MDA

[KBW75] M. Kim, K. Bahrami, and K. B. Woo. Mathematical description and analysis of cell cycle kinetics and the application to Ehrlich ascites tumor. Journal of Theoretical Biology, 50(2): 437–459, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900922.

Karlin:1978:EAS

[KC78] S. Karlin and D. Carmelli. Evolutionary aspects and sensitivity studies of some major gene models. Journal of Theoretical Biology, 75(2):197–222, November 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890231X.

Kalsow:1972:FDM

[KD72] Carolyn M. Kalsow and R. J. Doyle. Facilitated diffusion of monosaccharides in saccharomyces cerevisiae I. Theoretical considerations of asymmetry. Journal of Theoretical Biology, 34(1):125–130, IN1, 131–133, January 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900598.

Kudritskaya:1976:QMS

[KD76] Z. G. Kudritskaya and V. I. Danilov. Quantum mechanical study of bases interactions in various associates in atomic dipole approximation. Journal of Theoretical Biology, 59(2):303–318, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901727.

Kingsolver:1979:MEN

[KD79] J. G. Kingsolver and T. L. Daniel. On the mechanics and energetics of nectar feeding in butterflies. Journal of Theoretical Biology, 76(2):167–168, IN1, 169–179, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903680.

Kier:1977:SAS

[KDH77] Lemont B. Kier, Thérèse Di Paolo, and Lowell H. Hall. Structure-activity studies on odor molecules using molecular connectivity. Journal of Theoretical Biology, 67(3):585–595, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900595.

Kurganov:1976:TAKa

[KDKY76a] B. I. Kurganov, A. I. Dorozhko, Z. S. Kagan, and V. A. Yakovlev. The theoretical analysis of kinetic behaviour of “hysteretic” allosteric enzymes. I. The kinetic manifestations of slow conformational change of an oligomeric enyzme in the Monod, Wyman and changeux model. Journal of Theoretical Biology, 60(2):247–269, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690059X.

Kurganov:1976:TAKb

[KDKY76b] B. I. Kurganov, A. I. Dorozhko, Z. S. Kagan, and V. A. Yakovlev. The theoretical analysis of kinetic behaviour of “hysteretic” allosteric enzymes. II. The dissociating and associating enzymic systems in which the rate of installation of equilibrium between the oligomeric forms is small in comparison with that of enzymatic reaction. Journal of Theoretical Biology, 60(2):271– 286, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900606.

Kurganov:1976:TAKc

[KDKY76c] B. I. Kurganov, A. I. Dorozhko, Z. S. Kagan, and V. A. Yakovlev. The theoretical analysis of kinetic behaviour of “hysteretic” allosteric enzymes III. Dissociating and associating enzyme systems in which the rate of installation of equilibrium between the oligomeric forms is comparable to that of enzymatic reaction. Journal of Theoretical Biology, 60(2):287–299, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900618.

Keizer:1975:TCCa

[Kei75a] Joel Keizer. Thermodynamic coupling in chemical reactions. Journal of Theoretical Biology, 49(1):323–335, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800373.

Keizer:1975:TCCb

[Kei75b] Joel Keizer. Thermodynamic coupling in chemical reactions. Journal of Theoretical Biology, 49(2):323–335, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901757.

Keleti:1971:SMM

[Kel71] T. Keleti. Simple mechanisms for the modification of homotetrameric proteins. Journal of Theoretical Biology, 30(3):545– 551, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900075.

Keller:1977:MFM

[Kel77] Stuart R. Keller. Mechanics of flagellar motion with an application to a conical spiral flagellate. Journal of Theoretical Biology, 68(1):73–94, September 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902296.

Kemeny:1974:CAC

[Kem74] Gabor Kemeny. Collective aspects of conformons and the electron transfer chain. Journal of Theoretical Biology, 48(1):231– 241, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901933.

Kember:1979:PCL

[Kem79] N. F. Kember. Proliferation controls in a linear growth system: Theoretical studies of cell division in the cartilage growth plate. Journal of Theoretical Biology, 78(3):365–374, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903369.

Kerr:1975:GMP

[Ker75] A. Kerr. A genetic model for pathogenicity in agrobacterium and for tumour induction in plants. Journal of Theoretical Biology, 51(2):409–417, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900703.

Kitayama:1976:MOT

[KF76] Hajime P. Kitayama and Hideo Fukutome. A molecular orbital theoretical study on the acid-base property and reactivity of the charge relay system in α-chymotrypsin. Journal of Theoretical Biology, 60(1):1–18, July 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901521.

Kang:1974:CDM

[KFH74] Sungzong Kang, Mark Froimowitz, and David Hankins. Conformation of 2,3-diformylglycerol, a model compound of membrane phospholipids. Journal of Theoretical Biology, 44(2):337– 347, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901647.

Kemeny:1973:PC

[KG73] Gabor Kemeny and Indur M. Goklany. Polarons and conformons. Journal of Theoretical Biology, 40(1):107–123, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901677.

Kemeny:1974:QMM

[KG74] Gabor Kemeny and Indur M. Goklany. Quantum mechanical model for conformons. Journal of Theoretical Biology, 48(1):23– 38, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901775.

Klopman:1979:MCC

[KGH79] G. Klopman, H. Grinberg, and A. J. Hopfinger. MINDO/3 calculations of the conformation and carcinogenicity of epoxymetabolites of aromatic hydrocarbons; 7,8-dihydroxy-9,10-oxy-7,8,9,10-tetrahydrobenzo(a)pyrene. Journal of Theoretical Biology, 79(3):355–366, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379903527.

King:1975:EDP

[KGL75] Charles E. King, Edward E. Gallaher, and Donald A. Levin. Equilibrium diversity in plant-pollinator systems. Journal of Theoretical Biology, 53(2):263–275, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800038.

Kapur:1976:SCO

[KGN76] P. Kapur, P. Ghosh, and N. G. Nath. Stability, controllability and observability of arterial circulation. Journal of Theoretical Biology, 61(1):15–19, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901004.

Kaneti:1970:SSO

[KGYG70] Jose Kaneti, Evgeny Golovinsky, Ivan Yukhnovsky, and Dobri Genchev. A study of some orotic acid derivatives and analogues by Hückel’s method of molecular orbitals: I. A theoretical study of the active site of orotidine-5’-phosphate pyrophosphorylase. Journal of Theoretical Biology, 26(1):19–24, IN5, 25– 27, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800285.

Kier:1975:SMRa

[KH75a] Lemont B. Kier and Hans-Dieter Höltje. A stochastic model of the remote recognition of preferred conformation in a drugreceptor interaction. Journal of Theoretical Biology, 49(1):401– 416, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800439.

Kier:1975:SMRb

[KH75b] Lemont B. Kier and Hans-Dieter Höltje. A stochastic model of the remote recognition of preferred conformation in a drugreceptor interaction. Journal of Theoretical Biology, 49(2):401– 416, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901812.

Kirkwood:1975:CSM

[KH75c] T. B. L. Kirkwood and R. Holliday. Commitment to senescence: A model for the finite and infinite growth of diploid and transformed human fibroblasts in culture. Journal of Theoretical Biology, 53(2):481–496, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580018X.

King-Hele:1979:AAS

[KH79a] J. A. King-Hele. Approximate analytical solutions for water and solute flow in intercellular spaces with a leaky tight junction. Journal of Theoretical Biology, 80(4):451–465, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900882.

Kohler:1979:UFS

[KH79b] Hans-Helmut Kohler and Klaus Heckmann. Unidirectional fluxes in saturated single-file pores of biological and artificial membranes I. Pores containing no more than one vacancy. Journal of Theoretical Biology, 79(3):381–401, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903540.

Kikuchi:1979:NTS

[KHK79] O. Kikuchi, A. J. Hopfinger, and G. Klopman. A new type of semi-empirical molecular orbital method for large molecules. Journal of Theoretical Biology, 77(1):129–139, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901437.

King-Hele:1977:ILT

[KHP77] J. A. King-Hele and R. W. Paulson. On the influence of a leaky tight junction on water and solute transport in epithelia. Journal of Theoretical Biology, 67(1):61–84, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901850.

Kibby:1979:MRP

[Kib79] Michael R. Kibby. The matrix representation of pharmacokinetic models. Journal of Theoretical Biology, 77(3):333–348, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903606.

Kiefer:1971:TTS

[Kie71] Jürgen Kiefer. Target theory and survival curves. Journal of Theoretical Biology, 30(2):307–317, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900579.

Kier:1973:CSC

[Kie73] Lemont B. Kier. Chlorpromazine and serotonin: Conformational similarities correlating with activities. Journal of Theoretical Biology, 40(2):211–217, August 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901288.

Kiger:1973:BCM

[Kig73] John A. Kiger. The bithorax complex — a model for cell determination in Drosophila. Journal of Theoretical Biology, 40(3):455–467, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900040.

Kilkson:1970:DEM

[Kil70] R. Kilkson. Domain effects in membrane transitions. Journal of Theoretical Biology, 26(3):497–499, March 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900986.

Kilmer:1972:SRC

[Kil72] W. L. Kilmer. On some realistic constraints in prey-predator mathematics. Journal of Theoretical Biology, 36(1):9–22, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901737.

Kirkman:1979:RTRa

[Kir79a]  Henry N. Kirkman.  Rectification of tandemly repeti-

tious DNA I. Influences on rate. Journal of Theoretical Biology, 81(2):259–274, November 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901656.

Kirkman:1979:RTRb

[Kir79b] Henry N. Kirkman. Rectification of tandemly repetitious DNA II. Clustering and periodicity. Journal of Theoretical Biology, 81(2):275–288, November 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901668.

Kitney:1975:ANB

[Kit75] R. I. Kitney. An analysis of the nonlinear behaviour of the human thermal vasomotor control system. Journal of Theoretical Biology, 52(1):231–248, July 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900545.

Karlsson:1972:FMB

[KJ72] H. G. Karlsson and A. Johnsson. A feedback model for biological rhythms: II. Comparisons with experimental results, especially on the petal rhythm of Kalanchoë. Journal of Theoretical Biology, 36(1):175–194, July 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901865.

Ketley:1973:SFR

[KK73] J. N. Ketley and A. D. Ketley. A suggested function of RNA-linked fragments in DNA replication. Journal of Theoretical Biology, 40(3):607–609, August 22, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900143.

Kijima:1978:COR

[KK78] Hiromasa Kijima and Shôko Kijima. Co-operative response of chemically excitable membrane: I. Formulation: Unified theory of co-operativity. Journal of Theoretical Biology, 71(4):567–585, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903259.

Kurganov:1974:KMA

[KKDY74] B. I. Kurganov, Z. S. Kagan, A. I. Dorozhko, and V. A. Yakovlev. Kinetic manifestations of allosteric interactions in models of regulatory enzymes with “indirect” co-operativity. Journal of Theoretical Biology, 47(1):1–41, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900964.

Krishnan:1977:TTD

[KKJ77] Leela Krishnan, Engikolai C. Krishnan, and William R. Jewell. Theoretical treatment of the distribution and degradation of vascular, interstitial and intracellular albumin. Journal of Theoretical Biology, 67(4):609–623, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902491.

Keilson:1978:CMH

[KKW78] J. Keilson, A. Kester, and C. Waterhouse. A circulatory model for human metabolism. Journal of Theoretical Biology, 74(4):535–547, October 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902400.

Kurganov:1979:REA

[KL79] B. I. Kurganov and N. I. Loboda. Regulation of enzyme activity in adsorptive enzyme systems. Journal of Theoretical Biology, 79(3):281–301, August 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903473.

Klambt:1975:MMR

[Klä75] Dieter Klämbt. A model for messenger RNA sequences maximizing secondary structure due to code degeneracy. Journal of Theoretical Biology, 52(1):57–65, July 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900399.

Kleiber:1972:BSC

[Kle72] Max Kleiber. Body size, conductance for animal heat flow and Newton’s law of cooling. Journal of Theoretical Biology, 37(1):139–150, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901208.

Kleiber:1975:MTR

[Kle75a] Max Kleiber. Metabolic turnover rate: A physiological meaning of the metabolic rate per unit body weight. Journal of Theoretical Biology, 53(1):199–204, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901101.

Klein:1975:MIT

[Kle75b] Gerd Klein. Monitoring intracellular transport. Journal of Theoretical Biology, 52(2):481–491, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900144.

Kitasato:1975:ACK

[KM75] H. Kitasato and K. Murayama. An analysis of the change in kpermeability on depolarization in terms of affinity and numbers of total channels. Journal of Theoretical Biology, 51(1):181–200, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901472.

Krooth:1977:TCM

[KMS77] Robert S. Krooth, S. Randolph May, and Harvey J. Stern. A theory for the chemical mediation of the excitability of the brain, with special reference to natural and drug-induced sleep. Journal of Theoretical Biology, 66(4):595–652, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902351.

Kaufman:1974:CAY

[KMSP74] Gary E. Kaufman, Morton W. Miller, John R. K. Savage, and D. G. Papworth. Chromosome aberration yields from multiple fixation regimens. Journal of Theoretical Biology, 44(1):91– 103, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800305.

Kuchel:1974:SSK

[KNJ74] P. W. Kuchel, L. W. Nichol, and P. D. Jeffrey. Steady state kinetics of consecutive enzyme catalysed reactions involving single substrates: Procedures for the interpretation of coupled assays. Journal of Theoretical Biology, 48(1):39–49, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901787.

Knowlton:1974:NEG

[Kno74] Nancy Knowlton. A note on the evolution of gamete dimorphism. Journal of Theoretical Biology, 46(1):283–285, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901532.

Kunisawa:1978:COL

[KO78] Takashi Kunisawa and Jinya Otsuka. Co-operative ligand binding in a protein composed of subunits. Journal of Theoretical Biology, 74(4):559–578, October 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902424.

Kobatake:1977:ENP

[Kob77] Yonosuke Kobatake. On the excitability of naked protoplasmic droplets of nitella. Journal of Theoretical Biology, 68(3):453–455, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900728.

Koch:1970:OCB

[Koc70] Arthur L. Koch. Overall controls on the biosynthesis of ribosomes in growing bacteria. Journal of Theoretical Biology, 28(2):203–231, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900536.

Koch:1971:ERBa

[Koc71a] Arthur L. Koch. Evaluation of the rates of biological processes from tracer kinetic data: III. The net synthesis lemma and exchangeable pools. Journal of Theoretical Biology, 32(3):429–450, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901500.

Koch:1971:ERBb

[Koc71b] Arthur L. Koch. Evaluation of the rates of biological processes from tracer kinetic data: IV. Digital simulation of nucleic acid metabolism in bacteria. Journal of Theoretical Biology, 32(3):451–469, September 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901512.

Koch:1972:DHD

[Koc72] Arthur L. Koch. Deviations from hyperbolic dependency of transport processes. Journal of Theoretical Biology, 36(1):23–40, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901749.

Koch:1974:CRA

[Koc74a] Arthur L. Koch. Coexistence resulting from an alternation of density dependent and density independent growth. Journal of Theoretical Biology, 44(2):373–386, April 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901684.

Koch:1974:CCT

[Koc74b] Arthur L. Koch. Competitive coexistence of two predators utilizing the same prey under constant environmental conditions. Journal of Theoretical Biology, 44(2):387–395, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901696.

Kang:1970:DIP

[KOH70] Sungzong Kang, John F. Olsen, and Jon R. Hamann. Dipolar ion of phenylalanine: Conformational analysis and electronic structure. Journal of Theoretical Biology, 28(2):195–201, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900524.

Koide:1973:DKP

[Koi73] Frank T. Koide. Derivation of k parameter of smooth muscle bridge resistance equations. Journal of Theoretical Biology, 42(1):181–183, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901574.

Kominz:1971:RSM

[Kom71] David R. Kominz. Role of swelling in muscle contraction. Journal of Theoretical Biology, 31(2):255–267, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190186X.

Kootsey:1977:SSF

[Koo77] J. Mailen Kootsey. The steady-state finite cable: Numerical method for non-linear membrane. Journal of Theoretical Biology, 64(3):413–420, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790279X.

Koppenhofer:1977:INP

[KOR^{+}77] E. Koppenhöfer, A. Ode, Chr. Rimmel, M. Schramm, and … H. Schumann. Isolated nitella protoplasm is not excitable. Journal of Theoretical Biology, 68(3):449–451, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900716.

Kounosu:1973:NRO

[Kou73] Shigeru Kounosu. A note on the resonant oscillator model of the chemicomechanical energy conversion in muscle contraction. Journal of Theoretical Biology, 41(2):397–399, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901264.

Kowal:1971:MEA

[Kow71] N. E. Kowal. Models of elemental assimilation by invertebrates. Journal of Theoretical Biology, 31(3):469–474, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900221.

Kirk:1970:SCA

[KOWG70] J. Kirk, J. S. Orr, T. E. Wheldon, and W. M. Gray. Stress cycle analysis in the biocybernetic study of blood cell populations. Journal of Theoretical Biology, 26(2):265–276, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800170.

Kozlov:1979:ELO

[Koz79] A. P. Kozlov. Evolution of living organisms as a multilevel process. Journal of Theoretical Biology, 81(1):1–17, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900766.

Kuhn:1972:PSA

[KP72] Jonathan Kuhn and Yair Parag. Protein-subunit aggregation model for self-incompatibility in higher fungi: Schizophyllum commune. Journal of Theoretical Biology, 35(1):77–91, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901932.

Katz:1977:GMS

[KP77a] D. F. Katz and L. Pedrotti. Geotaxis by motile spermatozoa hydrodynamic reorientation. Journal of Theoretical Biology, 67(4):723–732, August 22, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902570.

Kim:1977:MMD

[KP77b] M. Kim and Seymour Perry. Mathematical methods for determining cell DNA synthesis rate and age distribution utilizing flow microfluorometry. Journal of Theoretical Biology, 68(1):27–42, September 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902259.

Keen:1979:DED

[KP79] Robert Keen and David L. Parker. Determining expected duration of development under conditions of alternating temperatures. Journal of Theoretical Biology, 81(3):599–607, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900572.

Kharkyanen:1978:DAM

[KPU78] V. N. Kharkyanen, E. G. Petrov, and I. I. Ukrainskii. Donoracceptor model of electron transfer through proteins. Journal of Theoretical Biology, 73(1):29–50, July 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901789.

Klambt:1976:CPS

[KR76] Dieter Klämbt and Otto Richter. Computer programs for studying conformations in ribonucleic acids. Journal of Theoretical Biology, 58(2):319–324, ???? 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801221.

Krasnow:1978:MLG

[Kra78] Richard A. Krasnow. Mass, length and growth rate in single cells. Journal of Theoretical Biology, 72(4):659–699, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902783.

Kremer:1970:IAC

[Kre70] Mordechai L. Kremer. Independent activity of catalase haematins. Journal of Theoretical Biology, 29(3):387–394, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901049.

Kremen:1971:TMP

[Kre71] Alexandr Kremen. Thermodynamic model of the production of antibiotics of the tetracycline type. Journal of Theoretical Biology, 31(2):363–370, May 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901949.

Koschel:1977:ACK

[KRH77] K. W. Koschel, J. M. Radley, and G. S. Hodgson. The analysis of cell kinetics using cohort fraction labelled mitosis (COFLM) curves. Journal of Theoretical Biology, 64(2):355–361, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903629.

Krishnamurthy:1977:FPO

[Kri77] L. Krishnamurthy. Functional properties and organization of grazing lands ecosystem. Journal of Theoretical Biology, 68(1):65–72, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902284.

Kumamoto:1971:TBA

[KRL71] Junh Kumamoto, John K. Raison, and James M. Lyons. Temperature “breaks” in Arrhenius plots: A thermodynamic consequence of a phase change. Journal of Theoretical Biology, 31(1): 47–51, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901202.

Krooth:1974:DMC

[Kro74] Robert S. Krooth. A deterministic mechanism for cellular and organismic ageing. Journal of Theoretical Biology, 46(2):501– 507, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900113.

Kroes:1977:NSE

[Kro77] H. W. Kroes. The niche structure of ecosystems. Journal of Theoretical Biology, 65(2):317–326, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903277.

Keller:1970:ISM

[KS70] Evelyn F. Keller and Lee A. Segel. Initiation of slime mold aggregation viewed as an instability. Journal of Theoretical Biology, 26(3):399–415, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900925.

Keller:1971:MC

[KS71a] Evelyn F. Keller and Lee A. Segel. Model for chemotaxis. Journal of Theoretical Biology, 30(2):225–234, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900506.

Keller:1971:TBC

[KS71b] Evelyn F. Keller and Lee A. Segel. Traveling bands of chemotactic bacteria: A theoretical analysis. Journal of Theoretical Biology, 30(2):235–248, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900518.

Kirschbaum:1971:OEM

[KS71c] Thomas H. Kirschbaum and Norman Z. Shapiro. Oxygen equilibration of maternal and fetal blood. Journal of Theoretical Biology, 31(2):375, May 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901962.

Karfunkel:1972:RIE

[KS72] Heinrich R. Karfunkel and Friedrich F. Seelig. Reversal of inhibition of enzymes and the model of a spike oscillator. Journal of Theoretical Biology, 36(2):237–253, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900951.

Kashyap:1974:SEP

[KS74] R. L. Kashyap and S. Subas. Statistical estimation of parameters in a phylogenetic tree using a dynamic model of the substitutional process. Journal of Theoretical Biology, 47(1):75–101, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901003.

Kennedy:1977:ECS

[KS77a] C. M. Kennedy and E. R. Smith. Effects of charge structure on long-range interactions between cylindrical polyelectrolyte molecules. Journal of Theoretical Biology, 65(1):203–207, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900844.

Kometani:1977:MDC

[KS77b] K. Kometani and H. Shimizu. On the molecular dynamics of the chemo-mechanical conversion in muscle contraction. Journal of Theoretical Biology, 69(3):415–428, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901497.

Korostyshevsky:1974:SPE

[KSR74] M. A. Korostyshevsky, M. R. Schtabnoy, and V. A. Ratner. On some principles of evolution viewed as a stochastic process. Journal of Theoretical Biology, 48(1):85–103, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901817.

Kittredge:1972:ESP

[KT72a] J. S. Kittredge and F. T. Takahashi. The evolution of sex pheromone communication in the arthropoda. Journal of Theoretical Biology, 35(3):467–471, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901452.

Kuba:1972:ENM

[KT72b] K. Kuba and T. Tomita. Effects of noradrenaline on miniature end-plate potentials and on end-plate potential. Journal of Theoretical Biology, 36(1):81–88, July 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901786.

Koushanpour:1971:MSN

[KTS71] Esmail Koushanpour, Ronald R. Tarica, and William F. Stevens. Mathematical simulation of normal nephron function in rat and man. Journal of Theoretical Biology, 31(2):177–214, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901822.

Kubitschek:1970:EGL

[Kub70] H. E. Kubitschek. Evidence for the generality of linear cell growth. Journal of Theoretical Biology, 28(1):15–29, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900615.

Kubo:1976:VEP

[Kub76] Kazumi Kubo. A view on the equivalence point in the precipitin test. Journal of Theoretical Biology, 63(2):453–478, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900461.

Kumbar:1974:SDH

[Kum74] Mahadevappa Kumbar. Supercoiled DNA: Hydrodynamic model. Journal of Theoretical Biology, 43(2):253–263, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800585.

Kumbar:1975:DMS

[Kum75a] Mahadevappa Kumbar. Dynamic model for sedimentation of supercoiled molecules. Journal of Theoretical Biology, 51(2): 307–312, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900624.

Kumbar:1975:SCH

[Kum75b] Mahadevappa Kumbar. Stable conformations of histamine by empirical method. Journal of Theoretical Biology, 53(2):333– 340, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800087.

Kurjian:1971:LSA

[Kur71] J. G. Kurjian. Logical structure of adaptive systems: I. Microevolution, conditioning and their decision problems. Journal of Theoretical Biology, 32(2):299–330, August 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901688.

Kurganov:1977:TAK

[Kur77] B. I. Kurganov. The theoretical analysis of kinetic behaviour of “hysteretic” allosteric enzymes: IV. Kinetics of dissociation-association processes of allosteric enzymes. Journal of Theoretical Biology, 68(4):521–543, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901035.

Kim:1974:AEP

[KV74] Y. J. Kim and A. J. Valleron. Analytical expressions for PLM(t), CL(t) and CLM(t) without use of Laplace transforms. Journal of Theoretical Biology, 44(2):353–365, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901660.

Klein:1977:CMS

[KV77] B. Klein and A.-J. Valleron. A compartmental model for the study of diurnal rhythms in cell proliferation. Journal of Theoretical Biology, 64(1):27–42, January 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901126.

Kharkats:1977:TIE

[KVB77] Yurij I. Kharkats, Aleksandr G. Volkov, and Leonid I. Boguslavsky. Transfer of ions and electrons across the interface between two immiscible liquids in functioning enzyme membrane systems. Journal of Theoretical Biology, 65(2):379–391, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903320.

Kauffman:1975:MOP

[KW75] Stuart Kauffman and John J. Wille. The mitotic oscillator in Physarum polycephalum. Journal of Theoretical Biology, 55(1):47–82, IN1–IN2, 83–93, November 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801081.

Kohn:1979:IFC

[KWG79] Michael C. Kohn, Lucy M. Whitley, and David Garfinkel. Instantaneous flux control analysis for biochemical systems. Journal of Theoretical Biology, 76(4):437–452, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900122.

Koons:1979:MDV

[KZ79] Stephen J. Koons and C. Richard Zobel. A modified distance variable for the hill formalism for kinetic models of muscle contraction. Journal of Theoretical Biology, 81(3):505–513, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900481.

Leake:1976:ECA

[LA76] Barbara Leake and Photios A. Anninos. Effect of connectivity on the activity of neural net models. Journal of Theoretical Biology, 58(2):337–363, ???? 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801245.

Ladde:1976:SME

[Lad76] G. S. Ladde. Stability of model ecosystems with timedelay. Journal of Theoretical Biology, 61(1):1–13, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900990.

Ladde:1977:CPI

[Lad77] G. S. Ladde. Competitive processes II. Stability of random systems. Journal of Theoretical Biology, 68(3):331–354, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900649.

Laing:1975:SAR

[Lai75] Richard Laing. Some alternative reproductive strategies in artificial molecular machines. Journal of Theoretical Biology, 54(1): 63–84, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800555.

Laing:1977:AMR

[Lai77] Richard Laing. Automaton models of reproduction by selfinspection. Journal of Theoretical Biology, 66(3):437–456, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902946.

Laki:1974:AET

[Lak74] K. Laki. Actin as an energy transducer. Journal of Theoretical Biology, 44(1):117–130, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800329.

Lakshminarayanaiah:1975:CEN

[Lak75] N. Lakshminarayanaiah. Comment on “evidence for a new concept of membrane potential” by ohki. Journal of Theoretical Biology, 50(2):497–500, April 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900958.

Lang:1975:NRS

[Lan75] Alexander Lang. A new role for the sieve plate. Journal of Theoretical Biology, 51(2):303–306, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900612.

Lanks:1977:PMH

[Lan77] Karl W. Lanks. A prediction of mRNA hybridization kinetics based on polypeptide abundances. Journal of Theoretical Biology, 69(1):1–9, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903848.

Lancaster:1978:SEM

[Lan78] Jack R. Lancaster. Site-exposure model for proton-lactose symport in Escherichia coli. Journal of Theoretical Biology, 75(1):35–50, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902011.

Lapidus:1978:TPM

[Lap78] I. Richard Lapidus. Time periods of morphological stages in populations of differentiating micro-organisms. Journal of Theoretical Biology, 73(2):319–327, July 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901935.

Laplante:1979:ISS

[Lap79] Jean-Pierre Laplante. Inhomogeneous steady state distributions of species in predator–prey systems. A specific example. Journal of Theoretical Biology, 81(1):29–45, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990078X.

Latimer:1975:IPD

[Lat75] Paul Latimer. The influence of photometer design on opticalconformational changes. Journal of Theoretical Biology, 51(1): 1–12, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901356.

Lauria:1972:MAS

[Lau72] F. E. Lauria. Mathematical approach to the study of the cerebral cortex: IV. The spread of excitation in the space and time domains. Journal of Theoretical Biology, 35(1):145–158, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901993.

Lavenda:1971:COE

[Lav71] B. Lavenda. Co-operative effects and charge fluctuation forces. Journal of Theoretical Biology, 31(3):395–404, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190018X.

Lavorel:1976:MAO

[Lav76] J. Lavorel. Matrix analysis of the oxygen evolving system of photosynthesis. Journal of Theoretical Biology, 57(1):171– 185, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800112.

Lavender:1977:KSR

[Lav77] James A. Lavender. Kauffman’s square-root law: Possible correlatives. Journal of Theoretical Biology, 65(4):791–793, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900261.

Laxton:1978:MD

[Lax78] R. R. Laxton. The measure of diversity. Journal of Theoretical Biology, 70(1):51–67, January 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903028.

Layzer:1978:MAP

[Lay78] David Layzer. A macroscopic approach to population genetics. Journal of Theoretical Biology, 73(4):769–788, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901364.

Lubliner:1971:MBP

[LB71] J. Lubliner and J. J. Blum. Model for bend propagation in flagella. Journal of Theoretical Biology, 31(1):1–24, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901172.

Levine:1972:WHE

[LB72a] Arnold J. Levine and Max M. Burger. A working hypothesis explaining the maintenance of the transformed state by SV40 and polyoma. Journal of Theoretical Biology, 37(3):435–446, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900847.

Lubliner:1972:MFW

[LB72b] J. Lubliner and J. J. Blum. Model of flagellar waves. Journal of Theoretical Biology, 34(3):515–534, March 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901397.

Lubliner:1977:ABI

[LB77] J. Lubliner and J. J. Blum. Analysis of bend initiation in cilia according to a sliding filament model. Journal of Theoretical Biology, 69(1):87–99, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903897.

Lespinasse:1976:FDH

[LBCV76] Jean-Noël Lespinasse, Henri Broch, Raphaël Cornillon, and Dane Vasilescu. Flexibility of deoxyguanosine 5’monophosphate (5’dGMP): an E.H.T. conformational analysis. Journal of Theoretical Biology, 57(1):225–230, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800148.

Lee:1976:FMM

[LBGK76] Kwang Yun Lee, Robert O. Barr, Stuart H. Gage, and Anil N. Kharkar. Formulation of a mathematical model for insect pest ecosystems — the cereal leaf beetle problem. Journal of Theoretical Biology, 59(1):33–76, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800239.

Leibovic:1970:BSR

[LBM70] K. N. Leibovic, E. Balslev, and T. A. Mathieson. Binocular space and its representation by cellular assemblies in the brain. Journal of Theoretical Biology, 28(3):513–529, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937090086X.

Loew:1972:MOCa

[LCC72] G. H. Loew, M. S. Chadha, and S. Chang. A molecular orbital and chemical study of aminoacetonitrile: A possible prebiotic peptide precursor. Journal of Theoretical Biology, 35(2):359– 373, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900446.

Levin:1978:CPE

[LCH78] R. L. Levin, E. G. Cravalho, and C. E. Huggins. The concentration polarization effect in a multicomponent electrolyte solution — the human erythrocyte. Journal of Theoretical Biology, 71(2):225–254, March 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902692.

Laffort:1973:APC

[LD73] Paul Laffort and Andrew Dravnieks. An approach to a physicochemical model of olfactory stimulation in vertebrates by single compounds. Journal of Theoretical Biology, 38(2):335–345, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901781.

Lovely:1975:SMB

[LD75] Peter S. Lovely and F. W. Dahlquist. Statistical measures of bacterial motility and chemotaxis. Journal of Theoretical Biology, 50(2):477–496, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900946.

Lebovitz:1973:SME

[Leb73] Robert M. Lebovitz. Significance of microthermal effects derived from low level UHF-microwave irradiation of the head: Indirect caloric vestibular stimulation. Journal of Theoretical Biology, 41(2):209–221, September 21, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901148.

Lefebvre:1978:TDB

[Lef78] Gerald M. Lefebvre. The time dependence of bacterial spore germination. Journal of Theoretical Biology, 75(3):307–326, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903375.

Leibovic:1970:IRC

[Lei70] K. N. Leibovic. Interactions of retinal cells in visual adaptation. Journal of Theoretical Biology, 28(3):507–512, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900858.

Leibel:1977:BRS

[Lei77] Rudolph L. Leibel. A biologic radar system for the assessment of body mass the model of a geometry sensitive endocrine system is presented. Journal of Theoretical Biology, 66(2):297–306, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901734.

Leith:1978:DLL

[Lei78] A. Leith. Detection of limited life span in small clones. Journal of Theoretical Biology, 71(3):453–464, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901728.

Leopold:1971:TSE

[Leo71] Luna B. Leopold. Trees and streams: The efficiency of branching patterns. Journal of Theoretical Biology, 31(2):339–354, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901925.

Leon:1975:LCP

[Leó75a] Jesús Alberto León. Limit cycles in populations with separate generations. Journal of Theoretical Biology, 49(1):241–244, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800312.

Leon:1975:FLC

[Leó75b] Pedro E. León. Function of lampbrush chromosomes: A hypothesis. Journal of Theoretical Biology, 55(2):481–490, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800956.

Leon:1976:LHA

[Leó76] Jesús Alberto León. Life histories as adaptive strategies. Journal of Theoretical Biology, 60(2):301–335, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690062X.

Lerman:1978:BER

[Ler78] M. I. Lerman. The biological essence of resting cells in cell populations. Journal of Theoretical Biology, 73(4):615–629, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901261.

Lerner:1979:IF

[Ler79]  Michael R. Lerner. Immortal fibroblasts?  Journal of

Theoretical Biology, 77(2):213–216, March 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903084.

Lesk:1970:OGC

[Les70] Arthur M. Lesk. On the origin of the genetic code: Photochemical interaction between amino acids and nucleic acids not requiring adaptors. Journal of Theoretical Biology, 27(1):171– 173, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901359.

Lesk:1973:HSP

[Les73] Arthur M. Lesk. On hypothesized selective pressure by u.v. on DNA base compositions. Journal of Theoretical Biology, 40(1): 201–202, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901720.

Lesk:1974:CSE

[Les74] Arthur M. Lesk. A combinatorial study of the effects of admitting non-watson-crick base pairings and of base composition on the helix-forming potential of polynucleotides of random sequence. Journal of Theoretical Biology, 44(1):7–17, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800251.

Lesk:1977:ETF

[Les77] Arthur M. Lesk. An encoding technique to facilitate the detection of homologies in biopolymer sequences. Journal of Theoretical Biology, 69(4):767–769, December 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903824.

Lestienne:1978:SAT

[Les78] Patrick Lestienne. The specificity of aminoacylation: a tRNA-tRNA interaction model. Journal of Theoretical Biology, 73(1):159–180, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901844.

Levitt:1972:CTE

[Lev72] David G. Levitt. Capillary-tissue exchange kinetics: An analysis of the krogh cylinder model. Journal of Theoretical Biology, 34(1):103–124, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900586.

Levitzki:1974:NCO

[Lev74a] Alexander Levitzki. Negative co-operativity in clustered receptors as a possible basis for membrane action. Journal of Theoretical Biology, 44(2):367–372, April 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901672.

Levy:1974:FEE

[Lev74b] Elinor Miller Levy. Flagellar elongation: An example of controlled growth. Journal of Theoretical Biology, 43(1):133–149, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800494.

Levine:1977:EIS

[Lev77] Stephen H. Levine. Exploitation interactions and the structure of ecoystems. Journal of Theoretical Biology, 69(2):345–355, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901424.

Lewin:1970:PBS

[Lew70a] S. Lewin. Possible basis for specificity in attachment of histones to DNA in relation to genetic control in mRNA synthesis. Journal of Theoretical Biology, 29(1):1–11, IN1–IN4, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901141.

Lewin:1970:WEB

[Lew70b] S. Lewin. Water extrusion in biological reactions. Journal of Theoretical Biology, 26(3):481–486, IN3–IN8, 487–497, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900974.

Lewinson:1971:EWS

[Lew71] T. M. Lewinson. Enumerating ways for the stepwise synthesis of macromolecules. Journal of Theoretical Biology, 31(3):557– 559, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900300.

Lewin:1972:PFH

[Lew72] Sherry Lewin. Probabilities of formation of α-helices in polyl-glutamic acid and in poly-l-aspartic acid. Journal of Theoretical Biology, 36(1):1–2, IN1–IN3, 3–7, July 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901725.

Lewis:1973:TCM

[Lew73] Julian Lewis. The theory of clonal mixing during growth. Journal of Theoretical Biology, 39(1):47–54, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390204X.

Lewis:1976:MGM

[Lew76] R. W. Lewis. Mucus globule membrane: An hypothesis concerning its role in determining the viscosity of mucus. Journal of Theoretical Biology, 61(1):21–26, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901016.

Leung:1977:STR

[LF77] L. Stan Leung and Walter J. Freeman. The spatial temporal response of the Limulus eye neural set. Journal of Theoretical Biology, 69(1):41–56, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903873.

Leith:1971:SMC

[LG71] Ardean G. Leith and Narendra S. Goel. Simulation of movement of cells during self-sorting. Journal of Theoretical Biology, 33(1):171–188, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371902232.

Levine:1976:VIN

[LG76] Daniel S. Levine and Stephen Grossberg. Visual illusions in neural networks: Line neutralization, tilt after effect, and angle expansion. Journal of Theoretical Biology, 61(2):477–504, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937690031X.

Lacalli:1978:DOA

[LH78a] T. C. Lacalli and L. G. Harrison. Development of ordered arrays of cell wall pores in desmids: a nucleation model. Journal of Theoretical Biology, 74(1):109–138, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902928.

Lacalli:1978:RCT

[LH78b] T. C. Lacalli and L. G. Harrison. The regulatory capacity of Turing’s model for morphogenesis, with application to slime moulds. Journal of Theoretical Biology, 70(3):273–295, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903776.

Lacalli:1979:TCA

[LH79a] T. C. Lacalli and L. G. Harrison. Turing’s conditions and the analysis of morphogenetic models. Journal of Theoretical Biology, 76(4):419–436, February 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900110.

Lyons:1979:SRB

[LH79b] N. I. Lyons and K. Hutcheson. A stopping rule based on Simpson’s index of diversity. Journal of Theoretical Biology, 77(4):497–503, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379900225.

Levy:1976:MPR

[LHC76] William B. Levy, John W. Haycock, and Carl W. Cotman. Membrane permeability and receptor function. Journal of Theoretical Biology, 60(1):109–130, July 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901582.

Lindenmayer:1971:DSC

[Lin71] Aristid Lindenmayer. Developmental systems without cellular interactions, their languages and grammars. Journal of Theoretical Biology, 30(3):455–484, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900026.

Lindenmayer:1975:DAM

[Lin75] Aristid Lindenmayer. Developmental algorithms for multicellular organisms: A survey of l-systems. Journal of Theoretical Biology, 54(1):3–22, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800518.

Lin:1976:ODS

[Lin76] S. H. Lin. Oxygen diffusion in a spherical cell with nonlinear oxygen uptake kinetics. Journal of Theoretical Biology, 60(2):449– 457, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900710.

Lin:1979:MMP

[Lin79a] S. H. Lin. A modified model for predicting the performance of a compact artificial kidney. Journal of Theoretical Biology, 77(4):441–451, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379900195.

Linkens:1979:MAF

[Lin79b] D. A. Linkens. Modulation analysis of forced non-linear oscillators for biological modelling. Journal of Theoretical Biology, 77(3):235–251, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379903564.

Linkens:1979:TAB

[Lin79c] D. A. Linkens. Theoretical analysis of beating and modulation phenomena in weakly inter-coupled van der Pol oscillator systems for biological modelling. Journal of Theoretical Biology, 79(1):31–54, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902558.

Litwin:1970:KDR

[LK70] Samuel Litwin and Neville R. Kallenbach. Kinetics of DNA replication in bacteria with heterogeneous generation times. Journal of Theoretical Biology, 27(3):357–376, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800029.

Lin:1976:AMP

[LK76] Juan Lin and Peter B. Kahn. Averaging methods in predator–prey systems and related biological models. Journal of Theoretical Biology, 57(1):73–102, March 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800069.

Lin:1977:QBP

[LK77a] Juan Lin and Peter B. Kahn. Qualitative behavior of predator– prey communities. Journal of Theoretical Biology, 65(1):101– 132, March 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900790.

Litwin:1977:LOD

[LK77b] Samuel Litwin and Neville R. Kallenbach. On the localization of the origin of DNA replication in E. coli. Journal of Theoretical Biology, 64(4):747–760, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902739.

Lauffenburger:1979:ELR

[LK79] Douglas Lauffenburger and Kenneth H. Keller. Effects of leukocyte random motility and chemotaxis in tissue inflammatory response. Journal of Theoretical Biology, 81(3):475–503, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937990047X.

Lange:1970:MID

[LL70] Yvonne Lange and Robert V. Lange. Molecular interactions and diffusion permeability. Journal of Theoretical Biology, 28 (1):47–58, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900639.

Lakshminarayanan:1979:MLS

[LL79] A. V. Lakshminarayanan and A. Lent. Methods of least squares and SIRT in reconstruction. Journal of Theoretical Biology, 76(3):267–295, February 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903138.

Lloyd:1977:GPM

[Llo77] David G. Lloyd. Genetic and phenotypic models of natural selection. Journal of Theoretical Biology, 69(3):543–560, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901552.

Lin:1974:OFA

[LM74a] C. H. Lin and L. V. McIntire. Oscillatory finite amplitude supercritical equilibria induced by active stress. Journal of Theoretical Biology, 46(1):129–142, July 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901441.

Lundstrom:1974:PNM

[LM74b] Ingemar Lundström and Douglas McQueen. A proposed 1f noise mechanism in nerve cell membranes. Journal of Theoretical Biology, 45(2):405–409, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901210.

Lindner:1975:ESOa

[LM75a] Peter Lindner and Olle Mårtensson. On the electronic structure and odours of esters of the isothiocyanic acid. Journal of Theoretical Biology, 49(1):377–383, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800415.

Lindner:1975:ESOb

[LM75b] Peter Lindner and Olle Mårtensson. On the electronic structure and odours of esters of the isothiocyanic acid. Journal of Theoretical Biology, 49(2):377–383, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901794.

Lefever:1971:CIS

[LN71] R. Lefever and G. Nicolis. Chemical instabilities and sustained oscillations. Journal of Theoretical Biology, 30(2):267– 284, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900543.

Lavenda:1971:CIR

[LNHK71] B. Lavenda, G. Nicolis, and M. Herschkowitz-Kaufman. Chemical instabilities and relaxation oscillations. Journal of Theoretical Biology, 32(2):283–292, August 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901664.

Loew:1971:CHC

[Loe71] Gilda H. Loew. Conformation of hydrogen cyanide dimer and its role in chemical evolution. Journal of Theoretical Biology, 33(1):121–130, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371902207.

Longmuir:1976:RLE

[Lon76] I. S. Longmuir. Reply to a letter to the editor by A. N. Stokes. Journal of Theoretical Biology, 60(2):479–480, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900758.

Lakshmanan:1975:NES

[LP75] M. Lakshmanan and P. K. Ponnuswamy. Nonlinear effects on the synaptic basis of memory. Journal of Theoretical Biology, 53(1):239–242, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901149.

Lapidus:1970:SSR

[LR70] I. Richard Lapidus and Bernard Rosen. Secondary structure of 5 s ribosomal RNA. Journal of Theoretical Biology, 27(3): 417–423, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800066.

Lou:1974:CVS

[LR74] Ru-Shuoh Lou and Robert Rosen. Comments on the validity of the simultaneous use of the “conservation of edge” equations in the model proposed by goel et al. (1970). Journal of Theoretical Biology, 46(2):587–589, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900204.

Lamb:1976:ICO

[LR76] C. J. Lamb and P. H. Rubery. Inhibition of co-operative enzymes by substrate-analogues: Possible implications for the physiological significance of negative co-operativity illustrated by phenylalanine metabolism in higher plants. Journal of Theoretical Biology, 60(2):441–447, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900709.

Lebowitz:1971:TCK

[LRB71] J. L. Lebowitz, S. I. Rubinow, and Shlomo Burstein. Theorem in chemical kinetics. Journal of Theoretical Biology, 32(2):335– 339, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901706.

Lapidus:1977:SSQ

[LRH77] I. Richard Lapidus, Bernard Rosen, and Robert Hepperle. Secondary structure of QβRNA. Journal of Theoretical Biology, 64(3):587–595, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902909.

Levy:1979:MCM

[LRS79] Harvey M. Levy, Fausto Ramirez, and Kamal K. Shukla. Muscle contraction: a mechanism of energy transduction. Journal of Theoretical Biology, 81(2):327–332, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901693.

Lieb:1971:RCS

[LS71] W. R. Lieb and W. D. Stein. Rejection criterion for some forms of the conventional carrier. Journal of Theoretical Biology, 30(1):219–222, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900488.

Larson:1972:MRB

[LS72a] Kenneth B. Larson and Donald L. Snyder. Measurement of relative blood flow, transit-time distributions and transportmodel parameters by residue detection when radiotracer recirculates. Journal of Theoretical Biology, 37(3):503–529, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900884.

Lieb:1972:IUL

[LS72b] W. R. Lieb and W. D. Stein. The influence of unstirred layers on the kinetics of carrier-mediated transport. Journal of Theoretical Biology, 36(3):641–645, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900161.

Lapidus:1975:BCF

[LS75] I. Richard Lapidus and Ralph Schiller. Bacterial chemotaxis in a fixed attractant gradient. Journal of Theoretical Biology, 53(1):215–222, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901125.

Lardner:1976:DLC

[LS76] T. J. Lardner and N. Solomon. The determination of local cell membrane diffusion coefficients. Journal of Theoretical Biology, 60(2):433–440, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900692.

Lieb:1977:THA

[LS77] W. R. Lieb and W. D. Stein. Is there a high affinity site for sugar transport at the inner face of the human red cell membrane? Journal of Theoretical Biology, 69(2):311–319, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901400.

Lundstrom:1978:EIB

[LS78] I. Lundström and M. Stenberg. Electrostatic interaction between gating charges in nerve membranes. Journal of Theoretical Biology, 70(2):229–244, January 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903491.

Levin:1978:CFC

[LSB78] Gilbert V. Levin, Patricia Ann Straat, and William D. Benton. Color and feature changes at Mars Viking lander site. Journal of Theoretical Biology, 75(3):381–382, IN1–IN4, 383–390, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903429.

Lopez-Saez:1975:MCC

[LSGFdlT^{+}75] Jorge F. López-Sáez, Aurora González-Fernández, Consuelo de la Torre, José Luis Díez, Ma Encarnacion Fernández-Gomez, Matilde H. Navarrete, Gregorio García-Herdugo, and Gonzalo Giménez-Martín. A model for cell cycle and growth kinetics in roots. Journal of Theoretical Biology, 53(2):463–473, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800166.

Laki:1977:CTL

[LSL77] K. Laki, M. Seel, and J. Ladik. Conformation of triphosphate and lysine-triphosphate-arginine complex. Journal of Theoretical Biology, 67(3):489–498, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900510.

Lardner:1974:NMO

[LST74a] T. J. Lardner, W. J. Shack, and P. Y. Tam. A note on the mechanics of ovulation. Journal of Theoretical Biology, 48(2): 481–483, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800160.

Lumsden:1974:QME

[LST74b] Charles J. Lumsden, Melvin Silverman, and L. E. H. Trainor. Quantum mechanical effects in molecule-membrane interactions. Journal of Theoretical Biology, 48(2):325–343, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800044.

Lewis:1977:TD

[LSW77] J. Lewis, J. M. W. Slack, and L. Wolpert. Thresholds in development. Journal of Theoretical Biology, 65(3):579–590, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902168.

Loew:1972:MOCb

[LT72] Gilda H. Loew and David D. Thomas. Molecular orbital calculations of the catalytic effect of lysozyme: 1. Glu 35 as general acid catalyst. Journal of Theoretical Biology, 36(1):89–104, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901798.

Leon:1975:CBT

[LT75] Jesús Alberto León and Daniel B. Tumpson. Competition between two species for two complementary or substitutable resources. Journal of Theoretical Biology, 50(1):185–201, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900326.

Lassalles:1977:DSM

[LT77] J. P. Lassalles and M. Thellier. Discussion on the symmetry of macroscopic coefficients in the formulation of the cellular fluxes: Possible application to a “symmetry criterion” for the study of active and passive transports. Journal of Theoretical Biology, 68(1):53–63, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902272.

Leguizamon:1978:CFM

[LTP78] Carlos A. Leguizamón, Elma A. A. Tenreyro, and Victorio Pecorini. Coronary flow. Method of returning coronary bolus. Journal of Theoretical Biology, 73(3):563–574, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901595.

Lubliner:1973:AIS

[Lub73] J. Lubliner. An analysis of interfilament shear in flagella. Journal of Theoretical Biology, 41(1):119–125, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901926.

Luck:1975:EBR

[Lüc75] Hermann B. Lück. Elementary behavioural rules as a foundation for morphogenesis. Journal of Theoretical Biology, 54(1):23– 34, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937580052X.

Luchnik:1979:MSR

[Luc79] Andrey N. Luchnik. On the mechanism of SOS-repair and prophage induction: Relaxation hypothesis. Journal of Theoretical Biology, 77(2):229–231, March 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903102.

Lundstrom:1974:MWP

[Lun74] Ingemar Lundström. Mechanical wave propagation on nerve axons. Journal of Theoretical Biology, 45(2):487–499, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901271.

Lewis:1976:PNE

[LW76] J. H. Lewis and L. Wolpert. The principle of non-equivalence in development. Journal of Theoretical Biology, 62(2):479–490, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901326.

Lewis:1979:DES

[LW79a] Julian Lewis and Lewis Wolpert. Diploidy, evolution and sex. Journal of Theoretical Biology, 78(3):425–438, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903412.

Lurie:1979:NET

[LW79b] D. Lurié and J. Wagensberg. Non-equilibrium thermodynamics and biological growth and development. Journal of Theoretical Biology, 78(2):241–250, May 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902674.

Lymn:1974:AAM

[Lym74a] Richard W. Lymn. Actin activation of the myosin ATPase: A kinetic analysis. Journal of Theoretical Biology, 43(2):313– 328, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800639.

Lymn:1974:EMT

[Lym74b] Richard W. Lymn. Effect of modifier on three-step enzymic process: General rate equation. Journal of Theoretical Biology, 43(2):305–312, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800627.

Lymn:1975:SSAa

[Lym75a] Richard W. Lymn. The steady-state actomyosin ATPase: A further note. Journal of Theoretical Biology, 49(1):425–429, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800452.

Lymn:1975:SSAb

[Lym75b] Richard W. Lymn. The steady-state actomyosin ATPase: A further note. Journal of Theoretical Biology, 49(2):425–429, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901836.

Mylrea:1971:MAD

[MA71] Kenneth C. Mylrea and Peter H. Abbrecht. Mathematical analysis and digital simulation of the control of erythropoiesis. Journal of Theoretical Biology, 33(2):279–297, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900671.

Matthews:1978:FKD

[MA78] Richard H. Matthews and James O. Alben. Fitting kinetic data for two independent saturable terms by Multifit II, a general purpose curve fitting program in FORTRAN IV. Journal of Theoretical Biology, 72(2):219–230, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900905.

Mackay:1973:HOT

[Mac73] A. L. Mackay. How to organize a typists’ pool — an operational analysis of an old metaphor. Journal of Theoretical Biology, 40(1):203–204, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901732.

Mackay:1975:MME

[Mac75] A. L. Mackay. A metaphor for molecular evolution. Journal of Theoretical Biology, 54(2):399–401, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801433.

Mackay:1976:GCE

[Mac76]  A. L. Mackay. A gyro-chemical effect?  Journal of

Theoretical Biology, 63(1):229–232, November 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900953.

MacDonald:1977:BTA

[Mac77a] N. MacDonald. Birfurcation theory applied to a simple model of a biochemical oscillator. Journal of Theoretical Biology, 65(4):727–734, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377900182.

MacDonald:1977:PCO

[Mac77b] N. MacDonald. A polar co-ordinate system for positional information in the vertebrate neural retina. Journal of Theoretical Biology, 69(1):153–165, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903939.

MacDonald:1977:TLM

[Mac77c] N. MacDonald. Time lag in a model of a biochemical reaction sequence with end product inhibition. Journal of Theoretical Biology, 67(3):549–556, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790056X.

Maclean:1977:AMM

[Mac77d] F. I. Maclean. Analysis of metabolism by mathematical programming techniques. Journal of Theoretical Biology, 65(3): 513–522, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902119.

Macnair:1978:ESR

[Mac78] Mark R. Macnair. An ESS for the sex ratio in animals, with particular reference to the social hymenoptera. Journal of Theoretical Biology, 70(4):449–459, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902527.

MacDonald:1979:SAF

[Mac79] N. MacDonald. Simple aspects of foodweb complexity. Journal of Theoretical Biology, 80(4):577–588, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900948.

Maddy:1976:CEA

[Mad76] A. H. Maddy. A critical evaluation of the analysis of membrane proteins by polyacrylamide gel electrophoresis in the presence of dodecyl sulphate. Journal of Theoretical Biology, 62(2):315–326, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901235.

Maden:1977:RPI

[Mad77a] M. Maden. The regeneration of positional information in the amphibian limb. Journal of Theoretical Biology, 69(4):735–753, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903794.

Madsen:1977:TTE

[Mad77b] Ernest L. Madsen. Theory for a test of the electric cable model of the myelinated axon and saltatory conduction. Journal of Theoretical Biology, 67(2):203–212, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901941.

Magar:1970:MWD

[Mag70] Magar E. Magar. Molecular weight distribution functions and virial coefficients from sedimentation equilibrium data. Journal of Theoretical Biology, 27(1):127–169, April 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901347.

Magar:1971:PDS

[Mag71] Magar E. Magar. On the possibility of determining the secondary structure of proteins in solution. Journal of Theoretical Biology, 33(1):105–119, October 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902190.

Maggiora:1973:TAS

[Mag73] Gerald M. Maggiora. Theoretical aspects of the structural chemistry of biotin I. The electronic structure of biotin and protonated biotins. Journal of Theoretical Biology, 41(3):523–534, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900593.

Maggiora:1974:TAS

[Mag74a] Gerald M. Maggiora. Theoretical aspects of the structural chemistry of biotin II. Carboxylated biotins. Journal of Theoretical Biology, 43(1):51–64, January 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800445.

Maguire:1974:NCB

[Mag74b] Marjorie P. Maguire. The need for a chiasma binder. Journal of Theoretical Biology, 48(2):485–487, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800172.

Makowski:1976:PTM

[Mak76a] Lee Makowski. A phase transition model for hemolysis. Journal of Theoretical Biology, 61(1):47–53, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690103X.

Makowski:1976:PBR

[Mak76b] Lee Makowski. A physical basis for red cell membrane elasticity. Journal of Theoretical Biology, 61(1):27–45, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901028.

Malcolm:1976:TSE

[Mal76] Alan D. B. Malcolm. On transition states in enzyme catalysed reactions. Journal of Theoretical Biology, 60(1):239–240, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901673.

Malcolm:1979:AAS

[Mal79] Alan D. B. Malcolm. Amino acid sequence homology. Journal of Theoretical Biology, 77(2):233–234, March 21, 1979.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903114.

McCabe:1979:ETU

[MAM79] M. McCabe, K. Adam, and D. Maguire. The effect of temperature upon the combined diffusional and kinetic parameters of tissue respiration. Journal of Theoretical Biology, 78(1):51– 59, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903254.

Maniloff:1975:TCD

[Man75a] Jack Maniloff. Thermodynamic considerations of the deoxyribonucleic acid helix-cruciform transition. Journal of Theoretical Biology, 53(2):497–501, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800191.

Manning:1975:GDC

[Man75b] J. T. Manning. Gamete dimorphism and the cost of sexual reproduction: Are they separate phenomena? Journal of Theoretical Biology, 55(2):393–395, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800889.

Manning:1975:OIS

[Man75c] J. T. Manning. The optimization of immune system “decisions”. Journal of Theoretical Biology, 52(1):219–222, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900521.

Manning:1975:SRP

[Man75d] J. T. Manning. Sexual reproduction and parent-offspring conflict in the RNA “tumour” virus/host relationship: Implications for vertebrate oncogene evolution. Journal of Theoretical Biology, 55(2):397–413, December 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800890.

Manabe:1978:TAC

[Man78] Tomohiro Manabe. Time average in chemical reaction and its application to the biological system. Journal of Theoretical Biology, 74(1):143–148, September 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902941.

Mandecki:1979:KMI

[Man79a] W. Mandecki. A kinetic model for interaction of regulatory proteins and RNA polymerase with the control region of the lac operon of Escherichia coli. Journal of Theoretical Biology, 81(1):105–122, November 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900845.

Maniloff:1979:MGM

[Man79b] Jack Maniloff. Mycoplasma gliding may be a biased Brownian motion. Journal of Theoretical Biology, 81(3):617–620, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900602.

Margulis:1970:RNC

[Mar70] Lynn Margulis. Recombination of non-chromosomal genes in Chlamydomonas: Assortment of mitochondria and chloroplasts? Journal of Theoretical Biology, 26(2):337–342, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800236.

Markowitz:1971:CMW

[Mar71a] D. Markowitz. Collective modes and wavelike solutions to cellular control equations. Journal of Theoretical Biology, 31(3): 475–492, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900233.

Markowttz:1971:GCN

[Mar71b] David Markowttz. Is generalized complementarity necessary? Journal of Theoretical Biology, 30(1):211–213, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900464.

Markowitz:1972:COM

[Mar72a] David Markowitz. Co-operativity model of cellular control: Organization and complexity. Journal of Theoretical Biology, 35(1):27–53, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901907.

Martinez:1972:MCD

[Mar72b] Hugo M. Martinez. Morphogenesis and chemical dissipative structures: A computer simulated case study. Journal of Theoretical Biology, 36(3):479–501, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900045.

Markowitz:1973:GCR

[Mar73] David Markowitz. Generalized complementarity re-entered. Journal of Theoretical Biology, 40(2):399–402, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901422.

Maroudas:1975:ASCa

[Mar75a] N. G. Maroudas. Adhesion and spreading of cells on charged surfaces. Journal of Theoretical Biology, 49(1):417–424, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800440.

Maroudas:1975:ASCb

[Mar75b] N. G. Maroudas. Adhesion and spreading of cells on charged surfaces. Journal of Theoretical Biology, 49(2):417–424, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901824.

Martynov:1975:MMI

[Mar75c] L. A. Martynov. A morphogenetic mechanism involving instability of initial forth. Journal of Theoretical Biology, 52(2):471– 480, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900132.

Margineanu:1976:WDT

[Măr76a] Doru-Georg Mărgineanu. Water distribution in tissues: A compartmental analysis and some testing criteria. Journal of Theoretical Biology, 61(2):377–391, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900254.

Martin:1976:EMC

[Mar76b] A. R. Martin. The effect of membrane capacitance on nonlinear summation of synaptic potentials. Journal of Theoretical Biology, 59(1):179–187, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800318.

Marshall:1977:OFE

[Mar77] E. A. Marshall. The osmotic flow of an electrolyte through a charged porous membrane. Journal of Theoretical Biology, 66(1):107–135, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903150.

Markvart:1978:ETP

[Mar78a] T. Markvart. Exciton transport in the photosynthetic unit. Journal of Theoretical Biology, 72(1):91–116, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900206.

Markvart:1978:IMC

[Mar78b] T. Markvart. Infinite model of connected photosynthetic units. Journal of Theoretical Biology, 72(1):117–130, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900218.

Maroudas:1979:LAF

[Mar79a] N. G. Maroudas. On the low adhesiveness of fluid phospholipid substrata. Journal of Theoretical Biology, 79(1):101–116, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902595.

Marshall:1979:EAS

[Mar79b] E. A. Marshall. On the exact analytical solution of the Poisson–Boltzmann equation. Journal of Theoretical Biology, 81(3):613–616, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379900596.

Matsuno:1978:EDS

[Mat78] Koichiro Matsuno. Evolution of dissipative system: A theoretical basis of Margalef’s principle on ecosystem. Journal of Theoretical Biology, 70(1):23–31, January 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903004.

May:1973:MEF

[May73] Robert M. May. Mass and energy flow in closed ecosystems: A comment. Journal of Theoretical Biology, 39(1):155–163, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902105.

Mayhew:1974:CBC

[May74] E. Mayhew. Complementarity between cell surfaces. Journal of Theoretical Biology, 47(2):483–484, October 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902100.

May:1975:BPO

[May75] Robert M. May. Biological populations obeying difference equations: Stable points, stable cycles, and chaos. Journal of Theoretical Biology, 51(2):511–524, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900788.

Mazanov:1973:MSP

[Maz73] A. Mazanov. A multi-stage population model. Journal of Theoretical Biology, 39(3):581–587, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900696.

Mazanov:1976:SMP

[Maz76] Alex Mazanov. Stability of multi-pool models with lags. Journal of Theoretical Biology, 59(2):429–442, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901818.

Mazanov:1978:ACM

[Maz78] Alex Mazanov. Acceptor control in model ecosystems. Journal of Theoretical Biology, 71(1):21–38, March 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902114.

Mura:1978:PIE

[MB78] Umberto Mura and Carlo Bauer. PH influence on enzymic activity: The involvement of two active ionized forms of either substrate or enzyme in the reaction. Journal of Theoretical Biology, 75(2):181–188, November 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902291.

Moore:1973:MCM

[MBG73] G. William Moore, John Barnabas, and Morris Goodman. A method for constructing maximum parsimony ancestral amino acid sequences on a given network. Journal of Theoretical Biology, 38(3):459–485, March 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390252X.

Markin:1972:PME

[MC72] V. S. Markin and Yu. A. Chismadjev. Properties of a multicomponent excitable medium. Journal of Theoretical Biology, 36(1):

61–80, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901774.

Mauzerall:1973:NCA

[MC73] D. Mauzerall and A. Chivvis. A novel cyclical approach to the oxygen producing mechanism of photosynthesis. Journal of Theoretical Biology, 42(2):387–395, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900969.

Meunier:1977:AAA

[MCB77] Danielle Meunier and Paule Chevillotte-Brivet. Action de l’antimycine a sur l’activité succinoxydase des mitochondries du tubercule de pomme de terre, et comportement des cytochromes de type b. Journal of Theoretical Biology, 64(1):137–163, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901175.

McClare:1971:CMM

[McC71] C. W. F. McClare. Chemical machines, Maxwell’s demon and living organisms. Journal of Theoretical Biology, 30(1):1–34, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900336.

McClare:1972:DHE

[McC72a] C. W. F. McClare. In defence of the high energy phosphate bond. Journal of Theoretical Biology, 35(2):233–246, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900367.

McClare:1972:MEM

[McC72b] C. W. F. McClare. A “molecular energy” muscle model. Journal of Theoretical Biology, 35(3):569–595, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901518.

McClendon:1975:E

[McC75a] John H. McClendon. Efficiency. Journal of Theoretical Biology, 49(1):213–218, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800294.

McCutchen:1975:DMC

[McC75b] C. W. McCutchen. Do mineral crystals stiffen bone by straitjacketing its collagen? Journal of Theoretical Biology, 51(1): 51–58, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901381.

McCoy:1977:COE

[McC77] J. Wynne McCoy. Complexity in organic evolution. Journal of Theoretical Biology, 68(3):457–458, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790073X.

McCoy:1978:NAP

[McC78] J. Wynne McCoy. New approaches to the problem of macronuclear assortment. Journal of Theoretical Biology, 74(4):475–489, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902357.

McElwain:1978:REO

[McE78] D. L. S. McElwain. A re-examination of oxygen diffusion in a spherical cell with Michaelis–Menten oxygen uptake kinetics. Journal of Theoretical Biology, 71(2):255–263, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902709.

McGavin:1970:ARS

[McG70] Stewart McGavin. Axes of rotational symmetry in protein structures. Journal of Theoretical Biology, 26(2):335–336, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800224.

McGavin:1979:RPS

[McG79] Stewart McGavin. A reconsideration of the possibility of the specific pairing of base pairs. Journal of Theoretical Biology, 77(1):83–99, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901395.

McKean:1973:EHC

[McK73] T. A. McKean. An electromechanical hypothesis for ciliary coordination in ciliated protozoa. Journal of Theoretical Biology, 39(2):315–319, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901008.

McMurtrie:1975:DSL

[McM75] Ross E. McMurtrie. Determinants of stability of large randomly connected systems. Journal of Theoretical Biology, 50(1): 1–11, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900193.

McElwain:1979:MVC

[MCM79] D. L. S. McElwain, R. Callcott, and L. E. Morris. A model of vascular compression in solid tumours. Journal of Theoretical Biology, 78(3):405–415, June 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903394.

McNair:1979:MTF

[McN79] James N. McNair. A model of tentacle function in certain suctorians. Journal of Theoretical Biology, 78(4):593–610, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901929.

Medina:1977:QMP

[MCO77] J. R. Medina and E. Cerdá-Olmedo. A quantitative model of phycomyces phototropism. Journal of Theoretical Biology, 69(4):709–719, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903770.

Medici:1972:INN

[Med72] Michael A. Medici. The immunoprotective niche — a new pathogenic mechanism for syphilis, the systemic mycoses and other infectious diseases. Journal of Theoretical Biology, 36(3):617–625, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900124.

Meganathan:1979:WNP

[Meg79] R. Meganathan. Why has the “nitrogenase plasmid” snubbed the pseudomonads? Journal of Theoretical Biology, 80(2): 301–304, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902133.

Meinhardt:1978:SDC

[Mei78] Hans Meinhardt. Space-dependent cell determination under the control of a morphogen gradient. Journal of Theoretical Biology, 74(2):307–321, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900784.

Menzel:1976:EUM

[Men76] E. Roland Menzel. On the energy upconversion mechanism of light utilization by chlorophyll in photosynthesis. Journal of Theoretical Biology, 56(2):401–416, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800823.

Metzner:1975:WDP

[Met75] Helmut Metzner. Water decomposition in photosynthesis? A critical reconsideration. Journal of Theoretical Biology, 51(1): 201–231, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901484.

Meyer:1973:EOM

[Mey73] Ralph R. Meyer. On the evolutionary origin of mitochondrial DNA. Journal of Theoretical Biology, 38(3):647–663, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902646.

Matela:1979:TEM

[MF79] Raymond J. Matela and Robert J. Fletterick. A topological exchange model for cell self-sorting. Journal of Theoretical Biology, 76(4):403–414, February 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900092.

Mann:1977:RPP

[MFS77] James E. Mann, Elizabeth Foley, and Nick Sperelakis. Resistance and potential profiles in the cleft between two myocardial cells: Electrical analog and computer simulations. Journal of Theoretical Biology, 68(1):1–15, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902235.

Mar:1972:KMO

[MG72] Ted Mar and Govindjee. Kinetic models of oxygen evolution in photosynthesis. Journal of Theoretical Biology, 36(3):427–446, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937290001X.

Morel:1977:SWV

[MG77a] Jean E. Morel and Marcel P. Gingold. On a simple way to verify the experimental agreement of any two-state mathematical model of muscle contraction. Journal of Theoretical Biology, 68(3):437–447, October 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900704.

Mulholland:1977:TAM

[MG77b] Robert J. Mulholland and Carolyne M. Gowdy. Theory and application of the measurement of structure and determination of function for laboratory ecosystems. Journal of Theoretical Biology, 69(2):321–344, November 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901412.

Moore:1973:IAS

[MGB73] G. William Moore, M. Goodman, and J. Barnabas. An iterative approach from the standpoint of the additive hypothesis to the dendrogram problem posed by molecular data sets. Journal of Theoretical Biology, 38(3):423–457, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902518.

Meiske:1978:IRP

[MGNR78] W. Meiske, M. Glende, G. Nürnberg, and J. G. Reich. On the influence of rapid periodic parameter oscillations on the long-time behaviour of cell metabolism. Journal of Theoretical Biology, 71(1):11–19, March 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902102.

Maxham:1974:SDU

[MH74] J. V. Maxham and H. J. Hickman. Substrate diffusion and uptake within spherical bacterial flocs. Journal of Theoretical Biology, 43(2):229–239, February 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800561.

Moore:1979:CLN

[MHB79] G. William Moore, Grover M. Hutchins, and Bernadine H. Bulkley. Certainty levels in the nullity method of symbolic logic application to the pathogenesis of congenital heart malformations. Journal of Theoretical Biology, 76(1):53–81, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903758.

Miyanishi:1979:GLS

[MHC79] K. Miyanishi, A. R. Hoy, and P. B. Cavers. A generalized law of self-thinning in plant populations (self-thinning in plant populations). Journal of Theoretical Biology, 78(3):439–442, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903424.

MacDonald:1975:APK

[MHMS75] R. L. MacDonald, D. Hsu, J. E. Mann, and N. Sperelakis. An analysis of the problem of K^{+} accumulation in the intercalated disk clefts of cardiac muscle. Journal of Theoretical Biology, 51(2):455–473, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900740.

Matsudaira:1977:TMF

[MHY77] Y. Matsudaira, K. Hieda, and T. Yamasaki. On two models for the frequency distribution of lethals in Drosophila observed in F_{3} s derived from a single F_{1} gonadal mosaic female, and the transmissibility. Journal of Theoretical Biology, 68(4):501–513, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901011.

Michod:1979:GAK

[Mic79] Richard Michod. Genetical aspects of kin selection: Effects of inbreeding. Journal of Theoretical Biology, 81(2):223–233, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901620.

Mikhaltsev:1972:GCM

[Mik72] Igor E. Mikhaltsev. On a general concept of memory. Journal of Theoretical Biology, 37(3):587–595, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900926.

Mikelsaar:1975:CAA

[Mik75a] H. N. Mikelsaar. A concept of amino acid archaeorelation: Origin of life and the genetic code. Journal of Theoretical Biology, 50(1):203–212, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900338.

Mikhaltsev:1975:PIC

[Mik75b] Igor E. Mikhaltsev. On the perception of information in CNS memory mechanism. Journal of Theoretical Biology, 49(1):245– 247, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800324.

Mikulecky:1977:SNTb

[Mik77] Donald C. Mikulecky. A simple network thermodynamic method for series-parallel coupled flows II. The non-linear theory, with applications to coupled solute and volume flow in a series membrane. Journal of Theoretical Biology, 69(3):511–541, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901540.

Mikhaltsev:1978:PCM

[Mik78] Igor E. Mikhaltsev. On the physics of CNS memory. Journal of Theoretical Biology, 70(1):33–49, January 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903016.

Miller:1970:TSC

[Mil70] D. R. Miller. Theoretical survival curves for radiation damage in bacteria. Journal of Theoretical Biology, 26(3):383–398, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900913.

Mild:1973:TAT

[Mil73] Kjell Hansson Mild. On the theoretical aspects of tracer exchange in red blood cells. Journal of Theoretical Biology, 39(1): 183–189, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902142.

Miller:1974:SAV

[Mil74] D. R. Miller. Sensitivity analysis and validation of simulation models. Journal of Theoretical Biology, 48(2):345–360, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800056.

Milch:1978:DSI

[Mil78a] James R. Milch. The dynamical stability of isometrically contracting muscle and its relation to transient force response. Journal of Theoretical Biology, 73(4):755–767, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901352.

Miller:1978:PGM

[Mil78b] Judith E. Miller. A proposal for the genetic manipulation of plant metabolism. Journal of Theoretical Biology, 74(1): 153–154, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902965.

Miller:1979:MLM

[Mil79] Rodman G. Miller. A model for lamellar myelin. Journal of Theoretical Biology, 80(1):101–113, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901826.

Mittenthal:1975:SFM

[Mit75] Jay E. Mittenthal. A sliding filament model for skeletal muscle: dependence of isometric dynamics on temperature and sarcomere length. Journal of Theoretical Biology, 52(1):1–16, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900375.

Mitchell:1976:PMM

[Mit76] Peter Mitchell. Possible molecular mechanisms of the protonmotive function of cytochrome systems. Journal of Theoretical Biology, 62(2):327–367, October 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901247.

Mittenthal:1979:MGC

[Mit79a] Jay E. Mittenthal. A model for the genetic control of differential adhesion in insect epidermis. Journal of Theoretical Biology, 81(1):129–149, November 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900869.

Mitz:1979:CBN

[Mit79b] M. A. Mitz. CO 2 biodynamics: a new concept of cellular control. Journal of Theoretical Biology, 80(4):537–551, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900924.

Mizuhashi:1977:PDI

[Miz77] Seiji Mizuhashi. Out-of-plane distortion of the iron in ferrohemoglobin due to pseudo-Jahn–Teller effect. Journal of Theoretical Biology, 66(1):13–20, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903083.

Moore:1971:ISP

[MJ71] L. E. Moore and E. Jakobsson. Interpretation of the sodium permeability changes of myelinated nerve in terms of linear relaxation theory. Journal of Theoretical Biology, 33(1):77–89, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902177.

Mirsky:1976:EAW

[MJ76] I. Mirsky and R. F. Janz. The effect of age on the wall stiffness of the human thoracic aorta: A large deformation “anisotropic” elastic analysis. Journal of Theoretical Biology, 59(2):467–484, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901831.

Mulholland:1974:ALC

[MK74] Robert J. Mulholland and Marvin S. Keener. Analysis of linear compartment models for ecosystems. Journal of Theoretical Biology, 44(1):105–116, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800317.

McMahon:1976:TSD

[MK76] Thomas A. McMahon and Richard E. Kronauer. Tree structures: Deducing the principle of mechanical design. Journal of Theoretical Biology, 59(2):443–466, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690182X.

Morand:1976:SOP

[MKLKV76] Peter Morand, Mathew Kalapurackal, Ludmila Lompa-Krzymien, and Alex Van Tongerloo. Steroid oxides as potential precursors in the biosynthesis of estrogens. Journal of Theoretical Biology, 56(2):503–506, February 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800902.

Mazat:1977:DSE

[MLM77] Jean-Pierre Mazat, Joél Langla, and Françoise Mazat. Doublesite enzymes and squatting a study of the regulation by one or several ligands binding at two different classes of site. Journal of Theoretical Biology, 68(3):365–383, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900662.

Mimura:1978:DPP

[MM78a] M. Mimura and J. D. Murray. On a diffusive preypredator model which exhibits patchiness. Journal of Theoretical Biology, 75(3):249–262, December 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903326.

Mulas:1978:ESC

[MM78b] Pablo Mulás and Donald A. McQuarrie. Effect of surface charge density on the kinetics of transmembrane lipid migration in asymmetric bilayers. Journal of Theoretical Biology, 72(3):475– 486, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903132.

Murakami:1979:COA

[MM79] Hideo Murakami and Hiromasa Mori. On the codings and origins of t and t antigens in SV40. Journal of Theoretical Biology, 81(4):809–822, December 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902832.

Miles:1974:MOS

[MME74] Daniel W. Miles, Douglas L. Miles, and H. Eyring. A molecular orbital study of the conformation of formycin. Journal of Theoretical Biology, 45(2):577–583, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901313.

Miles:1977:CBS

[MMRE77] Douglas L. Miles, Daniel W. Miles, Patrick Redington, and Henry Eyring. A conformational basis for the selective action of ara-adenine. Journal of Theoretical Biology, 67(3):499–514, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900522.

Macdonald:1974:DGE

[MMS74] Robert L. Macdonald, James E. Mann, and Nick Sperelakis. Derivation of general equations describing tracer diffusion in any two-compartment tissue with application to ionic diffusion in cylindrical muscle bundles. Journal of Theoretical Biology, 45(1):107–130, May 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900460.

Murakami:1973:ACM

[MMT73] Hideo Murakami, Hiromasa Mori, and Sadako Taira. Active centre model on the structure and function of the bleomycin molecule as the prototype of esteratic enzymes. Journal of Theoretical Biology, 42(3):443–460, December 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902403.

Murakami:1976:MEB

[MMT76] Hideo Murakami, Hiromasa Mori, and Sadako Taira. On the molecular embrace of bleomycin with helical DNA. Journal of Theoretical Biology, 59(1):1–23, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800215.

Markowitz:1973:COB

[MN73] D. Markowitz and R. M. Nisbet. Co-operativity in biological machines. Journal of Theoretical Biology, 39(3):653–657, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900763.

Mitchell:1972:VWF

[MNR72] D. J. Mitchell, B. W. Ninham, and P. Richmond. Van der Waals forces between thin anisotropic cylinders. Journal of Theoretical Biology, 37(2):251–259, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900203.

MacGregor:1973:GPE

[MO73] R. J. MacGregor and R. M. Oliver. A general-purpose electronic model for arbitrary configurations of neurons. Journal of Theoretical Biology, 38(3):527–534, IN1–IN8, 535–538, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902543.

Maruyama:1975:ODGa

[MO75a] Mutsumi Maruyama and Fumio Oosawa. Orientation distribution of globular protein molecules in a two-dimensional lattice: II. Thermal effect. Journal of Theoretical Biology, 49(1):249–254, IN1, 255–262, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800336.

Maruyama:1975:ODGb

[MO75b] Mutsumi Maruyama and Fumio Oosawa. Orientation distribution of globular protein molecules in a two-dimensional lattice: II. Thermal effect. Journal of Theoretical Biology, 49(2):249–254, IN1, 255–262, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590171X.

Mcminn:1976:CEP

[MO76] C. Linda Mcminn and J. H. Ottaway. On the control of enzyme pathways. Journal of Theoretical Biology, 56(1):57–73, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800458.

Mobley:1973:SAE

[Mob73] Curtis D. Mobley. A systematic approach to ecosystems analysis. Journal of Theoretical Biology, 42(1):119–136, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901525.

Moffat:1970:QMS

[Mof70] J. B. Moffat. A quantum-mechanical study of the peptide molecule using Gaussian orbitals. Journal of Theoretical Biology, 26(3):437–443, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900950.

Moffat:1973:PPN

[Mof73] J. B. Moffat. The protonation of peptides: A non-empirical theoretical study. Journal of Theoretical Biology, 40(2):247–258, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901306.

Molz:1976:WTT

[Mol76] Fred J. Molz. Water transport through plant tissue: the apoplasm and symplasm pathways. Journal of Theoretical Biology, 59(2):277–292, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901703.

Moore:1971:CMA

[Moo71] G. E. Moore. Cytological model of antibody production. Journal of Theoretical Biology, 32(3):625–629, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901597.

Moore:1974:DUD

[Moo74] David Moore. Dynamic unwinding of DNA helices: A mechanism for genetic recombination. Journal of Theoretical Biology, 43(1):167–186, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800512.

Moore:1977:PPP

[Moo77] G. William Moore. Proof of the populous path algorithm for missing mutations in parsimony trees. Journal of Theoretical Biology, 66(1):95–106, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903149.

Moraal:1972:MEB

[Mor72] H. Moraal. A model for the effects of binding and interaction energies on the kinetics of multivalent enzymes. Journal of Theoretical Biology, 37(2):283–290, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900239.

Morgan:1977:ARM

[Mor77] Richard S. Morgan. Actin rotates as myosin translates. Journal of Theoretical Biology, 67(4):769–771, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902612.

Morel:1978:FVR

[Mor78] Jean E. Morel. Force-velocity relationship in single muscle fibres. Journal of Theoretical Biology, 73(3):445–451, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901522.

Mountford:1971:PSV

[Mou71] M. D. Mountford. Population survival in a variable environment. Journal of Theoretical Biology, 32(1):75–79, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901366.

Moynihan:1970:CSD

[Moy70] M. Moynihan. Control, suppression, decay, disappearance and replacement of displays. Journal of Theoretical Biology, 29(1):85–112, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901207.

McGinness:1973:IFM

[MP73] John McGinness and Peter Proctor. The importance of the fact that melanin is black. Journal of Theoretical Biology, 39(3): 677–678, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900817.

MacWilliams:1978:MGI

[MP78] Harry K. MacWilliams and Spyros Papageorgiou. A model of gradient interpretation based on morphogen binding. Journal of Theoretical Biology, 72(3):385–406, IN2, 407–411, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903089.

Medina:1979:LHH

[MP79] J. R. Medina and C. Petit. L’effet hitch-hiking comme processus dispersif. Journal of Theoretical Biology, 81(2):235–246, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901632.

Maigret:1970:MOCa

[MPD70] B. Maigret, B. Pullman, and M. Dreyfus. Molecular orbital calculations on the conformation of polypeptides and proteins: I. Preliminary investigations and simple dipeptides. Journal of Theoretical Biology, 26(2):321–333, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800212.

Malone:1972:SMD

[MPH72] George H. Malone, Stephen Prager, and Thos. E. Hutchinson. A simple model for diffusion in independent, temporally fluctuating pores. Journal of Theoretical Biology, 36(2):379–395, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901063.

Morel:1976:MCC

[MPHG76] Jean E. Morel, Ingrid Pinset-Härström, and Marcel P. Gingold. Muscular contraction and cytoplasmic streaming: A new general hypothesis. Journal of Theoretical Biology, 62(1):17–26, IN1, 27–51, October 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900503.

Maigret:1970:MOCb

[MPP70] Bernard Maigret, David Perahia, and Bernard Pullman. Molecular orbital calculations on the conformation of polypeptides and proteins IV. The conformation of the prolyl and hydroxyprolyl residues. Journal of Theoretical Biology, 29(2):275–291, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900226.

Maigret:1971:MOC

[MPP71] Bernard Maigret, Bernard Pullman, and David Perahia. Molecular orbital calculations on the conformation of polypeptides and proteins: VII. refined calculations on the alanyl residue. Journal of Theoretical Biology, 31(2):269–285, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901871.

Mar:1974:KMP

[MR74a] Ted Mar and Guy Roy. A kinetic model of the primary back reaction in photosynthesis of green plants. Journal of Theoretical Biology, 48(2):257–281, ???? 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800019.

Moore:1974:NIH

[MR74b] John W. Moore and Fidel Ramon. On numerical integration of the Hodgkin and Huxley equations for a membrane action potential. Journal of Theoretical Biology, 45(1):249–273, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490054X.

Mrosovsky:1973:TLP

[Mro73] N. Mrosovsky. Temperature and learning in poikilotherms. Journal of Theoretical Biology, 39(3):659–663, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900775.

Moore:1977:AQN

[MRS77] G. William Moore, U. N. Riede, and W. Sandritter. Application of Quine’s nullities to a quantitative organelle pathology. Journal of Theoretical Biology, 65(4):633–651, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900133.

Mishra:1979:SIA

[MRTG79] R. K. Mishra, G. Subba Rao, R. S. Tyagi, and S. Gandhi. A study of the intermolecular association in living systems. IV. Studies on a model molecular fragment (pentane). Journal of Theoretical Biology, 80(3):435–440, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901036.

Magar:1971:ECM

[MS71] Magar E. Magar and Robert F. Steiner. Equivalence of certain models in protein ligand equilibria and the possibility of distinguishing between them. Journal of Theoretical Biology, 32(3):495–506, September 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901536.

Magar:1971:APL

[MSF71] Magar E. Magar, Robert F. Steiner, and John E. Fletcher. Analysis of protein ligand equilibria. Journal of Theoretical Biology, 32(1):59–74, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901354.

Maranon:1978:SCC

[MSG^{+}78] Julio Marañon, Oscar M. Sorarrain, Horacio Grinberg, Samuel Lamdan, and Carlos H. Gaozza. Semiempirical CNDO/2 calculation of the electronic structure of the DNA molecule I. Ground state potential curves; tunneling and tautomeric equilibrium in the N − H⋯N and O⋯H − N bonds of the adenine–thymine base pair. Journal of Theoretical Biology, 74(1):11–22, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902862.

Metcalf:1979:POC

[MSK79] Robert A. Metcalf, Judy A. Stamps, and V. V. Krishnan. Parent-offspring conflict that is not limited by degree of kinship. Journal of Theoretical Biology, 76(1):99–107, January 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903771.

Muller:1973:CSE

[MT73] F. J. Muller and W. K. Taylor. A comparative study of electronic and neural networks involved in pattern recognition. Journal of Theoretical Biology, 41(1):97–118, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901914.

Marshall:1974:MMU

[MT74a] E. A. Marshall and E. A. Trowbridge. A mathematical model of the ultra-filtration process in a single glomerular capillary. Journal of Theoretical Biology, 48(2):389–412, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800081.

Mazanov:1974:TSE

[MT74b] Alex Mazanov and Keith P. Tognetti. Taylor series expansion of delay differential equations — a warning. Journal of Theoretical Biology, 46(1):271–282, July 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901520.

Moffat:1976:TSR

[MT76] J. B. Moffat and K. F. Tang. A theoretical study of the reactive dimerization of HCN. Journal of Theoretical Biology, 58(1): 83–95, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901405.

McGuire:1978:REO

[MT78a] James B. McGuire and Colin J. Thompson. On the reconstruction of an evolutionary order. Journal of Theoretical Biology, 75(2):141–147, November 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902266.

Mikulecky:1978:SNT

[MT78b] D. C. Mikulecky and S. R. Thomas. A simple network thermodynamic method for series-parallel coupled flows III. Application to coupled solute and volume flows through epithelial membranes. Journal of Theoretical Biology, 73(4):697–710, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901315.

Murakami:1979:GTE

[MT79] Chikara Murakami and Kazuhisa Tomita. Giant transitory excitation — a thermokinetic model. Journal of Theoretical Biology, 79(2):203–222, July 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902467.

Miller:1976:TEC

[MTL76] C. E. Miller, S. Thompson, and J. Lozar. A theoretical evaluation of cardiac output as a function of mean arterial pressure in the human cardiovascular system. Journal of Theoretical Biology, 63(1):89–98, November 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900850.

Murakami:1977:MDR

[MTM77] Hideo Murakami, Sadako Taira, and Hiromasa Mori. Modes of DNA replication primed with the oligomer of palindrome. Journal of Theoretical Biology, 68(2):183–197, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901588.

Murray:1974:RMM

[Mur74a] J. D. Murray. On the role of myoglobin in muscle respiration. Journal of Theoretical Biology, 47(1):115–126, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901027.

Murray:1974:CTO

[Mur74b] M. J. Murray. Conductile trees: Obtaining structural and functional information from endpoint measurements. Journal of Theoretical Biology, 43(1):113–132, January 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800482.

Murray:1975:NEW

[Mur75] J. D. Murray. Non-existence of wave solutions for the class of reaction-diffusion equations given by the Volterra interacting-population equations with diffusion. Journal of Theoretical Biology, 52(2):459–469, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900120.

Murray:1976:TWS

[Mur76] J. D. Murray. On travelling wave solutions in a model for the Belousov–Zhabotinskii reaction. Journal of Theoretical Biology, 56(2):329–353, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800781.

Marbach:1975:MME

[MV75] G. Marbach and P. M. Vignais. A molecular mechanism of energy transduction at a cytochrome level. Journal of Theoretical Biology, 54(2):335–343, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801354.

Michel-Villaz:1976:FPP

[MV76] M. Michel-Villaz. Fluorescence polarization: Pigment orientation and energy transfer in photosynthetic membranes. Journal of Theoretical Biology, 58(1):113–129, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901429.

Marshall:1978:ITD

[MW78] E. A. Marshall and D. W. Windle. Ion transfer during electrolyte osmosis through a charged porous membrane. Journal of Theoretical Biology, 73(3):383–415, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901492.

Miller:1973:PSI

[MWH73] D. R. Miller, D. E. Weidhaas, and R. C. Hall. Parameter sensitivity in insect population modeling. Journal of Theoretical Biology, 42(2):263–274, November 15, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900891.

MacDonald:1976:RNC

[MWM76] Ian A. MacDonald, C. Noel Williams, and David E. Mahony. A rapid non-chromatographic analysis of individual bile acids in human bile extracts. Journal of Theoretical Biology, 57(2): 385–389, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900096.

Mikulecky:1977:SNTa

[MWS77] D. C. Mikulecky, W. A. Wiegand, and J. S. Shiner. A simple network thermodynamic method for modeling seriesparallel coupled flows I. The linear case. Journal of Theoretical Biology, 69(3):471–510, December 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901539.

Macdonald:1977:SPN

[MYF77] R. D. Macdonald, T. Yamamoto, and P. Fedorak. Statistical prediction of the number of cells surviving infection by an autointerfering virus using the Poisson distribution. Journal of Theoretical Biology, 68(3):355–363, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900650.

Mangoubi:1979:TAP

[MZ79] S. S. Mangoubi and Y. Y. Zeevi. Theoretical analysis of photoreceptor noise. Journal of Theoretical Biology, 81(4):621–632, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937990273X.

Nagylaki:1976:MES

[Nag76] Thomas Nagylaki. A model for the evolution of self-fertilization and vegetative reproduction. Journal of Theoretical Biology, 58(1):55–58, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901387.

Nagy:1978:MHA

[Nag78] Imre Zs. Nagy. A membrane hypothesis of aging. Journal of Theoretical Biology, 75(2):189–195, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902308.

Nanjundiah:1973:CSR

[Nan73] Vidyanand Nanjundiah. Chemotaxis, signal relaying and aggregation morphology. Journal of Theoretical Biology, 42 (1):63–105, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901495.

Nanjundiah:1976:SRS

[Nan76] Vidyanand Nanjundiah. Signal relay by single cells during wave propagation in a cellular slime mold. Journal of Theoretical Biology, 56(2):275–282, February 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800756.

Nazarea:1977:ICL

[Naz77] Apolinario D. Nazarea. Intrinsic critical length associated with oscillatory reactive processes in complex solutions: Hydrodynamic estimates. Journal of Theoretical Biology, 64(2):311–322, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903599.

Nelson:1972:SPA

[NB72] S. C. Nelson and J. D. Barchas. Sub-picogram assay of biogenic materials: Theoretical application of interactions with isotopes that yield heavy ionizing particles upon neutron exposure and use of radiographic dielectric track registration techniques. Journal of Theoretical Biology, 34(1):73–85, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900550.

Nicolis:1976:MRN

[NB76] John S. Nicolis and Marianna Benrubi. A model on the role of noise at the neuronal and the cognitive levels. Journal of Theoretical Biology, 59(1):77–96, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800240.

Nelson:1975:ECA

[NCM75] A. P. Nelson, P. Colonomos, and D. A. McQuarrie. Electrostatic coupling across a membrane with titratable surface groups. Journal of Theoretical Biology, 50(2):317–325, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900843.

Naora:1979:SBT

[NDF79] H. Naora, N. J. Deacon, and K. E. Fry. Similarity between 5’- and 3’-terminal nucleotide sequences and doublestranded RNA-derived sequences of eukaryotic mRNA. Journal of Theoretical Biology, 80(2):205–221, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902066.

Neal:1972:AMK

[Nea72] Jack L. Neal. Analysis of Michaelis kinetics for two independent, saturable membrane transport functions. Journal of Theoretical Biology, 35(1):113–118, April 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901968.

Nedelcu:1975:TKS

[Ned75] Constantin Nedelcu. Thermodynamic and kinetic study of blood filtration. Journal of Theoretical Biology, 55(2):505–517, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937580097X.

Nelson:1975:GSF

[Nel75] Jeremiah I. Nelson. Globality and stereoscopic fusion in binocular vision. Journal of Theoretical Biology, 49(1):1–88, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800208.

Nelson:1977:PCA

[Nel77] Jeremiah I. Nelson. The plasticity of correspondence: Aftereffects, illusions and horopter shifts in depth perception. Journal of Theoretical Biology, 66(2):203–266, May 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901709.

Nelson:1978:EHC

[Nel78] Clifford V. Nelson. The effects of heart-cavity blood resistivity and volume on endocardial potentials: theoretical predictions. Journal of Theoretical Biology, 71(3):387–399, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901674.

Newman:1970:NCS

[New70] Stuart A. Newman. Note on complex systems. Journal of Theoretical Biology, 28(3):411–413, September 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900780.

Newman:1972:SSM

[New72] Stuart A. Newman. A source of stability in metabolic networks. Journal of Theoretical Biology, 35(2):227–232, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900355.

Newman:1979:BPS

[New79] Stuart A. Newman. The un body: A possible structure for transcriptionally inactivated chromatin subunits. Journal of Theoretical Biology, 79(1):55–66, July 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990256X.

Nisbet:1976:PDP

[NG76] R. M. Nisbet and W. S. C. Gurney. Population dynamics in a periodically varying environment. Journal of Theoretical Biology, 56(2):459–475, February 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800860.

Nigon:1977:GAC

[NG77] Victor Nigon and Jacqueline Godet. Genetic analysis of cell differentiation: Possible extension of the haemoglobin differon model to Drosophila morphogenesis and immunoglobulin determination. Journal of Theoretical Biology, 64(1):97–111, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901151.

Nisbet:1977:ELM

[NGP77] R. M. Nisbet, W. S. C. Gurney, and M. A. Pettipher. An evaluation of linear models of population fluctuations. Journal of Theoretical Biology, 68(1):143–160, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902338.

Nisbet:1978:EFT

[NGP78] R. M. Nisbet, W. S. C. Gurney, and M. A. Pettipher. Environmental fluctuations and the theory of the ecological niche. Journal of Theoretical Biology, 75(2):223–237, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902321.

Nickerson:1973:BFM

[Nic73] Kenneth W. Nickerson. Biological functions of multistable proteins. Journal of Theoretical Biology, 40(3):507–515, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900076.

Nichols:1976:SCB

[Nic76] David E. Nichols. Structural correlation between apomorphine and LSD: Involvement of dopamine as well as serotonin in the actions of hallucinogens. Journal of Theoretical Biology, 59(1): 167–177, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800306.

Nieder:1971:SPC

[Nie71] P. C. Nieder. Suggested principle of cochlear organization: The addressed exponential delay line. Journal of Theoretical Biology, 31(2):371–374, May 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901950.

Ninio:1979:AGL

[Nin79] Jacques Ninio. An algorithm that generates a large number of geometric visual illusions. Journal of Theoretical Biology, 79(2):167–201, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902455.

Nir:1975:LRI

[Nir75] S. Nir. Long range intermolecular forces between macroscopic bodies: Macroscopic and microscopic approaches. Journal of Theoretical Biology, 53(1):83–100, September 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901046.

Niven:1977:MMA

[Niv77] B. S. Niven. A mathematical model for animal ecology with special reference to the ecology of the tribolium beetle. Journal of Theoretical Biology, 69(2):357–369, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901436.

Neville:1972:MCS

[NKB72] Margaret C. Neville, Robert B. Kelman, and Franklin W. Briese. Monte Carlo simulation of co-operative adsorption of ligands onto polymers. Journal of Theoretical Biology, 35(1): 129–144, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901981.

Noordergraaf:1979:MHN

[NLC79] Abraham Noordergraaf, John K.-J. Li, and Kenneth B. Campbell. Mammalian hemodynamics: a new similarity principle. Journal of Theoretical Biology, 79(4):485–489, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990239X.

Nelson:1972:EHC

[NM72] R. L. Nelson and H. S. Mason. An explicit hypothesis for chemical carcinogenesis. Journal of Theoretical Biology, 37(1):197– 200, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901270.

Nelson:1975:EDC

[NM75] A. P. Nelson and D. A. McQuarrie. The effect of discrete charges on the electrical properties of a membrane. I. Journal of Theoretical Biology, 55(1):13–27, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801068.

Noy-Meir:1978:SSG

[NM78] I. Noy-Meir. Stability in simple grazing models: Effects of explicit functions. Journal of Theoretical Biology, 71(3):347– 380, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901650.

Nicholls:1978:MMP

[NMK78] P. Nicholls, Judith M. Martin, and J. Kahn. A mathematical model predicting chromosome loss in cultured cells of young adults. Journal of Theoretical Biology, 73(2):237–245, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901881.

Nishikaw:1974:CHT

[NO74] Ken Nishikaw and Tatsuo Ooi. Comparison of homologous tertiary structures of proteins. Journal of Theoretical Biology, 43(2):351–374, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800664.

Nichol:1972:BDL

[NOB72] L. W. Nichol, K. O’Dea, and P. A. Baghurst. Binding of dissimilar ligand molecules to an interacting acceptor: A model for the action of effectors. Journal of Theoretical Biology, 34(2):255– 263, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901592.

Noldus:1973:PMA

[Nol73] E. Noldus. A perturbation method for the analysis of impulse propagation in a mathematical neuron model. Journal of Theoretical Biology, 38(2):383–395, February 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901811.

Noldus:1976:OCA

[Nol76] Erik J. Noldus. Optimal control aspects of left ventricular ejection dynamics. Journal of Theoretical Biology, 63(2):275–309, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900357.

Norwich:1971:CDT

[Nor71] K. H. Norwich. Convective diffusion of tracers. Journal of Theoretical Biology, 32(1):47–57, July 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901342.

Norman:1975:DSI

[Nor75a] Richard S. Norman. Diffusional spread of iontophoretically injected ions. Journal of Theoretical Biology, 52(1):159–162, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900478.

Norwich:1975:DPF

[Nor75b] Kenneth H. Norwich. Determination of pulsatile flow rates by indicator-dilution methods. Journal of Theoretical Biology, 50(2):353–361, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900867.

Nossal:1977:MAC

[Nos77] Ralph Nossal. Mathematical analysis of a capillary migration assay for cellular immune sensitivity: Simple theory for dose-response. Journal of Theoretical Biology, 64(4):703–704, IN1, 705–722, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902697.

Novotny:1973:GIP

[Nov73] Jirí Novotný. Genealogy of immunoglobulin polypeptide chains: A consequence of amino acid interactions, conserved in their tertiary structures. Journal of Theoretical Biology, 41(1):171–180, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901975.

Novotny:1975:IPS

[Nov75] Jirí Novotný. Immunoglobulin phenotype as a solution of the antibody dilemma. Journal of Theoretical Biology, 53(1):67–82, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901034.

Novotny:1976:NSJ

[Nov76] Jirí Novotný. A note on the stability of Jerne’s immune network. Journal of Theoretical Biology, 60(2):487–491, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900771.

Ninham:1971:EPB

[NP71] Barry W. Ninham and V. Adrian Parsegian. Electrostatic potential between surfaces bearing ionizable groups in ionic equilibrium with physiologic saline solution. Journal of Theoretical Biology, 31(3):405–428, June 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900191.

Nir:1979:TMA

[NP79] Avinoam Nir and Robert Pfeffer. Transport of macromolecules across arterial wall in the presence of local endothial injury. Journal of Theoretical Biology, 81(4):685–711, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902777.

Nicolis:1977:CMC

[NPV77] J. S. Nicolis, E. N. Protonotarios, and I. Voulodemou. Control Markov chain models for biological hierarchies. Journal of Theoretical Biology, 68(4):563–581, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901060.

Nevo:1970:DAS

[NR70] Abraham C. Nevo and Robert Rikmenspoel. Diffusion of ATP in sperm flagella. Journal of Theoretical Biology, 26(1):11– 18, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519370800273.

Noyes:1971:ULM

[NR71] David H. Noyes and Warren S. Rehm. Unstirred-layer model for the long time-constant transient voltage response to current in epithelial tissue. Journal of Theoretical Biology, 32(1):25–45, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901330.

Nir:1972:ACA

[NRW72] S. Nir, Robert Rein, and L. Weiss. On the applicability of certain approximations of the Lifshitz theory to thin films. Journal of Theoretical Biology, 34(1):135–153, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900604.

Nir:1973:FCA

[NRW73] S. Nir, R. Rein, and L. Weiss. Further comments on the applicability of certain approximations of the Lifshitz theory to thin films. Journal of Theoretical Biology, 41(3):561–566, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900623.

Ninham:1976:RIA

[NS76a] B. W. Ninham and R. A. Sammut. Refractive index of arrays of spheres and cylinders. Journal of Theoretical Biology, 56(1):125– 149, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800495.

Novak:1976:PSM

[NS76b] Bohuslav Novak and Friedrich F. Seelig. Phase-shift model for the aggregation of amoebae: A computer study. Journal of Theoretical Biology, 56(2):301–327, February 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937680077X.

Norman:1978:MPT

[NS78] Jack C. Norman and Paul E. Sager. Modelling phosphorus transfer rates in lake water. Journal of Theoretical Biology, 71(3):381–385, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901662.

Nastansky:1974:ISG

[NSS74] L. Nastansky, S. M. Selkow, and N. F. Stewart. An improved solution to the generalized camin-sokal model for numerical cladistics. Journal of Theoretical Biology, 48(2):413–424, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800093.

Norman:1977:SES

[NTR77] Richard F. Norman, Peter D. Taylor, and Raleigh J. Robertson. Stable equilibrium strategies and penalty functions in a game of attrition. Journal of Theoretical Biology, 65(3):571–578, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902156.

Nossal:1973:ADA

[NW73] Ralph Nossal and George H. Weiss. Analysis of a densitometry assay for bacterial chemotaxis. Journal of Theoretical Biology, 41(1):143–147, September 14, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390194X.

Nossal:1974:DTC

[NW74] Ralph Nossal and George H. Weiss. A descriptive theory of cell migration on surfaces. Journal of Theoretical Biology, 47(1):103–113, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901015.

Nichol:1979:EFS

[NWW79] L. W. Nichol, P. R. Wills, and D. J. Winzor. An examination of the forms of scatchard plots of binding results outside domains of sigmoidality. Journal of Theoretical Biology, 80(1):39–50, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901784.

Nyberg:1975:ESP

[Nyb75] Dennis Nyberg. Evolution of sex I. Primitive sex. Journal of Theoretical Biology, 50(2):429–436, April 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900910.

Nyholm:1978:DPL

[Nyh78] Niels Nyholm. Dynamics of phosphate limited algal growth: Simulation of phosphate shocks. Journal of Theoretical Biology, 70(4):415–425, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902503.

Ostrander:1973:FBM

[OB73] Lee E. Ostrander and Joe E. Brand. Functional Boolean models for systems with continuous variables. Journal of Theoretical Biology, 38(3):665–675, March 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902658.

Oster:1971:TTT

[OD71] G. F. Oster and C. A. Desoer. Tellegen’s theorem and thermodynamic inequalities. Journal of Theoretical Biology, 32(2):219– 241, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901627.

Otten:1973:ESS

[OD73] H. A. Otten and L. N. M. Duysens. An extension of the steady-state approximation of the kinetics of enzyme-containing systems. Journal of Theoretical Biology, 39(2):387–396, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901070.

Odell:1976:NMC

[Ode76] Garrett M. Odell. A new mathematical continuum theory of axoplasmic transport. Journal of Theoretical Biology, 60(1):223– 237, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901661.

Odutola:1977:TCM

[Odu77] Akintola B. Odutola. A theoretical consideration of a model simulating the functional organization of the thoracolumbar sympathetic subsystem of the spinal cord. Journal of Theoretical Biology, 67(1):1–10, July 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901813.

Odell:1975:CTM

[OF75] Garrett M. Odell and H. L. Frisch. A continumn theory of the mechanics of amoeboid pseudopodium extension. Journal of Theoretical Biology, 50(1):59–86, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900247.

Ordaq:1977:BMB

[OF77] George W. Ordaq and Russell B. Fields. A biochemical mechanism for bacterial chemotaxis. Journal of Theoretical Biology, 68(4):491–500, October 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790100X.

Offner:1971:NPP

[Off71] Franklin F. Offner. Nernst–Planck–Poisson diffusion equation: Numerical solution of the boundary value problem. Journal of Theoretical Biology, 31(2):215–227, May 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901834.

Offner:1974:STD

[Off74] Franklin F. Offner. Solution of the time-dependent ionic diffusion equation. Journal of Theoretical Biology, 45(1):81–91, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900447.

Ortega:1974:PHR

[OG74] Joseph K. E. Ortega and R. Igor Gamow. The problem of handedness reversal during the spiral growth of phycomyces. Journal of Theoretical Biology, 47(2):317–324, IN3–IN5, 325–332, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490201X.

Ohki:1970:PLB

[Ohk70] Shinpei Ohki. Properties of lipid bilayer membranes membrane thickness. Journal of Theoretical Biology, 26(2):277–287, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800182.

Ohki:1973:ENC

[Ohk73] Shinpei Ohki. Evidence for a new concept of membrane potential. Journal of Theoretical Biology, 42(3):593–596, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902506.

Ohshima:1977:MEI

[Ohs77] Hiroyuki Ohshima. A model for the electrostatic interaction of cells. Journal of Theoretical Biology, 65(3):523–530, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902120.

Odell:1976:TBC

[OK76a] G. M. Odell and E. F. Keller. Traveling bands of chemotactic bacteria revisited. Journal of Theoretical Biology, 56(1):243– 247, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800550.

Offner:1976:RIA

[OK76b] Franklin F. Offner and S. H. Kim. Role of ionic adsorption in the excitable membrane. Journal of Theoretical Biology, 61(1):113–127, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901089.

Offner:1976:CEA

[OK76c] Franklin F. Offner and San Hyung Kim. Conformational effects on the activation free energy of diffusion through membranes as influenced by electric fields. Journal of Theoretical Biology, 61(1):97–112, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901077.

Olien:1973:TCF

[Oli73] C. R. Olien. Thermodynamic components of freezing stress. Journal of Theoretical Biology, 39(1):201–210, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902178.

Oliver:1977:MYM

[Oli77] Stephen G. Oliver. On the mutability of the yeast mitochondrial genome. Journal of Theoretical Biology, 67(2):195–201, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790193X.

Olovnikov:1973:TMI

[Olo73] A. M. Olovnikov. A theory of marginotomy: The incomplete copying of template margin in enzymic synthesis of polynucleotides and biological significance of the phenomenon.

Journal of Theoretical Biology, 41(1):181–190, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901987.

Olsen:1975:FPMa

[Ols75a] Richard W. Olsen. Filamentous protein model for cyclic AMPmediated cell regulatory mechanisms. Journal of Theoretical Biology, 49(1):263–287, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800348.

Olsen:1975:FPMb

[Ols75b] Richard W. Olsen. Filamentous protein model for cyclic AMP-mediated cell regulatory mechanisms. Journal of Theoretical Biology, 49(2):263–287, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901721.

Olson:1979:LCT

[Ols79] Kenneth R. Olson. The linear cable theory as a model of gill blood flow. Journal of Theoretical Biology, 81(3):377–388, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900419.

Oosawa:1972:ODG

[OMF72] Fumio Oosawa, Mutsumi Maruyama, and Satoru Fujime. Orientation distribution of globular protein molecules in a two-dimensional lattice: Computer simulation. Journal of Theoretical Biology, 36(2):203–214, IN1, 215–221, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900938.

Oosawa:1977:BMO

[ON77] Fumio Oosawa and Yasuo Nakaoka. Behavior of microorganisms as particles with internal state variables. Journal of Theoretical Biology, 66(4):747–761, June 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902430.

Onda:1979:NHM

[Ond79] Hiromasa Onda. A new hypothesis on mitotic control mechanism in eukaryotic cells: Cell-specific mitosis-inhibiting protein excretion hypothesis. Journal of Theoretical Biology, 77(3):367– 377, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937990362X.

Ohki:1976:MOW

[OO76] Shinpei Ohki and Catherine B. Ohki. Monolayers at the oil/water interface as a proper model for bilayer membranes. Journal of Theoretical Biology, 62(2):389–407, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901260.

Oobatake:1977:ANB

[OO77] Motohisa Oobatake and Tatsuo Ooi. An analysis of non-bonded energy of proteins. Journal of Theoretical Biology, 67(3):567– 584, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900583.

Oonuki:1976:FPN

[Oon76] Makoto Oonuki. Firing probability of a neural network: First-order differential equation. Journal of Theoretical Biology, 58(1):59–81, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901399.

Oosawa:1970:SDP

[Oos70] Fumio Oosawa. Size distribution of protein polymers. Journal of Theoretical Biology, 27(1):69–86, April 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901293.

Oosawa:1973:FFI

[Oos73] Fumio Oosawa. Field fluctuation in ionic solutions and its biological significance. Journal of Theoretical Biology, 39(2):373– 386, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901069.

Oosawa:1975:EFF

[Oos75] Fumio Oosawa. The effect of field fluctuation on a macromolecular system. Journal of Theoretical Biology, 52(1):175–186, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900491.

Oplatka:1972:MMC

[Opl72] Avraham Oplatka. On the mechanochemistry of muscular contraction. Journal of Theoretical Biology, 34(2):379–403, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901695.

Ord:1976:INC

[Ord76] M. J. Ord. The interaction of nuclear and cytoplasmic damage after treatments with toxic chemicals. Journal of Theoretical Biology, 62(2):369–378, IN1–IN5, 379–387, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901259.

Orgel:1977:GDO

[Org77] L. E. Orgel. Gene-duplication and the origin of proteins with novel functions. Journal of Theoretical Biology, 67(4):773, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902624.

Oritsland:1974:WSR

[Øri74] N. A. Øritsland. A windchill and solar radiation index for homeotherms. Journal of Theoretical Biology, 47(2):413–420, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374902070.

Orloci:1970:AVS

[Orl70] Laszlo Orloci. Analysis of vegetation samples based on the use of information. Journal of Theoretical Biology, 29(2):173–189, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900184.

Orlove:1975:MKSa

[Orl75a] M. J. Orlove. A model of kin selection not invoking coefficients of relationship. Journal of Theoretical Biology, 49(1):289–310, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580035X.

Orlove:1975:MKSb

[Orl75b] M. J. Orlove. A model of kin selection not invoking coefficients of relationship. Journal of Theoretical Biology, 49(2):289–310, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375901733.

Orlove:1975:SFI

[Orl75c] M. J. Orlove. Some further insight into kin selection. Journal of Theoretical Biology, 55(2):547–551, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801019.

Orlove:1977:KSC

[Orl77] M. J. Orlove. Kin selection and cancer. Journal of Theoretical Biology, 65(3):605–607, April 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790220X.

Orlove:1979:PDE

[Orl79a] M. J. Orlove. Putting the diluting effect into inclusive fitness. Journal of Theoretical Biology, 78(3):449–450, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903448.

Orlove:1979:RIF

[Orl79b] M. J. Orlove. A reconciliation of inclusive fitness and personal fitness approaches: A proposed correcting term for the inclusive fitness formula. Journal of Theoretical Biology, 81(3):577–586, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900559.

Orr:1979:CAS

[Orr79] J. S. Orr. Continuous administration of short-lived radioisotope tracers and the analogous Laplace transform. Journal of Theoretical Biology, 78(1):101–111, May 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990328X.

Othmer:1971:IDP

[OS71] H. G. Othmer and L. E. Scriven. Instability and dynamic pattern in cellular networks. Journal of Theoretical Biology, 32(3):507–537, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901548.

Othmer:1974:NLA

[OS74] H. G. Othmer and L. E. Scriven. Non-linear aspects of dynamic pattern in cellular networks. Journal of Theoretical Biology, 43(1):83–112, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800470.

Osborn:1974:CTR

[Osb74] J. W. Osborn. On the control of tooth replacement in reptiles and its relationship to growth. Journal of Theoretical Biology, 46(2):509–527, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900125.

Otto:1979:RPS

[OSLM79] P. Otto, M. Seel, J. Ladik, and R. Müller. From retrospective to predictive structure activity correlations. Journal of Theoretical Biology, 78(2):197–210, May 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902649.

Ostroumov:1977:PCM

[Ost77] S. A. Ostroumov. Participation of chloroplasts and mitochondria in virus reproduction and the evolution of the eukaryotic cell. Journal of Theoretical Biology, 67(2):287–297, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902016.

Otten:1974:MAR

[Ott74] H. A. Otten. A mathematical analysis of the relation between (bacterio) chlorophyll fluorescence yield and the concentration of reaction centre traps in photosynthesis. Journal of Theoretical Biology, 46(1):75–100, July 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901404.

Ostroumov:1978:MPS

[OV78] S. A. Ostroumov and L. N. Vorobiev. Membrane potential and surface charge densities as possible generalized regulators of membrane protein activities. Journal of Theoretical Biology, 75(3):289–297, December 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903351.

Overton:1977:TIS

[Ove77] Jane Overton. Is there an instance of Steinberg’s “site frequency model”? Journal of Theoretical Biology, 65(4):787–789, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790025X.

OBrien:1973:APM

[OW73] James J. O’Brien and J. S. Wroblewski. On advection in phytoplankton models. Journal of Theoretical Biology, 38(1):197– 202, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937390235X.

Orlove:1978:CRC

[OW78] M. J. Orlove and Constance L. Wood. Coefficients of relationship and coefficients of relatedness in kin selection: A covariance form for the RHO formula. Journal of Theoretical Biology, 73(4):679–686, August 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901297.

Okuda:1974:DBA

[OYT74] Masami Okuda, Akihiko Yoshida, and Koichi Takahashi. A dynamical behaviour of active regions in randomly connected neural networks. Journal of Theoretical Biology, 48(1):51–73, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901799.

Poste:1971:MFR

[PA71] G. Poste and A. C. Allison. Membrane fusion reaction: A theory. Journal of Theoretical Biology, 32(1):165–184, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901445.

Polgar:1974:SCS

[PA74] László Polgár and Bence Asbóth. On the stereochemistry of catalysis by serine proteases. Journal of Theoretical Biology, 46(2):543–558, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374900149.

Paau:1978:PME

[Paa78] Alan S. Paau. A possible mechanism for the establishment of nitrogen fixation in bacteroids of fast-growing, acid-producing rhizobium. Journal of Theoretical Biology, 74(1):139–142, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937890293X.

Paillotin:1972:TCE

[Pai72] Guy Paillotin. Transport and capture of electronic excitation energy in the photosynthetic apparatus. Journal of Theoretical Biology, 36(2):223–235, August 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290094X.

Paillotin:1976:CFE

[Pai76a] Guy Paillotin. Capture frequency of excitations and energy transfer between photosynthetic units in the photosystem II. Journal of Theoretical Biology, 58(1):219–235, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901491.

Paillotin:1976:MEP

[Pai76b] Guy Paillotin. Movement of excitations in the photosynthetic domains of photosystem II. Journal of Theoretical Biology, 58(1):237–252, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901508.

Painter:1976:MCB

[Pai76c] P. R. Painter. A model for the control of the batch culture growth rate of enteric bacteria. Journal of Theoretical Biology, 56(2):477–489, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800872.

Palmer:1970:BDL

[Pal70] Lars S. Palmer. Binomial distribution of lactic dehydrogenase isoenzymes from multicellular sources. Journal of Theoretical Biology, 27(3):455–470, ???? 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800091.

Paltridge:1973:ST

[Pal73] G. W. Paltridge. On the shape of trees. Journal of Theoretical Biology, 38(1):111–137, January 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902282.

Pande:1978:ETS

[Pan78] L. K. Pande. Ecosystems with three species: One-prey-and-two-predator system in an exactly solvable model. Journal of Theoretical Biology, 74(4):591–598, October 21, 1978.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902448.

Papageorgiou:1971:ALN

[Pap71] George Papageorgiou. Absorption of light by non-refractive spherical shells. Journal of Theoretical Biology, 30(2):249–254, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937190052X.

Papaphilis:1973:RRT

[Pap73a] Anastasios D. Papaphilis. On the rate of ribosome translocation anisotropy and ribosome saturation in the polysome. Journal of Theoretical Biology, 38(3):613–625, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902610.

Papentin:1973:DESa

[Pap73b] Frank Papentin. A Darwinian evolutionary system: I. Definition and basic properties. Journal of Theoretical Biology, 39(2):397– 415, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901082.

Papentin:1973:DESb

[Pap73c] Frank Papentin. A Darwinian evolutionary system: II. Experiments on protein evolution and evolutionary aspects of the genetic code. Journal of Theoretical Biology, 39(2):417–430, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901094.

Papentin:1973:DESc

[Pap73d] Frank Papentin. A Darwinian evolutionary system: III. Experiments on the evolution of feeding patterns. Journal of Theoretical Biology, 39(2):431–445, May 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901100.

Papahadjopoulos:1974:CCM

[Pap74] Demetrios Papahadjopoulos. Cholesterol and cell membrane function: A hypothesis concerning the etiology of atherosclerosis. Journal of Theoretical Biology, 43(2):329–337, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800640.

Parry:1970:PCF

[Par70] D. A. D. Parry. A proposed conformation for α-fibrous proteins. Journal of Theoretical Biology, 26(3):429–430, IN1, 431– 435, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900949.

Parks:1973:SMA

[Par73a] J. R. Parks. A stochastic model of animal growth. Journal of Theoretical Biology, 42(3):505–518, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902439.

Parks:1973:DSR

[Par73b] John R. Parks. Diet space and response surfaces. Journal of Theoretical Biology, 42(2):349–358, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900945.

Park:1974:ADN

[Par74a] David J. M. Park. An algorithm for detecting non-steady state moieties in steady state subnetworks. Journal of Theoretical Biology, 48(1):125–131, November 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901830.

Park:1974:HSM

[Par74b] David J. M. Park. The hierarchical structure of metabolic networks and the construction of efficient metabolic simulators. Journal of Theoretical Biology, 46(1):31–74, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901398.

Parker:1974:ASE

[Par74c] G. A. Parker. Assessment strategy and the evolution of fighting behaviour. Journal of Theoretical Biology, 47(1):223–243, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901118.

Park:1975:FIB

[Par75a] David J. M. Park. Flux induced biochemical differentiation. Journal of Theoretical Biology, 54(2):363–379, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801378.

Park:1975:LSMa

[Par75b] David J. M. Park. Local stability in metabolic networks with conserved moieties and steady state subnetworks. Journal of Theoretical Biology, 49(1):431–437, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800464.

Park:1975:LSMb

[Par75c] David J. M. Park. Local stability in metabolic networks with conserved moieties and steady state subnetworks. Journal of Theoretical Biology, 49(2):431–437, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901848.

Parks:1975:AGP

[Par75d] John R. Parks. The animal growth phase plane. Journal of Theoretical Biology, 55(2):371–380, December 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800865.

Parks:1975:TAW

[Par75e] John R. Parks. A theory of animal weight response to controlled feeding. Journal of Theoretical Biology, 55(2):381–391, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800877.

Parkhurst:1977:TDM

[Par77] David F. Parkhurst. A three-dimensional model for CO 2 uptake by continuously distributed mesophyll in leaves. Journal of Theoretical Biology, 67(3):471–488, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900509.

Parker:1978:SNR

[Par78] G. A. Parker. Selection on non-random fusion of gametes during the evolution of anisogamy. Journal of Theoretical Biology, 73(1):1–28, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901777.

Pasternak:1976:CSR

[Pas76] C. A. Pasternak. The cell surface in relation to the growth cycle. Journal of Theoretical Biology, 58(2):365–382, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801257.

Patton:1970:CRC

[Pat70] Stuart Patton. Correlative relationship of cholesterol and sphingomyelin in cell membranes. Journal of Theoretical Biology, 29(3):489–491, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901116.

Pavlidis:1971:PBO

[Pav71] Theodosios Pavlidis. Populations of biochemical oscillators as circadian clocks. Journal of Theoretical Biology, 33(2):319–338, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900701.

Papadopoullos:1979:MTN

[PAZ79] Maximos S. Papadopoullos, Photios A. Anninos, and Silvio Zenone. Memory topology of neural nets. Journal of Theoretical Biology, 80(4):505–535, October 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900912.

Perkins:1977:TMA

[PB77] Donald L. Perkins and Stanley E. Babb. A thermodynamic model of ameboid locomotion based on the pushpull nature of gel-to-sol interconversions. Journal of Theoretical Biology, 69(2):293–299, November 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901382.

Poltev:1978:MMN

[PB78] V. I. Poltev and V. I. Bruskov. On molecular mechanisms of nucleic acid synthesis fidelity aspects 1. Contribution of base interactions. Journal of Theoretical Biology, 70(1):69–83, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890303X.

Petty:1979:PMS

[PB79] Howard R. Petty and James K. Baird. On the polarizability of macromolecules in solution. Journal of Theoretical Biology, 80(2):295–299, September 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902121.

Parker:1972:OEG

[PBS72] G. A. Parker, R. R. Baker, and V. G. F. Smith. The origin and evolution of gamete dimorphism and the male-female phenomenon. Journal of Theoretical Biology, 36(3):529–553, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900070.

PenningDeVries:1974:PRE

[PBV74] F. W. T. Penning De Vries, A. H. M. Brunsting, and H. H. Van Laar. Products, requirements and efficiency of biosynthesis a quantitative approach. Journal of Theoretical Biology, 45(2): 339–377, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901192.

Pohl:1972:DF

[PC72] Herbert A. Pohl and Joe S. Crane. Dielectrophoretic force. Journal of Theoretical Biology, 37(1):1–13, October 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901129.

Parnas:1976:OSM

[PC76] Hanna Parnas and Dan Cohen. The optimal strategy for the metabolism of reserve materials in micro-organisms. Journal of Theoretical Biology, 56(1):19–55, January 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800446.

Pullman:1972:MOS

[PCC72] B. Pullman, J. L. Coubeils, and Ph. Courrare. A molecular orbital study of the conformation of barbiturates. Journal of Theoretical Biology, 35(2):375–385, May 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900458.

Parrish:1976:CST

[PCT76] W. Parrish, F. Chittenden, and E. Triplett. A computer simulation of transcription and processing of eucaryotic mRNA. Journal of Theoretical Biology, 56(2):355–361, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800793.

Paltridge:1974:PYS

[PD74] G. W. Paltridge and J. V. Denholm. Plant yield and the switch from vegetative to reproductive growth. Journal of Theoretical Biology, 44(1):23–34, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800275.

Perram:1973:RTB

[Per73] J. W. Perram. Relaxation times in bacteriological culture and the approach to steady state. Journal of Theoretical Biology, 38(3):571–578, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902579.

Percus:1976:TCL

[Per76a] J. K. Percus. Tissue correlations in linear samples. Journal of Theoretical Biology, 57(2):281–299, April 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900023.

Perelson:1976:RCM

[Per76b] Alan S. Perelson. Remarks on conservation of mass in open chemical reaction systems. Journal of Theoretical Biology, 63(1):233–237, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900965.

Pedley:1978:DOF

[PF78] T. J. Pedley and J. Fischbarg. The development of osmotic flow through an unstirred layer. Journal of Theoretical Biology, 70(4):427–447, February 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902515.

Pfaff:1971:SES

[Pfa71] Donald W. Pfaff. Statistical effects of sensitivity differences among neurophysiological preparations. Journal of Theoretical Biology, 31(1):159–160, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901287.

Pfaffelhuber:1973:MLM

[Pfa73] E. Pfaffelhuber. Mathematical learning models and neuronal networks. Journal of Theoretical Biology, 40(1):63–76, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901653.

Pomerantz:1979:CCC

[PG79] M. J. Pomerantz and M. E. Gilpin. Community covariance and coexistence. Journal of Theoretical Biology, 79(1):67–81, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902571.

Prado:1978:CDP

[PGPP78] Fernando Ribas Prado, Claude Giessner-Prettre, and Bernard Pullman. On the conformational dependence of the proton chemical shifts in nucleosides and nucleotides III. Proton chemical shifts of 5’-nucleotides as a function of different conformational parameters. Journal of Theoretical Biology, 74(2):259–277, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900759.

Pielou:1972:CTE

[Pie72] E. C. Pielou. 2^{k} contingency tables in ecology. Journal of Theoretical Biology, 34(2):337–352, February 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290166X.

Pielou:1974:VZR

[Pie74] E. C. Pielou. Vegetation zones: Repetition of zones on a monotonic environmental gradient. Journal of Theoretical Biology, 47(2):485–489, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374902112.

Pincus:1970:ES

[Pin70] Martin Pincus. An evolutionary strategy. Journal of Theoretical Biology, 28(3):483–488, September 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900822.

Pinder:1972:SHR

[Pin72] D. N. Pinder. Shape of human red cells. Journal of Theoretical Biology, 34(3):407–408, IN1–IN2, 409–410, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901312.

Pitt:1978:WHB

[Pit78] Roger Pitt. Warfare and hominid brain evolution. Journal of Theoretical Biology, 72(3):551–575, June 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890317X.

Ponnuswamy:1977:NLM

[PL77] P. K. Ponnuswamy and M. Lakshmanant. A non-linear model for the synaptic basis of neuronal memory. Journal of Theoretical Biology, 68(4):483–489, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900996.

Ponnuswamy:1976:PBC

[PM76] P. K. Ponnuswamy and P. Manavalan. Preferred backbone conformations of amino acid residues for solvent interaction. Journal of Theoretical Biology, 60(2):481–486, August 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690076X.

Pearson:1977:CCM

[PM77] M. J. Pearson and D. I. McLaren. A criticism of catastrophe modelling of the differentiative process in amphibian development. Journal of Theoretical Biology, 69(4):721–734, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903782.

Proctor:1974:HPD

[PMC74] Peter Proctor, John McGinness, and Peter Corry. A hypothesis on the preferential destruction of melanized tissues. Journal of Theoretical Biology, 48(1):19–22, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901763.

Perez:1978:QAO

[PMC78] J. F. Perez, C. P. Malta, and F. A. B. Coutinho. Qualitative analysis of oscillations in isolated populations of flies. Journal of Theoretical Biology, 71(4):505–514, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903211.

Pinter:1973:ILS

[PMN73] R. B. Pinter, R. D. Martin, and P. O. Neudorfer. Invariance of least squared-error estimation under some commonly used neurophysiological data pre-processing methods. Journal of Theoretical Biology, 41(2):323–330, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901215.

Parsegian:1973:VWF

[PN73] V. Adrian Parsegian and Barry W. Ninham. Van der Waals forces in many-layered structures: Generalizations of the Lifshitz result for two semi-infinite media. Journal of Theoretical Biology, 38(1):101–109, January 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902270.

Perelson:1979:TSC

[PO79] Alan S. Perelson and George F. Oster. Theoretical studies of clonal selection: Minimal antibody repertoire size and reliability of self-non-self discrimination. Journal of Theoretical Biology, 81(4):645–670, December 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902753.

Pogar:1971:MPT

[Pog71] L. Pogár. On the mechanism of proton transfer in the catalysis by serine proteases. Journal of Theoretical Biology, 31(1):165– 169, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901305.

Polak:1973:MPM

[Pol73] Ernest H. Polak. Multiple profile-multiple receptor site model for vertebrate olfaction. Journal of Theoretical Biology, 40 (3):469–484, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900052.

Poole:1975:DEM

[Poo75] Brian Poole. Diffusion effects in the metabolism of hydrogen peroxide by rat liver peroxisomes. Journal of Theoretical Biology, 51(1):149–167, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375901459.

Popel:1979:DTS

[Pop79] Aleksander S. Popel. Diffusion in tissue slices with metabolism obeying Michaelis–Menten kinetics. Journal of Theoretical Biology, 80(3):325–332, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379900973.

Power:1975:MHS

[Pow75] Harry W. Power. A model of how the sickle-cell gene produces malaria resistance. Journal of Theoretical Biology, 50(1):121– 127, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900272.

Prahhakaran:1979:SDP

[PP79] M. Prahhakaran and P. K. Ponnuswamy. The spatial distribution of physical, chemical, energetic and conformational properties of amino acid residues in globular proteins. Journal of Theoretical Biology, 80(4):485–504, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900900.

Polozov:1975:INB

[PPS75] R. V. Polozov, V. I. Poltev, and B. I. Sukhorukov. The influence of nitrous bases interactions on the DNA secondary structure formation. Journal of Theoretical Biology, 55(2):491–503, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800968.

Prieve:1976:SPD

[PR76] Dennis C. Prieve and Eli Ruckenstein. The surface potential of and double-layer interaction force between surfaces characterized by multiple ionizable groups. Journal of Theoretical Biology, 56(1):205–228, January 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800537.

Pyshnov:1977:DMS

[PR77] M. B. Pyshnov and S. A. Reggirt. Deterministic model of steady-state cell proliferation. Journal of Theoretical Biology, 68(2):247–257, September 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901631.

Pratt:1972:NTC

[Pra72] Vernon Pratt. Numerical taxonomy — a critique. Journal of Theoretical Biology, 36(3):581–592, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900100.

Pratt:1974:NTI

[Pra74] Vernon Pratt. Numerical taxonomy: On the incoherence of its rationale. Journal of Theoretical Biology, 48(2):497–499, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800202.

Previc:1970:BDB

[Pre70] Edward P. Previc. Biochemical determination of bacterial morphology and the geometry of cell division. Journal of Theoretical Biology, 27(3):471–497, ???? 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800108.

Prentice:1973:DSC

[Pre73] R. L. Prentice. A design for studying the clustering of plant or animal species using quadrat sizes in geometric progression. Journal of Theoretical Biology, 39(3):601–608, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900714.

Parrish:1970:ICP

[PS70] J. D. Parrish and S. B. Saila. Interspecific competition, predation and species diversity. Journal of Theoretical Biology, 27(2): 207–220, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901384.

Parnas:1978:CSP

[PS78] Hanna Parnas and Lee A. Segel. A computer simulation of pulsatile aggregation in Dictyostelium discoideum. Journal of Theoretical Biology, 71(2):185–207, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902667.

Presber:1976:EMR

[PSK76] Wolfgang Presber, Cornelia Schroeder, and Detlev H. Krueger. The evolution of mutation rates via DNA base composition and mutator genes: A proposition. Journal of Theoretical Biology, 59(2):353–360, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901752.

Palatt:1970:TPR

[PSM70] P. J. Palatt, G. M. Saidel, and M. Macklin. Transport processes in the renal cortex. Journal of Theoretical Biology, 29(2):251– 274, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900214.

Palatt:1974:HMP

[PST74] Paul J. Palatt, Henry Sackin, and Roger I. Tanner. A hydrodynamic model of a permeable tubule. Journal of Theoretical Biology, 44(2):287–303, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519374901611.

Prothero:1975:MTP

[PT75] John W. Prothero and Ruth W. Tyler. A model of thymocyte proliferation. Journal of Theoretical Biology, 51(2):357– 382, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900661.

Post:1979:QSM

[PT79] Wilfred M. Post and C. C. Travis. Quantitative stability in models of ecological communities. Journal of Theoretical Biology, 79(4):547–553, August 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990242X.

Pruitt:1974:KMQ

[PTB74] Kenneth M. Pruitt, Malcolm E. Turner, and Robert J. Boackle. A kinetic model for the quantitative analysis of complement. Journal of Theoretical Biology, 44(2):207–217, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490157X.

Pulliam:1973:AF

[Pul73] H. Ronald Pulliam. On the advantages of flocking. Journal of Theoretical Biology, 38(2):419–422, February 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901847.

Purves:1976:CFP

[Pur76a] R. D. Purves. Current flow and potential in a three-dimensional syncytium. Journal of Theoretical Biology, 60(1):147–162, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901600.

Purves:1976:MSC

[Pur76b] R. D. Purves. Microelectrodes in spherical cells. Journal of Theoretical Biology, 63(1):225–228, November 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900941.

Purves:1977:RDI

[Pur77a] R. D. Purves. The release of drugs from iontophoretic pipettes. Journal of Theoretical Biology, 66(4):789–798, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902466.

Purves:1977:TCC

[Pur77b] R. D. Purves. The time course of cellular responses to iontophoretically applied drugs. Journal of Theoretical Biology, 65(2):327–344, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377903289.

Perelson:1979:CNI

[PW79] Alan S. Perelson and Frederik W. Wiegel. A calculation of the number of IgG molecules required per cell to fix complement. Journal of Theoretical Biology, 79(3):317–332, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903497.

Plesset:1976:RTI

[PWW76] Milton S. Plesset, Christopher G. Whipple, and Howard Winet. Rayleigh–Taylor instability of surface layers as the mechanism for bioconvection in cell cultures. Journal of Theoretical Biology, 59(2):331–351, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901740.

Richards:1975:CFP

[RAH75] W. G. Richards, D. G. Aschman, and Jane Hammond. Conformational flexibility in physiologically active amines. Journal of Theoretical Biology, 52(1):223–230, July 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900533.

Ramon:1976:MPA

[RAJM76] Fidel Ramón, Nels C. Anderson, Ronald W. Joyner, and John W. Moore. A model for propagation of action potentials in smooth muscle. Journal of Theoretical Biology, 59(2):381– 408, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901788.

Rakitzis:1977:KIE

[Rak77] Emmanuel T. Rakitzis. Kinetics of irreversible enzyme inhibition: Co-operative effects. Journal of Theoretical Biology, 67(1): 49–59, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901849.

Rakitzis:1978:KIEb

[Rak78a] Emmanuel T. Rakitzis. Kinetics of irreversible enzyme inhibition by an unstable inhibitor: Co-operative effects. Journal of Theoretical Biology, 75(2):239–243, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902333.

Rakitzis:1978:KIEa

[Rak78b] Emmanuel T. Rakitzis. Kinetics of irreversible enzyme inhibition: The tsou plot in irreversible binding co-operativity. Journal of Theoretical Biology, 70(4):461–465, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902539.

Raman:1976:ACD

[Ram76] R. K. Raman. The autonomous cell in Dictyostelium discoideum. Journal of Theoretical Biology, 59(2):491–495, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901855.

Raman:1977:CAT

[Ram77] Rajagopalan K. Raman. Cellular aggregation towards steady point sources of attractant. Journal of Theoretical Biology, 64(1):43–69, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377901138.

Rand:1973:SMT

[Ran73] William M. Rand. A stochastic model of the temporal aspect of breeding strategies. Journal of Theoretical Biology, 40(2):337–351, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901367.

Ransom:1975:CAD

[Ran75] Robert Ransom. Computer analysis of division patterns in the Drosophila head disc. Journal of Theoretical Biology, 53(2): 445–462, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800154.

Randolph:1976:DDH

[Ran76] M. L. Randolph. Degradation of DNA in Haemophilus influenzae cells after X-ray irradiation: II. Comparison with theoretical models. Journal of Theoretical Biology, 60(1):59–91, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901569.

Ransom:1977:CAC

[Ran77] Robert Ransom. Computer analysis of cell division in Drosophila imaginal discs: Model revision and extension to simulate leg disc growth. Journal of Theoretical Biology, 66(2):361–377, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901771.

Rapport:1971:GIT

[Rap71] Ellen W. Rapport. Genes and immunological tolerance. Journal of Theoretical Biology, 30(3):597–601, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900129.

Rappaport:1976:ABR

[Rap76] H. P. Rappaport. An analysis of the binding of related ligands. Journal of Theoretical Biology, 58(2):253–268, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376801191.

Rappaport:1977:ITS

[Rap77] H. P. Rappaport. Interpretation of transition state and substrate analog binding. Journal of Theoretical Biology, 65(4):781, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900236.

Rapoport:1978:MMV

[Rap78] Stanley I. Rapoport. A mathematical model for vasogenic brain edema. Journal of Theoretical Biology, 74(3):439–467, October 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902242.

Rappaport:1979:EGC

[Rap79] H. P. Rappaport. Evaluation of group contributions to ligand binding. Journal of Theoretical Biology, 79(2):157–165, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902443.

Rapp:1977:OCC

[RB77] P. E. Rapp and M. J. Berridge. Oscillations in calciumcyclic AMP control loops form the basis of pacemaker activity and other high frequency biological rhythms. Journal of Theoretical Biology, 66(3):497–525, June 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902995.

Rutledge:1976:ESI

[RBM76] Robert W. Rutledge, Bennett L. Basore, and Robert J. Mulholland. Ecological stability: An information theory viewpoint. Journal of Theoretical Biology, 57(2):355–371, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900072.

Roomans:1978:CTA

[RBP78] G. M. Roomans and G. W. F. H. Borst-Pauwels. Co-transport of anions and neutral solutes with cations across charged biological membranes. Effects of surrface potential on uptake kinetics. Journal of Theoretical Biology, 73(3):453–468, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901534.

Rehm:1970:MEU

[RBSS70] W. S. Rehm, C. F. Butler, S. G. Spangler, and S. S. Sanders. A model to explain uphill water transport in the mammalian stomach. Journal of Theoretical Biology, 27(3):433–453, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937080008X.

Rooney:1977:SBM

[RC77] David W. Rooney and Joseph P. Costello. Set-back model of division-synchronization: Optimal spacing of thermal shocks that accelerate cell cycling. Journal of Theoretical Biology, 69(4):597–611, December 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903691.

Reichert:1973:AIT

[RCW73] Thomas A. Reichert, Donald N. Cohen, and Andrew K. C. Wong. An application of information theory to genetic mutations and the matching of polypeptide sequences. Journal of Theoretical Biology, 42(2):245–261, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390088X.

Reshodko:1977:BSC

[RD77a] L. V. Reshodko and Z. Drska. Biological systems of cellular organization and their computer models. Journal of Theoretical Biology, 69(4):563–566, IN1–IN2, 567–569, IN4, 570–579, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903678.

Russev:1977:COS

[RD77b] George Russev and Kalin Dudov. Chromatin organization studied by computer simulation of models of growing polymer chains. Journal of Theoretical Biology, 64(2):221–230, IN1, 231–235, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903538.

Rein:1976:RBG

[RE76] Robert Rein and John T. Egan. The relation between genetic recombination and base sequence complementarity. Journal of Theoretical Biology, 62(2):409–427, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901272.

Reanney:1974:OP

[Rea74] D. C. Reanney. On the origin of prokaryotes. Journal of Theoretical Biology, 48(1):243–251, November 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901945.

Reanney:1975:RRVa

[Rea75a] D. C. Reanney. A regulatory role for viral RNA in eukaryotes. Journal of Theoretical Biology, 49(1):461–492, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800488.

Reanney:1975:RRVb

[Rea75b] D. C. Reanney. A regulatory role for viral RNA in eukaryotes. Journal of Theoretical Biology, 49(2):461–492, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901861.

Reanney:1977:ATE

[Rea77] D. C. Reanney. Aminoacyl thiol esters and the origins of genetic specificity. Journal of Theoretical Biology, 65(3):555–569, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902144.

Reed:1973:AAM

[Ree73] F. B. Reed. Amino acids in meteorites. Journal of Theoretical Biology, 39(3):683, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900830.

Reiff:1970:COH

[Rei70] Theodore R. Reiff. Colloid osmotic homeostasis in humans I. Theoretical aspects and background. Journal of Theoretical Biology, 28(1):1–14, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900603.

Reijnders:1976:OEN

[Rei76] L. Reijnders. The origin of endosymbiotic nitrogen fixation by rhizobium. Journal of Theoretical Biology, 61(1):245–248, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937690117X.

Reijnders:1978:AGT

[Rei78] L. Reijnders. On the applicability of game theory to evolution. Journal of Theoretical Biology, 75(2):245–247, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902345.

Reiff:1979:COH

[Rei79] Theodore R. Reiff. Colloid osmotic homeostasis in humans: Intravascular oncotic energy under varying gravitational potential. Journal of Theoretical Biology, 80(1):15–20, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901759.

Rescigno:1973:TTT

[Res73] Aldo Rescigno. On transfer times in tracer experiments. Journal of Theoretical Biology, 39(1):9–27, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902026.

Resnick:1976:RDS

[Res76] Michael A. Resnick. The repair of double-strand breaks in DNA: A model involving recombination. Journal of Theoretical Biology, 59(1):97–106, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800252.

Resnick:1978:SRI

[Res78] Michael A. Resnick. Similar responses to ionizing radiation of fungal and vertebrate cells and the importance of DNA doublestrand breaks. Journal of Theoretical Biology, 71(3):339–346, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901649.

Rudolph:1973:KTS

[RF73] Frederick B. Rudolph and Herbert J. Fromm. Kinetics of three substrate enzyme systems: Treatment of partially random mechanisms using equilibrium assumptions. Journal of Theoretical Biology, 39(2):363–371, May 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901057.

Rogers:1978:CSC

[RG78] Gary Rogers and Narendra S. Goel. Computer simulation of cellular movements: Cell-sorting, cellular migration through a mass of cells and contact inhibition. Journal of Theoretical Biology, 71(1):141–166, March 7, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902163.

Rosenberger:1978:SGR

[RGZW78] R. F. Rosenberger, N. B. Grover, A. Zaritsky, and C. L. Woldringh. Surface growth in rod-shaped bacteria. Journal of Theoretical Biology, 73(4):711–721, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901327.

Richards:1975:BIR

[RHA75] W. G. Richards, Jane Hammond, and D. G. Aschman. Barriers to internal rotation in histamine and 4-methyl histamine. Journal of Theoretical Biology, 51(1):237–239, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901502.

Rao:1979:ICD

[RHR79] Gita Subba Rao, Zehra Hamid, and J. Subba Rao. The information content of DNA and evolution. Journal of Theoretical Biology, 81(4):803–807, December 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902820.

Ribeiro:1979:PMT

[Rib79] J. A. Ribeiro. Purinergic modulation of transmitter release. Journal of Theoretical Biology, 80(2):259–270, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902108.

Richards:1975:MMC

[Ric75] Whitman Richards. Mosaic model for color vision. Journal of Theoretical Biology, 53(1):177–197, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901095.

Rigney:1979:NKS

[Rig79a] D. R. Rigney. Note on the kinetics and stochastics of induced protein synthesis as influenced by various models for messenger RNA degradation. Journal of Theoretical Biology, 79(2):247– 257, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902509.

Rigney:1979:SMC

[Rig79b] David R. Rigney. Stochastic model of constitutive protein levels in growing and dividing bacterial cells. Journal of Theoretical Biology, 76(4):453–480, February 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900134.

Riley:1973:AAS

[Ril73] W. T. Riley. Amino acid sequences and double-stranded messages — a means of directing the site of mutation? Journal of Theoretical Biology, 40(2):285–300, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901331.

Riley:1974:NAS

[Ril74] W. T. Riley. Nucleic acid sizes — a role for histones? Journal of Theoretical Biology, 44(1):141–143, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800342.

Riley:1979:EES

[Ril79] John G. Riley. Evolutionary equilibrium strategies. Journal of Theoretical Biology, 76(2):109–123, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903655.

Rubinow:1972:FVF

[RK72] S. I. Rubinow and Joseph B. Keller. Flow of a viscous fluid through an elastic tube with applications to blood flow. Journal of Theoretical Biology, 35(2):299–313, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900410.

Rubin:1976:IEF

[RL76] Howard A. Rubin and John J. Lee. Informational energy flow as an aspect of the ecological efficiency of marine ciliates. Journal of Theoretical Biology, 62(1):69–91, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900527.

Ramamoorthy:1977:FPC

[RL77] S. Ramamoorthy and Gary G. Leppard. Fibrillar pectin and contact cation exchange at the root surface. Journal of Theoretical Biology, 66(3):527–530, IN1, 531–540, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903009.

Richardson:1974:MSW

[RLB74] I. W. Richardson, Vojtech Licko, and Ettore Bartoli. Mechanisms of salt and water transport in proximal tubule as measured by split drop. Journal of Theoretical Biology, 47(1):57– 64, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374900988.

Roberts:1973:IMT

[RLS73] George W. Roberts, Kenneth B. Larson, and Edmund E. Spaeth. The interpretation of mean transit time measurements for multiphase tissue systems. Journal of Theoretical Biology, 39(2):447–475, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901112.

Rescigno:1973:DTE

[RM73] Aldo Rescigno and Lester D. Michels. On dispersion in tracer experiments. Journal of Theoretical Biology, 41(3):451–460, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900556.

Rajaraman:1978:MNT

[RMF78] R. Rajaraman, J. M. Macsween, and R. A. Fox. Models of normal and transformed cell adhesion, and capping and locomotion in vitro. Journal of Theoretical Biology, 74(2):177–201, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937890070X.

Ruckenstein:1976:CDR

[RMG76] Eli Ruckenstein, Abraham Marmur, and William N. Gill. Coverage dependent rate of cell deposition. Journal of Theoretical Biology, 58(2):439–454, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376801300.

Ruckenstein:1977:GKP

[RMG77] E. Ruckenstein, A. Marmur, and W. N. Gill. Growth kinetics of platelet thrombi. Journal of Theoretical Biology, 66(1):147– 168, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903174.

Ross:1976:PMB

[RN76] George G. Ross and Paul Nival. Plankton modeling in the bay of villefranche. Journal of Theoretical Biology, 56(2):381–399, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800811.

Rein:1971:PSP

[RNS71] Robert Rein, S. Nir, and Mary N. Stamatiadou. Photochemical survival principle in molecular evolution. Journal of Theoretical Biology, 33(2):309–318, November 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900695.

Rafanelli:1974:ISE

[RO74] Kenneth Rafanelli and Stephen Orenstein. On the inclusion of spontaneous emission in a neuronic equation of state. Journal of Theoretical Biology, 43(2):381–383, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800688.

Roberts:1970:MPS

[Rob70] A. M. Roberts. Motion of paramecium in static electric and magnetic fields. Journal of Theoretical Biology, 27(1):97–106, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901311.

Roberts:1977:CPM

[Rob77] D. W. Roberts. A contact pressure model for semi-decussate and related phyllotaxis. Journal of Theoretical Biology, 68 (4):583–597, October 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377901072.

Roberts:1978:OFP

[Rob78] David W. Roberts. The origin of Fibonacci phyllotaxis — an analysis of Adler’s contact pressure model and Mitchison’s expanding apex model. Journal of Theoretical Biology, 74 (2):217–233, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378900735.

Romesburg:1976:UNG

[Rom76] H. Charles Romesburg. Use of the normal generated distribution for estimating population survival. Journal of Theoretical Biology, 61(2):447–457, September 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900291.

Rosen:1970:SSP

[Ros70] Robert Rosen. Subunit and subassembly processes. Journal of Theoretical Biology, 28(3):415–422, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900792.

Rosen:1971:TSA

[Ros71] Philip Rosen. Theoretical significance of arsenic as a carcinogen. Journal of Theoretical Biology, 32(2):425–426, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901780.

Rosen:1972:DSO

[Ros72a] Robert Rosen. Are the dynamics of a system operationally determinable? Journal of Theoretical Biology, 36(3):635–638, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900148.

Ross:1972:DDM

[Ros72b] George G. Ross. A difference-differential model in population dynamics. Journal of Theoretical Biology, 37(3):477–492, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900860.

Rosen:1973:FTA

[Ros73a] Gerald Rosen. Fundamental theoretical aspects of bacterial chemotaxis. Journal of Theoretical Biology, 41(2):201–208, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901136.

Ross:1973:MCG

[Ros73b] George G. Ross. A model for the competitive growth of two diatoms. Journal of Theoretical Biology, 42(2):307–331, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900921.

Ross:1973:PML

[Ros73c] George G. Ross. A population model for limited food competition. Journal of Theoretical Biology, 42(2):333–347, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900933.

Rosen:1975:ASIa

[Ros75a] Gerald Rosen. Analytical solution to the initial-value problem for traveling bands of chemotactic bacteria. Journal of Theoretical Biology, 49(1):311–321, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800361.

Rosen:1975:ASIb

[Ros75b] Gerald Rosen. Analytical solution to the initial-value problem for traveling bands of chemotactic bacteria. Journal of Theoretical Biology, 49(2):311–321, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901745.

Rosen:1975:MNA

[Ros75c] Gerald Rosen. Molecule numbers associated with the chemotactic migration of neutrophils at various temperatures. Journal of Theoretical Biology, 54(2):381–382, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580138X.

Rosen:1975:NSD

[Ros75d] Gerald Rosen. On the nature of solutions to the dreitlein-smoes model for oscillatory kinetic systems. Journal of Theoretical Biology, 54(2):391–394, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937580141X.

Rosen:1975:DRA

[Ros75e] Robert Rosen. On the dynamical realization of automata. Journal of Theoretical Biology, 54(1):109–120, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800579.

Rosen:1976:CTT

[Ros76a] Gerald Rosen. Chemotactic transport theory for neutrophil leukocytes. Journal of Theoretical Biology, 59(2):371–380, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901776.

Rosen:1976:ENB

[Ros76b] Gerald Rosen. Existence and nature of band solutions to generic chemotactic transport equations. Journal of Theoretical Biology, 59(1):243–246, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800380.

Rosen:1976:NCR

[Ros76c] Robert Rosen. A note on the cross-reactivity of antibodies. Journal of Theoretical Biology, 60(1):177–182, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901624.

Rosen:1976:RSC

[Ros76d] Robert Rosen. On rate-sensitive chemoreceptor systems and the origin of the immune response. Journal of Theoretical Biology, 62(2):429–445, October 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901284.

Rosen:1976:SSA

[Ros76e] Robert Rosen. Structural stability, alternate descriptions and information. Journal of Theoretical Biology, 63(1):19–31, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900813.

Rossler:1976:PRB

[Rös76f] O. E. Rössler. Prescriptive relational biology and bacterial chemotaxis. Journal of Theoretical Biology, 62(1):141–157, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900552.

Rosen:1977:PES

[Ros77a] Gerald Rosen. On the persistence of ecological systems. Journal of Theoretical Biology, 65(4):795–799, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900273.

Rosen:1977:VC

[Ros77b] Philip Rosen. Viral carcinogenesis. Journal of Theoretical Biology, 64(2):215–220, January 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903526.

Rose:1978:CEG

[Ros78a] Michael R. Rose. Cheating in evolutionary games. Journal of Theoretical Biology, 75(1):21–34, November 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890200X.

Rosen:1978:XRL

[Ros78b] Philip Rosen. X- or γ-ray leukemogenesis in humans. Journal of Theoretical Biology, 74(4):603–606, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902461.

Rosen:1978:FGS

[Ros78c] Robert Rosen. Feedforwards and global system failure: A general mechanism for senescence. Journal of Theoretical Biology, 74(4):579–590, October 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902436.

Rosenthal:1978:CNC

[Ros78d] Gerald A. Rosenthal. Is canavanine a natural component of mammalian metabolism? Journal of Theoretical Biology, 71(2):265–273, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902710.

Rosin:1978:HBL

[Ros78e] R. Rosin. The honey bee “language” controversy. Journal of Theoretical Biology, 72(4):589–602, June 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902734.

Rosen:1979:CSM

[Ros79a] Robert Rosen. A comment on surface modulating assemblies. Journal of Theoretical Biology, 77(3):383–384, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903643.

Rosen:1979:PCC

[Ros79b] Robert Rosen. Patching and capping of cell membrane receptors as examples of morphogenetic movement. Journal of Theoretical Biology, 80(1):149–153, September 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901863.

Rotenberg:1972:TPT

[Rot72] Manuel Rotenberg. Theory of population transport. Journal of Theoretical Biology, 37(2):291–305, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900240.

Rothman:1973:MBM

[Rot73] James E. Rothman. The molecular basis of mesomorphic phase transitions in phospholipid systems. Journal of Theoretical Biology, 38(1):1–16, January 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390221X.

Rotenberg:1975:ESP

[Rot75] Manuel Rotenberg. Equilibrium and stability in populations whose interactions are age-specific. Journal of Theoretical Biology, 54(2):207–224, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375801263.

Rotenberg:1977:MDH

[Rot77a] Manuel Rotenberg. Mathematical description of holometabolous life cycles. Journal of Theoretical Biology, 64(2):333–353, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903617.

Rotenberg:1977:SSC

[Rot77b] Manuel Rotenberg. Selective synchrony of cells of differing cycle times. Journal of Theoretical Biology, 66(2):389–398, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901795.

Rotenberg:1979:SAT

[Rot79] Manuel Rotenberg. Stability analysis of the two compartment model of cell proliferation. Journal of Theoretical Biology, 77(1):51–63, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901371.

Routledge:1979:DIW

[Rou79] R. D. Routledge. Diversity indices: Which ones are admissible? Journal of Theoretical Biology, 76(4):503–515, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900158.

Rotshenker:1973:CFC

[RP73] S. Rotshenker and Y. Palti. Changes in fraction of current penetrating an axon as a function of duration of stimulating pulse. Journal of Theoretical Biology, 41(3):401–407, October 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900507.

Ruckenstein:1975:DCD

[RP75] Eli Ruckenstein and Dennis C. Prieve. Dynamics of cell deposition on surfaces. Journal of Theoretical Biology, 51(2):429– 438, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900727.

Rotshenker:1976:LPA

[RP76] S. Rotshenker and Y. Palti. The latent period of anode break excitation in myelinated and giant axons. Journal of Theoretical Biology, 59(2):293–302, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901715.

Roti:1977:EMR

[RP77] Joseph L. Roti Roti and Robert B. Painter. Equations for measuring the rate of DNA chain growth and replicon size by density labeling techniques. Journal of Theoretical Biology, 64(4):681–696, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902673.

Richardson:1979:AMT

[RR79] I. W. Richardson and Robert Rosen. Aging and the metrics of time. Journal of Theoretical Biology, 79(4):415–423, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902340.

Rahmann:1976:FMS

[RRB76] Hinrich Rahmann, Harald Rösner, and Heinz Breer. A functional model of sialo-glyco-macromolecules in synaptic transmission and memory formation. Journal of Theoretical Biology, 57(1):231–237, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251937680015X.

Rein:1972:QEA

[RRS72] Robert Rein, James R. Rabinowitz, and T. J. Swissler. Quantitative examination of the approximations in the monopole and dipole theories of intermolecular interactions. Journal of Theoretical Biology, 34(2):215–218, February 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901567.

Rikmenspoel:1970:BME

[RS70] Robert Rikmenspoel and Michael A. Sleigh. Bending moments and elastic constants in cilia. Journal of Theoretical Biology, 28(1):81–100, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370900652.

Raven:1976:ECE

[RS76] J. A. Raven and F. A. Smith. The evolution of chemiosmotic energy coupling. Journal of Theoretical Biology, 57(2):301– 312, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900035.

Rigney:1977:SML

[RS77] David R. Rigney and William C. Schieve. Stochastic model of linear, continuous protein synthesis in bacterial populations. Journal of Theoretical Biology, 69(4):761–766, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903812.

Rohlfing:1978:EPP

[RS78] Duane L. Rohlfing and Mary A. Saunders. Evolutionary processes possibly limiting the kinds of amino acids in protein to twenty: A review. Journal of Theoretical Biology, 71(4):487– 503, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937890320X.

Ross:1979:PPV

[RS79] George G. Ross and Norman A. Slade. A predator– prey viewpoint of a single species population. Journal of Theoretical Biology, 77(4):513–522, April 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900249.

Ramirez:1979:MAS

[RSL79] F. Ramirez, K. K. Shukla, and H. M. Levy. A model for the active site of skeletal muscle myosin. Journal of Theoretical Biology, 76(3):351–354, IN5, 355–357, February 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903175.

Rehm:1973:PCC

[RSS^{+}73] W. S. Rehm, S. S. Sanders, R. L. Shoemaker, J. O’Callaghan, and … E. A. Friday. Proton conductance of cell membranes. Journal of Theoretical Biology, 39(1):131–153, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902099.

Rapport:1970:DPF

[RT70] D. J. Rapport and J. E. Turner. Determination of predator food preferences. Journal of Theoretical Biology, 26(3):365– 372, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900895.

Rapport:1975:PPI

[RT75] David J. Rapport and James E. Turner. Predator–prey interactions in natural communities. Journal of Theoretical Biology, 51(1):169–180, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375901460.

Reay:1977:HPT

[RT77] D. A. Reay and F. A. J. Thiele. Heat pipe theory applied to a biological system: Quantification of the role of the “resting” eccrine sweat gland in thermoregulation. Journal of Theoretical Biology, 64(4):789–803, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902776.

Rubin:1977:TMP

[Rub77] B. T. Rubin. A theoretical model of the pinocytotic vesicular transport process in endothelial cells. Journal of Theoretical Biology, 64(4):619–647, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902648.

Rudin:1979:SCM

[Rud79] Donald O. Rudin. The second cell membrane: The basal lamina and the tissue cell theory of metazoa. Journal of Theoretical Biology, 78(1):143–152, IN3, 153–165, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903321.

Russell:1972:BPO

[Rus72] C. S. Russell. Biosynthesis of porphyrins and the origin of the “mauve factor”. Journal of Theoretical Biology, 35(2):277–283, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900392.

Ruse:1973:SIN

[Rus73] Michael Ruse. On the supposed incoherence of numerical taxonomy. Journal of Theoretical Biology, 40(3):603–605, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900131.

Russell:1974:BPI

[Rus74] C. S. Russell. Biosynthesis of porphyrins. II. Journal of Theoretical Biology, 47(1):145–151, September 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901052.

Robinson:1979:CEI

[RV79] James V. Robinson and Walter D. Valentine. The concepts of elasticity, invulnerability and invadability. Journal of Theoretical Biology, 81(1):91–104, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900833.

Rudich:1976:TSF

[RWD76] S. W. Rudich, J. B. Wade, and V. A. DiScala. Theoretical study of flow, pressure and solute concentration in double membrane systems. Journal of Theoretical Biology, 60(1):163–175, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901612.

Ryan:1972:IEV

[Rya72] James Philip F. J. Ryan. Information, entropy and various systems. Journal of Theoretical Biology, 36(1):139–146, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290183X.

Rodenbeck:1975:CAT

[RZMF75] M. Rödenbeck, A. Ziegler, H. Marek, and Th. Friedrich. Computer analysis of a tracer-kinetic model of the collagenic metabolism of the rat’s aorta. Journal of Theoretical Biology, 49(1):191–200, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519375800270.

Schumaker:1970:VLD

[SA70] V. N. Schumaker and G. H. Adams. Very low density lipoproteins: Surface-volume changes during metabolism. Journal of Theoretical Biology, 26(1):89–91, January 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800340.

Sachs:1972:PBA

[Sac72] T. Sachs. A possible basis for apical organization in plants. Journal of Theoretical Biology, 37(2):353–361, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900288.

Sachs:1975:PTR

[Sac75] T. Sachs. Plant tumors resulting from unregulated hormone synthesis. Journal of Theoretical Biology, 55(2):445–453, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800920.

Sagan:1973:USP

[Sag73] Carl Sagan. Ultraviolet selection pressure on the earliest organisms. Journal of Theoretical Biology, 39(1):195–200, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902166.

Salu:1974:CSE

[Sal74] Y. Salu. Computer simulation of electrical activity of the heart. Journal of Theoretical Biology, 43(1):43–49, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800433.

Samish:1971:KSE

[Sam71] Yochai B. Samish. Kinetics of the substrate for the evolution of CO 2 in light by photosynthesizing organs. Journal of Theoretical Biology, 33(3):557–564, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900968.

Samuels:1977:TLN

[Sam77] Stanley Samuels. Transport of the large-neutral amino acids by the γ-glutamyl cycle: A proposal. Journal of Theoretical Biology, 64(4):729–738, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902715.

Sandler:1974:DTD

[San74] Sheldon S. Sandler. Direct three-dimensional reconstruction of a corneal stromal lamella from electron micrographs. Journal of Theoretical Biology, 48(1):207–212, IN1–IN3, 213, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490191X.

Shuler:1972:DEE

[SAT72] M. L. Shuler, R. Aris, and H. M. Tsuchiya. Diffusive and electrostatic effects with insolubilized enzymes. Journal of Theoretical Biology, 35(1):67–76, April 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901920.

Sattler:1974:NCS

[Sat74] Rolf Sattler. A new conception of the shoot of higher plants. Journal of Theoretical Biology, 47(2):367–382, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902045.

Smith:1979:AIA

[SATT79] Truett L. Smith, Juan J. Angulo, Janice O. Tsokos, and Chris P. Tsokos. An analysis of the influence of age and school-attendance status on the spread of variola minor. Journal of Theoretical Biology, 76(2):157–165, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903679.

Savageau:1970:BSA

[Sav70] Michael A. Savageau. Biochemical systems analysis: III. Dynamic solutions using a power-law approximation. Journal of Theoretical Biology, 26(2):215–226, February 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800133.

Savageau:1979:FIBa

[Sav79] Michael A. Savageau. Feedforward inhibition in biosynthetic pathways: inhibition of the aminoacyl-tRNA synthetase by the penultimate product. Journal of Theoretical Biology, 77(4):385– 404, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937990016X.

Stuart:1971:QTV

[SB71] R. N. Stuart and E. W. Branscomb. Quantitative theory of in vivo lac regulation: Significance of repressor packaging: I. Equilibrium considerations. Journal of Theoretical Biology, 31(2): 313–329, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901901.

Szabadi:1973:SDT

[SB73] E. Szabadi and C. M. Bradshaw. Spread of drags in tissue: Diffusion or iontophoresis? Journal of Theoretical Biology, 42(1):185–187, November 5, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901586.

Saijnders:1975:SFC

[SB75a] P. T. Saijnders and M. J. Bazin. On the stability of food chains. Journal of Theoretical Biology, 52(1):121–142, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900442.

Strain:1975:MCM

[SB75b] George M. Strain and William H. Brockman. A modified cable model for neuron processes with non-constant diameters. Journal of Theoretical Biology, 51(2):475–494, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900752.

Solimano:1976:TGT

[SB76] F. Solimano and E. Beretta. The tree graphs theory for enzymatic reactions. Journal of Theoretical Biology, 59(1):159– 166, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251937680029X.

Sabesan:1977:EMT

[SB77] M. N. Sabesan and R. Balasubramanian. An energy minimization technique using the least-squares method. Journal of Theoretical Biology, 68(2):241–246, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790162X.

Sokolove:1978:CSP

[SB78a] Phillip G. Sokolove and Wayne N. Bushell. The chi square periodogram: Its utility for analysis of circadian rhythms. Journal of Theoretical Biology, 72(1):131–160, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890022X.

Stein:1978:AFC

[SB78b] Robert B. Stein and J. J. Blum. On the analysis of futile cycles in metabolism. Journal of Theoretical Biology, 72(3):487– 522, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903144.

Sonnenberg:1979:SCV

[SB79] A. Sonnenberg and A. L. Blum. Simple calculations of volume flow and flow rates across the pylorus. Journal of Theoretical Biology, 79(2):235–242, July 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902480.

Smythies:1974:MSS

[SBB^{+}74a] J. R. Smythies, F. Benington, R. J. Bradley, W. F. Bridgers, and R. D. Morin. The molecular structure of the sodium channel. Journal of Theoretical Biology, 43(1):29–32, IN1–IN8, 33– 42, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800421.

Smythies:1974:MMA

[SBB^{+}74b] J. R. Smythies, F. Benington, R. J. Bradley, W. F. Bridgers, and R. D. Morin. On the molecular mechanism of action of scorpion neurotoxin: II. From androctonus australis hector. Journal of Theoretical Biology, 43(1):65–72, IN9, 69–72, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800457.

Smythies:1974:MIBa

[SBB^{+}74c] J. R. Smythies, F. Benington, R. J. Bradley, R. D. Morin, and W. O. Romine. On the mechanism of interaction between histone I and DNA and histone III and DNA. Journal of Theoretical Biology, 47(2):309–315, October 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902008.

Smythies:1974:MIBb

[SBB^{+}74d] J. R. Smythies, F. Benington, R. J. Bradley, R. D. Morin, and W. O. Romine. On the mechanism of interaction between histone II B 1 and DNA and histone II B 2 and DNA. Journal of Theoretical Biology, 47(2):383–395, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374902057.

Smythies:1975:MSR

[SBB^{+}75] J. R. Smythies, F. Benington, R. J. Bradley, W. F. Bridgers, and … W. O. Romine. The molecular structure of the receptorionophore complex at the neuromuscular junction. Journal of Theoretical Biology, 51(1):111–126, May 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901423.

Smythies:1976:MMI

[SBB^{+}76] J. R. Smythies, J. M. Beaton, F. Benington, R. J. Bradley, and R. F. Morin. On the molecular mechanism of interaction between the neurophysins and oxytocin and vasopressin. Journal of Theoretical Biology, 63(1):33–46, IN1, 47–48, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900825.

Sorarrain:1979:ABE

[SBC79] Oscar M. Sorarrain, Rafael Boggio, and Cristina Cordo. An actual Bellman–0Harris epidemic process. The infection of septoria in barley or wheat. Journal of Theoretical Biology, 80(3):425–433, October 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901024.

Smythies:1972:MIH

[SBM72] J. R. Smythies, F. Benington, and R. D. Morin. A mechanism for the interaction of a histone and DNA. Journal of Theoretical Biology, 37(1):151–154, IN3, 155–158, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290121X.

Solimano:1977:TGT

[SBP77] F. Solimano, E. Beretta, and E. Piatti. The tree graphs theory for enzymatic reactions: A theorem for the reactions among the non-enzymatic species. Journal of Theoretical Biology, 64(3):401–412, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902788.

Simon:1977:MDR

[SBR77a] Zeno Simon, Ilie Badilescu, and Traian Racovitan. Mapping of dihydrofolate-reductase receptor site by correlation with minimal topological (steric) differences. Journal of Theoretical Biology, 66(3):485–495, June 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902983.

Srinivasan:1977:APC

[SBR77b] R. Srinivasan, R. Balasubramanian, and S. S. Rajan. Analysis of protein crystallographic structaral data: II analysis of θ and δ parameters on some proteins. Journal of Theoretical Biology, 67(2):299–312, July 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902028.

Soucek:1975:FPR

[SC75a] B. Soucek and A. D. Carlson. Flash pattern recognition in fireflies. Journal of Theoretical Biology, 55(2):339–352, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800841.

Steward:1975:CAG

[SC75b] E. G. Steward and G. R. Clarke. Conformational adaptability at GABA-sensitive inhibitory synapses. Journal of Theoretical Biology, 52(2):493–498, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900156.

Spencer:1977:ATA

[SC77] N. Spencer and N. Carter. Amino terminal acetylation and the levels of some erythrocyte proteins. Journal of Theoretical Biology, 67(2):279–286, July 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902004.

Sharpe:1977:DMO

[SCDC77] Peter J. H. Sharpe, Guy L. Curry, Don W. DeMichele, and Charles L. Cole. Distribution model of organism development times. Journal of Theoretical Biology, 66(1):21–38, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903095.

Scheie:1970:ELC

[Sch70] Paul O. Scheie. Environmental limits of cellular existence. Journal of Theoretical Biology, 28(3):315–325, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900743.

Schaap:1971:DID

[Sch71] Tamar Schaap. Dual information in DNA and the evolution of the genetic code. Journal of Theoretical Biology, 32(2):293– 298, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901676.

Schotz:1972:CLI

[Sch72a] W. E. Schotz. Continuous labeling indices: CLI(t) and CLMII(t). Journal of Theoretical Biology, 34(1):29–45, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900525.

Schotz:1972:DLI

[Sch72b] W. E. Schotz. Double label indices. Journal of Theoretical Biology, 36(2):397–412, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901075.

Schwarz:1972:CRC

[Sch72c] Gerhard Schwarz. Chemical relaxation of co-operative conformational transitions of linear biopolymers. Journal of Theoretical Biology, 36(3):569–580, September 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900094.

Schleidt:1973:TCC

[Sch73a] Wolfgang M. Schleidt. Tonic communication: Continual effects of discrete signs in animal communication systems. Journal of Theoretical Biology, 42(2):359–386, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900957.

Schreiner:1973:GRC

[Sch73b] Kristian E. Schreiner. Geometrical restrictions on crossbridge motion and sliding filament dynamics. Journal of Theoretical Biology, 40(3):591–601, August 22, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390012X.

Schotz:1974:DLE

[Sch74] W. E. Schotz. Double label estimation of the mean duration of the s-phase. Journal of Theoretical Biology, 46(2):353–368, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900046.

Schulman:1976:SGS

[Sch76a] L. S. Schulman. Stable generation of simple forms. Journal of Theoretical Biology, 57(2):453–468, April 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900151.

Schwabauer:1976:EN

[Sch76b] Robert Schwabauer. Enzymatic neurons. Journal of Theoretical Biology, 59(1):223–230, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519376800343.

Schobert:1977:TOR

[Sch77a] Brigitte Schobert. Is there an osmotic regulatory mechanism in algae and higher plants? Journal of Theoretical Biology, 68(1):17–26, September 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902247.

Schwartz:1977:DSV

[Sch77b] Eric L. Schwartz. The development of specific visual connections in the monkey and the goldfish: Outline of a geometric theory of receptotopic structure. Journal of Theoretical Biology, 69(4):655–664, IN7, 665–683, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903745.

Schaup:1978:PRF

[Sch78] H. W. Schaup. A paradigm for ribosome fidelity modulation in protein synthesis. Journal of Theoretical Biology, 71(2):215– 224, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902680.

Scheie:1979:FCO

[Sch79a] Paul O. Scheie. Forces causing osmotic pressure. Journal of Theoretical Biology, 77(1):47–50, March 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990136X.

Schenzle:1979:EPB

[Sch79b] D. Schenzle. On estimating parameters of the binomial model for transmitter release at synapses. Journal of Theoretical Biology, 80(3):441–443, October 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901048.

Schnakenberg:1979:SCR

[Sch79c] J. Schnakenberg. Simple chemical reaction systems with limit cycle behaviour. Journal of Theoretical Biology, 81(3):389–400, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900420.

Schoen:1979:AAM

[Sch79d] Daniel J. Schoen. An angiosperm analogue to megagametophyte analysis. Journal of Theoretical Biology, 79(4):543–546, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902418.

Scott:1974:MPT

[Sco74] H. L. Scott. A model for phase transitions in lipid bilayers and biological membranes. Journal of Theoretical Biology, 46(1): 241–253, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901507.

Scott:1979:HEP

[Sco79a] J. E. Scott. Hydrated electrons and prebiotic evolution. Journal of Theoretical Biology, 81(1):19–27, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900778.

Scott:1979:CCTa

[Sco79b] P. D. Scott. Conditional control transfer mechanisms in the neocortex: 1. The basic model. Journal of Theoretical Biology, 81(3):423–452, December 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900456.

Scott:1979:CCTb

[Sco79c] P. D. Scott. Conditional control transfer mechanisms in the neocortex: 2. Cytoarchitectural and electrophysiological evidence and predictions. Journal of Theoretical Biology, 81(3):453–473, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900468.

Sullivan:1974:MGC

[SD74] Jerome L. Sullivan and A. Gib DeBusk. Membrane glycoproteins and cellular ageing. Journal of Theoretical Biology, 46(1): 291–294, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901556.

Somogyi:1975:RBL

[SD75] B. Somogyi and S. Damjanovich. Relationship between the lifetime of an enzyme-substrate complex and the properties of the molecular environment. Journal of Theoretical Biology, 51(2): 393–401, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900685.

Stubblefield:1976:STK

[SD76] Elton Stubblefield and Cheryl M. Dennis. The simulation of thymidine kinase enzyme kinetics in cultured cells using the computer language cellsim. Journal of Theoretical Biology, 61(1):171–184, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901119.

Sharpe:1977:RKP

[SD77] Peter J. H. Sharpe and Don W. DeMichele. Reaction kinetics of poikilotherm development. Journal of Theoretical Biology, 64(4):649–670, February 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790265X.

Silk:1979:KPG

[SE79] Wendy Kuhn Silk and Ralph O. Erickson. Kinematics of plant growth. Journal of Theoretical Biology, 76(4):481–501, February 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900146.

Seaton:1976:DPI

[Sea76] B. Seaton. The detection of periodicity in irregular data. Journal of Theoretical Biology, 63(2):311–324, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900369.

Sebba:1979:SCB

[Seb79] Felix Sebba. A surface-chemical basis for cell motility. Journal of Theoretical Biology, 78(3):375–388, IN1–IN10, 389–391, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903370.

Seelig:1970:USO

[See70] Friedrich Franz Seelig. Undamped sinusoidal oscillations in linear chemical reaction systems. Journal of Theoretical Biology, 27(2):197–206, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901372.

Seelig:1971:STM

[See71a] Friedrich F. Seelig. Systems-theoretic model for the spontaneous formation of optical antipodes in strongly asymmetric yield. Journal of Theoretical Biology, 31(2):355–361, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901937.

Seelig:1971:AEC

[See71b] Friedrich Franz Seelig. Activated enzyme catalysis as a possible realization of the stable linear chemical oscillator model. Journal of Theoretical Biology, 30(3):497–514, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190004X.

Seelig:1971:MBB

[See71c] Friedrich Franz Seelig. Mono- or bistable behaviour in a weakly or strongly open chemical reaction system. Journal of Theoretical Biology, 32(1):93–106, July 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190138X.

Seelig:1972:SAS

[See72] Friedrich F. Seelig. On spontaneous asymmetric synthesis. Journal of Theoretical Biology, 34(1):197–198, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900641.

Seely:1973:ESV

[See73a] G. R. Seely. Effects of spectral variety and molecular orientation on energy trapping in the photosynthetic unit: A model calculation. Journal of Theoretical Biology, 40(1):173–187, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901707.

Seely:1973:ETM

[See73b] G. R. Seely. Energy transfer in a model of the photosynthetic unit of green plants. Journal of Theoretical Biology, 40(1):189– 199, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901719.

Segel:1970:SGF

[Seg70] Lee A. Segel. Standing-gradient flows driven by active solute transport. Journal of Theoretical Biology, 29(2):233–250, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900202.

Segel:1976:IRK

[Seg76] Lee A. Segel. Incorporation of receptor kinetics into a model for bacterial chemotaxis. Journal of Theoretical Biology, 57(1): 23–42, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800045.

Seghers:1979:QSI

[Seg79] M. Seghers. A qualitative study of an idiotypic cyclic network. Journal of Theoretical Biology, 80(4):553–576, October 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900936.

Sejnowski:1977:SCS

[Sej77] T. J. Sejnowski. Statistical constraints on synaptic plasticity. Journal of Theoretical Biology, 69(2):385–389, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901461.

Seshadri:1971:ETS

[Ses71] M. S. Seshadri. Evolution towards steady state in membrane transport problems. Journal of Theoretical Biology, 31(2):245– 253, May 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371901858.

Seshagiri:1972:CMA

[Ses72] N. Seshagiri. Chart method for the analysis of enzyme kinetic reactions. Journal of Theoretical Biology, 34(3):469–486, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901361.

Seshadri:1974:RDC

[Ses74] M. S. Seshadri. A reaction-diffusion coupled system with multiple steady states. Journal of Theoretical Biology, 47(2):351– 365, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374902033.

Seymour:1972:CHT

[Sey72] Roger S. Seymour. Convective heat transfer in the respiratory systems of panting animals. Journal of Theoretical Biology, 35(1):119–127, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937290197X.

Seyfarth:1977:MSG

[Sey77] Robert M. Seyfarth. A model of social grooming among adult female monkeys. Journal of Theoretical Biology, 65(4):671–698, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900157.

Stayton:1979:CAV

[SF79] Mark M. Stayton and Herbert J. Fromm. A computer analysis of the validity of the integrated Michaelis–Menten equation. Journal of Theoretical Biology, 78(3):309–323, June 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903333.

Schultz:1977:EEC

[SFN77] Stanley G. Schultz, Raymond A. Frizzell, and Hugh N. Nellans. An equivalent electrical circuit model for “sodiumtransporting” epithelia in the steady-state. Journal of Theoretical Biology, 65(2):215–229, March 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903228.

Sineriz:1975:CUA

[SFT75] Fausmo Siñeriz, Ricardo N. Farias, and Raul E. Trucco. The convenience of the use of allosteric “probes” for the study of lipid-protein interactions in biological membranes: Thermodynamic considerations. Journal of Theoretical Biology, 52(1): 113–120, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900430.

Saran:1971:QCS

[SG71a] Anil Saran and Girjesh Govil. Quantum chemical studies on the conformational structure of nucleic acids III. Calculation of backbone structure by extended Hückel theory. Journal of Theoretical Biology, 33(2):407–418, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900749.

Seelig:1971:SLR

[SG71b] Friedrich F. Seelig and Friedrich Göbber. Stable linear reaction oscillator. Journal of Theoretical Biology, 30(3):485–496, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900038.

Singh:1971:PUM

[SG71c] U. N. Singh and R. S. Gupta. Polyribosomes and unstable messenger RNA II. Some further implications of the tape theory of protein synthesis. Journal of Theoretical Biology, 30(3):603– 619, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900130.

Saran:1972:CCC

[SG72] Anil Saran and Girjesh Govil. CNDO calculations on the conformational structure of acetylcholine. Journal of Theoretical Biology, 37(1):181–185, October 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901245.

Strebel:1973:IMA

[SG73] Donald E. Strebel and Narendra S. Goel. On the isocline methods for analyzing prey-predator interactions. Journal of Theoretical Biology, 39(1):211–239, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390218X.

Smeach:1975:SDMa

[SG75a] Stephen C. Smeach and Harvey J. Gold. Stochastic and deterministic models for the kinetic behavior of certain structured enzyme systems i: One enzyme-one substrate systems. Journal of Theoretical Biology, 51(1):59–78, May 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901393.

Smeach:1975:SDMb

[SG75b] Stephen C. Smeach and Harvey J. Gold. Stochastic and deterministic models for the kinetic behavior of certain structured enzyme systems II: Consecutive two enzyme systems. Journal of Theoretical Biology, 51(1):79–96, May 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590140X.

Shymko:1976:CGC

[SG76] R. M. Shymko and Leon Glass. Cellular and geometric control of tissue growth and mitotic instability. Journal of Theoretical Biology, 63(2):355–374, December 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900394.

Scheufens:1972:UGD

[SH72a] Ernst Edvard Scheufens and Niels Rudolph Hartmann. Use of gamma distributed transit times and the Laplace transform method in theoretical cell kinetics. Journal of Theoretical Biology, 37(3):531–543, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900896.

Silvester:1972:AHF

[SH72b] N. R. Silvester and M. E. J. Holwill. An analysis of hypothetical flagellar waveforms. Journal of Theoretical Biology, 35(3):505–506, IN3, 507–523, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901488.

Stusnick:1972:NDM

[SH72c] Eric Stusnick and R. P. Hurst. Numerical determination of membrane permeability parameters. Journal of Theoretical Biology, 37(2):261–271, November 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900215.

Saunders:1976:ICE

[SH76] P. T. Saunders and M. W. Ho. On the increase in complexity in evolution. Journal of Theoretical Biology, 63(2):375–384, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900400.

Saunders:1977:DC

[SH77] P. T. Saunders and M. W. Ho. In defense of complexity. Journal of Theoretical Biology, 69(1):235–237, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903976.

Schauer:1979:AQS

[SH79] M. Schauer and R. Heinrich. Analysis of the quasi-steady-state approximation for an enzymatic one-substrate reaction. Journal of Theoretical Biology, 79(4):425–442, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902352.

Shackney:1973:CMH

[Sha73] Stanley E. Shackney. A cytokinetic model for heterogeneous mammalian cell populations: I. Cell growth and cell death. Journal of Theoretical Biology, 38(2):305–333, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390177X.

Shackney:1974:CMH

[Sha74] Stanley E. Shackney. A cytokinetic model for heterogeneous mammalian cell populations: II. Tritiated thymidine studies: The per cent labeled mitosis (PLM) curve. Journal of Theoretical Biology, 44(1):49–90, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800299.

Shackney:1977:CMH

[Sha77] Stanley E. Shackney. A cytokinetic model for heterogeneous mammalian cell populations. III. tritiated thymidine studies: Correlations among multiple kinetic parameters in human tumors. Journal of Theoretical Biology, 65(3):421–464, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902077.

Shahinpoor:1978:RPS

[Sha78] M. Shahinpoor. The role of parametric self-excitation in DNA self-replication. Journal of Theoretical Biology, 70(1):17–22, January 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902990.

Stuck:1971:AOD

[SHC71] J. D. Stuck, J. A. Howell, and H. T. Cullinan. Analysis of oxygen diffusion to a measuring probe in animal tissue. Journal of Theoretical Biology, 31(3):509–532, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900269.

Shear:1970:EFM

[She70] David B. Shear. Electrostatic forces in muscle contraction. Journal of Theoretical Biology, 28(3):531–546, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900871.

Sheppard:1971:SMT

[She71] C. W. Sheppard. Stochastic models for tracer experiments in the circulation III. The lumped catenary system. Journal of Theoretical Biology, 33(3):491–515, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900932.

Shear:1973:CPS

[She73a] David B. Shear. The chemical potential in solutions and the free energy of reaction. Journal of Theoretical Biology, 39(1): 171–177, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902129.

Shear:1973:GBD

[She73b] David B. Shear. The generalized Boltzmann distribution. Journal of Theoretical Biology, 39(1):165–169, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902117.

Strehler:1971:CRT

[SHG^{+}71] B. Strehler, G. Hirsch, D. Gusseck, R. Johnson, and M. Bick. Codon-restriction theory of aging and development. Journal of Theoretical Biology, 33(3):429–474, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900919.

Shiba:1971:HBC

[Shi71] Haruo Shiba. Heaviside’s “bessel cable” as an electric model for flat simple epithelial cells with low resistive junctional membranes. Journal of Theoretical Biology, 30(1):59–68, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900361.

Shinagawa:1977:DPS

[Shi77a] Yoshiya Shinagawa. Donnan potential in the stationary state. Journal of Theoretical Biology, 64(3):551–556, February 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902879.

Shinagawa:1977:ENC

[Shi77b] Yoshiya Shinagawa. Evidence for a new concept of Donnan potential in the stationary state. Journal of Theoretical Biology, 69(1):231–234, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903964.

Shinagawa:1978:ASP

[Shi78a] Yoshiya Shinagawa. Analytical solution of the Poisson– Boltzmann equation for membrane potential. Journal of Theoretical Biology, 72(4):603–610, June 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902746.

Shine:1978:PSP

[Shi78b] Richard Shine. Propagule size and parental care: The “safe harbor” hypothesis. Journal of Theoretical Biology, 75(4):417–424, December 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903533.

Shinagawa:1979:NVC

[Shi79] Yoshiya Shinagawa. Note on the validity of constant field assumption in relation to the exact solution of Nernst–Planck equations. Journal of Theoretical Biology, 81(2):333–340, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937990170X.

Shotkin:1974:MLL

[Sho74a] Louis M. Shotkin. A model for LH levels in the recentlycastrated adult rat and its comparison with experiment. Journal of Theoretical Biology, 43(1):1–14, January 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800408.

Shotkin:1974:MED

[Sho74b] Louis M. Shotkin. A model for the effect of daily injections of gonadal hormones on LH levels in recently-castrated adult rats and its comparison with experiment. Journal of Theoretical Biology, 43(1):15–28, January 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937480041X.

Shuter:1979:MPA

[Shu79] Brian Shuter. A model of physiological adaptation in unicellular algae. Journal of Theoretical Biology, 78(4):519–552, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901899.

Sibatani:1978:BMD

[Sib78] A. Sibatani. A bifurcation model for developmental processes. Journal of Theoretical Biology, 75(2):149–166, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902278.

Siekevitz:1972:TPU

[Sie72] Philip Siekevitz. The turnover of proteins and the usage of information. Journal of Theoretical Biology, 37(2):321–334, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900264.

Silver:1976:DDG

[Sil76] Ira Lee Silver. The dynamics of a discrete geotropic sensor subject to rotation-induced gravity compensation. Journal of Theoretical Biology, 61(2):353–362, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900230.

Simon:1973:BCM

[Sim73a] Z. Simon. Bacterial cell model: Cell cycle parameters and macromolecular synthesis. Journal of Theoretical Biology, 38(1):39–49, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902233.

Simons:1973:CFA

[Sim73b] R. Simons. The constant field assumption for symmetric membrane systems. Journal of Theoretical Biology, 39(1):191– 194, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902154.

Simon:1979:IPF

[Sim79] I. Simon. Investigation of protein folding: Uneven distribution of point mutations along polypeptide chains. Journal of Theoretical Biology, 81(2):247–258, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901644.

Singh:1973:PUM

[Sin73] U. N. Singh. Polyribosomes and unstable messenger RNA III. Temporal relationship between the syntheses of mRNA and inducible enzymet. Journal of Theoretical Biology, 40(3):553–571, August 22, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373900106.

Sinha:1975:AMM

[Sin75] A. S. C. Sinha. Analysis of a mathematical model for transport of 131 i-albumin. Journal of Theoretical Biology, 50(2):259– 267, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900806.

Singh:1976:AMO

[Sin76] U. N. Singh. Adaptation in micro-organisms: Variation in macromolecular composition with growth rate. Journal of Theoretical Biology, 59(1):107–126, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800264.

Singh:1978:AMO

[Sin78] U. N. Singh. Adaptation in micro-organisms II: Transient behaviour following shift-up and shift-down. Journal of Theoretical Biology, 72(3):459–474, June 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903120.

Sirs:1970:SE

[Sir70] John A. Sirs. Structure of the erythrocyte. Journal of Theoretical Biology, 27(1):107–110, IN1–IN2, 111–115, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901323.

Seno:1978:IOB

[SIS78] Manabu Seno, Kazutoshi Iwamoto, and Kiyoshi Sawada. Instability and oscillatory behavior of membrane-chemical reaction systems. Journal of Theoretical Biology, 72(3):577–588, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903181.

Segel:1972:DSE

[SJ72] Lee A. Segel and Julius L. Jackson. Dissipative structure: An explanation and an ecological example. Journal of Theoretical Biology, 37(3):545–559, December 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900902.

Savageau:1979:FIBb

[SJ79] Michael A. Savageau and Gerald Jacknow. Feedforward inhibition in biosynthetic pathways: inhibition of the aminoacyltRNA synthetase by intermediates of the pathway. Journal of Theoretical Biology, 77(4):405–425, April 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900171.

Skadhauge:1972:ACS

[SK72] E. Skadhauge and K. Kristensen. An analogue computer simulation of cloacal resorption of salt and water from ureteral urine in birds. Journal of Theoretical Biology, 35(3):473–487, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901464.

Stasiak:1979:FSD

[SK79] A. Stasiak and T. Klopotowski. Four-stranded DNA structure and DNA base methylation in the mechanism of action of restriction endonucleases. Journal of Theoretical Biology, 80(1):65–82, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901802.

Sanyal:1977:EAU

[SKS77] Nitish K. Sanyal, U. Kumar, and Ravi P. Sinha. Energy of association of uracil with Watson-Crick a-t pair. Journal of Theoretical Biology, 66(1):39–46, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903101.

Shigesada:1979:SSI

[SKT79] Nanako Shigesada, Kohkichi Kawasaki, and Ei Teramoto. Spatial segregation of interacting species. Journal of Theoretical Biology, 79(1):83–99, July 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902583.

Somogyi:1978:EVA

[SKTC78] Béla Somogyi, Frank E. Karasz, Lajos Trón, and Peter R. Couchma. The effect of viscosity on the apparent decomposition rate on enzyme-ligand complexes. Journal of Theoretical Biology, 74(2):209–216, September 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900723.

Slatkin:1978:SPD

[Sla78] Montgomery Slatkin. Spatial patterns in the distributions of polygenic characters. Journal of Theoretical Biology, 70 (2):213–228, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937890348X.

Saidel:1976:SDM

[SLK76] Gerald M. Saidel, Lance A. Liotta, and Jerome Kleinerman. System dynamics of a metastatic process from an implanted tumor. Journal of Theoretical Biology, 56(2):417–434, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800835.

Surkinshaw:1971:AEP

[SM71] L. Surkinshaw and D. H. Marshall. Accuracy of estimation of the parameters in a model of calcium tracer kinetics. Journal of Theoretical Biology, 30(2):255–265, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900531.

Smith:1976:NSC

[SM76] Warren D. Smith and R. R. Mohler. Necessary and sufficient conditions in the tracer determination of compartmental system order. Journal of Theoretical Biology, 57(1):1–21, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800033.

Sperelakis:1977:EEF

[SM77] Nick Sperelakis and James E. Mann. Evaluation of electric field changes in the cleft between excitable cells. Journal of Theoretical Biology, 64(1):71–96, January 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790114X.

Skagen:1978:ATQ

[SM78] Dankert W. Skagen and Lars Mørkrid. An approach to the theory of quantitative and double label autoradiography. Journal of Theoretical Biology, 70(2):185–197, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903466.

Smith:1979:PES

[SM79] R. H. Smith and R. Mead. On predicting extinction in simple population models I. Stochastic linearization. Journal of Theoretical Biology, 80(2):189–203, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902054.

Smart:1976:STB

[Sma76] J. S. Smart. Statistical tests of the broken-stick model of speciesabundance relations. Journal of Theoretical Biology, 59(1):127– 139, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800276.

Smerage:1976:MEF

[Sme76] Glen H. Smerage. Matter and energy flows in biological and ecological systems. Journal of Theoretical Biology, 57(1):203– 223, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519376800136.

Smith:1971:WUS

[Smi71]  J. Maynard Smith. What use is sex?  Journal of

Theoretical Biology, 30(2):319–335, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900580.

Smith:1973:DIF

[Smi73] P. G. Smith. Dependence of ion flux across a membrane on ionic concentration. Journal of Theoretical Biology, 41(2):269–286, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901197.

Smith:1974:TGE

[Smi74] J. Maynard Smith. The theory of games and the evolution of animal conflicts. Journal of Theoretical Biology, 47(1):209–221, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901106.

Smith:1976:STA

[Smi76a] J. Maynard Smith. A short-term advantage for sex and recombination through sib-competition. Journal of Theoretical Biology, 63(2):245–258, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900333.

Smith:1976:SSH

[Smi76b] John Maynard Smith. Sexual selection and the handicap principle. Journal of Theoretical Biology, 57(1):239–242, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800161.

Smith:1977:ASD

[Smi77] Geoffrey D. Smith. Active site-directed and allosteric effectors of regulatory enzymes: The activation of aspartate transcarbamylase by substrate and transition state analogues. Journal of Theoretical Biology, 69(2):275–285, November 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901369.

Smith:1978:PDS

[Smi78a] H. J. Smith. Perspectives in the design of small molecule enzyme inhibitors as useful drugs. Journal of Theoretical Biology, 73(3):531–538, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937890156X.

Smith:1978:HPC

[Smi78b] J. Maynard Smith. The handicap principle — a comment. Journal of Theoretical Biology, 70(2):251–252, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890351X.

Smith:1973:VWF

[SMN73] E. R. Smith, D. J. Mitchell, and B. W. Ninham. Van der Waals forces between cylinders: Arrays of thin cylinders and three body forces. Journal of Theoretical Biology, 41(1):149–160, September 14, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901951.

Smoes:1976:TMD

[Smo76] Marie-Luce Smoes. Toward a mathematical description of chemical phase waves. Journal of Theoretical Biology, 58(1):1–14, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901351.

Smythies:1972:MMA

[Smy72] J. R. Smythies. Molecular mechanisms of adrenergic receptors. Journal of Theoretical Biology, 35(1):93–96, IN5–IN12, 97–101, April 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901944.

Smythies:1979:PMS

[Smy79] J. R. Smythies. The possible molecular structure of binding sites for tumor-promoting phorbol esters, 2’-deacyl phospholipids and PGE 1 on the molecule of phospholipase a 2. Journal of Theoretical Biology, 81(3):401–405, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900432.

Saito:1974:LON

[SO74] Nobuhiko Saitô and Masaaki Oosima. Localized oscillations in neural networks I. General considerations with nonlinear differential equations. Journal of Theoretical Biology, 47(2):421– 437, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374902082.

Sauve:1979:IDC

[SO79a] Rémy Sauvé and Shinpei Ohki. Interactions of divalent cations with negatively charged membrane surfaces. I. Discrete charge potential. Journal of Theoretical Biology, 81(2):157–179, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901589.

So:1979:NGS

[So79b] Joseph W. H. So. A note on the global stability and bifurcation phenomenon of a Lotka–Volterra food chain. Journal of Theoretical Biology, 80(2):185–187, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902042.

Steiner:1972:ABT

[SOF72] R. F. Steiner, J. Osborne, and K. Flanigan. The analysis of the binding of two ligands by an allosteric protein. Journal of Theoretical Biology, 36(3):459–477, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900033.

Soucek:1971:CMS

[Sou71] Branko Soucek. Complete model for the statistical composition of the end-plate potential. Journal of Theoretical Biology, 30(3):631–645, March 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900154.

Soucek:1975:MAA

[Sou75] Branko Soucek. Model of alternating and aggressive communication with the example of katydid chirping. Journal of Theoretical Biology, 52(2):399–417, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900090.

Stein:1971:PAF

[SP71] R. B. Stein and K. G. Pearson. Predicted amplitude and form of action potentials recorded from unmyelinated nerve fibres. Journal of Theoretical Biology, 32(3):539–558, September 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190155X.

Savage:1973:EVG

[SP73a] John R. K. Savage and D. G. Papworth. The effect of variable G_{2} duration upon the interpretation of yield-time curves of radiation-induced chromatid aberrations. Journal of Theoretical Biology, 38(1):17–38, January 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902221.

Silver:1973:EAN

[SP73b] W. S. Silver and J. R. Postgate. Evolution of asymbiotic nitrogen fixation. Journal of Theoretical Biology, 40(1):1–10, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901604.

Savage:1975:SPS

[SP75] John R. K. Savage and D. G. Papworth. Some problems of sampling for chromosomal aberrations from synchronous populations. Journal of Theoretical Biology, 54(2):129–152, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801196.

Spencer:1970:PBH

[Spe70] Richard P. Spencer. Physical basis for Haase’s law of fetal growth. Journal of Theoretical Biology, 26(3):507–508, March 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901001.

Spertell:1976:TIR

[Spe76] Robert B. Spertell. A theoretical inquiry into the role of phospholipids in membrane fusion. Journal of Theoretical Biology, 60(1):197–208, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901648.

Spertell:1978:DLP

[Spe78] Robert B. Spertell. On the decay of lateral phase separations in biological membranes. Journal of Theoretical Biology, 71(1): 1–9, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902096.

Steward:1972:IXR

[SR72] V. W. Steward and S. Rudofsky. Integrating X-ray absorption for mass determination of cells: A new calibration method: Practical and theoretical. Journal of Theoretical Biology, 37(1):115–116, IN1, 117–124, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290118X.

Streilein:1976:MMI

[SR76] J. Wayne Streilein and Campbell B. Read. A mathematical model in immunoregulation where the spleen has a pivotal role. Journal of Theoretical Biology, 61(2):363–376, September 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900242.

Sitaramam:1977:HMA

[SR77] V. Sitaramam and N. J. Rao. Hierarchical modelling of acclimatory processes. Journal of Theoretical Biology, 67(1):25–47, July 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901837.

Srebro:1974:REE

[Sre74] Zbigniew Srebro. Relative electron excess and relative electron deficiency: An explanation for differences in the response of the organism to various factors. Journal of Theoretical Biology, 46(2):573–578, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900162.

Shukla:1979:MHM

[SRML79] K. K. Shukla, F. Ramirez, J. F. Marecek, and H. M. Levy. A mechanism for the hydrolysis of MgATP by actomyosin of skeletal muscle. Journal of Theoretical Biology, 76(3):359–367, February 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903187.

Sabah:1970:RRH

[SS70] N. H. Sabah and R. A. Spangler. Repetitive response of the Hodgkin–Huxley model for the squid giant axon. Journal of Theoretical Biology, 29(2):155–171, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900172.

Segel:1972:ILC

[SS72] Lee A. Segel and Beth Stoeckly. Instability of a layer of chemotactic cells, attractant and degrading enzyme. Journal of Theoretical Biology, 37(3):561–585, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900914.

Shoemaker:1973:MSO

[SS73a] E. M. Shoemaker and L. M. Srivastava. The mechanics of stomatal opening in corn (Zea mays L.) leaves. Journal of Theoretical Biology, 42(2):219–220, IN1–IN2, 221–225, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900866.

Smeach:1973:CSD

[SS73b] Stephen C. Smeach and Woollocott Smith. A comparison of stochastic and deterministic models for cell membrane transport. Journal of Theoretical Biology, 42(1):157–167, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901549.

Shall:1979:MTC

[SS79] Sydney Shall and Wilfred D. Stein. A mortalization theory for the control of cell proliferation and for the origin of immortal cell lines. Journal of Theoretical Biology, 76(2):219–231, January 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903710.

Scribner:1974:NSF

[SSR74] Terry L. Scribner, Lee A. Segel, and Edwin H. Rogers. A numerical study of the formation and propagation of traveling bands of chemotactic bacteria. Journal of Theoretical Biology, 46(1): 189–219, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901477.

Sinha:1972:HSE

[SSZ^{+}72] Arabinda K. Sinha, Harvey Smythe, Vincent P. Zarcone, Jack D. Barchas, and William C. Dement. Human sleepelectroencephalogram: A damped oscillatory phenomenon. Journal of Theoretical Biology, 35(2):387–393, May 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290046X.

Sato:1976:SPT

[ST76a] Shunsuke Sato and Nakaakira Tsukahara. Some properties of the theoretical membrane transients in Rall’s neuron model. Journal of Theoretical Biology, 63(1):151–163, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900904.

Starzak:1976:OSK

[ST76b] Michael E. Starzak and Daniel A. Tranchina. An open system kinetic transport model for the Hodgkin–Huxley equations. Journal of Theoretical Biology, 56(2):283–299, February 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800768.

Stevenson:1976:FRK

[ST76c] Duncan R. Stevenson and Colin J. Thompson. Fairy ring kinetics. Journal of Theoretical Biology, 58(1):143–163, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901442.

Sepulveda:1977:SSS

[ST77] A. Sepúlveda and J. C. Toha. Secondary structure in synthetic messenger RNAs from immunoglobulins. Journal of Theoretical Biology, 65(3):591–596, April 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790217X.

Shindler:1978:E

[ST78a] Julian S. Shindler and Keith F. Tipton. Erratum. Journal of Theoretical Biology, 73(1):i, July 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901868.

Shindler:1978:QPD

[ST78b] Julian S. Shindler and Keith F. Tipton. A quantitative plot for the determination of negatively co-operative ligand binding sites in proteins. Journal of Theoretical Biology, 70(3):315–321, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937890379X.

Silver:1979:LAT

[ST79] F. H. Silver and R. L. Trelstad. Linear aggregation and the turbidimetric lag phase: Type I collagen fibrillogenesis in vitro. Journal of Theoretical Biology, 81(3):515–526, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900493.

Staff:1970:SDR

[Sta70] P. J. Staff. A stochastic development of the reversible Michaelis– Menten mechanism. Journal of Theoretical Biology, 27(2):221– 232, May 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901396.

Shuler:1973:DEE

[STA73a] M. L. Shuler, H. M. Tsuchiya, and R. Aris. Diffusive and electrostatic effects with insolubilized enzymes subject to substrate inhibition. Journal of Theoretical Biology, 41(2):347–356, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901239.

Starzak:1973:MCCa

[Sta73b] Michael E. Starzak. A model for conductance changes in the squid giant axon: I. Interactive relaxation. Journal of Theoretical Biology, 39(3):487–504, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900647.

Starzak:1973:MCCb

[Sta73c] Michael E. Starzak. A model for conductance changes in the squid giant axon: II. Channel kinetics. Journal of Theoretical Biology, 39(3):505–522, June 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900659.

Starzak:1976:ACT

[Sta76] Michael E. Starzak. Analysis of clamps with time dependent voltages in the squid axon using a kinetic transport model. Journal of Theoretical Biology, 57(1):153–169, March 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800100.

Stauffer:1979:DWD

[Sta79a] H. B. Stauffer. A derivation for the Weibull distribution. Journal of Theoretical Biology, 81(1):55–63, November 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900808.

Stauffer:1979:GPI

[Sta79b] Howard B. Stauffer. A generalization of Pielou’s index of nonrandomness. Journal of Theoretical Biology, 77(1):19–25, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901334.

Somogyi:1977:PAM

[STD77] Béla Somogyi, Lajos Trón, and Sándor Damjanovich. Physical analysis of the molecular motions during transcription. Journal of Theoretical Biology, 67(1):175–180, July 7, 1977.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901916.

Steiner:1970:AAT

[Ste70] R. F. Steiner. Analysis of the association of two different monomers by combined colligative and solubility measurements. Journal of Theoretical Biology, 28(2):185–193, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900512.

Stewart:1972:ARC

[Ste72] Charles R. Stewart. Asymmetric replication and cellular differentiation. Journal of Theoretical Biology, 36(3):639–640, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937290015X.

Stevens:1973:EE

[Ste73] E. D. Stevens. The evolution of endothermy. Journal of Theoretical Biology, 38(3):597–611, March 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902609.

Steiner:1974:BLA

[Ste74] R. F. Steiner. The binding of ligand by an associating protein system. Journal of Theoretical Biology, 45(1):93–106, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900459.

Steinberg:1975:AGM

[Ste75] Malcolm S. Steinberg. Adhesion-guided multicellular assembly: a commentary upon the postulates, real and imagined, of the differential adhesion hypothesis, with special attention to computer simulations of cell sorting. Journal of Theoretical Biology, 55(2):431–432, IN1–IN3, 433–443, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800919.

Steele:1976:ATM

[Ste76a] John Steele. Application of theoretical models in ecology. Journal of Theoretical Biology, 63(2):443–451, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690045X.

Stein:1976:AWF

[Ste76b] W. D. Stein. An algorithm for writing down flux equations for carrier kinetics, and its application to co-transport. Journal of Theoretical Biology, 62(2):467–478, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901314.

Stein:1978:CGP

[Ste78] T. P. Stein. Cachexia, gluconeogenesis and progressive weight loss in cancer patients. Journal of Theoretical Biology, 73(1): 51–59, July 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901790.

Stillwell:1977:OP

[Sti77] William Stillwell. On the origin of photophosphorylation. Journal of Theoretical Biology, 65(3):479–497, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902090.

Storey:1970:CHE

[Sto70] Bayard T. Storey. Chemical hypothesis for energy conservation in the mitochondrial respiratory chain. Journal of Theoretical Biology, 28(2):233–259, August 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900548.

Storey:1971:CHE

[Sto71] Bayard T. Storey. Chemical hypothesis for energy conservation in the mitochondrial respiratory chain II. Redox properties of the coupling site enzymes at sites I and II. Journal of Theoretical Biology, 31(3):533–552, June 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900270.

Stokfs:1975:FDE

[Sto75] A. N. Stokfs. Facilitated diffusion: The elasticity of oxygen supply. Journal of Theoretical Biology, 52(2):285–297, August 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900028.

Stokes:1976:ROE

[Sto76] A. N. Stokes. Rapid oxygen entry into unperfused tissue slices: An experimental artefact? Journal of Theoretical Biology, 60(2):473–475, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900734.

Strunk:1971:HLN

[Str71] Thomas H. Strunk. Heat loss from a Newtonian animal. Journal of Theoretical Biology, 33(1):35–61, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902153.

Straub:1974:SST

[Str74] Karl David Straub. A solid state theory of oxidative phosphorylation. Journal of Theoretical Biology, 44(2):191–206, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901568.

Strandberg:1976:APG

[Str76] M. W. P. Strandberg. Action potential in the giant axon of loligo: a physical model. Journal of Theoretical Biology, 58(1): 33–53, May 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901375.

Stroeve:1977:DIC

[Str77] Pieter Stroeve. Diffusion with irreversible chemical reaction in heterogeneous media: Application to oxygen transport in respiring tissue. Journal of Theoretical Biology, 64(2):237–251, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937790354X.

Stuart:1978:SNL

[STU78] C. I. J. M. Stuart, Y. Takahashi, and H. Umezawa. On the stability and non-local properties of memory. Journal of Theoretical Biology, 71(4):605–618, April 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903272.

Subirana:1970:HMA

[Sub70] Juan A. Subirana. Hydrodynamic model of amoeboid movement. Journal of Theoretical Biology, 28(1):111–120, July 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900676.

Subrahmanyam:1977:MMR

[Sub77] N. C. Subrahmanyam. Mechanism of meiotic recombination in eukaryotes — a theory. Journal of Theoretical Biology, 64(4):739–745, February 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902727.

Schreiber:1976:PFO

[SUE76] Gerhard Schreiber, Jörg Urban, and Kaylene Edwards. Possible functions of the oligopeptide extension in the albumin precursor. Journal of Theoretical Biology, 60(1):241–245, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901685.

Suezaki:1978:SMA

[Sue78] Yukio Suezaki. Statistical mechanical analysis of interfacial tension of black lipid membrane. Journal of Theoretical Biology, 71(3):279–294, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901601.

Sugita:1975:FAC

[Sug75] Motoyosi Sugita. Functional analysis of chemical systems in vivo using a logical circuit equivalent: V. Molecular biological interpretation of the self-reproducing automata theory and chemico-physical interpretation of information in biological systems. Journal of Theoretical Biology, 53(1):223–237, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901137.

Summerton:1979:IIS

[Sum79a] James Summerton. Intracellular inactivation of specific nucleotide sequences: A general approach to the treatment of viral diseases and virally-mediated cancers. Journal of Theoretical Biology, 78(1):77–99, May 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903278.

Summerton:1979:SSC

[Sum79b] James Summerton. Sequence-specific crosslinking agents for nucleic acids: Design and functional group testing. Journal of Theoretical Biology, 78(1):61–75, May 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903266.

Sungchul:1977:PMM

[Sun77] J. I. Sungchul. A possible molecular mechanism of free energy transfer in oxidative phosphorylation. Journal of Theoretical Biology, 68(4):607–612, October 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901096.

Soucek:1975:BCS

[SV75] B. Soucek and F. Vencl. Bird communication study using digital computer. Journal of Theoretical Biology, 49(1):147–172, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800245.

Stein:1974:AMA

[SW74] R. B. Stein and E. Y.-M. Wong. Analysis of models for the activation and contraction of muscle. Journal of Theoretical Biology, 46(2):307–327, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900022.

Sparling:1978:MAA

[SW78] Donald W. Sparling and John D. Williams. Multivariate analysis of avian vocalizations. Journal of Theoretical Biology, 74(1):83–107, September 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902916.

Swanson:1974:ATS

[Swa74] C. J. Swanson. Application of thin shell theory to helicallywound fibrous cuticles. Journal of Theoretical Biology, 43 (2):293–304, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519374800615.

Shimizu:1975:SET

[SY75] H. Shimizu and T. Yamada. The synergetic enzyme theory of muscular contraction: A two-headed myosin model. Journal of Theoretical Biology, 49(1):89–109, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580021X.

Synder:1979:TBT

[Syn79] G. Synder. Thermodynamics and body temperature. Journal of Theoretical Biology, 80(1):145–147, September 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901851.

Shimizu:1976:SET

[SYNY76] H. Shimizu, T. Yamada, K. Nishiyama, and M. Yano. The synergetic enzyme theory of muscular contraction: II relation between Hill’s equations and functions of two-headed myosin. Journal of Theoretical Biology, 63(1):165–189, November 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900916.

Schotz:1971:ELS

[SZ71] W. E. Schotz and M. Zelen. Effect of length sampling bias on labeled mitotic index waves. Journal of Theoretical Biology, 32(2):383–404, August 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371901755.

Szabadi:1977:MTF

[Sza77] E. Szabadi. A model of two functionally antagonistic receptor populations activated by the same agonist. Journal of Theoretical Biology, 69(1):101–112, November 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903903.

Tagami:1978:PSM

[Tag78] Yukiko Tagami. A perm-selective model membrane exhibiting oscillatory behavior. Journal of Theoretical Biology, 73 (4):739–754, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378901340.

Tansky:1978:SEP

[Tan78] Mumay Tansky. Switching effect in prey-predator system. Journal of Theoretical Biology, 70(3):263–271, February 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903764.

Tarin:1972:TIM

[Tar72] David Tarin. Tissue interactions in morphogenesis, morphostasis and carcinogenesis. Journal of Theoretical Biology, 34(1):61– 72, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900549.

Taschdjian:1970:ALS

[Tas70] Edgar Taschdjian. Asymmetry of living systems. Journal of Theoretical Biology, 26(3):417–428, March 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900937.

Taylor:1972:SBN

[Tay72] J. G. Taylor. Spontaneous behaviour in neural networks. Journal of Theoretical Biology, 36(3):513–528, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900069.

Taylor:1976:AS

[Tay76] James Taylor. The advantage of spacing-out. Journal of Theoretical Biology, 59(2):485–490, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901843.

Taylor:1978:FBA

[Tay78] Fritz Taylor. Foraging behavior of ants: Theoretical considerations. Journal of Theoretical Biology, 71(4):541–565, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903247.

Taylor:1979:AMS

[Tay79] Peter D. Taylor. An analytical model for a short-term advantage for sex. Journal of Theoretical Biology, 81(3):407–421, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379900444.

Thompson:1972:GFH

[TB72] Colin J. Thompson and Robert Ter Bush. General formulation of Haldane’s laws for competing ligands. Journal of Theoretical Biology, 34(3):441–449, March 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901348.

Theuvenet:1976:ISC

[TBP76] A. P. R. Theuvenet and G. W. F. H. Borst-Pauwels. The influence of surface charge on the kinetics of ion-translocation across biological membranes. Journal of Theoretical Biology, 57(2): 313–329, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900047.

Testa:1978:RFM

[TBP78] Bernard Testa, Jean-Claude Bünzli, and William P. Purcell. Regioelectronic factors in metabolic hydroxylation of aliphatic carbon atoms. Journal of Theoretical Biology, 70(4):339–344, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902473.

Thompson:1979:PBM

[TC79] R. W. Thompson and D. A. Cauley. A population balance model for fish population dynamics. Journal of Theoretical Biology, 81(2):289–307, November 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990167X.

Trincher:1972:SLI

[TD72] K. S. Trincher and A. G. Dudoladov. Spin-lattice interaction of water and protein membranes in cell metabolism. Journal of Theoretical Biology, 34(3):557–576, March 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901415.

Teague:1976:IPC

[TD76a] Rodney B. Teague and Michael A. B. Deakin. Inversion polymorphism: Character and stability of equilibria. Journal of Theoretical Biology, 56(1):75–94, January 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937680046X.

Treisman:1976:CMT

[TD76b] Michel Treisman and Richard Dawkins. The “cost of meiosis”: is there any? Journal of Theoretical Biology, 63(2):479–484, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900473.

Thomson:1977:DED

[TD77a] Alan B. R. Thomson and John M. Dietschy. Derivation of the equations that describe the effects of unstirred water layers on the kinetic parameters of active transport processes in the intestine. Journal of Theoretical Biology, 64(2):277–294, January 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903575.

Troitsky:1977:QRF

[TD77b] N. A. Troitsky and S. E. Dromashko. On the question of the recombination frequencies of proximal markers in bacterial conjugation. Journal of Theoretical Biology, 67(4):671–676, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902533.

Thames:1976:ESO

[TE76] Howard D. Thames and Allen D. Elster. Equilibrium states and oscillations for localized two-enzyme kinetics: A model for circadian rhythms. Journal of Theoretical Biology, 59(2):415– 427, July 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901806.

Teague:1976:RSR

[Tea76] R. Teague. A result on the selection of recombination altering mechanisms. Journal of Theoretical Biology, 59(1):25–32, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800227.

Temussi:1975:ACS

[Tem75] P. A. Temussi. Automatic comparison of the sequences of calf thymus histones. Journal of Theoretical Biology, 50(1):25– 33, March 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900211.

Teodorescu:1973:HRI

[Teo73] M. Teodorescu. A hypothesis on the role of immune RNA in antibody variability. Journal of Theoretical Biology, 42 (2):189–218, November 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373900854.

Tiwari:1973:GRF

[TF73] Jawahar Tiwari and Alex Fraser. Genetic regulation by feedback repression. Journal of Theoretical Biology, 39(3):679–681, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900829.

Trubatch:1974:CPP

[TF74] S. L. Trubatch and A. Franco. Canonical procedures for population dynamics. Journal of Theoretical Biology, 48(2):299– 324, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800032.

Tiwari:1974:GFR

[TFB74] Jawahar Tiwari, Alex Fraser, and Richard Beckman. Genetical feedback repression I. Single locus models. Journal of Theoretical Biology, 45(2):311–326, June 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901179.

Thames:1973:DSI

[Tha73] Howard D. Thames. Dependence of stability on interenzyme distance in cooperative systems. Journal of Theoretical Biology, 41(2):331–345, September 21, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901227.

Thellier:1971:NET

[The71] M. Thellier. Non-equilibrium thermodynamics and electrokinetic interpretation of biological systems. Journal of Theoretical Biology, 31(3):389–393, June 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900178.

Thibault:1978:NAS

[Thi78] P. Thibault. A new attempt to study the oxygen evolving system of photosynthesis: Determination of transition probabilities of a state i. Journal of Theoretical Biology, 73(2):271–284, July 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890190X.

Thompson:1971:RAE

[Tho71] J. S. Thompson. Role of autolytic enzymes in the growth and morphogenesis of bacterial cell walls. Journal of Theoretical Biology, 33(1):63–75, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371902165.

Thomas:1973:NUC

[Tho73a] Hugh B. G. Thomas. “noise” as a unifying concept in the theory of schizophrenia. Journal of Theoretical Biology, 41(2):223–242, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937390115X.

Thomas:1973:BFG

[Tho73b] René Thomas. Boolean formalization of genetic control circuits. Journal of Theoretical Biology, 42(3):563–585, December 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902476.

Thoma:1974:SFR

[Tho74a] John A. Thoma. Separation of factors responsible for lysozyme catalysis. Journal of Theoretical Biology, 44(2):305–317, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901623.

Thomas:1974:BFG

[Tho74b] René Thomas. Boolean formalization of genetic control circuits. Journal of Theoretical Biology, 44(2):i, April 1974.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901726.

Thorne:1977:TAD

[Tho77a] M. C. Thorne. Theoretical apects of the distribution and retention of radionuclides in biological systems. Journal of Theoretical Biology, 65(4):743–754, April 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900200.

Thornley:1977:MAB

[Tho77b] J. H. M. Thornley. A model of apical bifurcation applicable to trees and other organisms. Journal of Theoretical Biology, 64(1):165–176, January 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901187.

Thomas:1978:LAS

[Tho78] R. Thomas. Logical analysis of systems comprising feedback loops. Journal of Theoretical Biology, 73(4):631–656, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901273.

Tomita:1977:SMB

[TIO77] Kazuhisa Tomita, Akira Ito, and Takao Ohta. Simplified model for Belousov-zhabotinsky reaction. Journal of Theoretical Biology, 68(4):459–481, October 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900984.

Tansey:1975:MMM

[TJ75] Michael R. Tansey and Michele A. Jack. Moonlight, mushrooms and moulds. Journal of Theoretical Biology, 51(2):403–407, June 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900697.

Tai:1974:OTH

[TkC74] Ronald C. Tai and Hsin kang Chang. Oxygen transport in heterogeneous tissue. Journal of Theoretical Biology, 43(2):265– 276, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800597.

Tawada:1974:AEM

[TKO74] K. Tawada, S. Kounosu, and F. Oosawa. An “active enzyme” muscle model. Journal of Theoretical Biology, 45(1):219–233, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900526.

Teramoto:1979:SEP

[TKS79] Ei Teramoto, Kohkichi Kawasaki, and Nanako Shigesada. Switching effect of predation on competitive prey species. Journal of Theoretical Biology, 79(3):303–315, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903485.

Thingstad:1974:DCC

[TL74] Tron Frede Thingstad and Tor I. Langeland. Dynamics of chemostat culture: The effect of a delay in cell response. Journal of Theoretical Biology, 48(1):149–159, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901866.

Tallarida:1976:PDR

[TLJ76] Ronald J. Tallarida, Oscar L. Laskin, and Leonard S. Jacob. Perturbation of drug receptor equilibrium in the presence of competitive blocking agents. Journal of Theoretical Biology, 61(1):211–219, September 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901156.

Thornton:1974:LSE

[TM74] Kent W. Thornton and Robert J. Mulholland. Lagrange stabffity and ecological system. Journal of Theoretical Biology, 45(2):

473–485, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937490126X.

Tewari:1974:QCS

[TNG74] Ravindra Tewari, Rabindra K. Nanda, and Girjesh Govil. Quantum chemical studies on the conformational structure of nucleic acids IV. Calculation of backbone structure by CNDO method. Journal of Theoretical Biology, 46(1):229–239, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901490.

Todd:1970:KFC

[Tod70] N. B. Todd. Karyotypic fissioning and canid phylogeny. Journal of Theoretical Biology, 26(3):445–480, March 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900962.

Torda:1970:PMI

[Tor70] Clara Torda. Proposed model for information processing, including initiation of memory storage. Journal of Theoretical Biology, 28(3):497–505, September 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900846.

Torre:1976:TSH

[Tor76] Vincent Torre. A theory of synchronization of heart pace-maker cells. Journal of Theoretical Biology, 61(1):55–71, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901041.

Tornheim:1979:OGP

[Tor79] Keith Tornheim. Oscillations of the glycolytic pathway and the purine nucleotide cycle. Journal of Theoretical Biology, 79(4):491–541, August 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379902406.

Travis:1979:DCS

[TP79] C. C. Travis and Wilfred M. Post. Dynamics and comparative statics of mutualistic communities. Journal of Theoretical Biology, 78(4):553–571, June 21, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901905.

Toivonen:1978:MCS

[TR78a] H. Toivonen and T. Rytömaa. Monte Carlo simulation of malignant growth. Examples of the behavior of chloroleukemia after specific perturbations. Journal of Theoretical Biology, 72(2):257–267, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378900929.

Tuckwell:1978:NIT

[TR78b] Henry C. Tuckwell and Wolfgang Richter. Neuronal interspike time distributions and the estimation of neurophysiological and neuroanatomical parameters. Journal of Theoretical Biology, 71(2):167–183, March 20, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902655.

Thakur:1979:GAI

[TR79] Ajit K. Thakur and David Rodbard. Graphical aids to interpretation of scatchard plots and dose-response curves. Journal of Theoretical Biology, 80(3):383–403, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901000.

Trehub:1975:FOM

[Tre75] Arnold Trehub. First-occupancy model for signal transfer in the brain. Journal of Theoretical Biology, 50(2):387–395, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900880.

Treisman:1976:ESR

[Tre76] Michel Treisman. The evolution of sexual reproduction: A model which assumes individual selection. Journal of Theoretical Biology, 60(2):421–431, August 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900680.

Trehub:1977:NMC

[Tre77] Arnold Trehub. Neuronal models for cognitive processes: Networks for learning, perception and imagination. Journal of Theoretical Biology, 65(1):141–169, March 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900819.

Trehub:1978:NMS

[Tre78] Arnold Trehub. Neuronal model for stereoscopic vision. Journal of Theoretical Biology, 71(4):479–484, IN1–IN2, 485–486, April 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903193.

Trehub:1979:ASR

[Tre79] Arnold Trehub. Associative sequential recall in a synaptic matrix. Journal of Theoretical Biology, 81(3):569–576, December 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900547.

Tributsch:1975:PEC

[Tri75] Helmut Tributsch. Parametric energy conversion — a possible, universal approach to bioenergetics in biological structures. Journal of Theoretical Biology, 52(1):17–56, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900387.

Theodoridis:1971:CRS

[TS71a] George C. Theodoridis and Lawrence Stark. Central role of solar information flow in pregenetic evolution. Journal of Theoretical Biology, 31(3):377–388, June 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900166.

Tsanev:1971:PMM

[TS71b] R. Tsanev and Bl. Sendov. Possible molecular mechanism for cell differentiation in multicellular organisms. Journal of Theoretical Biology, 30(2):337–393, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900592.

Tuljapurkar:1977:DEU

[TS77] S. D. Tuljapurkar and J. S. Semura. Dynamic equilibrium under periodic perturbations in simple ecosystem models. Journal of Theoretical Biology, 66(2):327–343, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901758.

Texter:1978:DPT

[TS78a] John Texter and John Clark Sutherland. The distribution of pyrimidine tracts in DNA can be calculated from nearest neighbor frequencies. Journal of Theoretical Biology, 75(3):263–269, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378903338.

Tyson:1978:NDP

[TS78b] J. Tyson and W. Sachsenmaier. Is nuclear division in Physarum controlled by a continuous limit cycle oscillator? Journal of Theoretical Biology, 73(4):723–738, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901339.

Tyson:1979:DMC

[TS79] J. Tyson and W. Sachsenmaier. Derepression as a model for control of the DNA-division cycle in eukaryotes. Journal of Theoretical Biology, 79(3):275–280, August 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903461.

Tapia:1978:EEH

[TSP78] O. Tapia, F. Sussman, and E. Poulain. Environmental effects on H-bond potentials: A SCRF MO CNDO/2 study of some model systems. Journal of Theoretical Biology, 71(1):49–72, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902138.

Tsukamoto:1974:VAI

[Tsu74] Yoshihiko Tsukamoto. The valve action of intensity receptor. Journal of Theoretical Biology, 47(2):253–268, October 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901969.

Tsukamoto:1979:ITG

[Tsu79] Yoshihiko Tsukamoto. An information theory of the genetic code. Journal of Theoretical Biology, 78(4):451–498, June 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901875.

Tawada:1975:COR

[TT75] Yuriko Tawada and Katsuhisa Tawada. Co-operative regulation mechanism of muscle contraction: Inter-tropomyosin cooperation model. Journal of Theoretical Biology, 50(2):269– 283, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519375900818.

Tuckwell:1977:SMA

[Tuc77] Henry C. Tuckwell. On stochastic models of the activity of single neurons. Journal of Theoretical Biology, 65(4):783–785, April 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900248.

Tuckwell:1979:STM

[Tuc79] Henry C. Tuckwell. Synaptic transmission in a model for stochastic neural activity. Journal of Theoretical Biology, 77 (1):65–81, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379901383.

Tullock:1971:BE

[Tul71] Gordon Tullock. Biological externalities. Journal of Theoretical Biology, 33(3):565–576, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://

www.sciencedirect.com/science/article/pii/002251937190097X.

Trubatch:1972:SII

[TV72] Janett Trubatch and A. Van Harreveld. Spread of iontophoretically injected ions in a tissue. Journal of Theoretical Biology, 36(2):355–358, IN3, 359–366, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290104X.

Thompson:1975:AMC

[TV75] William A. Thompson and Ilan Vertinsky. Application of Markov chains to analysis of a simulation of birds’ foraging. Journal of Theoretical Biology, 53(2):285–307, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800051.

Torbet:1976:XRD

[TW76] J. Torbet and M. H. F. Wilkins. X-ray diffraction studies of lecithin bilayers. Journal of Theoretical Biology, 62(2):447–458, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376901296.

Thames:1977:SVM

[TW77] H. D. Thames and R. A. White. State-vector models of the cell cycle 1. Parametrization and fits to labeled mitoses data. Journal of Theoretical Biology, 67(4):733–756, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902582.

Tyree:1970:SCG

[Tyr70] M. T. Tyree. The symplast concept a general theory of symplastic transport according to the thermodynamics of irreversible processes. Journal of Theoretical Biology, 26(2):181–214, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800121.

Tyson:1979:PES

[Tys79] John J. Tyson. Periodic enzyme synthesis: Reconsideration of the theory of oscillatory repression. Journal of Theoretical Biology, 80(1):27–38, September 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901772.

Ulanowicz:1971:TTH

[UF71] Robert E. Ulanowicz and George C. Frazier. Transport through hemoglobin solutions: A sensitivity analysis. Journal of Theoretical Biology, 32(1):185–190, July 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901457.

Urist:1973:PBM

[UI73] Marshall R. Urist and Hisashi Iwata. Preservation and biodegradation of the morphogenetic property of bone matrix. Journal of Theoretical Biology, 38(1):155–167, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902312.

Umeyama:1974:MOS

[UIN74] Hideaki Umeyama, Akira Imamura, and Chikayoshi Nagata. Molecular orbital study on the interaction between coenzymes and enzymes; NAD + or NADH and dehydrogenases. Journal of Theoretical Biology, 46(1):1–19, July 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901374.

Umeyama:1973:MOS

[UINH73] Hideaki Umeyama, Akira Imamura, Chikayoshi Nagata, and Manabu Hanano. A molecular orbital study on the enzymic reaction mechanism of α-chymotrypsm. Journal of Theoretical Biology, 41(3):485–502, October 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390057X.

Ulanowicz:1972:MEF

[Ula72] Robert E. Ulanowicz. Mass and energy flow in closed ecosystems. Journal of Theoretical Biology, 34(2):239–253, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901580.

Uttley:1970:INA

[Utt70] A. M. Uttley. The informon: A network for adaptive pattern recognition. Journal of Theoretical Biology, 27(1):31–67, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901281.

Uttley:1975:ICCa

[Utt75a] A. M. Uttley. The informon in classical conditioning. Journal of Theoretical Biology, 49(1):355–376, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800403.

Uttley:1975:ICCb

[Utt75b] A. M. Uttley. The informon in classical conditioning. Journal of Theoretical Biology, 49(2):355–376, February 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901782.

Uttley:1977:MSB

[Utt77a] A. M. Uttley. Methods of simulating the behaviour of granule cells in hippocampus based on informon theory. Journal of Theoretical Biology, 69(3):391–399, December 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901473.

Uttley:1977:SBG

[Utt77b] A. M. Uttley. Similarity of behaviour of granule cells in hippocampus to that predicted by informon theory. Journal of Theoretical Biology, 69(3):401–413, December 7, 1977.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901485.

Venkatesan:1971:IRM

[VAA71] Natarajan Venkatesan, Joseph C. Arcos, and Mary F. Argus. Induction and repression of microsomal drug-metabolizing enzymes by polycyclic hydrocarbons and phenobarbital: Theoretical models. Journal of Theoretical Biology, 33(3):517–537, December 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900944.

Villarreal:1976:TIM

[VAC76] Ernesto Villarreal, Ziya Akcasu, and Raymond P. Canale. A theory of interacting microbial populations: Multigroup approach. Journal of Theoretical Biology, 58(2):285–317, ???? 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937680121X.

Vaidhyanathan:1971:ICRb

[Vai71a] V. S. Vaidhyanathan. Influence of chemical reactions on fluxes and a theory of active and facilitated types of transport. Journal of Theoretical Biology, 33(1):1–27, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937190213X.

Vaidhyanathan:1971:ICRa

[Vai71b]  V. S. Vaidhyanathan.  Influence of chemical reactions

on fluxes of non-electrolytes across membranes. Journal of Theoretical Biology, 31(1):53–67, April 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901214.

VanDerSteen:1972:EEE

[Van72] W. J. Van Der Steen. Ecology, evolution and explanatory patterns in biology. Journal of Theoretical Biology, 36(3):593–616, September 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372900112.

Vandermeer:1973:RSL

[Van73] John H. Vandermeer. On the regional stabilization of locally unstable predator–prey relationships. Journal of Theoretical Biology, 41(1):161–170, September 14, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901963.

VanValen:1974:MSS

[Van74a] Leigh Van Valen. Multivariate structural statistics in natural history. Journal of Theoretical Biology, 45(1):235–247, May 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900538.

VanValen:1974:PSD

[Van74b] Leigh Van Valen. Predation and species diversity. Journal of Theoretical Biology, 44(1):19–21, March 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800263.

Vance:1976:RLR

[Van76] Dennis E. Vance. The rate-limiting reaction catalyzed by a multi-enzyme complex — the fatty acid synthetases. Journal of Theoretical Biology, 59(2):409–413, July 7, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690179X.

VanSickle:1977:ADP

[Van77] John VanSickle. Analysis of a distributed-parameter population model based on physiological age. Journal of Theoretical Biology, 64(3):571–586, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902892.

Vassent:1974:MRB

[Vas74] Gérard Vassent. Multi-random billing molecular interactions: A general kinetic model. Journal of Theoretical Biology, 44(2): 241–270, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901593.

Vergonet:1970:ESS

[VB70] G. Vergonet and H. J. C. Berendsen. Evaluation of the steady-state approximation in a system of coupled rate equations. Journal of Theoretical Biology, 28(2):155–173, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900494.

Vandermeer:1978:VMI

[VB78] John H. Vandermeer and Douglas H. Boucher. Varieties of mutualistic interaction in population models. Journal of Theoretical Biology, 74(4):549–558, October 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902412.

Vasilescu:1976:CAA

[VBCL76] D. Vasilescu, H. Broch, R. Cornillon, and J.-N. Lespinasse. Conformational analysis of aminoacyladenylates I. General backbone structure. Journal of Theoretical Biology, 57(2):395– 405, April 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900114.

VanCauter:1978:CIM

[VD78] E. Van Cauter and J. E. Dumont. Cross inhibition models for the transmission of hormonal signals. Journal of Theoretical Biology, 73(4):657–677, August 21, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901285.

vandenDriessche:1974:SLP

[vdD74] P. van den Driessche. Stability of linear population models. Journal of Theoretical Biology, 48(2):473–476, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800147.

vanDrongelen:1978:COS

[vDHD78] Wim van Drongelen, Andre Holley, and Kjell B. Døving. Convergence in the olfactory system: quantitative aspects of odour sensitivity. Journal of Theoretical Biology, 71(1):39–48, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902126.

vanderSteen:1971:BCA

[vdSJ71] W. J. van der Steen and J. C. Jager. Biology, causality and abstraction, with illustrations from a behavioural study of chemoreception. Journal of Theoretical Biology, 33(2):265–278, November 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251937190066X.

Venzl:1976:SFA

[Ven76] G. Venzl. Statistical fluctuations of activity in localized neural populations. Journal of Theoretical Biology, 63(2):259–274, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519376900345.

Vologodskii:1975:TSD

[VFK75] A. V. Vologodskii and M. D. Frank-Kamenetskii. Theoretical study of DNA unwinding under the action of formaldehyde. Journal of Theoretical Biology, 55(1):153–166, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801147.

Vologodskii:1974:TCO

[VFKL74] A. V. Vologodskii, M. D. Frank-Kamenetskii, and A. V. Lukashin. Theory of co-operative formation of defects in DNA molecule under ultraviolet irradiation. Journal of Theoretical Biology, 44(2):271–286, April 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937490160X.

Vaccaro:1979:IPE

[VG79] S. R. Vaccaro and H. S. Green. Ionic processes in excitable membranes. Journal of Theoretical Biology, 81(4):771–802, December 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379902819.

vonHeijne:1976:KAD

[vHB76] Gunnar von Heijne and Clas Blomberg. Kinetic aspects of the DNA helix-cruciform transition. Journal of Theoretical Biology, 63(2):347–353, December 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376900382.

vonHeijne:1979:CDE

[vHB79] Gunnar von Heijne and Clas Blomberg. The concentration dependence of the error frequencies and some related quantities in protein synthesis. Journal of Theoretical Biology, 78(1):113– 120, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903291.

Viale:1970:SDK

[Via70] Richard O. Viale. Similarities and differences in the kinetics of the Michaelis scheme and the Henri scheme. Journal of Theoretical Biology, 27(3):377–385, ???? 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800030.

Viale:1971:RTS

[Via71] Richard O. Viale. Resolution of two simple protein-ligand binding schemes with kinetic measurements. Journal of Theoretical Biology, 31(3):501–507, June 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371900257.

Vig:1975:CNM

[Vig75] B. K. Vig. Chromatin-nuclear membrane attachment in relation to DNA replication and chromosome aberrations: A new hypothesis. Journal of Theoretical Biology, 54(2):191–199, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937580124X.

Vine:1971:RVD

[Vin71] Ian Vine. Risk of visual detection and pursuit by a predator and the selective advantage of flocking behaviour. Journal of Theoretical Biology, 30(2):405–422, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900610.

Vine:1973:DPF

[Vin73] Ian Vine. Detection of prey flocks by predators. Journal of Theoretical Biology, 40(2):207–210, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901276.

Vine:1974:RGF

[Vin74] I. Vine. Reply to G. F. Estabrook and H. H. Robinson. Journal of Theoretical Biology, 47(1):249–250, September 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901131.

Visvader:1977:NAM

[Vis77] John Visvader. A note on apter machines. Journal of Theoretical Biology, 69(4):685–692, December 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903757.

VanHam:1978:RML

[VL78] P. Van Ham and I. Lasters. Reduction methods for logical control networks. Journal of Theoretical Biology, 72(2):269–281, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378900930.

VonHeijne:1978:ENF

[VLB78] Gunnar Von Heijne, Olof Leimar, and Clas Blomberg. On the emergence of new function in primitive proteins. Journal of Theoretical Biology, 75(2):167–180, November 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890228X.

VonHeijne:1977:TMR

[VNB77] Gunnar Von Heijne, Lennart Nilsson, and Clas Blomberg. Translation and messenger RNA secondary structure. Journal of Theoretical Biology, 68(3):321–329, October 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900637.

Vogel:1977:GHF

[Vog77] Richard I. Vogel. Gingival hyperplasia and folic acid deficiency from anticonvulsive drug therapy a theoretical relationship. Journal of Theoretical Biology, 67(2):269–278, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901990.

Volkenstein:1972:C

[Vol72a] M. V. Volkenstein. The conformon. Journal of Theoretical Biology, 34(1):193–195, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937290063X.

Vollmer:1972:TCS

[Vol72b] Robin T. Vollmer. Twin concordance: A set theoretic and probability theory approach. Journal of Theoretical Biology, 36(2):367–378, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901051.

Volkenstein:1979:MVI

[Vol79] M. V. Volkenstein. Mutations and the value of information. Journal of Theoretical Biology, 80(2):155–169, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902029.

Vener:1977:USC

[VP77] Kirt J. Vener and Donald M. Perrill. Urea and sickle cell anemia. Journal of Theoretical Biology, 66(3):457–460, June 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902958.

VanDruten:1978:RAA

[VPBD78] J. A. M. Van Druten, P. G. M. Peer, A. B. H. Bos, and W. W. De Jong. Reciprocal amino acid substitutions in the evolution of homologous peptides. Journal of Theoretical Biology, 73(3):549– 561, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378901583.

Varela:1977:CCM

[VRM77] Francisco J. Varela, J. Carter Rowley, and David T. Moran. The control of ciliary movements: An application of the cusp catastrophe. Journal of Theoretical Biology, 65(3):531–553, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902132.

Verhoff:1971:TAC

[VS71] F. H. Verhoff and F. E. Smith. Theoretical analysis of a conserved nutrient ecosystem. Journal of Theoretical Biology, 33(1):131–147, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519371902219.

Vogt:1979:CBT

[VSW^{+}79] H.-P. Vogt, W. Strassburger, A. Wollmer, J. Fleischhauer, and … Dan Mercola. Calcium binding by troponin-C and homologs is correlated with the position and linear density of “β-turn forming” residues. Journal of Theoretical Biology, 76(3):297–302, 303–310, February 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937990314X.

VanderSteen:1979:TSA

[VtM79] W. J. Van der Steen and A. ter Maat. Theoretical studies on animal orientation I. Methodological appraisal of classifications. Journal of Theoretical Biology, 79(2):223–234, July 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902479.

vWangenheim:1976:MEC

[vW76] K.-H. v. Wangenheim. A mechanism for the endocellular control of cell differentiation and cell proliferation. Journal of Theoretical Biology, 59(1):205–210, IN1, 211–222, June 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800331.

Wu:1978:BEC

[WA78] T. M. Wu and Steven Austin. Bose–Einstein condensation in biological systems. Journal of Theoretical Biology, 71(2):209– 214, March 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902679.

Walter:1970:OSL

[Wal70] Charles F. Walter. The occurrence and the significance of limit cycle behavior in controlled biochemical systems. Journal of Theoretical Biology, 27(2):259–272, May 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901414.

Wallach:1973:PPI

[Wal73a] Donald F. Hoelzl Wallach. Possible processes in the immune surveillance of the intracellular space. Journal of Theoretical Biology, 39(2):321–324, May 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937390101X.

Walz:1973:PMM

[Wal73b] Frederick G. Walz. A proposed mechanism for myosin magnesium adenosine triphosphatase activity. Journal of Theoretical Biology, 41(2):357–364, IN1–IN5, 365–373, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901240.

Walter:1974:VUQ

[Wal74a] Charles Walter. The validity of using a quasi-steady state approximation for the reversible Michaelis–Menten mechanism of enzyme action. Journal of Theoretical Biology, 44(1):1–5, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251937480024X.

Walter:1974:SDP

[Wal74b] Charles F. Walter. Some dynamic properties of linear, hyperbolic and sigmoidal multi-enzyme systems with feedback control. Journal of Theoretical Biology, 44(2):219–240, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901581.

Waller:1979:MMF

[Wal79] Donald M. Waller. Models of mast fruiting in trees. Journal of Theoretical Biology, 80(2):223–232, September 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379902078.

Warner:1970:AAH

[War70a] Donald T. Warner. Attempted application of the hexagonal conformation to proteins containing cystine and cysteine. Journal of Theoretical Biology, 27(3):393–416, ???? 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800054.

Warner:1970:PMM

[War70b] Donald T. Warner. A proposal for the mechanism of muscle contraction at the molecular level. Journal of Theoretical Biology, 26(2):289–306, IN1–IN4, 307–313, February 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800194.

Warner:1974:AAH

[War74] D. T. Warner. Additional applications of the hexagonal conformation to calcitonin, myoglobin and other proteins of known sequence. Journal of Theoretical Biology, 46(2):329–351, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900034.

Watson:1972:RET

[Wat72] M. R. Watson. Rate effectors and their role in metabolic control. Journal of Theoretical Biology, 36(1):195–202, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901877.

Wedler:1973:MEC

[WB73] Frederick C. Wedler and Paul D. Boyer. Mechanisms of enzyme control as probed by equilibrium exchange rates: Patterns of modifier effects with a two substrate, two product system. Journal of Theoretical Biology, 38(3):539–558, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902555.

Williams:1974:EQU

[WB74a] James D. Williams and John M. Bowen. Effects of quantal unit latency on statistics of Poisson and binomial neurotransmitter release mechanisms. Journal of Theoretical Biology, 43(1):151– 165, January 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800500.

Williams:1974:MCC

[WB74b] James D. Williams and John M. Bowen. Membrane capacitance and correction of end-plate potentials for bonlinear summation of quantal units. Journal of Theoretical Biology, 44(2): 349–351, April 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901659.

Warner:1975:ECC

[WB75] D. Warner and P. W. Borthwick. The efficient computation of conformational energy surfaces using SCF molecular orbital methods. Journal of Theoretical Biology, 54(2):389–390, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801408.

Waltman:1977:TMA

[WB77a] Paul Waltman and E. Butz. A threshold model of antigenantibody dynamics. Journal of Theoretical Biology, 65(3):499– 512, April 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902107.

Williams:1977:ICE

[WB77b] James D. Williams and John M. Bowen. Integral correction of end-plate potentials: A useful alternative to peak amplitude correction. Journal of Theoretical Biology, 67(2):313–317, July 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790203X.

Willard:1975:CRC

[WC75] William K. Willard and Donald S. Cherry. Comparative radiosensitivity in the class insecta. Journal of Theoretical Biology, 52(1):149–158, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375900466.

Williamson:1976:EAF

[WC76] John A. Williamson and Brian Charlesworth. The effect of age of founder on the probability of survival of a colony. Journal of Theoretical Biology, 56(1):175–190, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800513.

Weaver:1972:PSV

[WD72] Charles S. Weaver and Eugene Dong. A phase-sensitive vagal heart-rate control model. Journal of Theoretical Biology, 36(2):291–311, August 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900999.

Will:1973:RBL

[WD73a] H. Will and G. Damaschun. Relationship between ligand binding and conformational change of macromolecules in the twostate allosteric model. Journal of Theoretical Biology, 38(3):579– 585, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902580.

Woolley:1973:ADC

[WD73b] Warren H. Woolley and Andrew G. De Rocco. Asynchronous division in cell colonies: I. General considerations and a linear control model. Journal of Theoretical Biology, 39(1):73–89, April 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902063.

Wolf:1978:ABK

[WD78] Matthew B. Wolf and Edward C. Deland. The association between K^{+} and H^{+} distribution in skeletal muscle: Implication to linked transport. Journal of Theoretical Biology, 72(4):729– 741, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902801.

Wilby:1975:MGP

[WE75] O. K. Wilby and D. A. Ede. A model generating the pattern of cartilage skeletal elements in the embryonic chick limb. Journal of Theoretical Biology, 52(1):199–217, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590051X.

Weber:1970:PEM

[Web70] C. E. Weber. Potassium etiology for mesenchyme degeneration. Journal of Theoretical Biology, 29(2):327–328, November 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900251.

Webb:1974:SRA

[Web74] D. J. Webb. The statistics of relative abundance and diversity. Journal of Theoretical Biology, 43(2):277–291, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800603.

Weir:1971:TLI

[Wei71] B. S. Weir. Two locus inbreeding functions. Journal of Theoretical Biology, 30(2):423–425, February 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900622.

Weissbluth:1972:JTE

[Wei72] M. Weissbluth. The Jahn–Teller effect in hemoglobin. Journal of Theoretical Biology, 35(3):597–600, June 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290152X.

Weihs:1974:EAB

[Wei74] D. Weihs. Energetic advantages of burst swimming of fish. Journal of Theoretical Biology, 48(1):215–229, November 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374901921.

Welch:1977:FEC

[Wel77a] G. Rickey Welch. On the free energy “cost of transition” in intermediary metabolic processes and the evolution of cellular infrastructure. Journal of Theoretical Biology, 68(2):267–291, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377901655.

Weller:1977:FTT

[Wel77b] Malcolm Weller. A further theoretical treatment of the turnover of protein bound phosphate in the presence of both protein kinase and phosphatase activities. Journal of Theoretical Biology, 64(2):391–399, January 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903666.

Went:1977:CRT

[Wen77] Hans A. Went. Can a reverse transcriptase be involved in centriole duplication? Journal of Theoretical Biology, 68(1):95–100, September 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377902302.

Werber:1976:RMI

[Wer76] Moshe M. Werber. The role of metal ions in the mechanism of action of hydrolytic metalloenzymes: Carbonic anhydrase. Journal of Theoretical Biology, 60(1):51–58, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901557.

Westall:1973:EDS

[Wes73] Fred C. Westall. An explanation for the determination of “self” versus “non-self” proteins. Journal of Theoretical Biology, 38(1):139–141, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902294.

Westall:1974:MVA

[Wes74a] Fred C. Westall. A mechanism for viral activation of multiple sclerosis. Journal of Theoretical Biology, 48(2):469–471, ???? 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800135.

Westall:1974:PAS

[Wes74b] Fred C. Westall. A possible affect of serotonin on the basic protein of myelin. Journal of Theoretical Biology, 45(2):591– 592, June 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519374901337.

Westall:1974:SVN

[Wes74c] Fred C. Westall. “self” versus “non-self”: In situ proteins. Journal of Theoretical Biology, 46(2):579–580, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900174.

West:1975:TAR

[Wes75] J. D. West. A theoretical approach to the relation between patch size and clone size in chimaeric tissue. Journal of Theoretical Biology, 50(1):153–160, March 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900296.

Westall:1978:PCR

[Wes78] Fred C. Westall. Possible change of a recognition mechanism for determining encephalitogenicity. Journal of Theoretical Biology, 71(3):465–468, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937890173X.

Wever:1972:VST

[Wev72] R. Wever. Virtual synchronization towards the limits of the range of entrainment. Journal of Theoretical Biology, 36(1): 119–132, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901816.

Wilbur:1975:KMD

[WFW75] R. Wilbur, E. Robert Fernholz, and Catherine T. Wilson. Kinetic models of the drug-receptor interaction. Journal of Theoretical Biology, 52(2):335–341, August 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900053.

Wolpert:1970:SPM

[WG70] L. Wolpert and D. Gingell. Striping and the pattern of melanocyte cells in chimaeric mice. Journal of Theoretical Biology, 29(1):147–150, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901232.

Weiss:1972:SSR

[WH72] L. Weiss and J. P. Harlos. Some speculations on the rate of adhesion of cells to coverslips. Journal of Theoretical Biology, 37(1):169–179, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901233.

Weintraub:1973:CAS

[WH73a] H. J. R. Weintraub and A. J. Hopfinger. Conformational analysis of some phenethylamine molecules. Journal of Theoretical Biology, 41(1):53–75, September 14, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901896.

Wong:1973:SST

[WH73b] R. Wong and E. Harth. Stationary states and transients in neural populations. Journal of Theoretical Biology, 40(1):77– 106, July 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901665.

Wheldon:1975:MAN

[Whe75] T. E. Wheldon. Mitotic autoregulation of normal and abnormal cells: Alternative mechanisms for the derangement of growth control. Journal of Theoretical Biology, 53(2):421–433, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800130.

White:1971:TMC

[Whi71] D. C. S. White. Thermodynamics and muscle contraction: A reply to Banks and Vernon. Journal of Theoretical Biology, 32(3):631–632, September 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901603.

White:1978:UPP

[Whi78] R. Allen White. The use of population projection matrices in cell kinetics. Journal of Theoretical Biology, 74(1):49–67, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902898.

Whitehead:1979:SSS

[Whi79] E. P. Whitehead. The structure of steady-state enzyme kinetic equations: A graph-theoretical algorithm for obtaining conditions for reduction in degree by common-factor cancellation. Journal of Theoretical Biology, 80(3):355–381, October 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900997.

Wickler:1976:DSB

[Wic76] Wolfgang Wickler. Duetting songs in birds: Biological significance of stationary and non-stationary processes. Journal of Theoretical Biology, 60(2):493–497, August 7, 1976.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900783.

Wicken:1978:ITM

[Wic78] Jeffrey S. Wicken. Information transformations in molecular evolution. Journal of Theoretical Biology, 72(1):191–204, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900255.

Wicken:1979:GCE

[Wic79] Jeffrey S. Wicken. The generation of complexity in evolution: A thermodynamic and information-theoretical discussion. Journal of Theoretical Biology, 77(3):349–365, April 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379903618.

Wiegel:1979:HPP

[Wie79] F. W. Wiegel. Hydrodynamics of a permeable patch in the fluid membrane. Journal of Theoretical Biology, 77(2):189–193, March 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379903060.

Wiggins:1971:WSD

[Wig71] Philippa M. Wiggins. Water structure as a determinant of ion distribution in living tissue. Journal of Theoretical Biology, 32(1):131–146, July 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937190141X.

Wiggins:1972:IPS

[Wig72] Philippa M. Wiggins. Intracellular pH and the structure of cell water. Journal of Theoretical Biology, 37(2):363–371, November 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937290029X.

Wiggins:1973:IPD

[Wig73] Philippa M. Wiggins. Intracellular pH of deep and superficial muscle fibres. Journal of Theoretical Biology, 42(1):113–117, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901513.

Wiggins:1975:CFC

[Wig75] Philippa M. Wiggins. Cellular functions of a cell in a metastable equilibrium state. Journal of Theoretical Biology, 52(1):99–111, July 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900429.

Williams:1970:OTP

[Wil70a] John A. Williams. Origin of transmembrane potentials in nonexcitable cells. Journal of Theoretical Biology, 28(2):287–296, August 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370900561.

Williams:1970:DCE

[Wil70b] Mary B. Williams. Deducing the consequences of evolution: A mathematical model. Journal of Theoretical Biology, 29(3):343– 385, December 1970. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519370901037.

Williams:1975:DSS

[Wil75a] Mary B. Williams. Darwinian selection for self-limiting populations. Journal of Theoretical Biology, 55(2):415–430, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800907.

Wilson:1975:SMS

[Wil75b] Hugh R. Wilson. A synaptic model for spatial frequency adaptation. Journal of Theoretical Biology, 50(2):327–352, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900855.

Williams:1978:SSA

[Wil78] Mary B. Williams. Sexual selection, adaptation, and ornamental traits: The advantage of seeming fitter. Journal of Theoretical Biology, 72(3):377–383, June 6, 1978. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903077.

Winfree:1970:IVR

[Win70a] Arthur T. Winfree. Integrated view of resetting a circadian clock. Journal of Theoretical Biology, 28(3):327–374, September 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900755.

Winne:1970:FKW

[Win70b] D. Winne. Formal kinetics of water and solute absorption with regard to intestinal blood flow. Journal of Theoretical Biology, 27(1):1–18, April 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519370901268.

Winfree:1972:PCT

[Win72] Arthur T. Winfree. On the photosensitivity of the circadian time-sense in drosophila pseudoobscura. Journal of Theoretical Biology, 35(1):159–189, April 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372902007.

Winfree:1973:PPF

[Win73] Arthur T. Winfree. Polymorphic pattern formation in the fungus nectria. Journal of Theoretical Biology, 38(2):363–364, IN1–IN5, 365–382, February 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937390180X.

Winne:1975:IVC

[Win75] D. Winne. The influence of villous counter current exchange on intestinal absorption. Journal of Theoretical Biology, 53(1):145–176, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375901083.

Winkelhake:1976:DCM

[Win76] Jeffrey L. Winkelhake. A dynamic continuum model for molecular regulation of the humoral immune response. Journal of Theoretical Biology, 60(1):37–49, July 21, 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901545.

Wissel:1977:ASF

[Wis77] C. Wissel. On the advantage of the specialization of flowers on particular pollinator species. Journal of Theoretical Biology, 69(1):11–22, November 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790385X.

Witten:1978:FSL

[Wit78] Matthew Witten. Fitness and survival in logistic models. Journal of Theoretical Biology, 74(1):23–32, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902874.

Wang:1972:TLS

[WJ72] Chang-Yi Wang and Theodore L. Jahn. A theory for the locomotion of spirochetes. Journal of Theoretical Biology, 36(1): 53–60, July 1972. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519372901762.

Winet:1974:GPP

[WJ74] Howard Winet and Theodore L. Jahn. Geotaxis in protozoa I. A propulsion-gravity model for tetrahymena (ciliata). Journal of Theoretical Biology, 46(2):449–465, August 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900083.

Whisler:1978:EMI

[WJ78] John W. Whisler and Daniel Johnston. Epileptogenesis: A model for the involvement of slow membrane events and extracellular potassium. Journal of Theoretical Biology, 75 (3):271–288, December 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251937890334X.

Wheldon:1973:ECM

[WK73a] T. E. Wheldon and J. Kirk. An error cascade mechanism for tumour progression. Journal of Theoretical Biology, 42 (1):107–111, November 5, 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373901501.

Wheldon:1973:IDO

[WK73b] T. E. Wheldon and J. Kirk. Induction, differentiation and oncogenesis. Journal of Theoretical Biology, 41(2):261–268, September 21, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901185.

Warner:1975:MOM

[WK75] T. F. C. S. Warner and R. G. Krueger. The multicellular (as opposed to the monoclonal) origin of the murine myeloma cell. Journal of Theoretical Biology, 54(2):175–179, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801226.

Wilson:1978:ELD

[WK78] K. J. Wilson and D. L. Kilgore. The effects of location and design on the diffusion of respiratory gases in mammal burrows. Journal of Theoretical Biology, 71(1):73–101, March 7, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937890214X.

Wheldon:1973:MAG

[WKG73] T. E. Wheldon, J. Kirk, and W. M. Gray. Mitotic autoregulation, growth control and neoplasia. Journal of Theoretical Biology, 38(3):627–639, March 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902622.

White:1971:RDT

[WL71] Fred N. White and Robert C. Lasiewski. Rattlesnake denning: Theoretical considerations on winter temperatures. Journal of Theoretical Biology, 30(3):553–557, March 1971. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900087.

Wachman:1972:MAS

[WL72] Avraham Wachman and Alexander Levitzki. Mechanism for the assembly of spherical viruses: A geometrical approach. Journal of Theoretical Biology, 34(2):277–278, IN1, 279–287, February 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901610.

Wolfenden:1976:VPA

[WL76] Richard Wolfenden and Charles A. Lewis, Jr. A vapor phase analysis of the hydrophobic effect. Journal of Theoretical Biology, 59(1):231–235, June 1976. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800355.

Williams:1973:WRS

[WM73] George C. Williams and Jeffry B. Mitton. Why reproduce sexually? Journal of Theoretical Biology, 39(3):545–554, June 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373900672.

Wolkowski:1977:QSA

[WMHG77] Z. William Wolkowski, Dominique Moccatti, Françoise Heymans, and Jean-Jacques Godfroid. A quantitative structureactivity approach to chemoreception: Importance of lipophilic properties. Journal of Theoretical Biology, 66(1):181–193, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903198.

Weiss:1979:MTI

[WN79] Leonard Weiss and Shlomo Nir. On the mechanisms of transfilter induction of kidney tubules. Journal of Theoretical Biology, 78(1):11–20, May 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519379903217.

Welss:1975:LDI

[WNHS75] L. Welss, S. Nir, J. P. Harlos, and J. R. Subjeck. Longdistance interactions between Ehrlich ascites tumour cells. Journal of Theoretical Biology, 51(2):439–454, June 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375900739.

Wiegert:1971:TSA

[WO71] R. G. Wiegert and D. F. Owen. Trophic structure, available resources and population density in terrestrial vs. aquatic ecosystems. Journal of Theoretical Biology, 30(1):69–81, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900373.

Woese:1970:CRA

[Woe70] Carl R. Woese. Codon recognition: The allosteric ribosome hypothesis. Journal of Theoretical Biology, 26(1):83–88, January 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519370800339.

Woese:1971:EMC

[Woe71] Carl R. Woese. Evolution of macromolecular complexity. Journal of Theoretical Biology, 33(1):29–34, October 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371902141.

Woese:1973:RRG

[Woe73] C. R. Woese. The rotating ribosome: A gross mechanical model for translation. Journal of Theoretical Biology, 38(1):203–204, January 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373902361.

Wohlgemuth:1978:AIS

[Woh78] Andrew Wohlgemuth. Abstract immunogenetic systems. Journal of Theoretical Biology, 73(3):469–508, August 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901546.

Woolley:1970:MHC

[Woo70] Warren H. Woolley. Model for a homeostatic control system deriving from a hamiltonian. Journal of Theoretical Biology, 28(3):305–313, September 1970. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370900731.

Woolley:1972:NSL

[Woo72] Warren H. Woolley. A note on spatial localization and oscillatory cellular dynamics. Journal of Theoretical Biology, 37(1):205–206, October 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372901294.

Woodward:1978:DIU

[Woo78] David R. Woodward. Is the “difference index” useful for assessment of protein relatedness? Journal of Theoretical Biology, 72(4):743–749, June 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378902813.

Wolf:1971:FIE

[WP71] A. V. Wolf and Phoebe G. Prentiss. Fitness of the internal environment body fluids are potent buffers of concentrative properties such as osmotic pressure as well as pH. Journal of Theoretical Biology, 31(1):139–158, April 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371901275.

Wyatt:1972:SSB

[WP72] Philip J. Wyatt and David T. Phillips. Structure of single bacteria from light scattering. Journal of Theoretical Biology, 37(3):493–501, December 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519372900872.

Westhaus:1978:HBM

[WP78] Paul A. Westhaus and Herbert A. Pohl. Hydrogen bonds and masking in the sulfhydryl group in proteins. Journal of Theoretical Biology, 70(2):157–160, IN2, 161–169, January 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903442.

Wells:1977:KMG

[WPTF77] B. D. Wells, L. A. Parks, I. Tam, and J. R. Fisher. A kinetic model for glutamate dehydrogenase. Journal of Theoretical Biology, 66(1):81–93, May 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903137.

Wei:1977:CDT

[WR77] James Wei and Michael B. Russ. Convection and diffusion in tissues and tissue cultures. Journal of Theoretical Biology, 66(4):775–787, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519377902454.

Wright:1977:OMV

[Wri77] R. H. Wright. Odor and molecular vibration: Neural coding of olfactory information. Journal of Theoretical Biology, 64(3):473–474, IN1–IN2, 475–502, February 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902831.

White:1979:SVMb

[WRT79] R. Allen White, Shelley L. Rasmussen, and Howard D. Thames. State vector models of the cell cycle III: Continuous time cell cycle models. Journal of Theoretical Biology, 81(2):181–200, November 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519379901590.

Waterman:1978:SD

[WS78] M. S. Waterman and T. F. Smith. On the similarity of dendrograms. Journal of Theoretical Biology, 73(4):789–800, August 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901376.

Wolpert:1976:F

[WSB76] L. Wolpert, W. D. Stein, and E. A. Barnard. Frontispiece. Journal of Theoretical Biology, 62(2):i, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901181.

Wells:1976:INS

[WSF76] B. D. Wells, T. A. Stewart, and J. R. Fisher. Interpretation of nonlinear steady state enzyme kinetics-cyclic and mathematical properties of cooperative, second-site and random pathway models. Journal of Theoretical Biology, 60(1):209–221, July 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937690165X.

Warner:1979:APT

[WSL79] R. R. Warner, T. Strunk, and C. Lechene. Analysis of proximal tubule salt and water transport in standing droplets. Journal of Theoretical Biology, 77(4):453–471, April 21, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379900201.

Weinstein:1977:TMS

[WSMS77] Harel Weinstein, Shalom Srebrenik, Saul Maayani, and Mordechai Sokolovsky. A theoretical model study of the comparative effectiveness of atropine and scopolamine action in the central nervous system. Journal of Theoretical Biology, 64(2):295–309, January 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903587.

Waterman:1977:AET

[WSSB77] M. S. Waterman, T. F. Smith, M. Singh, and W. A. Beyer. Additive evolutionary trees. Journal of Theoretical Biology, 64(2):199–213, January 21, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903514.

Welch:1977:RWL

[WT77] William R. Welch and C. Richard Tracy. Respiratory water loss: A predictive model. Journal of Theoretical Biology, 65(2):253– 265, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377903241.

Walsh:1978:RSS

[WT78] John B. Walsh and Henry C. Tuckweli. Repetitive subthreshold synaptic excitation and transmitter depletion. Journal of Theoretical Biology, 70(4):467–469, February 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378902540.

White:1979:SVMa

[WT79] R. Allen White and H. D. Thames. State vector models of the cell cycle II: The first three moments of the transit time distribution. Journal of Theoretical Biology, 77(1):141–160, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901449.

Wilson:1975:RCR

[WTR75] H. R. Wilson, P. Tollin, and A. Rahman. RNA conformation in rod-shaped viruses: A possible molecular model for RNA in narcissus mosaic virus. Journal of Theoretical Biology, 53(2): 327–332, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800075.

Wu:1977:SEF

[Wu77] Lilian Shiao-Yen Wu. The stability of ecosystems — a finite-time approach. Journal of Theoretical Biology, 66(2):345–359, May 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790176X.

Wunder:1975:MEMa

[Wun75a] Bruce A. Wunder. A model for estimating metabolic rate of active or resting mammals. Journal of Theoretical Biology, 49(1): 345–354, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375800397.

Wunder:1975:MEMb

[Wun75b] Bruce A. Wunder. A model for estimating metabolic rate of active or resting mammals. Journal of Theoretical Biology, 49(2):345–354, February 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519375901770.

Wuilmart:1977:ESH

[WW77] Christian Wuilmart and Lode Wyns. An evolutionary scheme for the histones as derived from a study of internal repetitions and homologies among the different classes. Journal of Theoretical Biology, 65(2):231–252, March 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790323X.

Weiss:1975:DRT

[WY75] George H. Weiss and Stephen Yeandle. Distribution of response times in visual sense cells after weak stimuli. Journal of Theoretical Biology, 55(2):519–528, December 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800981.

Weiss:1972:HAS

[WZ72] Leonard Weiss and Robert Zeigel. Heterogeneity of anionic sites at the electrokinetic surfaces of fixed Ehrlich ascites tumor cells. Journal of Theoretical Biology, 34(1):21–27, January 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900513.

Yagil:1976:IEB

[Yag76] Gad Yagil. Isotope exchange in biochemical networks application of an electric circuit technique to a biological reaction system. Journal of Theoretical Biology, 61(1):73–80, September 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376901053.

Yager:1977:NMK

[Yag77] Paul Yager. A novel mechanism for the Na + –k + ATPase. Journal of Theoretical Biology, 66(1):1–4, IN1–IN3, 5–11, May 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903071.

Yamaguchi:1977:MOS

[YC77] Tamië Yamaguchi and Nagata Chikayoshi. Molecular orbital study on the effect of the binding of water to base pair. Journal of Theoretical Biology, 69(4):693–707, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377903769.

Ycas:1974:ESB

[Yca74] Martynas Ycas. On earlier states of the biochemical system. Journal of Theoretical Biology, 44(1):145–160, March 1974. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519374800354.

Ycas:1978:CTS

[YGJ78] Martynas Ycas, Narendra S. Goel, and Jacob W. Jacobsen. On the computation of the tertiary structure of globular proteins. Journal of Theoretical Biology, 72(3):443–457, June 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378903119.

Yamauchi:1971:SBI

[YK71] Akira Yamauchi and Hideo Kimizuka. Study of bi-ionic potentials. Journal of Theoretical Biology, 30(2):285–295, February 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900555.

Younkin:1978:IRR

[YM78] Burrows Younkin and Horace Martin. Insulin regulation of RNA synthesis. Journal of Theoretical Biology, 74(4):491–500, October 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902369.

Yang:1976:ALE

[YMM76] William Yang, Robert Minard, and Clifford Matthews. Azacyclopropenylidenimine: A low energy dimer of hydrogen cyanide. Journal of Theoretical Biology, 56(1):111–123, January 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519376800483.

Yockey:1974:AIT

[Yoc74] Hubert P. Yockey. An application of information theory to the central dogma and the sequence hypothesis. Journal of Theoretical Biology, 46(2):369–406, August 1974. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519374900058.

Yockey:1977:CPS

[Yoc77a] Hubert P. Yockey. A calculation of the probability of spontaneous biogenesis by information theory. Journal of Theoretical Biology, 67(3):377–398, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900443.

Yockey:1977:ICC

[Yoc77b] Hubert P. Yockey. On the information content of cytochrome c. Journal of Theoretical Biology, 67(3):345–376, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900431.

Yockey:1977:PWP

[Yoc77c] Hubert P. Yockey. A prescription which predicts functionally equivalent residues at given sites in protein sequences. Journal of Theoretical Biology, 67(3):337–343, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790042X.

Yockey:1978:CCD

[Yoc78] Hubert P. Yockey. Can the central dogma be derived from information theory? Journal of Theoretical Biology, 74(1):149–152, September 20, 1978. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519378902953.

Yockey:1979:DOG

[Yoc79] Hubert P. Yockey. Do overlapping genes violate molecular biology and the theory of evolution? Journal of Theoretical Biology, 80(1):21–26, September 7, 1979. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901760.

Yodzis:1978:ERT

[Yod78] P. Yodzis. Environmental randomness and the tenacity of equilibria. Journal of Theoretical Biology, 72(1):185–189, May 8, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378900243.

Yorke:1979:PMT

[Yor79] Ellen D. Yorke. A possible magnetic transducer in birds. Journal of Theoretical Biology, 77(1):101–105, March 7, 1979. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519379901401.

Yoshida:1977:TAM

[Yos77] Akihiko Yoshida. Theoretical approaches to the microphonics. Journal of Theoretical Biology, 68(2):215–224, September 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901606.

Young:1973:MTO

[You73] John H. Young. Membrane tubularization and osmotic swelling in mitochrondria. Journal of Theoretical Biology, 40(2):329–335, August 15, 1973. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519373901355.

Young:1974:LID

[You74] John H. Young. Light induced dipole layers in the thylakoid membrane. Journal of Theoretical Biology, 43(2):339– 350, February 1974. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519374800652.

Young:1976:CDV

[You76] J. Z. Young. Choice, determinism and value in the light of biological knowledge. Journal of Theoretical Biology, 62 (2):459–465, October 21, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519376901302.

Young:1978:DTP

[You78] David A. Young. On the diffusion theory of phyllotaxis. Journal of Theoretical Biology, 71(3):421–432, April 6, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519378901698.

Yamagata:1971:POH

[YS71a] Y. Yamagata and M. Sakamoto. Possible origin of the hot spot. Journal of Theoretical Biology, 30(1):215–217, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519371900476.

Yap:1971:AIM

[YS71b] W. T. Yap and H. A. Saroff. Application of the Ising model to hemoglobin. Journal of Theoretical Biology, 30(1):35–39, January 1971. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519371900348.

Yariv:1972:PRC

[YSK^{+}72] Joseph Yariv, Izchak Z. Steinberg, A. Joseph Kalb, Rachel Goldman, and Ephraim Katchalski. Permease as a rotatory carrier. Journal of Theoretical Biology, 35(3):459–465, June 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372901440.

Yu:1977:AFP

[Yu77] Richard Yu. The active form of phytochrome: A new hypothesis based on phytochrome pelletability studies. Journal of Theoretical Biology, 69(4):581–595, December 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937790368X.

Yakovlev:1977:KAI

[YZI77] Andrej Yu. Yakovlev, Aleksandr V. Zorin, and Nikita A. Isanin. The kinetic analysis of induced cell proliferation. Journal of Theoretical Biology, 64(1):1–25, January 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377901114.

Zahavi:1975:MSS

[Zah75] Amotz Zahavi. Mate selection — a selection for a handicap. Journal of Theoretical Biology, 53(1):205–214, September 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519375901113.

Zahavi:1977:CHF

[Zah77] Amotz Zahavi. The cost of honesty: Further remarks on the handicap principle. Journal of Theoretical Biology, 67(3):603– 605, August 7, 1977. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519377900613.

Zajicek:1977:IP

[Zaj77] Gershom Zajicek. The intestinal proliferon. Journal of Theoretical Biology, 67(3):515–521, August 7, 1977. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377900534.

Zamir:1976:OPA

[Zam76] M. Zamir. Optimality principles in arterial branching. Journal of Theoretical Biology, 62(1):227–251, October 7, 1976. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519376900588.

Zaritsky:1975:DDR

[Zar75] Arieh Zaritsky. On dimensional determination of rod-shaped bacteria. Journal of Theoretical Biology, 54(2):243–248, ???? 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375801299.

Zeigler:1977:PPP

[Zei77] Bernard P. Zeigler. Persistence and patchiness of predator–prey systems induced by discrete event population exchange mechanisms. Journal of Theoretical Biology, 67(4):687–713, August 22, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902557.

Zemlicka:1973:PRO

[Zem73] Jirí Zemlicka. A possible role of organic activators in the ribosome-catalyzed peptide and ester bond formation. Journal of Theoretical Biology, 39(1):241–248, April 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373902191.

Zervos:1975:VUS

[Zer75] C. Zervos. On the validity of using semilogarithmic plots to determine initial velocities of enzyme-catalyzed reactions. Journal of Theoretical Biology, 50(2):253–258, April 1975. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937590079X.

Zipser:1975:MCN

[ZG75]  David Zipser and Terrell Taylor Gibbs.  Mispair cor-

rection in nucleic acids with alternate forms. Journal of Theoretical Biology, 53(2):395–402, ???? 1975. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519375800117.

Zgierski:1975:MCC

[Zgi75] Marek Z. Zgierski. Mechanisms of conformon creation in αhelical structures. Journal of Theoretical Biology, 55(1):95–106, November 1975. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519375801093.

Zamir:1973:BFD

[ZR73]  M. Zamir and Margot R. Roach.  Blood flow down-

stream of a two-dimensional bifurcation. Journal of Theoretical Biology, 42(1):33–48, November 5, 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002251937390146X.

Zelman:1972:CFA

[ZS72] Allen Zelman and H. H. Shih. The constant field approximation: Numerical evaluation for monovalent ions migrating across a homogeneous membrane. Journal of Theoretical Biology, 37(2):373–383, November 1972. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900306.

Zucker:1973:TIS

[Zuc73] Robert S. Zucker. Theoretical implications of the size principle of motoneurone recruitment. Journal of Theoretical Biology, 38(3):587–596, March 1973. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519373902592.

Znojil:1977:TMR

[ZV77] V. Znojil and J. Vacha. A theoretical model of the red cell survival followed by means of random and cohort labeling. Journal of Theoretical Biology, 66(4):711–726, June 21, 1977. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519377902405.

Zeigler:1970:STA

[ZW70] Bernard P. Zeigler and Roger Weinberg. System theoretic analysis of models: Computer simulation of a living cell. Journal of Theoretical Biology, 29(1):35–56, October 1970. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519370901177.

Zotina:1972:TPT

[ZZ72] R. S. Zotina and A. I. Zotin. Towards a phenomenological theory of growth. Journal of Theoretical Biology, 35(2):213–225, May 1972. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519372900343.

Zhabotinsky:1973:APD

[ZZ73] A. M. Zhabotinsky and A. N. Zaikin. Autowave processes in a distributed chemical system. Journal of Theoretical Biology, 40(1):45–56, IN1–IN3, 57–61, July 1973. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022519373901641.